### v2

# Differentially Private Matchings

[…]
Differentially Private Matchings
Michael Dinitz Johns Hopkins University mdinitz@cs.jhu.edu George Z. Li Carnegie Mellon University gzli@andrew.cmu.edu Quanquan C. Liu Yale University quanquan.liu@yale.edu Felix Zhou Yale University felix.zhou@yale.edu
[…]
Motivated by the importance of computing matchings in sensitive graph data, we study differentially private algorithms for computing maximum matchings and $b$-matchings in general graphs. Our algorithms return implicit solutions for these problems. We provide a number of algorithms and lower bounds for these problems in different models and settings.
  • We first prove a lower bound showing that computing explicit solutions necessarily incurs large error, even if we try to obtain privacy by allowing ourselves to output non-edges.
  • We then consider implicit solutions, where at the end of the computation there is an ($\varepsilon$-differentially private) billboard and each node can determine its matched edge(s) based on what is written on this publicly visible billboard. For this solution concept, we provide tight upper and lower (bicriteria) bounds, where the degree bound is violated by a logarithmic factor, which we show is necessary via a novel lower bound.
  • We further show that our algorithm can be implemented in the local edge DP (LEDP) model, and can even be done in a logarithmic number of rounds if we further relax the degree bounds by logarithmic factors.
  • Moreover, we implement our algorithms in the node-DP setting by combining edge-DP algorithms with a novel use of arboricity sparsifiers. Interestingly, we prove an impossibility result for publicly releasing such sparsifiers under differential privacy, even though they form a key component of our node-DP algorithm. Our techniques also allow us to improve the bipartite results of [hsu2014private] by a polylogarithmic factor.
  • We demonstrate that all of our results can also be implemented in the continual release model under both edge- and node-privacy.
  • Finally, we extend our main algorithmic result to the $b$-matching setting, including an extension of our node-DP algorithm using a novel approximate $b$-matching arboricity sparsifier. Our sparsifier may be of independent interest even in the non-private graph setting.
1 Introduction
Maximum matching and its variants are central problems within graph algorithms, in both theory and practice. Theoretically, matching and its variants including maximal matching and $b$-matching have been studied in essentially every modern model of computation, including the sequential [edmonds1965maximum, HK71, MV80], streaming [feldman2024maximum, assadi2022semi, kapralov2013better], distributed [CHS09, Fischer20], dynamic [neiman2015simple, BDL21, GP13], and parallel [BFS12, FischerN18] models. In addition, a variety of lower bounds have also been proven [balliu2021lower, konrad2024unconditional].
Matching has been used in practice for many applications on sensitive data including ad allocation [mehta2013online], kidney exchange [BMW22], and online dating [wang2018privacy, yi2016practical]. For each of these applications, the user’s preferences and the existence of an edge between two individuals contain very sensitive information. So it is natural (and important) to study private versions of matching algorithms.
Differential privacy [DMNS06] is the gold standard for protecting the privacy of individuals. While differential privacy was originally introduced in the context of numerical and statistical databases, a growing number of works have placed its focus on sensitive graph data. Since many problems in graph algorithms are actually problems about graph analytics over graphs containing sensitive personal information, there has been a large amount of important work on differentially private graph algorithms. This includes differentially private versions of graph problems like triangle counting, shortest paths, graph cuts, densest subgraph, $k$-core decompositions, and many others [NRS07, GuptaLMRT10, BGM22, DMN23, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, RS16, raskhodnikova2016differentially, Upadhyay13]. In the graph setting, differential privacy is generally defined in terms of edge-neighboring graphs or node-neighboring graphs where the graphs differ in one edge or all edges incident to any one node, respectively; this intuitively corresponds to requiring that when we run our algorithm on neighboring graphs, the distribution of outputs is very similar (see Section 2 for formal definitions).
[…]
Thus, very natural and important questions remain: is it possible to give differentially private algorithms for maximum (or maximal) matchings and $b$-matchings in general graphs beyond bipartite graphs? In this paper, we introduce the study of privately computing matchings and $b$-matchings in general graphs. We give the first algorithms and lower bounds for this question in a variety of models including both edge- and node-DP, local privacy, and continual release via a variety of novel techniques, some of which may be of independent interest in both the privacy and non-private graph algorithms communities. As a corollary of our techniques, we improve the bounds of [hsu2014private] for the node-private bipartite setting.
1.1 Our Contributions
We now state our results and our main technical contributions in the various settings mentioned above.
1.1.1 Solution Types
Explicit Solutions.
It is easy to see that given an input graph, it is impossible to differentially privately output any subgraph (including a matching) containing only true edges; doing so would obviously reveal the existence of any edge that we output. So we must turn to other ways to output a matching that will not result in privacy violations.
One natural approach is to allow our algorithm to output “edges” that are not true edges; in other words, our matching algorithm returns a matching where only some of the edges actually exist. This corresponds to changing the input graph $G=(V,E)$ to a weighted complete graph, where every $\{u,v\}\not\in E$ is assigned weight 00 in the complete graph and actual edges have weight $1$, and asking for a max-weight matching in this graph. We can even allow our algorithm to be bicriteria and return a $b$-matching [^1] Recall that in a $b$-matching every node has degree at most $b$, thus a traditional matching is a $1$-matching. rather than a true matching. In this case, we would want our returned $b$-matching to contain an (approximate) maximum matching, while minimizing the magnitude of $b$.
As our first result, we show a strong lower bound in this model, even if we allow ourselves to output a $b$-matching and only compare its weight to the maximum matching. More carefully, we define a $(\gamma,\beta)$-approximation to be an algorithm that outputs a $b$-matching whose total weight is at least $(\text{OPT}/\gamma)-\beta$, where OPT is the size of the maximum matching in $G$.
Theorem 1.1 (Lower Bound on Explicit Solutions).
Let $\mathcal{A}$ be an algorithm which satisfies $(\varepsilon,\delta)$-edge DP and always outputs a graph $H$ with maximum degree at most $b$. If $\mathcal{A}$ outputs a $(\gamma,\beta)$-approximation (even just in expectation) of the maximum matching of a $0/1$-weighted (Definition 2.7) input graph $G=(V,E)$, then
$\gamma(e^{2\varepsilon}b+\delta n)+\beta\geq n/2.$
The proof of Theorem 1.1 is presented in Section 3. To interpret this theorem, think of $\delta\leq 1/n$ (so the $\delta$ term does not contribute anything meaningful) and small constant $\varepsilon$. Then a few meaningful regimes of $b,\gamma,\beta$ include the following:
  • If $\gamma=1$ (so we are bounding additive loss), then $\beta\geq n/2-O(b)$. So if $b$ is at most $n/c$ for some reasonably large constant $c$, then the additive loss is still linear even though we are allowing our solution to have a quadratic number of total edges!
  • If $\beta=0$ (so we are bounding multiplicative loss), then we get that $\gamma\geq\Omega(n/b)$.
  • If $\gamma\leq\frac{n}{cb}$ for some reasonably large constant $c$, then $\beta\geq\Omega(n)$. Thus we have linear additive loss even if we allow multiplicative loss of almost $n/b$.
We can actually strengthen this into a lower bound even on fractional matchings. A fractional $b$-matching is an assignment of values in $[0,1]$ to edges (i.e., a vector $X\in[0,1]^{\binom{V}{2}}$) so that $\sum_{v\in V}X_{\{u,v\}}\leq b$. As the proof of the following theorem proceeds similarly to to lower bound for explicit integral solutions, we defer it to Appendix B.
Theorem 1.2 (Lower Bound on Explicit Fractional Solutions).
Let $\mathcal{A}$ be an algorithm which satisfies $(\varepsilon,\delta)$-edge DP and always outputs a fractional $b$-matching $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$. If $\mathcal{A}$ outputs a $(\gamma,\beta)$-approximation (even just in expectation) of the maximum matching of a $0/1$-weighted (Definition 2.7) input graph $G=(V,E)$, then
$\gamma(e^{2\varepsilon}b+\delta n)+\beta\geq n/2.$
Implicit Solutions.
Theorem 1.1 suggests that we must find another way to return a subset of edges representing a matching. A natural approach, which we adopt through the rest of this paper, is to use “implicit” solutions in the sense of [GuptaLMRT10], which was later formalized by [hsu2014private] as the billboard model. In this model there is a billboard of public information released by the coordinator or the nodes themselves. Using the billboard of public information and its own privately stored information, each node can privately determine which of its incident edges are in the solution (if any). So in this model, we output something (to the billboard) that is “locally decodable” at every node (where each node can also use its private information in the decoding). That is, the information released to the billboard is $\varepsilon$-differentially private and each node computes its own part of the solution by combining the public information on the billboard with its own private adjacency list. Henceforth, we refer to the information on the billboard as an “implicit solution”. We formally define this implicit solution representation in Definition 2.4; such a solution has been used previously in the context of vertex cover, set cover, and maximum matching in bipartite graphs [GuptaLMRT10, hsu2014private]. It is important to note that the billboard model is not a privacy model, but a solution release model.
1.1.2 Edge Privacy (LEDP)
Our results for edge privacy are within the novel setting that is a combination of the billboard model of implicit solutions with a strong local version of privacy known as local edge differential privacy (LEDP) [IMC21locally, DLRSSY22]. In the LEDP model, nodes do not reveal their private information to anyone. Rather, nodes communicate with the curator over (possibly many) rounds, where in each round each node releases differentially private outputs which are accumulated in a global transcript. The transcript is publicly visible and must also be differentially private. It is very natural to combine the LEDP and billboard models by thinking of the billboard as this differentially private transcript, which is the approach we take.
Although both [hsu2014private] and our results are in the billboard model for solution release, their algorithm hinges on the “binary mechanism” [CSS11], a mechanism for releasing a stream of $n$ private counting queries with only $O(\log n)$ error, rather than the $O(\operatorname{poly}(n))$ error attained through naive composition. This mechanism releases noisy partial sums such that each query can be answered by combining $O(\log n)$ partial sums. However, this requires a centralized curator who can compute the true partial sums before adding noise, which seems incompatible with the local privacy model.
In comparison, we can guarantee LEDP rather than centralized differential privacy, and hence our solutions are in a much stronger privacy model (although only for edge-privacy rather than for node-privacy). We also later provide solutions for node-DP in the centralized setting.
As with our lower bound in Theorem 1.1, we will allow our algorithms to be bicriteria and output matchings that have up to $b$-degree. Interestingly, we present three main results for matchings depending on the size of $b$.
Our solutions are additionally guaranteed to (approximate) maximum matchings, with high probability. Given a large enough gap between $b$ and $b^{\prime}$, we also show that our solutions are guaranteed to (approximate) maximum $b^{\prime}$-matchings. When $b$ is sufficiently large, our returned matching will be a $(2+\eta)$-approximation of the maximum $b$-matching, and would no longer need to be a bicriteria approximation.
Our first result is the following, which we show in Section 4.
Theorem 1.3 (LEDP $2$-Approximate Maximum Matching).
For $\varepsilon\in(0,1)$, Algorithm 2 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$, in the billboard model, for $b=O(\log(n)/\varepsilon)$ that has the size of a $2$-approximate maximum matching.
We further extend our result to $b^{\prime}$-matchings where given a sufficient gap between $b$ and $b^{\prime}$, we can obtain an $\varepsilon$-LEDP algorithm that outputs a matching with degree at most $b$ that is a $2$-approximate maximum $b^{\prime}$-matching. We present these results in Section 10.
Theorem 1.4 (LEDP $2$-Approximate Maximum $b^{\prime}$-Matching).
For $\varepsilon\in(0,1)$ and $\eta\in(0,1)$, Algorithm 9 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$, in the billboard model, for $b\geq\frac{(1+\eta)^{2}}{1-\eta}\cdot b^{\prime}+O\left(\frac{\log(n)}{\eta^{2}\varepsilon}\right)$ that has the size of a $2$-approximate maximum $b^{\prime}$-matching.
Finally, when $b$ is sufficiently large, we no longer require a bicriteria approximation; instead, we are guaranteed to return an approximate maximum $b$-matching, with high probability. We present these results in detail in Section 10.
Theorem 1.5 (LEDP $2$-Approximate Maximum $b$-Matching).
For $\varepsilon\in(0,1)$ and $\eta\in(0,1/2)$, Algorithm 9 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) $b$-matching, in the billboard model, for $b=\Omega\left(\frac{\ln(n)}{\eta^{3}\varepsilon}\right)$ that has the size of a $(2+\eta)$-approximate maximum $b$-matching.
One downside of the algorithm used to prove Theorem 1.3 is its running time: it takes $\Omega(n)$ rounds. This is not ideal in the LEDP setting, where we have to wait for every node to communicate with the curator in each round before we can move onto the next, causing each round of the algorithm to be quite slow if there are straggler nodes. So we next show (in Section 5) an improved version that runs in only $O(\log n)$ rounds, based on combining the ideas of Theorem 1.3 with ideas from parallel and distributed algorithms. The price we pay for this efficiency is some extra loss in the $b$ parameter.
Theorem 1.6 (Efficient 2-Approximate Maximum Matching).
For $\varepsilon\in(0,1)$, Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) matching with degree at most $b$ in the billboard model for $b=O(\log^{2}(n)/\varepsilon)$ that has the size of a $2$-approximate maximum matching.
We complement our upper bounds with a lower bound which shows that even for the particular type of implicit solutions used in Theorem 1.3, our reliance on $b$ is necessary: if $b$ is sublogarithmic, then even an implicit solution cannot be a good approximation to the maximum matching in the graph. The following theorem is an informal version of this; see Section 11.1 and Theorem 11.1 for the precise statement.
Theorem 1.7 (Informal Lower Bound on Implicit Solutions; See Theorem 11.1).
Let $\mathcal{A}$ be an algorithm which satisfies $\varepsilon$-edge DP and outputs an implicit solution of the type used in Theorem 1.3 which is the size of a $\gamma$-approximate maximal matching [^2] Given a graph $G=(V,E)$, a $b$-matching $F$ is said to be a $\gamma$-approximate maximal matching if there exists a maximal matching with size $M$ such that $F$ has size at least $M/\gamma$. If there is a maximal $b$-matching of size $M$ where $F$ has size at least $M/\gamma$, then we say that $F$ is a $\gamma$-approximate $b$-matching. with probability at least $1-1/\mathrm{poly}(n)$. Then, when decoded from the billboard, with probability at least $1/2$, there is a node matched with at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes, regardless of the value of $\gamma>0$.
Interestingly, the value of $\eta$ does not actually affect this lower bound; a logarithmic loss in the degree is necessary for any multiplicative approximation.
1.1.3 Node Privacy
In the node differential privacy setting, two graphs are considered neighbors if one can be obtained from the other by replacing all edges incident on any one node. It is a significantly more challenging setting than edge privacy. We show in Section 6 that the ideas we developed for edge-DP in general graphs can be combined with a “matching sparsifier” [solomon2021local] to give the first implicit node-DP algorithms for matching in general graphs. Specifically, we show the following upper bound results in node-DP within Section 6, where $\alpha$ is the arboricity [^3] Recall that a graph has arboricity $\alpha$ if it can be partitioned into $\alpha$ edge-disjoint forests. of the input graph.
Theorem 1.8 (Node-DP Approximate Maximum Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}\right)$, and (ii) the implicit solution has the size of the $(2+\eta)$-approximate maximum matching.
Our node-DP algorithms rely on a stable implementation of a bounded arboricity sparsifier that sparsifies the graph (in a manner that doesn’t increase edge edit distance between neighboring graphs) such that each node’s degree is upper bounded (approximately) by the arboricity. Additional care is taken to ensure we use the privatized arboricity in our sparsifier.
We can alternatively present our approximation guarantees in terms of $b$-matchings, which are detailed in Section 10.
Theorem 1.9 (Node-DP Approximate Maximum $b$-Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that, with high probability, outputs an (implicit) $b$-matching in the billboard model, for $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\Omega\left(\frac{\alpha\log(n)}{\eta^{3}\varepsilon}+\frac{\log^{2}(n)}{\eta^{3}\varepsilon^{3}}+\frac{b^{\prime}\log(n)}{\eta^{2}\varepsilon}\right)$, that has the size of a $\left(2+\eta\right)$-approximate maximum $b^{\prime}$-matching.
Our techniques also lead to improved results for the setting of [hsu2014private]: bipartite graphs with node differential privacy in the billboard model. In this model, they view the nodes in the “left side” of the bipartition as goods and the “right side” nodes of the graph as bidders for these goods. Two such bipartite graphs are neighboring if one can be obtained from the other by adding/deleting any one bidder (and all its incident edges).
[hsu2014private] show guarantees of the following form. Suppose each item has supply $s$, meaning that each item can be matched with $s$ bidders (let us call these $s$-matchings); if $s=\Omega(\log^{3.5}{n}/\varepsilon)$, then they give an algorithm which (implicitly) outputs a $(1+\eta)$-maximum matching while guaranteeing differential privacy. [^4] They also consider the weighted version of the problem and further generalizations where each bidder has gross substitutes valuations. Our techniques also extend to these cases with improved results, but we limit our discussion here to matchings as that is the focus of this work. They also show that $s=\Omega(1/\sqrt{\eta})$ is necessary to obtain any guarantees, leaving open the question of tightening these bounds. We substantially tighten the upper bound in Section 7, showing that $s=\Theta(\log{n}/\varepsilon)$ suffices:
Theorem 1.10.
If the supply is at least $s\geq\Omega\left(\frac{\log(n)}{\eta^{4}\varepsilon}\right)$, Algorithm 5 outputs a $(1+\eta)$-approximate maximum (one-sided) $s$-matching with high probability.
1.1.4 Continual Release
Finally, all of our results translate to the continual release [DNPR10, CSS11] setting. In the graph continual release setting, edges are given as updates to the graph in a stream and the algorithm releases an output after each update in the stream. The continual release model requires the entire vector of outputs of the algorithm to be $\varepsilon$-differentially private. In this work, we focus on edge-insertion streams where updates contain edge insertions and edge-neighboring streams differ by one edge update.
We give results for two different types of streams in the continual release model. The first type of stream we consider is the arbitrary edge-order stream where the edges in the stream appear one by one in arbitrary order. In particular, in arbitrary edge-order streams, we show the following results for edge- and node-DP in Section 8. In our results below, we give bicriteria approximate solutions where an $(\psi,\phi)$-approximation is one with a $\psi$-multiplicative error and $\phi$-additive error. Our approximations hold, with high probability, for every update.
Theorem 1.11.
Let $\eta\in(0,1]$ and $\varepsilon\in(0,1)$. Algorithm 6 is an $\varepsilon$-edge DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution that has the size of a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
Theorem 1.12.
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{3}\varepsilon^{2}}\right)$and (ii) each implicit solution has the size of a
$\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta^{2}\varepsilon}\right)\right)$-approximate maximum matching.
The second type of stream we consider is the arbitrary adjacency-list order model [kallaugher2019complexity, goldreich2009proximity, czumaj2016relating, mcgregor2016better] in which nodes arrive in arbitrary order and once a node $v$ arrives as an update in the stream all edges adjacent to $v$ arrive in an arbitrary order as edge updates immediately after the arrival of $v$. In this model, we give the following results within Section 9.
Theorem 1.13.
For $\varepsilon\in(0,1)$ and $b=O(\log(n)/\varepsilon)$, Algorithm 8 is an $\varepsilon$-edge DP algorithm in the arbitrary adjacency-list continual release model that outputs, with high probability, an implicit $b$-matching that has the size of a $(2,1)$-approximate maximum matching.
We summarize all of our results in Table 1.
Model $b$ Approximation Bound Type Reference
$(\varepsilon,\delta)$-edge DP $O(1)$ $\Omega(n/b)$ Lower Bound Theorem 1.1
(explicit solutions)
$\varepsilon$-edge DP $\Omega(\log(n)/\varepsilon)$ $(2+\eta)$ Lower Bound Theorem 1.7
$\varepsilon$-LEDP ($n$ rounds) $O(\log(n)/\varepsilon)$ $2,(2+\eta)$ Upper Bound Theorem 1.3
Theorem 1.4
Theorem 1.5
$\varepsilon$-LEDP ($O(\log n)$ rounds) $O(\log^{2}(n)/\varepsilon)$ $2$ Upper Bound Theorem 1.6
$\varepsilon$-node DP $O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+{\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}}\right)$ $\left(2+\eta\right)$ Upper Bound Theorem 1.8
$\varepsilon$-node DP (bipartite) $s=\Omega\left(\frac{\log(n)}{{\eta^{4}}\varepsilon}\right)$ $(1+\eta)$, $s$-matching Upper Bound Theorem 1.10
Edge-Order $\varepsilon$-Edge DP $O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Theorem 1.11
Continual Release
Edge-Order $\varepsilon$-Node DP $O\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+{\frac{\log^{3}(n)}{\eta^{3}\varepsilon^{2}}}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Theorem 1.12
Continual Release
Adj-List Edge $O(\log(n)/\varepsilon)$ $(2,1)$ Upper Bound Theorem 1.13
Continual Release
Table 1: Summary of results on differentially private matchings. All approximations are given in terms of the optimum maximum matching in the input graph. For our node-DP results, $\alpha$ denotes the arboricity of the graph. All results except the first row are with respect to implicit solutions while the first one is a lower bound for explicit solutions.
1.2 Technical Contributions of Independent Interest.
Our algorithms not only use ideas from parallel and distributed algorithms but also use several novel algorithmic tools, which may be of independent interest outside of the scope of this paper. In this section, we describe such novel contributions. In Section 1.3, we describe our contributions with respect to differentially private matchings.
Our lower bound proof relies on a symmetrization argument that enables us to prove the lower bound for a restricted class of highly symmetric algorithms and extend the hardness to all algorithms. To the best of our knowledge, this is the first application of symmetrization arguments towards DP lower bounds and may be of use outside of private matchings.
Our Public Vertex Subset Mechanism (Algorithm 4) allows us to privately release a set of candidate vertices per node. Each node can then intersect the released set with its private adjacency list to produce implicit vertex subset solutions. Our mechanism is an independent contribution that may be useful for other implicit graph problems beyond matching.
In order to implement our algorithms in the node-DP setting, we make novel use of matching sparsifiers for bounded-arboricity graphs [solomon2021local] by providing a “stable” implementation in the sense of [wagaman2024time] that allows us to apply our edge-DP algorithms in a black-box fashion on the sparsified input, while incurring a factor of the arboricity $\alpha$ in the approximation guarantees. While sparsification has been studied in the context of node privacy [wagaman2024time], previous utility guarantees typically depend on the maximum degree, which can be significantly bigger than the arboricity. [^5] The star graph on $n$ vertices has maximum degree $n-1$ but arboricity $\alpha=1$. We also prove a novel extension of the sparsifier to $b$-matchings, which may be of interest beyond private graph algorithms to the non-private graph algorithm setting. In general, bounded arboricity sparsifiers may be of use in node-DP beyond matchings.
1.3 Technical Overview
We now give an overview for some of the main ideas behind our results.
Explicit Solution Lower Bound. (Section 3)
At a very high level, our lower bound for explicit solutions (Theorem 1.1) uses an argument similar to the “packing arguments” of [DMN23] for minimum cut. However, instead of using a packing/averaging argument, we use a symmetry argument reminiscent of symmetry arguments in distributed computing (e.g. [kuhn2016local]) and semidefinite programming [dam2015semidefinite]. Namely, we argue that any DP algorithm for a collection of input graphs must be highly symmetric. Highly symmetric is defined in terms of the marginal probability of each non-existing edge; we show that the marginal probability of choosing any of these non-existent edges is identical. Thus, any DP algorithm that gives a solution with decent utility cannot satisfy differential privacy. And so, any differentially private algorithm must give a poor approximation on at least one of the graphs in our collection of input graphs. Our paper is, to the best of our knowledge, the first use of this type of symmetry argument in DP lower bounds.
1.3.1 Edge Privacy
LEDP algorithms. (Sections 4 and 5)
Our main algorithms (Theorem 1.3 and Theorem 1.6) take inspiration from greedy algorithm and distributed algorithms for maximal matching, although our proofs also hold for $b$-matchings. In particular, the traditional greedy algorithm for maximal matching iterates through the nodes in the graph in an arbitrary order and matches the nodes greedily whenever there exists an unmatched node in the current iteration. However, this algorithm is clearly not private because each matched edge reveals the existence of a true edge in the graph. Thus, we must have some way of releasing public subsets of vertices that include both true edges and non-edges.
Instead, we introduce a novel DP proposal/receiver mechanism that allows nodes to propose a match set and receivers to accept the matches. We detail this proposal/receiver mechanism in Algorithm 2 and the guarantees are given in Theorem 1.3. Algorithm 2 proceeds in $n$ rounds using an arbitrary ordering of the nodes in the graph. We give as input a threshold $b$ which is a threshold for the number of edges in the matching each vertex is incident to. In the order provided by the ordering, each node $v$ privately checks using the Multidimensional AboveThreshold (MAT) technique introduced in [DLL23] whether it has reached its number of matched neighbors threshold. If it has not, then it asks all of its neighbors ordered after it whether they will match with it. This is the proposal phase of the algorithm. Each of these neighbors first determines using MAT whether they have exceeded their threshold for matching and then flips a coin (using progressively exponentially decreasing probability) to determine whether to be part of the set to match with $v$. Then, every vertex ordered after $v$ answers YES or NO depending on its coin flip. Intersecting the YES/NO responses with $v$’s private adjacency list results in $v$’s matches. This second portion of the algorithm is the receiver phase. All of these procedures can be done locally privately in the LEDP model.
Within this algorithm, we introduce the Public Vertex Subset Mechanism which may be of independent interest. In this mechanism, each node releases a public subset of nodes obtained via coin flips. Specifically, for each pair of nodes, we flip a set of coins with appropriate probability parameterized by subgraph indices $r$. For index $r$, each coin is flipped with probability $(1+\eta)^{-r}$. Then, a node $v$ publicly releases a subgraph index $r$ to determine a public subset of vertices consisting of coin flips which landed heads for each pair $\{v,u\}$ containing $v$. The value of $r$ determines the size of this public subset of vertices (in expectation). Then, each node $w$ takes the public subsets and intersects them with their private adjacency lists to determine a private subset of vertices only $w$ knows about. Thus, the public subset can be written on the billboard and the private subset is used to determine each node’s implicit answer. We believe this mechanism will be helpful for other graph problems that release implicit solutions.
To speed up our algorithm to run in polylogarithmic rounds (Theorem 1.6), we need to allow many vertices to simultaneously propose matches to their neighbors (rather than just one at a time). Intuitively, this can be done in ways similar to classical parallel and distributed algorithms for related problems like Boruvka’s parallel MST algorithm [Boruvka]. In these algorithms, nodes randomly choose whether to propose or to listen. Then, proposer nodes send proposals, receivers decide which of their proposals to accept, and receivers communicate the acceptances back to the proposers (see [DHIN19] for a recent example of this in distributed settings). However, the added LEDP constraint makes this process significantly more difficult. Note, for example, that even communicating “with neighbors” violates privacy, since the transcript would then reveal who the neighbors are! Hence, greater care must be taken regarding the messages each node transmits. Fortunately, by using techniques gleaned from our $n$-round algorithm, ideas from basic parallel algorithms, and improved analyses, our implementation goes through.
The crux of our $O(\log n)$ round distributed algorithm relies on our Public Vertex Subset Mechanism. We use this mechanism to release proposal sets on the part of the proposers and match sets on the part of the receivers. The proposers release public proposal sets that contain proposals to its neighbors to match. Using the Public Vertex Subset Mechanism, receivers can privately discern which nodes have proposed. Then, receivers also publish a public subset of nodes as their match sets. Using this mechanism, the proposers can then (privately) recognize which receivers accepted their proposals. Performing multiple rounds of this procedure leads to our desired Theorem 1.6. Our novel utility analyses depends on a novel noisy version of a charging argument, based on edge orientation, implicitly described in [khuller1994primal].
Implicit Solution Lower Bound.
Our lower bound on implicit solutions (Theorem 11.1) shows that we must violate the degree bounds in our returned matchings and so our bicriteria algorithms are reasonable. The proof is much more straightforward than the lower bound proof on explicit solutions. We start with a low-degree graph, and first use the DP guarantee over nearby graphs to argue that edges that do not exist must still be included with reasonably high probability (this is essentially a highly simplified version of the symmetry argument used in the previous lower bound). We then use group privacy to argue that there is a graph with logarithmic max degree (which is also logarithmically far from our starting graph) where many edges incident to the node of logarithmic degree must be included.
1.3.2 Node Privacy
General Graphs. (Section 6)
To achieve node privacy in the central model, we combine our edge privacy results with a “matching sparsifier” [solomon2021local]. Informally, a matching sparsifier of a graph $G$ is a subgraph $H$ such that 1) the maximum matching in $H$ is approximately as large as the maximum matching in $G$, and 2) the maximum degree of $H$ is at most $O(\alpha)$, where $\alpha$ is the arboricity of $G$. Note that the average degree in a graph with arboricity $\alpha$ is at most $O(\alpha)$, so such a sparsifier essentially turns the average degree into the maximum degree without harming the matching.
Given an algorithm to compute a matching sparsifier, it is obvious how to combine this with edge privacy to get node privacy: since the maximum degree in the sparsifier is at most $O(\alpha)$, basic group privacy implies that we can simply use our edge-private algorithm and incur an extra loss of at most an $O(\alpha)$ factor. In other words, if the maximum degree is small then node privacy is approximately the same as edge privacy, and a matching sparsifier guarantees small maximum degree.
So if we could compute and release a differentially private matching sparsifier, we would be done. Unfortunately, we show in Section 11.2 that this is impossible, by providing a strong lower bound against private matching sparsification (Theorem 11.2). However, we show that we can use these sparsifiers in a node-DP algorithm even though releasing them does not satisfy DP! Intuitively, this is because given two node-neighboring graphs, computing a sparsifier of them can bring the distributions of the outputs “closer together”. In particular, although these sparsifiers cannot be released publicly, they decrease the “edge-distance” between node-neighboring graphs, where the edge-distance is the number of edges that differ between neighboring graphs.
We extend our results to $b$-matchings using our novel arboricity sparsifiers for $b$-matchings in Section 10.2.
Improvement in Bipartite Graphs. (Section 7)
The starting point of our algorithm for node-private bipartite matchings is the same deferred acceptance type algorithm which [hsu2014private] base their private algorithm on. In their private implementation of the algorithm, they use continual counters [dwork2009complexity, CSS11] to keep track of the number of bidders each item is matched to. This is necessary for their implementation since their implicit solutions need to know the number of bidders an item is matched with at every iteration of the algorithm. We show that in a different private implementation of the algorithm, the counts of the number of bidders an item is matched to is only needed $O(1/\eta^{2})$ different times (where $\eta\in(0,1]$ is present in the $(1+\eta)$-approximation factor). Furthermore, a clever use of MAT to iteratively check if this count exceeds the threshold $s$ (where $s$ is the matching factor) suffices for the implementation, enabling us to substantially improve the error of the counts used.
1.3.3 Continual Release (Sections 8 and 9)
We extend all of our results to the continual release model under two different types of streams: edge-order streams and adjacency-list streams, both commonly seen in non-private streaming literature (see e.g. [kallaugher2019complexity, mcgregor2016better, mcgregor2005finding]). We consider edge-insertion streams where an update in the stream can either be $\bot$ (an empty update) or an insertion of an edge $\{u,v\}$. Edge-neighboring streams are two streams that differ in exactly one edge update. We also consider node-neighboring streams; node-neighboring streams are two streams that differ in all edges adjacent to any one node. We release a solution after every update over the course of $T=\operatorname{poly}(n)$ updates. [^6] Our algorithms extend to the case where $T=\omega(\operatorname{poly}(n))$ at the cost of factors of $\log(T)$. We focus on the case $T=\operatorname{poly}(n)$ as there are at most $O(n^{2})$ non-empty updates. The goal is to produce an accurate approximate solution for each update, with high probability.
The continual release setting is a more difficult setting than the static setting for several reasons. First, each piece of private data is used multiple times to produce multiple solutions, potentially leading to high error from composition. Second, depending on the algorithm, it may be possible to accumulate more errors as one releases more solutions (leading to compounding errors). Finally, for node-private algorithms, sparsification techniques need to be handled with more care since temporal edge updates can lead to sparsified solutions becoming unstable (causing the neighboring streams to become farther in edge-distance instead of closer). We solve all of these challenges to implement our matching algorithms in the continual release model.
We first adapt our LEDP and node-DP implicit $O(\log(n)/\varepsilon)$-matching algorithms to edge-order streams. These results are given in Theorem 1.11 and Theorem 1.12. The main idea behind our continual release algorithm is to use the sparse vector technique (SVT) to determine when to release a new solution at each timestep $t\in[T]$. Specifically, at timestep $t$, we check the current exact maximum matching size in the induced subgraph $G_{t}$ consisting of all updates up to and including $t$. If this matching size is greater than a $(1+\eta)$ factor (for some fixed $\eta\in(0,1]$) of our previously released solution, then we release a new solution. To release our new solution, we use our LEDP algorithm as a blackbox and pass $G_{t}$ into the algorithm. Then, we release the implicit solution that is the output of our LEDP algorithm. Since we can only increase our solution size $O(\log_{1+\eta}(n))$ times, we only accumulate an additional $O(\log(n)/\eta)$ factor in the error due to composition.
For our node-DP continual release algorithm for edge-order streams, we perform the same strategy as our edge-DP algorithm above except for one main change. We implement a stable version of our generalized matching sparsifier in Section 6 for edge-order streams in the continual release model. Then, for each update, we determine whether to keep it as part of our matching sparsifier. Using the sparsified set of edges, we run our SVT procedure to determine when to release a new matching and use our LEDP algorithm as a blackbox.
Finally, we adapt our LEDP matching algorithm to adjacency-list order streams in the continual release model. In adjacency-list order streams, updates consist of both vertices and edges where each edge shows up twice in the stream. Immediately following a vertex update, all edges adjacent to the vertex are given in an arbitrary order in the stream following the vertex update. Edge-neighboring adjacency-list order streams differ in exactly one edge update. We implement our LEDP algorithm in a straightforward manner in adjacency-list order streams. In particular, a node performs our proposal procedure once it sees all of its adjacent edges. Then, the node writes onto the blackboard the results of this proposal procedure. We maintain the same utility guarantees except for an additional additive error of $1$. Since a node must wait until it sees all of its adjacent edges before performing the proposal procedure, this error results from the most recent node update (where the node has yet to observe all of its adjacent edges).
1.4 Related Work
Private Allocations.
[…]
Now, auction-based methods and the implicit solution concept of [hsu2014private] are inherently tied to bipartite graphs. While it is possible to reduce general matchings to a bipartite instance via random bipartitions [liu2024pointwise], this transformation incurs a factor of 2 in the multiplicative approximation factor and the obtained approximation ratio is only in expectation. This would imply a node-DP $(2+\eta)$-approximate $s$-matching in expectation for general graphs. Our novel solution concept allows us to handle general graphs, under edge-DP or node-DP assuming low arboricity, and obtain a $(2+\eta)$-approximate $s$-matching with high probability while reducing the lower bound to $s=\Omega(\log(n)/\varepsilon)$. Moreover, specializing our techniques to the bipartite case (with arbitrary arboricity) recovers the node-DP $(1+\eta)$-approximation ratio of [hsu2014private], with the same reduced bound on $s$.
We remark that a followup work [hsu2016jointly] studies a class of private convex programming problems that capture fractional allocations with supply $s$ under approximate differential privacy. Similar to our algorithm, they also output an $s^{\prime}$-matching with $s^{\prime}>s$. Unlike our results, they compute a fractional allocation and guarantee that the sum of constraint violations ($\sum_{j}s_{j}^{\prime}-s_{j}$) is bounded, whereas our algorithm returns an (integral) solution while ensuring each constraint violation is bounded ($\max_{j}s_{j}^{\prime}-s_{j}$). Thus, neither the privacy nor the utility guarantees are comparable to ours.
Although [kannan2015pareto] also study private allocations, they do so in a relaxed notion of differential privacy called marginal differential privacy with the goal of computing approximate pareto optimal allocations. Thus both their privacy and utility guarantees differ from our setting.
Private Matching Size.
Differentially private estimations of the size of the maximum matching have been recently studied in various privacy models including the continual release model [FHO21, wagaman2024time, raskhodnikova2024fully, epasto2024sublinear] and the sublinear model [BGM22]. In the standard central DP model, computing an approximate solution for the size of the maximum matching in the input graph with $O(\log(n)/\varepsilon)$ additive error is trivial since the sensitivity of the size of the maximum matching is $1$ and so we can use the Laplace mechanism to obtain such an estimate.
Implicit Solutions.
Implicit solutions in the graph setting were first introduced by [GuptaLMRT10] for various problems such as set cover and vertex cover. An implicit solution is one where the released information from a differentially private algorithm is $\varepsilon$-DP or $\varepsilon$-LEDP but the released solution does not immediately solve the problem. In particular, the released solution is an implicit solution from which each individual node can compute the explicit solution for itself. In the case of our matching algorithms, the explicit solution computed by every node is the set of edges in the matching that are adjacent to it.
2 Preliminaries
2.1 Differential Privacy
We begin with the basic definitions of differential privacy for graphs. Two graphs $G$ and $G^{\prime}$ are said to be edge-neighboring if they differ in one edge. They are said to be node-neighboring if they differ in one node (and all edges incident to that node).
Definition 2.1 (Graph Differential Privacy [NRS07]).
Algorithm $\mathcal{A}(G)$ that takes as input a graph $G$ and outputs an object in $\text{Range}(\mathcal{A})$ is $(\varepsilon,\delta)$-edge (-node) differentially private ($(\varepsilon,\delta)$-edge (-node) DP) if for all $S\subseteq\text{Range}(\mathcal{A})$ and all edge- (node-)neighboring graphs $G$ and $G^{\prime}$,
$\Pr[\mathcal{A}(G)\in S]\leq e^{\varepsilon}\Pr[\mathcal{A}(G^{\prime})\in S]+\delta$
If $\delta=0$ in the above, then we drop it and simply refer to $\varepsilon$-edge (-node) differential privacy.
In this paper, we differentiate between the representation model which we will use to release solutions and the privacy model. The privacy model we use for our static algorithms is the local edge differential privacy (LEDP) model defined in [DLRSSY22]. We give the transcript-based definition, defined on $\varepsilon$-local randomizers, verbatim, below.
Definition 2.2 (Local Randomizer (LR) [DLRSSY22]).
An $\varepsilon$-local randomizer $R:\mathbf{a}\rightarrow\mathcal{Y}$ for node $v$ is an $\varepsilon$-edge DP algorithm that takes as input the set of its neighbors $N(v)$, represented by an adjacency list $\mathbf{a}=(b_{1},\dots,b_{|N(v)|})$. In other words,
$\frac{1}{e^{\varepsilon}}\leq\frac{\mathsf{Pr}\left[R(\mathbf{a}^{\prime})\in Y\right]}{\mathsf{Pr}\left[R(\mathbf{a})\in Y\right]}\leq e^{\varepsilon}$
for all $\mathbf{a}$ and $\mathbf{a}^{\prime}$ where the symmetric difference is $1$ and all sets of outputs $Y\subseteq\mathcal{Y}$. The probability is taken over the random coins of $R$ (but not over the choice of the input).
Definition 2.3 (Local Edge Differential Privacy (LEDP) [DLRSSY22]).
A transcript $\pi$ is a vector consisting of 5-tuples $(S^{t}_{U},S^{t}_{R},S^{t}_{\varepsilon},S^{t}_{\delta},S^{t}_{Y})$ – encoding the set of parties chosen, set of local randomizers assigned, set of randomizer privacy parameters, and set of randomized outputs produced – for each round $t$. Let $S_{\pi}$ be the collection of all transcripts and $S_{R}$ be the collection of all randomizers. Let $\mathtt{EOC}$ denote a special character indicating the end of computation. A protocol is an algorithm $\mathcal{A}:S_{\pi}\to(2^{[n]}\times 2^{S_{R}}\times 2^{\mathbb{R}^{\geq 0}}\times 2^{\mathbb{R}^{\geq 0}})\;\cup\{\mathtt{EOC}\}$ mapping transcripts to sets of parties, randomizers, and randomizer privacy parameters. The length of the transcript, as indexed by $t$, is its round complexity.
Given $\varepsilon\geq 0$, a randomized protocol $\mathcal{A}$ on (distributed) graph $G$ is $\varepsilon$-locally edge differentially private ($\varepsilon$-LEDP) if the algorithm that outputs the entire transcript generated by $\mathcal{A}$ is $\varepsilon$-edge differentially private on graph $G.$ If $t=1$, that is, if there is only one round, then $\mathcal{A}$ is called non-interactive. Otherwise, $\mathcal{A}$ is called interactive.
Now, we decouple the billboard model definition given in [hsu2014private] from the privacy model. Hence, the billboard model only acts as a solution release model. The billboard model takes as input a private graph and produces an implicit solution consisting of a public billboard. Then, each node can determine its own part of the solution using the public billboard and the private information about its adjacent neighbors. In this paper, our algorithms for producing the public billboard will be $\varepsilon$-LEDP.
Definition 2.4 (Billboard Model [hsu2014private]).
Given an input graph $G=(V,E)$, algorithms in the billboard model produces a public billboard. Then, each node $v\in V$ in the graph deduces the portion of the explicit solution that $v$ participates in.
In particular, one can easily show that if every node processes the information contained in the public billboard using a deterministic algorithm, then the explicit solution contained at every node will be $\varepsilon$-edge DP (with respect to edge-neighboring graphs) except for the two nodes adjacent to the edge that differs between neighboring graphs. Morally speaking, such a set of explicit solutions does not leak any additional private information since the nodes that are endpoints to the edge that differs already know that this edge exists. Hence, we decouple the privacy definition from the solution release model and say an algorithm is $\varepsilon$-LEDP in the billboard model if the public billboard is produced via a $\varepsilon$-LEDP algorithm and the explicit solutions obtained by every node is via a deterministic algorithm at each node.
Lemma 2.5.
Given a public billboard produced from the billboard model (Definition 2.4), if each node produces their individual explicit solution using a (predetermined) deterministic algorithm then the produced explicit solutions are $\varepsilon$-edge DP except for the solutions produced by the endpoints of the edge that differs between edge-neighboring graphs.
Proof.
Given identical adjacency lists and an identical billboard, a deterministic algorithm will output identical solutions. Hence, by the definition of $\varepsilon$-edge DP, every node will produce identical explicit solutions except for the endpoints of the edge that differs between edge-neighboring graphs. ∎
We now define the implicit solutions that our algorithms will post to the billboard. Note that by the definition of $\varepsilon$-LEDP, the entire transcript is public and so without loss of generality is posted to the billboard. But the particular information that each node will use to produce its explicit solution is the following.
Definition 2.6 (Implicit Solution).
Given a graph $G=(V,E)$, an implicit solution is a collection $\mathcal{S}=\{S_{v}\}_{v\in V}$ where each $S_{v}\subseteq V$. An implicit solution $\mathcal{S}$ defines an implicit graph $H(\mathcal{S})=(V,E^{\prime})$ where $E^{\prime}=\{\{u,v\}:u\in S_{v}\lor v\in S_{u}\}$. The degree of an implicit solution $\mathcal{S}$ is the maximum degree in the graph $(V,E\cap E(H(\mathcal{S})))$, i.e., the maximum degree in the graph which is the intersection of $G$ and $H(\mathcal{S})$.
In other words, for each vertex $v$ we have a subset of nodes $S_{v}$. Think of $S_{v}$ as “potential matches” for $v$. Then $H(\mathcal{S})$ is the graph obtained by adding an edge if either endpoint contains the other as a potential match. We get a third graph by intersecting the implicit graph with the true graph, and the maximum degree in this graph is what we call the degree of the solution. From a billboard/LEDP perspective, this means that any node $v$ can locally decode an explicit solution from the implicit solution since it can perform the intersection of $H(\mathcal{S})$ with its own neighborhood.
Now, we define an additional data input model which we use in our proof of our lower bound.
Definition 2.7 ($0/1$-Weight Model [hsu2014private]).
Given a complete graph $G=(V,E)$, each edge in the graph is given a binary weight of 00 or $1$. Edge-neighbors are two graphs $G$ and $G^{\prime}$ where the weight of exactly one edge differs. Node-neighbors are two graphs where the weights of all edges adjacent to exactly one node differ.
2.1.1 Continual Release
In this section, we define the continuous release model [DNPR10, CSS11]. We first define the concepts of edge-order and adjacency-list order streams and then edge-neighboring and node-neighboring streams.
Definition 2.8 (Edge-Order Graph Stream [wagaman2024time]).
In the edge-order continual release model, a graph stream $S\in\mathcal{S}^{T}$ of length $T$ is a $T$-element vector where the $t$-th element is an edge update $u_{t}=\{v,w,\mathtt{insert}\}$ (an edge insertion of edge $\{v,w\}$), or $\bot$ (an empty operation).
Definition 2.9 (Adjacency-List Order Graph Stream (adapted from [mcgregor2016better])).
In the adjacency-list order continual release model, a graph stream $S\in\mathcal{S}^{T}$ of length $T$ is a $T$-element vector where the $t$-th element is a node update $u_{t}=\{v\}$, an edge update $u_{t}=\{v,w,\mathtt{insert}\}$ (an edge insertion of edge $\{v,w\}$), or $\bot$ (an empty operation). Each node update is followed (in an arbitrary order) by all adjacent edges.
We use $G_{t}$ and $E_{t}$ to denote the graph induced by the set of updates in the stream $S$ up to and including update $t$. Now, we define neighboring streams as follows. Intuitively, two graph streams are edge neighbors if one can be obtained from the other by removing one edge update (replacing the edge update by an empty update in a single timestep); and they are node-neighbors if one can be obtained from the other via removing all edge updates incident to a particular vertex.
Definition 2.10 (Edge Neighboring Streams).
Two streams of updates, $S=[u_{1},\dots,u_{T}]$ and $S^{\prime}=[u^{\prime}_{1},\dots,u^{\prime}_{T}]$, are edge-neighboring if there exists exactly one timestamp $t^{*}\in[T]$ (containing an edge update in $S$ or $S^{\prime}$) where $u_{t^{*}}\neq u^{\prime}_{t^{*}}$ and for all $t\neq t^{*}\in[T]$, it holds that $u_{t}=u^{\prime}_{t}$. Streams may contain any number of empty updates, i.e. $u_{t}=\bot$. Without loss of generality, we assume for the updates $u_{t^{*}}$ and $u^{\prime}_{t^{*}}$, it holds that $u^{\prime}_{t^{*}}=\bot$ and $u_{t^{*}}=e_{t^{*}}$ is an edge insertion.
Definition 2.11 (Node Neighboring Streams).
Two streams of updates, $S=[u_{1},\dots,u_{T}]$ and $S^{\prime}=[u^{\prime}_{1},\dots,u^{\prime}_{T}]$, are node-neighboring if there exists exactly one vertex $v^{*}\in V$ where for all $t\in[T]$, $u_{t}\neq u^{\prime}_{t}$ only if $u_{t}$ or $u_{t}^{\prime}$ is an edge insertion of an edge adjacent to $v^{*}$. Streams may contain any number of empty updates, i.e. $u_{t}=\bot$. Without loss of generality, we assume for the updates $u_{t}\neq u^{\prime}_{t}$, it holds that $u^{\prime}_{t}=\bot$ and $u_{t}=e_{t}$ is an edge insertion of an edge adjacent to $v^{*}$.
We now define edge-privacy and node-privacy for edge-neighboring and node-neighboring streams, respectively.
Definition 2.12 (Edge Differential Privacy for Edge-Neighboring Streams).
Let $\varepsilon\in(0,1)$. An algorithm $\mathcal{A}(S):\mathcal{S}^{T}\rightarrow\mathcal{Y}^{T}$ that takes as input a graph stream $S\in\mathcal{S}^{T}$ is said to be $\varepsilon$-edge differentially private (DP) if for any pair of edge-neighboring graph streams $S,S^{\prime}$ (Definition 2.10) and for every $T$-sized vector of outcomes $Y\subseteq\text{Range}(\mathcal{A})$,
$\mathsf{Pr}\left[\mathcal{A}(S)\in Y\right]\leq e^{\varepsilon}\cdot\mathsf{Pr}\left[\mathcal{A}(S^{\prime})\in Y\right].$
Definition 2.13 (Node Differential Privacy for Node-Neighboring Streams).
Let $\varepsilon\in(0,1)$. An algorithm $\mathcal{A}(S):\mathcal{S}^{T}\rightarrow\mathcal{Y}^{T}$ that takes as input a graph stream $S\in\mathcal{S}^{T}$ is said to be $\varepsilon$-node differentially private (DP) if for any pair of node-neighboring graph streams $S,S^{\prime}$ (Definition 2.11) and for every $T$-sized vector of outcomes $Y\subseteq\text{Range}(\mathcal{A})$,
$\mathsf{Pr}\left[\mathcal{A}(S)\in Y\right]\leq e^{\varepsilon}\cdot\mathsf{Pr}\left[\mathcal{A}(S^{\prime})\in Y\right].$
2.2 Differential Privacy Tools
In this section, we state the privacy tools we use in our paper. The adaptive Laplace mechanism is a formalization of the Laplace mechanism for adaptive inputs that we employ in this work (and is used implicitly in previous works). For completeness, we include a proof of Lemma 2.14 in Section A.2.
Lemma 2.14 (Adaptive Laplace Mechanism; used implicitly in [wagaman2024time]).
Let $f_{1},\ldots,f_{k}$ with $f_{i}:\mathcal{G}\to\mathbb{R}$ be a sequence of adaptively chosen queries and let $f$ denote the vector $(f_{1},\ldots,f_{k})$. Suppose that the adaptive adversary gives the guarantee that the vector $f$ has $\ell_{1}$-sensitivity $\Delta$, regardless of the output of the mechanism. Then the Adaptive Laplace Mechanism $\mathcal{M}$ with vector-valued output $\tilde{f}(G)$ where $\tilde{f}_{i}(G)\coloneqq f_{i}(G)+\text{Lap}(\Delta/\varepsilon)$ for each query $f_{i}$ is $\varepsilon$-differentially private.
The Multidimensional AboveThreshold mechanism (Algorithm 1) is a generalization of the AboveThreshold mechanism [lyu2017understanding] which is traditionally used to privately answer sparse threshold queries.
Lemma 2.15 (Multidimensional AboveThreshold Mechanism [DLL23]).
Algorithm 1 is $\varepsilon$-LEDP.
1 Input: Graph $G$, adaptive queries $\{\vec{f}_{1},\dots,\vec{f}_{n}\}$, threshold vector $\vec{T}$, privacy $\varepsilon$, $\ell_{1}$-sensitivity $\Delta$.
2 Output: A sequence of responses $\{\vec{a}_{1},\dots,\vec{a}_{n}\}$ where $a_{i,j}$ indicates if $f_{i,j}(G)\geq\vec{T}_{j}$
1:  for $j=1,\ldots,d$ do
2:     $\hat{T}_{j}\leftarrow\vec{T}_{j}+\text{Lap}(2\Delta/\varepsilon)$
3:  end for
4:
5:  for each query $\vec{f}_{i}\in\{\vec{f}_{1},\dots,\vec{f}_{n}\}$ do
6:     for $j=1,\ldots,d$ do
7:        Let $\nu_{i,j}\leftarrow\text{Lap}(4\Delta/\varepsilon)$
8:        if $f_{i,j}(G)+\nu_{i,j}\geq\hat{T}_{j}$ then
9:           Output $a_{i,j}=$ “above”
10:           Stop answering queries for coordinate $j$
11:        else
12:           Output $a_{i,j}=$ “below”
13:        end if
14:     end for
15:  end for
Algorithm 1 Multidimensional AboveThreshold (MAT) [DLL23]
In the privacy analysis of our algorithms, we will often argue that each subroutine is DP and hence the whole algorithm is DP. However, since the access to private data is interactive, we will need some form of concurrent composition theorem, such as the one stated below.
Lemma 2.16 (Concurrent Composition Theorem [vadhan2021concurrent]).
If $k$ interactive mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are each $(\varepsilon,\delta)$-differentially private, then their concurrent composition is $\left(k\cdot\varepsilon,\frac{e^{k\varepsilon}-1}{e^{\varepsilon}-1}\cdot\delta\right)$-differentially private.
Finally, we use the following privacy amplification theorem for subsampling.
Lemma 2.17 (Privacy Amplification via Subsampling).
Let $\varepsilon\in(0,1]$. If elements from the private dataset are sampled with probability $p$ and we are given a $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ on the original dataset, then running $\mathcal{A}$ on the subsampled dataset gives a $(2p\varepsilon,p\cdot\delta)$-DP algorithm for $\varepsilon\in(0,1)$.
3 Lower Bound for Explicit Matchings
We now prove Theorem 1.1 which is restated below for convenience: in the basic 0/1 weight $(\varepsilon,\delta)$-edge DP model (Definition 2.7), the required error is essentially linear, even if we allow for the algorithm to output a $b$-matching for very large $b$.
See 1.1
Given a graph $G=(V,E)$, let $\nu(G)$ denote the size of a maximal matching in $G$ (while this is not fully well-defined, all graphs on which we use this definition will have the property that all maximal matchings have the same size, and so are actually maximum matchings). Given a (randomized) algorithm $\mathcal{A}$ which outputs some graph $H=(V,E_{H})$ when run on input graph $G=(V,E)$, let $\nu_{\mathcal{A}}(G)=\nu((V,E\cap E_{H}))$ denote the size of a maximal matching in the graph consisting of edges that are in both $G$ and $H$.
At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument we will instead use a symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that the marginal probability of choosing each non-edge is identical. This makes it easy to argue that high-quality algorithms cannot satisfy differential privacy.
3.1 Proof of Theorem 1.1
We begin with a description of our inputs and some more notation. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-differentially private algorithm which always returns a $b$-matching on $V$ (i.e. a graph with vertex set $V$ in which every vertex has degree at most $b$). For $G\in\mathcal{M}$ and $\{u,v\}\in\binom{V}{2}$, let $p(\mathcal{A},G,\{u,v\})$ denote the probability that $\mathcal{A}(G)$ contains $\{u,v\}$. Let $\nu_{\mathcal{A}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}(G)]$ denote the worst-case expected utility of $\mathcal{A}$ on $\mathcal{M}$. Note that for $G\in\mathcal{M}$, every subset of $E(G)$ is a matching. Hence $\text{E}[\nu_{\mathcal{A}}(G)]=\text{E}[|E(G)\cap\mathcal{A}(G)|]\geq\nu_{\mathcal{A}}$ for all $G\in\mathcal{M}$. By linearity of expectations, this is equivalent to saying that $\sum_{\{u,v\}\in E(G)}p(\mathcal{A},G,\{u,v\})\geq\nu_{\mathcal{A}}$ for all $G\in\mathcal{M}$.
With this notation in hand, we can now define and prove the main symmetry property.
Lemma 3.1.
If $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-DP and has expected utility at least $\nu_{\mathcal{A}}$ on $\mathcal{M}$, then there is an algorithm $\mathcal{A}^{\prime}$ which also satisfies $(\varepsilon,\delta)$-DP and has $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$, and moreover satisfies the following symmetry property: $p(\mathcal{A}^{\prime},G,\{u,v\})=p(\mathcal{A}^{\prime},G,\{u^{\prime},v^{\prime}\})$ for all $\{u,v\},\{u^{\prime},v^{\prime}\}\not\in E(G)$.
Proof.
Consider the following algorithm $\mathcal{A}^{\prime}$. We first choose a permutation $\pi$ of $V$ uniformly at random from the space $S_{V}$ of all permutations of $V$. Let $\pi(G)$ denote the graph isomorphic to $G$ obtained by applying this permutation, i.e., by creating an edge $\{\pi(u),\pi(v)\}$ for each edge $\{u,v\}\in E(G)$. We run $\mathcal{A}$ on $\pi(G)$, and let $\pi(H)$ denote the resulting $b$-matching. We then “undo” $\pi$ to get the $b$-matching $H=\{\{u,v\}:\{\pi(u),\pi(v)\}\in\pi(H)\}$. We return $H$.
We first claim that $\mathcal{A}^{\prime}$ is $(\varepsilon,\delta)$-DP. This is straightforward, since it simply runs $\mathcal{A}$ on a uniformly random permuted version of $G$, and then does some post-processing. Slightly more formally, for any neighboring graphs $G$ and $G^{\prime}$ and for any $\pi\in S_{V}$, the graphs $\pi(G)$ and $\pi(G^{\prime})$ are also neighboring graphs. Hence the differential privacy guarantee of $\mathcal{A}$ implies that running $\mathcal{A}$ on $\pi(G)$ and $\pi(G^{\prime})$ is $(\varepsilon,\delta)$-DP, and then “undoing” $\pi$ is simply post-processing. So $\mathcal{A}^{\prime}$ has the same privacy guarantees as $\mathcal{A}$.
Now we claim that $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$. Let $\alpha_{\mathcal{A}}(G)$ denote the expected utility of $\mathcal{A}$ on $G\in\mathcal{M}$, and similarly let $\alpha_{\mathcal{A}^{\prime}}(G)$ denote the expected utility of $\mathcal{A}^{\prime}$ on $G$. Note that by construction, there is a bijection between $\pi(H)\cap E(\pi(G))$ and $H\cap E(G)$, and hence $\nu_{\mathcal{A}^{\prime}}(G)=\nu_{\mathcal{A}}(\pi(G))$. Since $\pi$ was chosen uniformly at random, we know that $\pi(G)$ is distributed uniformly among $\mathcal{M}$. Hence for any $G\in\mathcal{M}$ we have that
$\text{E}[\nu_{\mathcal{A}^{\prime}}(G)]$ $=\frac{1}{|\mathcal{M}|}\sum_{G^{\prime}\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}(G^{\prime})]\geq\frac{1}{|\mathcal{M}|}\sum_{G^{\prime}\in\mathcal{M}}\nu_{\mathcal{A}}=\nu_{\mathcal{A}}.$
Sicne this is true for all $G\in\mathcal{M}$, we have that $\nu_{\mathcal{A}^{\prime}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}^{\prime}}(G)]\geq\nu_{\mathcal{A}}$, as claimed.
We now prove the symmetry property. Note that there are $n!$ permutations in $S_{V}$, but only $\frac{n!}{2^{n/2}(n/2)!}$ different perfect matchings (i.e., $|\mathcal{M}|=\frac{n!}{2^{n/2}(n/2)!}$). This is because $\pi(G)$ can equal $\pi^{\prime}(G)$ for many for many $\pi\neq\pi^{\prime}$. In particular, it is not hard to see that |$|\{\pi\in S_{V}:\pi(G)=G^{\prime}\}|=2^{n/2}(n/2)!$ for all $G,G^{\prime}\in\mathcal{M}$. Fix $G,G^{\prime}\in\mathcal{M}$, and let $\Pi(G,G^{\prime})=\{\pi\in S_{V}:\pi(G)=G^{\prime}\}$. It is not hard to see that if we draw a permutation $\pi$ uniformly at random from $\Pi(G,G^{\prime})$, then for any $\{u,v\}\not\in E(G)$, the pair $\{\pi(u),\pi(v)\}$ is uniformly distributed among the set $\{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})\}$. Thus we have that
$p(\mathcal{A}^{\prime},G,\{u,v\})$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}p(\mathcal{A},G^{\prime},\{\pi(u),\pi(v)\})$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\pi\in\Pi(G,G^{\prime})}p(\mathcal{A},G^{\prime},\{\pi(u),\pi(v)\})$
$=\frac{1}{|\mathcal{M}|}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})}p(\mathcal{A},G^{\prime},\{u^{\prime},v^{\prime}\}).$
Hence $p(\mathcal{A}^{\prime},G,\{u,v\})$ is actually independent of $\{u,v\}$, and thus for all $\{u,v\},\{u^{\prime},v^{\prime}\}\not\in E(G)$ we have that $p(\mathcal{A}^{\prime},G,\{u,v\})=p(\mathcal{A}^{\prime},G,\{u^{\prime},v^{\prime}\})$ as claimed. ∎
Lemma 3.1 implies that if we can prove an upper bound on the expected utility) for DP algorithms which obey the symmetry property, then the same upper bound applies to all DP algorithms. We we will now prove such a bound.
Lemma 3.2.
Let $\mathcal{A}$ be an algorithm which is $(\varepsilon,\delta)$-DP, always outputs a $b$-matching, and satisfies the symmetry property of Lemma 3.1. Then
$\nu_{\mathcal{A}}\leq e^{2\varepsilon}b+\delta n.$
Proof.
Fix some graph $G\in\mathcal{M}$, and let $\{u,v\},\{u^{\prime},v^{\prime}\}\in E(G)$. Let $G^{\prime}$ be the graph obtained by removing $\{u,v\}$ and $\{u^{\prime},v^{\prime}\}$, and adding $\{u,v^{\prime}\}$ and $\{u^{\prime},v\}$. Note that $G$ and $G^{\prime}$ are distance $2$ away from each other.
For all $v^{\prime\prime}\neq u,v^{\prime}$, it must be the case that $p(\mathcal{A},G^{\prime},\{u,v^{\prime\prime}\})\leq b/(n-2)$, since otherwise the symmetry property of Lemma 3.1 and linearity of expectations implies that the expected degree of $u$ in the output of $\mathcal{A}(G)$ is at least $\sum_{v^{\prime\prime}\neq u,v^{\prime}}p(\mathcal{A},G,\{u,v^{\prime\prime}\})>(n-2)\frac{b}{n-2}=b$, contradicting our assumption that $\mathcal{A}$ always outputs a $b$-matching. Thus $p(\mathcal{A},G^{\prime},\{u,v\})\leq b/(n-2)$. Since $G$ and $G^{\prime}$ are at distance $2$ from each other, this implies that $p(\mathcal{A},G,\{u,v\})\leq e^{2\varepsilon}\frac{b}{n-2}+2\delta$.
Since this argument applies to all $\{u,v\}\in E(G)$, we have that
$\text{E}[\nu_{\mathcal{A}}(G)]$ =$=\sum_{\{u,v\}\in E(G)}p(\mathcal{A},G,\{u,v\})$
$\leq\sum_{\{u,v\}\in E(G)}\left(e^{2\varepsilon}\frac{b}{n-2}+2\delta\right)$
$\leq e^{2\varepsilon}\frac{bn}{2(n-2)}+2\delta\frac{n}{2}.$
$\leq e^{2\varepsilon}b+\delta n.$
Since $\nu_{\mathcal{A}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}(G)]$, this implies that $\nu_{\mathcal{A}}\leq e^{2\varepsilon}b+\delta n$, as claimed. ∎
We can now use these two lemmas to prove Theorem 1.1.
Proof of Theorem 1.1.
The combination of Lemma 3.1 and Lemma 3.2 imply that any $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ which always outputs a $b$-matching has $\nu_{\mathcal{A}}\leq e^{2\varepsilon}b+\delta n$. We also know that $\nu(G)=n/2$ for all $G\in\mathcal{M}$. Thus by the definition of an $(\alpha,\beta)$-approximation, it must be the case that $\alpha(e^{2\varepsilon}b+\delta n)+\beta\geq n/2$. ∎
4 $\varepsilon$-Local Edge Differentially Private Implicit Matchings
In this section, we state our main algorithm for maximal matching (Algorithm 2) and prove its guarantees in Theorem 1.3, which is restated below for readibility. We modify Algorithm 2 in the later sections for various different models. See 1.3
Our LEDP algorithm is based on a simple procedure for maximal non-private $b$-matching, described as follows. We take an arbitrary ordering on the vertices, and process them one by one. When considering the $i^{th}$ vertex $v_{i}$, let $b^{\prime}$ be the number of additional vertices $v_{i}$ can match with. We wish to find some subset of vertices of size $b^{\prime}$ from later in the ordering with which to match $v_{i}$. Some of these vertices may have already matched with $b$ vertices from previous iterations. We choose an arbitrary subset of $b^{\prime}$ vertices which have not already reached their limit from later in the ordering. We say that vertices which have reached their limit satisfy the matching condition. If there are fewer than $b^{\prime}$ vertices satisfying the matching condition, we match all of them with $v_{i}$. Then we move on to the next vertex in the order. It is clear that this procedure yields a maximal $b$-matching. In the following algorithms, we analyze a suitable privatization of this algorithm and show that given sufficiently large $b$, we are guaranteed to produce an implicit maximal ($1$-)matching. [^7] We remark that all of our algorithms can be straightforwardly adjusted to guarantee maximal $b$-matchings.
High-Level Overview
Our approximation guarantees hold with probability $1-\frac{1}{n^{c}}$ for any constant $c\geq 3$. The constants in Algorithm 2 are given in terms of $c$ to achieve our high probability guarantee. At a high level, our $\varepsilon$-LEDP algorithm modifies our non-private procedure described above in the following way. The key part of the procedure that uses private information is the selection of the $b^{\prime}$ neighbors for each vertex $v$ to satisfy the matching condition. We cannot select an arbitrary set of these neighbors directly since this arbitrary selection would reveal the existence of an edge between $v$ and each selected neighbor. Thus, we must select an appropriate number of neighbors randomly and output a public vertex subset that also includes non-neighbors.
To solve the above challenge, we introduce a novel Public Vertex Subset Mechanism. We call a vertex that is proposing a set of vertices the proposer. The mechanism works by having a proposer who proposes a public set of vertices to match to. The public set of vertices contains vertices which are neighbors of the proposer and also non-neighbors of the proposer. To select a public subset of vertices, we flip a set of coins with appropriate probability for each pair of vertices in the graph. A total of $O(\log(n))$ coins are flipped for each pair; a coin is flipped with probability $1/(1+\eta)^{r}$ for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. Thus, these coins determine progressively smaller subsets of vertices. An edge is in the public set indexed by $r$ if the coin lands heads. It is necessary to have vertex subsets with different sizes since proposers need to choose an appropriately sized subset of vertices that simultaneously ensures it satisfies the matching condition and not does exceed the $b$ constraint, with high probability. The coin flips are public because they do not reveal the existence of any edge. Moreover, the intersection between the public subset of vertices with each node’s private knowledge of its adjacency list allows each node to know which of its neighbors are matched to it. This Public Vertex Subset Mechanism may be useful for other problems.
A proposer $v$ releases the proposal subset by releasing the index $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ that corresponds with the coin flips determining the set that $v$ wants to match to. We use the sparse vector technique on the size of the subset to determine which $r$ to release. In fact, we use a multidimensional version of the sparse vector technique, called the Multidimensional AboveThreshold (MAT) technique (Algorithm 1, [DLL23]) which is designed for SVT queries performed by all nodes of a graph. Once a proposer releases a public proposal subset, each of $v$’s neighbors, that have not met their matching conditions, can determine whether they are matched to $v$ using the public coin flips. We call $r$ the subgraph index that is released by each proposer. Thus, each node knows the set of vertices they are matched to using the transcript consisting of publicly released subgraph indices and the public releases of when each node satisfies their matching condition. Each node determines whether it has satisfied its matching condition using MAT.
4.1 Detailed Algorithm Description
We now describe our algorithm, Algorithm 2, in detail. The algorithm is provided with a private graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, a matching parameter $b\geq\frac{576c\ln(n)}{\eta\cdot\varepsilon}+1$ and a constant $c\geq 1$ which is used in the high probability guarantee. We first set the variables $\eta$ and $\varepsilon^{\prime}$ that we use in our algorithm (Algorithms 2 and 2). Then, we flip our public set of coins. We flip $O(\log(n))$ coins for each pair of vertices $u\neq v\in V\times V$ (Algorithm 2). Each of the coins flips for pair $\{u,v\}$ is flipped with probability $1/(1+\eta)^{r}$ for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. The result of the coin flip is released and stored in $coin(u,v,r)$ (Algorithm 2).
We now determine the matching condition for each node $u\in V$ (Algorithm 2). To do so, we draw a noise variable from $\text{Lap}(4/\varepsilon^{\prime})$ and add it to $b$. This noise is used as part of MAT (Algorithm 1, Multidimensional AboveThreshold) for a noisy threshold. Next, we subtract $36c\ln(n)/\varepsilon^{\prime}$ to ensure that the matching condition does not exceed $b$ even if a large positive noise was drawn from $\text{Lap}(4/\varepsilon^{\prime})$. In other words, Algorithm 2 ensures that the matching condition is satisfied when $v$ is actually matched to a sufficiently large number of vertices, with high probability. We also initialize the private data structure $M(u)$ stored by $u$ that contains the set of $u$’s neighbors that $u$ is matched to. Note that in practice, $M(u)$ is not centrally stored but can be decoded by $u$ using information posted to the public billboard.
We then initialize a set $A(u)\leftarrow\infty$ for every $u\in V$ (Algorithm 2) indicating that no vertices have yet to satisfy their matching condition. We now iterate through all of the vertices one by one in an arbitrary order (Algorithm 2). During the $i$-th iteration, if there is any node $u\in V$ which has not satisfied its matching condition (Algorithm 2), we use MAT (Algorithm 1) to check whether it now satisfies the matching condition (Algorithms 2 and 2). If node $u$ now satisfies the matching condition, then node $u$ releases $A(u)\leftarrow i$ to indicate that $u$ satisfied its matching condition at iteration $i$ (Algorithm 2).
If the current node in the iteration, $v$, has not satisfied its matching condition (Algorithm 2), then we determine its proposal set using the Public Vertex Subset Mechanism (Algorithms 2 to 2). For each subgraph index $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ (Algorithm 2), we determine the private subset of vertices $W_{r}(v)$ using the public coin flips. Specifically, $W_{r}(v)$ contains the set of neighbors $u$ of $v$ that are still active, come after $v$ in the ordering, and where $coin(u,v,r)=\textsc{Heads}$ (Algorithm 2). The public set of nodes determined by index $r$ are all nodes $x$ where $coin(v,x,r)=\textsc{Heads}$. We then determine the noisy size of $|W_{r}(v)|$ by drawing a noise variable from $\text{Lap}(2/\varepsilon^{\prime})$ to be used in the adaptive Laplace mechanism (Algorithm 2). Finally, we also determine the noisy number of nodes currently matched to $v$ (Algorithm 2). Then, $v$ determines the smallest $r$ that satisfies the SVT check in Algorithm 2. Intuitively, this means we are finding the largest subset of neighbors of $v$ that does not exceed the $b$ bound, with high probability. Then, $v$ releases $r_{v}$ (Algorithm 2) and for each neighbor $u$ of $v$ in $W_{r_{v}}(v)$ (Algorithm 2), $u$ (privately) adds $v$ to $M(u)$ (Algorithm 2) and $v$ (privately) adds $u$ to $M(v)$ (Algorithm 2).
If Algorithm 2 processed the vertices in order $v_{1},\dots,v_{n}$, then $v_{i}$ can decode its matched neighbors using the following formula
$M(v_{i})=\{v_{j}:j<i\land A(v_{j})>j\land v_{i}\in W_{r_{v_{j}}}(v_{j})\}\cup W_{r_{v_{i}}}(v_{i})\,.$
Here $W_{r_{v_{i}}}(v_{i})=\varnothing$ by convention if $A(v_{i})\leq i$ (i.e. if $v_{i}$ reaches its matching condition before the $i$-th iteration).
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching parameter $b\geq\frac{576c\ln(n)}{\varepsilon}+1$, and constant $c\geq 3$
Output: An $\varepsilon$-local edge differentially private implicit $b$-matching.
1 $\eta\leftarrow 1/2$
2 $\varepsilon^{\prime}\leftarrow\frac{\varepsilon}{2(1+2\eta)/\eta}$
3 for every pair of vertices $u\neq v\in V\times V$ and subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$  do
4       Flip and release coin $coin(u,v,r)$ which lands heads with probability $p_{r}=(1+\eta)^{-r}$
5
6for each node $u\in V$ do
7       $\tilde{b}(u)\leftarrow b-36c\ln(n)/\varepsilon^{\prime}+\text{Lap}(4/\varepsilon^{\prime})$
8       $M(u)\leftarrow\varnothing$
9       $A(u)\leftarrow\infty$
10
11for iteration $i=1$ to $n$, let $v=v_{i}$ do
       // Multidimensional-AboveThreshold for checking if each node has reached their matching threshold
12       for each node $u\in V$ such that $A(u)>i$  do
             // $u$ has not satisfied its matching condition
13             $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime})$
14             if $|M(u)|+\nu_{i}(u)\geq\tilde{b}(u)$  then
15                   $u$ releases $A(u)\leftarrow i$
16
17
      // If $v$ can still match with more edges, find an additional set to match $v$ with.
18       if $A(v)>i$  then
             // $v$ has not satisfied its matching condition
19             for subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$  do
20                   $W_{r}(v)=\{u:A(u)>i\wedge u\text{ later than $v$ in ordering }\land\{u,v\}\in E\wedge coin(u,v,r)=\textsc{Heads}\}$
21                   $|\widetilde{W}_{r}(v)|=|W_{r}(v)|+\text{Lap}(2/\varepsilon^{\prime})$
22
23            $|\widetilde{M}_{i}(v)|=|M(v)|+\text{Lap}(2/\varepsilon^{\prime})$
24             $v$ computes $r_{v}\leftarrow\min\{r:|\widetilde{M}_{i}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b\}$
25             $v$ releases $r_{v}$
26             for $u\in W_{r_{v}}(v)$ do
27                   $M(u)\leftarrow M(u)\cup\{v\}$
28                   $M(v)\leftarrow M(v)\cup\{u\}$
29
30
Algorithm 2 $\varepsilon$-LEDP Matching
4.2 Proof of Theorem 1.3
The pseudocode for Theorem 1.3 can be found in Algorithm 2. We prove its privacy and approximation guarantees separately. In Section 4.2.1, we first prove that Algorithm 2 is $\varepsilon$-DP. The utility proofs can then be found in Section 4.2.2.
4.2.1 Privacy Guarantees
We now prove that our algorithm is $\varepsilon$-LEDP using our privacy mechanisms given in Section 2.2 and show that Algorithm 2 can be implemented using local randomizers. In our proof, we implement three different types of local randomizers that use private data and perform various instructions of our algorithm. Note that not all of the local randomizers in our algorithm releases the computed information publicly but the computed information from these local randomizer algorithms satisfy edge-privacy.
Lemma 4.1.
Algorithm 2 is $\varepsilon$-locally edge differentially private.
Proof.
The algorithm consists of three different types of local randomizers which use the private information. The first type of local randomizer checks at each iteration whether or not each node has already reached its matching capacity. The second local randomizer computes the estimated number of neighbors, $|\widetilde{W}_{r}(v)|$, which are still active and are ordered after the current node for each subgraph index $r$. The third computes a noisy estimate $|\widetilde{M}(v)|$ for $v=v_{i}$ at iteration $i$. We will argue that each of these local randomizers are individually differentially private. Then, by the concurrent composition theorem (Lemma 2.16), the entire algorithm is differentially private as no other parts of the algorithm use the private information.
First, we prove that the local randomizers checking whether or not each node has reached their matching capacity (Algorithms 2, 2, 2, 2, 2 and 2) is an instance of the Multidimensional AboveThreshold (MAT) mechanism with sensitivity $\Delta=2$. The vector of queries at each iteration $i$ is the number of other nodes $|M(u)|$ each node $u$ has already been matched to before iteration $i$. Fixing the outputs of the previous iterations, the addition or removal of a single edge can only affect $|M(u)|$ for two nodes, each by at most $1$. Thus, Algorithms 2, 2, 2, 2, 2 and 2 indeed implement an instance of the Multidimensional AboveThreshold mechanism. But then these operations are $\varepsilon^{\prime}$-differentially private by Lemma 2.15.
Next, we prove that for each subgraph index $r$, the procedures that produce noisy estimates of the size of $W_{r}(v)$ in Algorithms 2, 2 and 2 is implemented via local randomizers using the Adaptive Laplace Mechanism with sensitivity $\Delta=2$. Fixing the outputs of the other parts of the algorithm (which concurrent composition allows us to do), the addition or removal of a single edge $\{u,w\}$ only affects $|W_{r}(v_{i})|$ if $u=v_{i}$ or $w=v_{i}$, and it affects each $|W_{r}(v_{i})|$ by 1. Thus, we have shown above that for a fixed subgraph index $r$, all estimates of $|W_{r}(v_{i})|$ over the iterations $i$ are $\varepsilon^{\prime}$-differentially private. Since we sample each edge with probability $p_{r}$, privacy amplification (Lemma 2.17) implies that this step is actually $2p_{r}\varepsilon^{\prime}$-differentially private since $\varepsilon^{\prime}\leq 1$.
Third, we prove that producing noisy estimates of $|M(v_{i})|$ in Algorithm 2 is an instance of the Adaptive Laplace Mechanism with sensitivity $\Delta=2$. Fixing the outputs of the other parts of the algorithm (again, which concurrent composition allows us to do), the additional and removal of an edge $\{u,w\}$ can again only affect $M(v_{i})$ for the iterations $i$ where $v_{i}=u$ and $v_{i}=w$, each by at most 1. Thus, the sensitivity is indeed $\Delta=2$ even with the adaptive choice of queries, and the noisy estimates of $|M(v_{i})|$ are $\varepsilon^{\prime}$-differentially private by Lemma 2.14.
Finally, applying the concurrent composition theorem (Lemma 2.16) proves that the entire algorithm is $2\varepsilon^{\prime}\cdot\left(1+\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}p_{r}\right)$-differentially private. We can upper bound this by an infinite geometric series:
$2\varepsilon^{\prime}\cdot\left(1+\textstyle\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}p_{r}\right)$ $\leq 2\varepsilon^{\prime}\cdot\left(1+\textstyle\sum_{r=0}^{\infty}(1+\eta)^{-r}\right)=2\varepsilon^{\prime}\cdot\left(1+\frac{1+\eta}{\eta}\right).$
By the choice of $\varepsilon^{\prime}=\frac{\varepsilon}{2(1+2\eta)/\eta}$ (Algorithm 2), we can conclude that the algorithm is $\varepsilon$-LEDP. ∎
4.2.2 Utility
We first prove that the implicitly output $b$-matching is at least half the size of a maximum matching, with high probability.
Lemma 4.2.
Let $M\subseteq E$ be an edge subset satisfying the following maximality condition: every vertex is either matched to at least $b^{\prime}$ edges of $M$, or all its unmatched neighbors are matched to at least $b^{\prime}$ edges of $M$. Then if $\operatorname{OPT}$ denotes the size of a maximum $b^{\prime}$-matching, we have
$|M|\geq\frac{1}{2}\operatorname{OPT}\,.$
We will mostly use Lemma 4.2 with $b^{\prime}=1$, but we state and prove the general version needed for Section 10.
Proof.
Let $M^{*}$ denote an arbitrary $b^{\prime}$-matching. We would like to show that $\frac{|M|}{|M^{*}|}\geq\frac{1}{2}$. Since $\frac{|M\setminus M^{*}|}{|M^{*}\setminus M|}\leq\frac{|M|}{|M^{*}|}$, we may assume without loss of generality that $M\cap M^{*}=\varnothing$ after adjusting the $b^{\prime}$ values per vertex.
Let $S\subseteq V$ denote the set of nodes $v$ with at least $b^{\prime}(v)$ matched edges in $M$. We claim that every edge of $M^{*}$ must have an endpoint in $S$. Otherwise, there is an edge $uv$ such that $u,v$ are matched to less than $b^{\prime}(u),b^{\prime}(v)$ neighbors, respectively, but is not in $M$, contradicting the maximality condition.
Now, each $v\in S$ is incident to at most $b^{\prime}(v)$ edges in $M^{*}$, while every edge of $M$ is incident to at most $2$ edges of $S$. Moreover, each node in $S$ is incident to at least $b^{\prime}(v)$ edges of $M$ by construction. Thus by a double counting argument, $|M^{*}|\leq\sum_{v\in S}b^{\prime}(v)\leq 2|M|$ as desired. ∎
Lemma 4.3.
For $b=O\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 outputs an implicit $b$-matching that is at least $\nicefrac{{1}}{{2}}$ the size of a maximum matching, with high probability.
Proof.
For each iteration $i$, we will show if node $v=v_{i}$ has not satisfied its matching condition and there exists at least $1$ node that comes after it in the order, then it is matched with at least one of the total number of nodes it can be matched with. That is, let $b^{*}(v)\in[0,b]$ denote the maximum number of nodes later than $v$ in the permutation which $v$ can match with (up to $b$); we wish to show that the number of nodes $v$ matches with after iteration $i$ is at least $1$ if $b^{*}(v)\geq 1$. By Lemma 4.2, this implies the utility guarantee. Finally, we show that with high probability, we do not match with more than $b$ nodes.
First, all of the Laplace noise (Algorithms 2, 2, 2 and 2) are drawn from distributions with expectation 00. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
  • $b-36c\log(n)/\varepsilon^{\prime}\leq\tilde{b}(u)\leq b-24c\log(n)/\varepsilon^{\prime}$ for all $u\in V$,
  • $|\nu_{i}(u)|\leq{24c\ln(n)/\varepsilon^{\prime}}$ for all $i\in[n],u\in V$,
  • $|W_{r}(v)|-{6}c\log(n)/\varepsilon^{\prime}\leq|\widetilde{W}_{r}(v)|\leq|W_{r}(v)|+{6}c\log(n)/\varepsilon^{\prime}$ for all $v\in V,r\in[\lceil\log_{1+\eta}(n)\rceil]$, and
  • $|M(v)|-{6}c\log(n)/\varepsilon^{\prime}\leq|\widetilde{M}_{i}(v)|\leq|M(v)|+{6}c\log(n)/\varepsilon^{\prime}$ for all $v_{i}\in V$.
Let $\mathcal{E}$ denote the intersection of all the events above. By a union bound, $\mathcal{E}$ occurs with probability at least $1-\frac{1}{n^{2c}}$ for $c\geq 3$. We condition on $\mathcal{E}$ for the rest of the proof.
Case I: Suppose that Algorithm 2 does not execute, i.e. $v$ has already satisfied its matching condition as evaluated on Algorithms 2, 2 and 2. Then we know that $|M(v)|+\nu_{i}(v)\geq\tilde{b}(v)$ by definition. We then have that
$|M(v)|$ $\geq b-{60}c\ln(n)/\varepsilon^{\prime}$
$\geq\frac{576c\ln(n)}{\varepsilon}+1-\frac{72c\ln(n)}{\varepsilon^{\prime}}$
$=\frac{576c\ln(n)}{\varepsilon}+1-\frac{72c\ln(n)\cdot(2(1+\eta)+2\cdot\eta)}{\varepsilon\cdot\eta}$ (by our choice of $\varepsilon^{\prime}$ in Algorithm 2)
$\geq\frac{576c\ln(n)}{\varepsilon}+1-\frac{4\cdot 2\cdot 72c\ln(n)}{\varepsilon}$ (by our choice of $\eta=0.5$)
$\geq\frac{576c\ln(n)}{\varepsilon}+1-\frac{576c\ln(n)}{\varepsilon}$
$\geq 1$
where the first line follows since we know that $b\geq\frac{576c\ln(n)}{\varepsilon}+1$. In other words, $v$ was already matched from a previous iteration.
Case II.a: Now, assume that Algorithm 2 does execute. Let $W_{0}(v)$ be the set of unmatched neighbors of $v$ which appear later in the ordering. First consider the case that $|W_{0}(v)|<\frac{8c\ln(n)}{\varepsilon^{\prime}}$. We know that either $|M(v)|=0$ or $|M(v)|\geq 1$. If $|M(v)|\geq 1$, then $v$ is already matched and there is nothing to prove. Otherwise, conditioned on $\mathcal{E}$, $|W_{0}(v)|+3\cdot\frac{12c\ln(n)}{\varepsilon^{\prime}}=\frac{8c\ln(n)}{\varepsilon^{\prime}}+\frac{36c\ln(n)}{\varepsilon^{\prime}}\leq\frac{44c\ln(n)}{\varepsilon^{\prime}}\leq\frac{8\cdot 44c\ln(n)}{\varepsilon}<\frac{576c\ln(n)}{\varepsilon}+1\leq b$ and all of $W_{0}(v)$ is matched. In particular, $v$ is matched to at least $b^{*}(v)$ neighbors.
Case II.b.1: For the remainder of the proof, we assume $k\coloneqq|W_{0}(v)|\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$. As a reminder, we are in the case that Algorithm 2 does execute. Let $r^{*}$ be the $r$ chosen in Algorithm 2. If the returned $r^{*}>\log_{1+\eta}(k\varepsilon^{\prime}/(8c\ln{n}))$, we know that the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b$ did not hold for $r=\log_{1+\eta}\left(\frac{k\varepsilon^{\prime}}{8c\ln{n}}\right)$. Thus, we have that $p_{r}=(1+\eta)^{r}=\frac{8c\ln(n)}{k\varepsilon^{\prime}}$ so the expected size of $W_{r}(v)$ is $\frac{8c\ln(n)}{\varepsilon^{\prime}}$; that is, $\mathbb{E}[|W_{r}(v)|]=\frac{8c\ln(n)}{\varepsilon^{\prime}}$. By a multiplicative Chernoff Bound (Theorem A.2), the true size of $W_{r}(v)$ is at most $16c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{4c}}$. Since the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\log(n)/\varepsilon^{\prime}\leq b$ didn’t hold for such $r$, we have that
$|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+\frac{12c\ln(n)}{\varepsilon^{\prime}}$ $>b$
$|M(v)|+|W_{r}(v)|+\frac{6c\ln(n)}{\varepsilon^{\prime}}+\frac{6c\ln(n)}{\varepsilon^{\prime}}+\frac{12c\ln(n)}{\varepsilon^{\prime}}$ $>b$ (conditioning on $\mathcal{E}$)
$|M(v)|+|W_{r}(v)|+\frac{24c\ln(n)}{\varepsilon^{\prime}}$ $>b\,.$
In other words,
$|M(v)|$ $>b-|W_{r}(v)|-\frac{24c\ln(n)}{\varepsilon^{\prime}}$
$|M(v)|$ $>b-\frac{16c\ln(n)}{\varepsilon^{\prime}}-\frac{24c\ln(n)}{\varepsilon^{\prime}}$ ( w.p. at least $1-\frac{1}{n^{4c}}$)
$|M(v)|$ $>b-\frac{40c\ln(n)}{\varepsilon^{\prime}}$
$|M(v)|$ $>b-\frac{4\cdot 40c\ln(n)}{\varepsilon}$ (by the choice of $\varepsilon^{\prime}$)
$|M(v)|$ $>1\,.$ (since $b\geq\frac{576c\ln(n)}{\varepsilon}+1$)
In other words, $v$ was already matched from a previous iteration.
Case II.b.2: If $r^{*}\leq\log_{1+\eta}\left(\frac{k\varepsilon^{\prime}}{8c\ln{n}}\right)$, there are two scenarios. Either $r^{*}=0$, then we match $v$ with all nodes which are still available so clearly the number of nodes $v$ is matched with is at least $b^{*}(v)$, or $r^{*}>1$. In the latter scenario, we know that the threshold is exceeded for $r^{*}-1$, so we have $|\widetilde{M}(v)|+|\widetilde{W}_{r^{*}-1}(v)|+12c\ln(n)/\varepsilon^{\prime}\geq b$. Conditioning on $\mathcal{E}$, $|M(v)|+|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Hence, by our assumption that $k\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$, if $|M(v)|=0$, it holds that $|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$ Thus, by a Chernoff bound, with probability at least $1-\frac{1}{n^{4c}}$
$|W_{r^{*}}(v)|\geq\frac{1-\eta}{1+\eta}\mathbb{E}[|W_{r^{*}-1}(v)|]\geq\frac{1-\eta}{1+\eta}(b-36c\ln(n)/\varepsilon^{\prime})\geq 1\,.$
It follows that $v$ is matched with at least 1 neighbor at the end of this iteration.
Taking a union bound over all $r$ and $v\in V$, all Chernoff bounds hold regardless of the value of $r^{*}$ with probability at least $1-\frac{1}{n^{2c}}$ for $c\geq 3$. Taking into account the conditioning on $\mathcal{E}$, our algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for $c\geq 3$.
In summary, if $b^{*}(v)>0$, then $v$ is matched with at least 1 neighbor at the end of the $i$-th iteration. If $b^{*}(v)=0$ and $v$ is not matched, then all of $v$’s neighbors are matched to at least one other node as we showed above. Thus the final output contains a maximal matching as required.
It remains only to check that we match each vertex to at most $b$ neighbors. To do so, it suffices to check that on the $i$-th iteration, each node which does not yet satisfy a matching condition can match with at least 1 more neighbor after Algorithm 2 and that $v=v_{i}$ is matched with at most $b$ neighbors at the end of the iteration (Algorithm 2). The first statement is guaranteed since any vertex $u$ not yet satisfying its matching condition satisfies
$|M(u)|$ $<\tilde{b}(u)-\nu_{i}(u)$
$|M(u)|$ $<b-24c\log(n)/\varepsilon^{\prime}+24c\log(n)$ (conditioning on $\mathcal{E}$)
$|M(u)|$ $<b.$
If $v$ already satisfies its matching condition after Algorithm 2, the second statement is guaranteed to hold. Otherwise, $|M(v)|<b$ and we choose $r$ such that
$b$ $\geq|\tilde{M}_{i}(v)|+|\tilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}$
$\geq|M(v)|+|W_{r}(v)|-12c\ln(n)/\varepsilon^{\prime}+12c\ln(n)/\varepsilon^{\prime}$ (conditioning on $\mathcal{E}$)
$=|M(v)|+|W_{r}(v)|.$
Thus in this case, $v$ is also matched to at most $b$ neighbors, as desired. ∎
Together, Lemma 4.1 and Lemma 4.3 proves Theorem 1.3.
5 $O(\log n)$-Round $\varepsilon$-LEDP Implicit Matchings
In this section, we present a distributed implementation of our matching algorithm that uses $O(\log n)$ rounds in the LEDP model. Specifically, we prove Theorem 1.6, restated below.
See 1.6
High-Level Overview
Our algorithm performs multiple rounds of matching where in each round some nodes are proposers and others are receivers. A node participates in a round if it has not satisfied its matching condition; a node that participates in the current round is an active node. The proposers are chosen randomly and propose to a set of nodes to match. We call the set of nodes each node proposes to as the proposal set. Receivers are active nodes that are not chosen as proposers and receive the proposers’ proposals. Then, each receiver chooses a subset of proposals to accept. The algorithm continues until all nodes satisfy the matching condition.
The algorithm we use to prove Theorem 1.6 modifies Algorithm 2. Our algorithm (the pseudocode is given in Algorithm 3) modifies Algorithm 2 in multiple ways and we briefly comment on the three main changes here. First, we perform several rounds of matching with multiple proposers in each round. Each round uses a fresh set of coin flips. A node may propose in more than one round because many proposers may propose to the same set of receivers; then, many of the proposers will remain unmatched and will need to propose again in a future round. Second, each node that has not satisfied its matching condition determines whether it is a proposer with $1/2$ probability. Third, when all proposers have released their proposal sets, the receivers release their match sets which chooses among the proposers who proposed to them. Then, only the pairs that exist in both the proposal and match sets will be matched.
We show that $O(\log(n))$ rounds are sufficient, with high probability, to obtain a matching that satisfies Theorem 1.6. Notably, we use our Public Vertex Subset Mechanism that we developed in Section 4 as the main routine for both our proposal and match sets.
We give some intuition about why we need to make these changes for the distributed version of our algorithm. We need a fresh set of coin flips for each round because a proposer may participate in multiple rounds. An intuitive reason for why this is the case is due to the fact that some (unlucky) proposers may not be matched to most of the receivers in their proposal sets, in the current round. Hence, a proposer would need to choose another proposal set in the next round. In order to ensure that this new proposal set does not depend on the matched vertices of the previous round, we must flip a new set of coins. Second, because proposers release their sets simultaneously, we cannot have a node simultaneously propose and receive; if a node simultaneously proposes and receives, we do not have a way to ensure that the matching thresholds are not exceeded. Thus, we have a two-round process where in the first synchronous round, proposers first propose and then in the next round, receivers decide the matches.
5.1 Detailed Algorithm Description
We now describe our pseudocode for Algorithm 3 in detail and then prove its properties. We use a number of constants that guarantees our $\Theta(\log n)$ round algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$. Thus, in our pseudocode, some of our constants depend on $c$. Our algorithm returns a maximal matching with high probability when we set $b=\Omega(\log^{2}(n)/\varepsilon)$.
In our pseudocode given in Algorithm 3, we define $M(u)$ to be the set of nodes currently matched to $u\in V$. For each round $i$, we let $V_{i}$ be the set of active vertices which have not reached their matching capacity in round $i$. Our algorithm uses a subroutine (Algorithm 4) which takes a vertex $v$, a set of vertices $S$ from which to select a proposal or matching set, and returns a subgraph index which determines a subset of vertices based on public i.i.d. coin flips. This algorithm implements a more specific version of our Public Vertex Subset Mechanism.
Algorithm 3 takes as input a graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, and a matching parameter $b$ where $b=\Omega(\log^{2}(n)/\varepsilon)$. The algorithm returns an $\varepsilon$-LEDP implicit $b$-matching. First, we set some additional approximation and privacy parameters in Algorithm 3. We then iterate through every node in Algorithm 3 simultaneously to determine the noisy threshold of every node (Algorithm 3). This threshold is used to determine if a vertex satisfies its matching condition. In particular, if the estimated number of possible matches for the node is greater than the threshold, then with high probability, the node is matched to at least one neighbor. The threshold is set in Algorithm 3 using Laplace noise. We add Laplace noise to the threshold as part of an instance of the Multidimensional AboveThreshold (MAT) technique (Lemma 2.15). This is in turn used to determine for every node whether the number of matches exceeds this threshold. We also initialize an empty set for each $u\in V$, denoted $M(u)$, that contains the set of nodes $v$ is matched to (Algorithm 3).
$V_{i}$ contains the set of remaining active nodes in round $i$. Initially, in the first round, all nodes are active (Algorithm 3). We proceed through $O(\log n)$ rounds of matching (Algorithm 3). For each node, we first check using the MAT (simultaneously in Algorithm 3) whether its matching condition has been met. To do so, we add Laplace noise from the appropriate distribution (Algorithm 3) to the size of $u$’s current matches (Algorithm 3). If this noisy size exceeds the noisy threshold, we output that node $u$ has satisfied its matching condition and we remove $u$ from $V_{i}$ (Algorithm 3). Next, we flip the coins for round $i$. These coin flips are used to determine the proposal and match sets. Recall that we produce coin flips for each pair of nodes in the graph. Then, the coin flips are used to determine an implicit set of edges used to match nodes. Although the coin flips are public, only the endpoints of each existing edge knows whether that edge is added to a proposal or match set. All coin flips are done simultaneously and are performed by the curator (Algorithm 3). We flip a coin for each unique pair of vertices and for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. The probability that the coin lands Heads is determined by $r$. Specifically, the coin for the $(i,j,r)$ tuple, denoted $coin(i,j,r)$, is Heads with probability $(1+\eta)^{-r}$ (Algorithm 3). This ensures that the $(r+1)$-th set is, in expectation, a factor of $\frac{1}{1+\eta}$ smaller than the $r$-th set.
We then flip another set of coins to determine which nodes are proposers (Algorithm 3) in round $i$. Each node is a proposer with $1/2$ probability; the result of the coin flip is stored in $a_{i}(v)$ for each $v\in V$. We only select proposers from the set of active vertices, $V_{i}$. For each selected proposer, we simultaneously call the procedure PrivateSubgraph with the inputs $G,\varepsilon^{\prime},\eta,w,V_{i}\setminus P_{i}$ and $coin$ (Algorithms 3 and 3). The pseudocode for PrivateSubgraph is given in Algorithm 4. The function takes as input a graph $G$, a privacy parameter $\varepsilon^{\prime}$, an approximation parameter $\eta$, a vertex subset $S$, a vertex $w\in V$, and all public coin flips $coin$. The function then iterates through all possible subgraph indices $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ (Algorithm 4). For each index, we determine the set of nodes $u$ that satisfy the following conditions: $u\in S$, the edge $\{w,u\}$ exists, and the coin flip $coin(u,w,r)$ is Heads. Since the set $W_{r}(w)$ is computed privately and not released, $w$ has access to the information about whether $\{w,u\}\in E$. This set of nodes is labeled $W_{r}(w)$ (Algorithm 4). We then add Laplace noise to the size of this set to obtain a noisy estimate for the size of $W_{r}(w)$ (Algorithm 4). We use the Adaptive Laplace Mechanism to determine the smallest $r$ (most number of nodes $w$ can propose to) that does not exceed $147c\log(n)/\varepsilon^{\prime}$. To do this, we add Laplace noise to the size of the set of nodes matched to $w$ (Algorithm 4) and find the smallest $r$ such that the sum of the noisy proposal set size, $|\widetilde{W}_{r}(w)|$, and the noisy matched set size, $|\widetilde{M}(w)|$, plus $27c\log(n)/\varepsilon^{\prime}$ does not exceed $147c\log(n)/\varepsilon^{\prime}$ (Algorithm 4). The term $27c\log(n)/\varepsilon^{\prime}$ is added to ensure that $|W_{r}(w)|+|M(w)|$ does not exceed our $b$ bound (by drawing negative noises), with high probability. The procedure releases the smallest subgraph index (Algorithm 4) satisfying Algorithm 4.
We save the released index from PrivateSubgraph in $r_{w}$. Then, $w$ releases $r_{w}$ (Algorithm 3). After the proposers release their subgraph indices, the receivers then determine their match sets. We iterate through all receivers simultaneously and for each receiver $w$, the receiver $w$ computes the set of proposers that proposed to it, denoted as $R$ (Algorithm 3). The receiver can privately compute this set since they know which edges are incident to it, which of their neighbors are proposers, the coin flips of each of the incident edges, and the released subgraph indices of their neighbors. Using the publicly released subgraph indices $r_{v}$ of each neighbor $v$, receiver $w$ can then check $coin(v,w,r_{v})$ to see if edge $\{v,w\}$ is included in proposer $v$’s proposal set. Using $R$, receiver $w$ then computes the match set by calling PrivateSubgraph (Algorithm 3). The receiver releases the subgraph index associated with the match set (Algorithm 3).
The final steps compute the new edges that are in the matching that each node stores privately. We iterate through all pairs of proposers and receivers simultaneously (Algorithm 3). For each pair, $v$ checks whether it is in $W_{r_{u}}(u)$ by checking if $coin(v,u,r_{u})=$ Heads and vice versa for $w$ (Algorithm 3). Then, if the pair is in both the proposal and match sets, $v,u$ matches with each other and $v$ adds $u$ to $M(v)$ and $u$ adds $v$ to $M(u)$ (Algorithm 3). The sets $M(u)$ and $M(v)$ are stored privately but they are computed using the public transcript; hence, the release transcript is the implicit solution that allows each node to know and privately store which nodes it is matched to.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching parameter $b\geq 147c\log^{2}(n)/\varepsilon$
Output: An $\varepsilon$-local edge differentially private ($\varepsilon$-LEDP) implicit $b$-matching
1
2Let $\eta\leftarrow 1/2,\varepsilon^{\prime}\leftarrow\varepsilon/(192c\log_{16/15}(n)),\varepsilon^{\prime\prime}\leftarrow\varepsilon/3$
3 for each node $u\in V$ (simultaneously) do
       $\tilde{b}(u)\leftarrow 38c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}+\text{Lap}(4/\varepsilon^{\prime\prime})$ // Noisy threshold
4       $M(u)\leftarrow\varnothing$
5 end for
[…]
7 for round $i=1$ to $\lceil 32c\cdot\log_{16/15}(n)\rceil$ do
       // MAT for checking matching capacity
8       for each node $u\in V_{i}$ (simultaneously) do
9             $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime\prime})$
10             if $|M(u)|+\nu_{i}(u)\geq\widetilde{b}(u)$ then
11                   Release: node $u$ has satisfied matching condition and remove $u$ from $V_{i}$
12
13             end if
14
15       end for
16
      // Curator flips coins for proposal and match sets; proposal set coins are stored in $coin_{p}$ and matching set coins in $coin_{m}$
17       for each tuple $(i,j,r)$ where $i\in[n]$, $j\in\{i+1,\dots,n\}$ and $r\in\{0,\ldots,\lceil\log_{1+\eta}(n)\rceil\}$ do
18             Flip and release $coin_{m}(i,j,r)$ and $coin_{p}(i,j,r)$ which each lands Heads with probability $p_{r}=(1+\eta)^{-r}$
19
20       end for
21      $P_{i}\leftarrow\{v:a_{i}(v)=\textsc{Heads}\wedge v\in V_{i}\}$ where $a_{i}(v)$ is Heads with probability $p=1/2$
22
23      for $w\in P_{i}$ simultaneously do
             // If $v$ is a proposer, find the proposal set
             $r_{w}\leftarrow$ PrivateSubgraph$(G,\varepsilon^{\prime},\eta,w,V_{i}\setminus P_{i},coin_{p})$   // Algorithm 4
24             Release $r_{w}$
25
26       end for
27
28      for $w\in V_{i}\setminus P_{i}$ simultaneously do
             // Determine proposers who proposed to $w$
29             $R\leftarrow\{u\mid w\in W_{r_{u}}(u)\wedge u\in P_{i}\}$
             // Determine the match set
             $r_{w}\leftarrow$ PrivateSubgraph$(G,\varepsilon^{\prime},\eta,w,R,coin_{m})$   // Algorithm 4
30             Release $r_{w}$
31
32       end for
33      for $v\in P_{i}$ and $u\in V_{i}\setminus P_{i}$ (simultaneously) do
34             if $v\in W_{r_{u}}(u)$ and $u\in W_{r_{v}}(v)$ then
35                   $v,u$ matches; add $u$ to $M(v)$ and $v$ to $M(u)$
36
37             end if
38
39       end for
40      $V_{i+1}\leftarrow V_{i}$
41 end for
Algorithm 3 $O(\log n)$-Round $\varepsilon$-LEDP Maximal Matching
Input: Graph $G=(V,E)$, privacy parameters $\varepsilon^{\prime}$, approximation parameter $\eta$, vertex set $S$, vertex $v\in V$, and public coin flips $coin$
Output: Release subgraph index $r$
1
2Function PrivateSubgraph($G,\varepsilon^{\prime},\eta,w,S,coin$) begin
3       for subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$ do
             // Determine sets with exponentially increasing size
4             $W_{r}(w)\leftarrow\{u\mid u\in S\wedge\{w,u\}\in E\wedge coin(u,w,r)=\textsc{Heads}\}$
5             $|\widetilde{W}_{r}(w)|\leftarrow|W_{r}(w)|+\text{Lap}(4/\varepsilon^{\prime})$
6       end for
      $|\widetilde{M}_{i}(w)|\leftarrow|M(w)|+\text{Lap}(2/\varepsilon^{\prime})$ // Noisy count of matched nodes
       $r\leftarrow\min\left(\{r:|\widetilde{M}_{i}(w)|+|\widetilde{W}_{r}(w)|+27c\log(n)/\varepsilon^{\prime}\leq 147c\log(n)/\varepsilon^{\prime}\}\right)$ // Smallest $r$
7       Return $r$
8 end
9
Algorithm 4 Private Subgraph Release
5.2 Proof of Theorem 1.6
The pseudocode for Theorem 1.6 is presented in Algorithm 3. We prove its privacy and utility guarantees separately. We first prove that it satisfies $\varepsilon$-DP in Section 5.2.1 and then prove its approximation guarantees in Section 5.2.2.
5.2.1 Privacy Guarantees
Our privacy proof follows a similar flavor to the $b$-matching privacy proof given in the previous section except for one main difference. While the adaptive Laplace mechanism was called twice per node in its use in the previous algorithm, this is not the case in our distributed algorithm. Suppose that given neighbor graphs $G$ and $G^{\prime}$ with edge $\{u,v\}$ that differs between the two. Nodes $u$ and/or $v$ can propose a set many times during the course of the algorithm since their proposed sets are not guaranteed to be matched. If they are particularly unlucky, they could propose a set during every round of the algorithm. This means that by composition, we privacy loss proportional to the number of rounds. We formally prove the privacy of our algorithm below.
Lemma 5.1.
Algorithm 3 is $\varepsilon$-LEDP.
Proof.
To prove that Algorithm 3 is $\varepsilon$-LEDP, we show that our algorithm is implemented using local randomizers. Each node $v$ in the algorithm only releases the following set of information in Algorithms 3, 3 and 3: when $v$ has satisfied its matching condition (Algorithm 3), when $v$ releases a proposal subgraph index (Algorithm 3), and when $v$ releases a match set index (Algorithm 3).
Each node implements three different local randomizers for releasing each of the aforementioned three types of information. Our local randomizers are created for each type of private information used to determine the released output. Namely, we use a local randomizer to determine when a node has satisfied the matching condition (Algorithm 3), a local randomizer for computing the noisy proposal or match sets in Algorithm 4 of Algorithm 4, and a local randomizer for determining the noisy set of matched edges in Algorithm 4. The releases in Algorithms 3 and 3 solely depend on the noisy proposal/match sets and the noisy set of matched edges. Then by the concurrent composition theorem, the entire algorithm is differentially private since no other parts of the algorithm use the private graph information.
First, before we dive into the privacy components, we note that the curator flip coins and releases the result of the coins (Algorithms 3 and 3); these coin flips do not lose any privacy since the coins are not tied to private information. The coin flips are performed for each public pair of distinct nodes. In other words, these coins only use the set of nodes which is public. Our proof follows the privacy proof of Lemma 4.1 except we need to account for multiple uses of the adaptive Laplace mechanism for each node.
The first set of local randomizers for all nodes $u\in V$ that releases whether $u$ has satisfied its matching condition can be implemented using an instance of the Multidimensional AboveThreshold (MAT) mechanism with sensitivity. We show how Algorithms 3, 3, 3, 3 and 3 can be implemented using MAT. First, our vectors of queries at each round $i$ is an $n$-length vector which contains the number of nodes each node $u\in V$ has already matched to. Specifically, the vector of queries at each round $i$ is the number of other nodes, $|M(u)|$, each node $u$ has already been matched to before round $i$. Conditioning on the outputs of the previous iterations, the addition or removal of a single edge can only affect $M(u)$ for two nodes, each by at most $1$. Hence, the sensitivity of the vector of queries is $2$; furthermore, the sensitivity of each vector of queries for all rounds is $2$. This means that MAT can be implemented with $\Delta_{M}=2$. Thus, the first set of local randomizers is $\varepsilon^{\prime\prime}$-differentially private by Lemma 2.15.
The second type of local randomizer computes the $|\widetilde{W}_{r}(w)|$ values for each $w\in V$ that is a proposer or receiver. As before, each call of Algorithm 4 can be implemented as a local randomizer using the Adaptive Laplace Mechanism with sensitivity $\Delta=1$; the sensitivity is $1$ because on neighboring adjacency lists, $|W_{r}(w)|$ differs by at most $1$. In fact, given that each edge is selected with probability $p_{r}=(1+\eta)^{-r}$, we can give the privacy guarantee for this local randomizer in terms of $p$ using Lemma 2.17. For a given randomizer and parameter $r$, by Lemma 2.17, the randomizer is $\frac{2}{(1+\eta)^{r}}\cdot\frac{\varepsilon^{\prime}}{4}=\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}$-DP. Then, the sum of the privacy parameters of all calls to this second type of local randomizer on node $w$ is
$\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}=\frac{\varepsilon(\eta n+n-1)}{2\eta\cdot n}\leq\frac{\varepsilon^{\prime}\cdot(\eta+1)}{2\eta}=\frac{3\varepsilon^{\prime}}{2}.$
The last local randomizer implementation is for estimating $|\widetilde{M}(w)|$ in Algorithm 4; as before it is an instance of the Adaptive Laplace Mechanism with sensitivity $1$. Given two neighboring adjacency lists and conditioning on all previous local randomizer outputs, the number of nodes matched to $w$ differs by $1$. Hence, each call of Algorithm 4 can be implemented using a $\frac{\varepsilon^{\prime}}{2}$-local randomizer by Lemma 2.14.
Finally, applying the concurrent composition theorem (Lemma 2.16) proves the privacy guarantee of the entire algorithm. All calls to our MAT local randomizers is $\varepsilon^{\prime\prime}$-DP which simplifies to $\varepsilon/3$-DP by our setting of $\varepsilon^{\prime\prime}$. Then, all calls to our second type of local randomizers incur a privacy loss of at most
$2\cdot\frac{3\varepsilon^{\prime}}{2}\cdot 32c\log_{16/15}(n)=3\cdot\frac{\varepsilon}{192c\log_{16/15}(n)}\cdot 32c\log_{16/15}(n)=\frac{\varepsilon}{2}\,.$
All calls to our third type of local randomizer incur a privacy loss of at most
$2\cdot\frac{\varepsilon^{\prime}}{2}\cdot 32c\log_{1+\eta}(n)=2\cdot\frac{\varepsilon}{2\cdot 192c\log_{1+\eta}(n)}\cdot 32c\log_{1+\eta}(n)=\frac{\varepsilon}{6}\,.$
Thus, all calls to all local randomizers give $\frac{\varepsilon}{3}+\frac{\varepsilon}{6}+\frac{\varepsilon}{2}=\varepsilon$-DP. Hence, our algorithm is $\varepsilon$-LEDP since we can implement our algorithm using local randomizers and our produced transcript preserves $\varepsilon$-differential privacy. ∎
5.2.2 Utility and Number of Rounds
The goal of this section is to prove the following lemma, which gives the formal approximation guarantees and round complexity of Algorithm 3.
Lemma 5.2.
For any $\varepsilon\in(0,1)$, Algorithm 3 returns a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum matching, with high probability, when $b=\Omega(\log^{2}(n)/\varepsilon)$, in $O(\log(n))$ rounds.
Towards proving Lemma 5.2, we first prove the following lemma. Recall that our algorithm randomly divides the population into (roughly) half proposers and half receivers. Then, the proposers first propose their proposal sets, and then the receivers respond with their match sets for each of $32c\cdot\lceil\log_{16/15}(n)\rceil$ rounds. The following crucial lemma gives the “progress” of the algorithm after each round. Namely, to ensure we produce a maximal matching in $O(\log n)$ rounds, with high probability, roughly a constant fraction of the nodes would need to exceed their matching capacity every round (hence, would not participate in future matching rounds). For convenience in the below proofs, when we write $\log(n)$, we mean $\log_{1+\eta}(n)$.
In the analysis below, we call the set of nodes that have not exceeded their matching capacity and have a non-zero number of neighbors who have not reached their matching capacity, the hopeful nodes. A node that is not hopeful is called unhopeful.
First, note that every active node in any round $i$ is either a proposer or receiver by definition. We first prove the following lemma, which states that a vertex $w$ will choose $r_{w}=0$ with high probability, as a proposer or receiver if its degree is less than $27c\log(n)/\varepsilon$.
Lemma 5.3.
If a node $w\in V_{i}$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$ among the hopeful nodes in round $i$, then it will choose $r_{w}=0$ with probability at least $1-\frac{1}{n^{5c}}$.
Proof.
Suppose $X\sim\text{Lap}(4/\varepsilon^{\prime})$ and $Y\sim\text{Lap}(2/\varepsilon^{\prime})$ are the noise random variables chosen in Algorithms 4 and 4, respectively. Since $w$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$, it will pick $r_{w}=0$ if $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$. By Lemma A.1, $X\leq 24c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$ and $Y\leq 12c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$. Then, $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{5c}}$ when $n\geq 2$. ∎
We now prove an additional lemma about the size of any proposal or match set an active vertex in round $i$ will choose. In particular, if a vertex is active in round $i$, then it will propose a proposal set or match set of size at least $32c\log_{16/15}(n)/\varepsilon^{\prime}$, with high probability.
Lemma 5.4.
If vertex $w\in V_{i}$ is active in round $i$, then $w$ will pick $r_{w}=0$ or an $r_{w}$ where $|W_{r_{w}}(w)|\geq 32c\log_{16/15}(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{2c}}$.
Proof.
If $w$ is still active in round $i$, this means that $w$ did not meet its matching condition in round $i-1$. By Lemma A.1, this means that with probability at least $1-\frac{1}{n^{4c}}$, it holds that $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ because the noise drawn in Algorithm 3 is at least $-12c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$ and the noise drawn in Algorithm 3 is at most $24c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since, $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}<<27c\log(N)/\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}>>\varepsilon^{\prime}$, by our proof for Lemma 5.3, node $w$ will choose a $|W_{r_{w}}(w)|$ can be as large as $50c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since we chose $\eta=0.5$, we will pick an $r_{w}$ where $33c\log(n)/\varepsilon^{\prime}<(50\cdot 2)/3c\log(n)/\varepsilon^{\prime}\leq\mathbb{E}[|W_{r}(w)|]\leq 50c\log(n)/\varepsilon^{\prime}$. Otherwise, if the expectation is less than $33c\log(n)/\varepsilon^{\prime}$, then $w$ will pick all available edges (and hence pick $r_{w}=0$). Now we show the concentration of the size with sufficiently high probability via the Chernoff bound. We can use the Chernoff bound since the edges picked for $W_{r_{w}}(n)$ are i.i.d. chosen at random using coins that are flipped heads with probability $(1+\eta)^{-r_{w}}$. Specifically, by Theorem A.2, if we take $\psi=0.4$, then, $|W_{r_{w}}(w)|<32c\log_{16/15}(n)/\varepsilon^{\prime}$ with probability at most $\exp\left(-\frac{0.4^{2}\cdot 33c\log(n)}{3\varepsilon^{\prime}}\right)<\frac{1}{n^{3c}}$. Hence, together, $w$ will pick an $r_{w}$ where $|W_{r_{w}}(w)|\geq 32c\log_{16/15}(n)/\varepsilon^{\prime}$ or it will pick $r_{w}=0$. ∎
Using our above lemmas, we will now prove our main lemma that enough progress is made in each round of our algorithm such that we find a maximal matching at the end of all of our rounds. Specifically, this lemma shows that with a large enough constant probability, a constant fraction of the remaining edges in the graph become unhopeful. An edge is unhopeful if at least one of its endpoints is unhopeful.
Lemma 5.5.
In each round $i$ of Algorithm 3 with $H$ hopeful edges, at least $H/16$ edges become unhopeful with probability at least $1/16$.
Proof.
Let $H$ be the set of hopeful edges in round $i$ where a hopeful edge is one where both endpoints of the edge are hopeful. We show that in each of $32c\log_{16/15}(n)$ rounds, a constant fraction of hopeful edges become unhopeful with high constant probability.
By definition of a hopeful edge, $e=\{u,v\}$, if either $u$ or $v$ becomes a proposer then they will propose a non-empty set (if they have at least one receiver neighbor). The marginal probability that any hopeful edge has one endpoint become the proposer and the other remains a receiver is $1/4$ by Algorithm 3. We now follow a charging scheme inspired by the edge orientation charging scheme given in [khuller1994primal]; however, our charging scheme is fundamentally different since we are performing charging in a noisy process for $b$-matching. Note that we do not actually perform any orientation algorithm; this orientation charging scheme is only used for the analysis. In this orientation charging scheme, we orient all hopeful edges from low to high degree endpoints in the induced graph consisting of $H$. We denote a vertex $v$ as good if at least $1/3$ of its degree is oriented into it. We call an edge good if it is oriented into a good vertex. We first show that at least $1/2$ of $H$ is good. We call a vertex or edge bad if it is not good.
A vertex is bad if more than $2/3$ of its degree is oriented out. These edges that are oriented out are bad edges if and only if they are oriented toward bad vertices. Furthermore, any bad vertex must either have degree $1$ or have at least $2$ outgoing edges for every bad incoming edge. A degree one vertex has a bad outgoing edge if it is paired with another vertex with at least two outgoing edges. Hence, for any node we can charge each incoming bad edge to a unique outgoing edge such that each outgoing edge receives at most one charge. Suppose the number of incoming bad edges charged to good edges is $H^{\prime}$, then there are at least $H^{\prime}$ good edges. Finally, there can be at most $\frac{1}{2}(H-H^{\prime})$ bad edges charged to bad edges, since any incoming bad edge must be charged to at most one outgoing edge and every bad vertex has at least two outgoing edges for each incoming edge. Hence, in total, at least $H/2$ edges are good edges.
We now show that at least a fixed constant fraction of the hopeful, good edges become unhopeful each round with high constant probability. For each good, hopeful edge $e=(u,v)$, oriented from $u$ to $v$, the marginal probability that the source of the edge is a proposer and the sink is a receiver is $1/4$. Let the degree of $v$ in the induced subgraph consisting of all hopeful edges be $\deg(v)$; since $e$ is oriented from $u$ to $v$, it means that $\deg(u)\leq\deg(v)$. The marginal probability that $u$’s proposed set contains $v$ is at least
$\min\left(1,\frac{32c\log_{16/15}(n)}{\varepsilon^{\prime}\cdot\deg(u)}\right)\geq\min\left(1,\frac{32c\log_{16/15}(n)}{\varepsilon^{\prime}\cdot\deg(v)}\right)$
by Lemma 5.4. Thus, the probability that $u$ is a proposer, $v$ is a receiver, and $u$ proposes $v$ is at least $\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$. Receiver $v$ can either choose to match with $u$ or not. If $v$ does not match with $u$, that means by Lemma 5.4, $v$ picked a set that did not contain $u$ and has reached its matching threshold, with probability at least $1-\frac{1}{n^{3c}}$.
The expected number of edges $v$ receives as part of all proposals is then
$\deg(v)\cdot\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(u)}\right)$ $\geq\deg(v)\cdot\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$
$=\min\left(\deg(v),\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}}\right)\,.$
Thus, in expectation, $v$ satisfies its matching condition. Now, we show the concentration on the probability that $v$ satisfies the matching condition. By the Chernoff bound and suppose $\psi=0.4$, the probability that $<10c\log(n)/(4\varepsilon^{\prime})$ proposers propose to $v$ is upper bounded by
$\exp\left(\frac{0.4^{2}\cdot 32c\log_{16/15}(n)/\varepsilon^{\prime}}{12}\right)<\exp\left(\frac{2c\log(n)}{12\varepsilon^{\prime}}\right)<\frac{1}{n^{4c}}\,.$
Hence, by the union bound over $n$ nodes, with probability at least $1-\frac{1}{n^{3c}}$, each good receiver satisfies its matching condition. Each hopeful, good edge has $1/4$ probability that it is going from a proposer to a receiver. The expected number of hopeful edges that are good and are going from proposers to receivers is then at least $\frac{H}{8}\cdot(1-1/n^{3})$; these are also the expected number of hopeful edges that become unhopeful. By the Markov bound, the probability that at most $1615H/$ 16 hopeful edges remain hopeful is $\frac{7/8+\frac{1}{n^{3}}}{15/16}=14/15+\frac{15}{16n^{3}}$. This means that with probability at least $1/16$ (assuming $n>10$—if $n\leq 10$, we can run our algorithm from Section 4 in $O(1)$ rounds), at least $H/16$ of the hopeful edges become unhopeful. ∎
Using all of the above lemmas, we prove Lemma 5.2, which states the final utility guarantee of our algorithm: It is guaranteed to return a maximal matching in $O(\log n)$ rounds, with high probability.
Proof of Lemma 5.2.
By Lemma 5.5, in each round $i$, at least $1/16$ of the hopeful edges become unhopeful with probability at least $1/16$. In $32\cdot 16c\log(n)$ rounds, the expected number of rounds for which at least $1/16$ of the hopeful edges become unhopeful is $32c\log(n)$ rounds. Because the Chernoff bound holds for binomials, by Theorem A.2, if we set $\psi=0.5$, then the probability that less than $16c\log(n)$ rounds are successful is at most $\exp\left(\frac{0.5^{2}\cdot 32c\log(n)}{3}\right)<\frac{1}{n^{8c}}$. Hence, in $O(\log n)$ rounds, all edges become unhopeful.
We now show that if all edges become unhopeful, then all nodes are either matched or all of their neighbors are matched. An edge becomes unhopeful if at least one of its endpoints satisfies the matching condition. By Lemma A.1, a node satisfies the matching condition if and only if $|M(u)|\geq 1$ with probability at least $1-\frac{1}{n^{2c}}$. Hence, if a node is adjacent to all unhopeful edges, then it either satisfies the matching condition and is matched to at least one neighbor with high probability, or it is adjacent to endpoints which are matched with high probability.
Finally, by Lemma A.1, $|M(u)|\geq 2\cdot 147c\log(n)/\varepsilon^{\prime}$ with probability at most $\frac{1}{n^{2c}}$. By Lemma 4.2 and a union bound over all probabilities of failure, with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$, we output a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum matching in $O(\log n)$ rounds when $b=\Omega(\log^{2}(n)/\varepsilon)$.
The number of rounds of our algorithm is determined by Algorithm 3 (which is $O(\log n)$) since round contains a constant number of synchronization points and all nodes (proposers and receivers) perform their instructions simultaneously. ∎
Combining Lemma 5.1 and Lemma 5.2 yields the proof of Theorem 1.6.
6 Node Differentially Private Implicit Matchings
Using sparsification techniques, we demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. In particular, we prove Theorem 1.8, which we restate below. See 1.8
The class of bounded arboricity [^8] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree $n-1$ and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. For simplicity of presentation, we first design an algorithm assuming a public bound $\tilde{\alpha}$. Then our algorithm is always private but the approximation guarantees hold when $\tilde{\alpha}$ upper bounds the arboricity of the input graph. When such a public bound is unavailable, we show that guessing the bound with powers of 2 achieves the same guarantees up to logarithmic factors.
The key idea is that a judicious choice of a sparsification algorithm reduces the edge edit distance between node-neighboring graphs to some factor $\Lambda=O(\alpha)$. Such sparsification is useful since, in the worst case, the edge edit distance pre-sparsification can be $\Omega(n)$. By group privacy, it then suffices to run any edge-DP algorithm with privacy parameter $\nicefrac{{\varepsilon}}{{\Lambda}}$ after sparsification to achieve node-privacy.
6.1 Bounded Arboricity Sparsifiers
[…]
Matching Sparsifier
[…]
Proposition 6.1.
Let $\pi\in P_{\binom{n}{2}}$ be a total ordering over unordered vertex pairs and $G\sim G^{\prime}$ be node neighboring graphs. Then the edge edit distance between $H\coloneqq\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ and $H^{\prime}\coloneqq\textsc{ContractionSparsify}_{\pi}(G^{\prime},\Lambda)$ is at most $2\Lambda$.
Proof.
Let $S_{G}$ and $S_{G^{\prime}}$ be the sparsified graphs of $G$ and $G^{\prime}$, respectively. Suppose without loss of generality that $G^{\prime}$ contains $E_{\operatorname{extra}}$ additional edges incident to vertex $v$ and $v$ has degree 00 in graph $G$. Then, for each edge $\{v,w\}\in E_{\operatorname{extra}}$, let $e^{w}_{\operatorname{last}}$ be the edge adjacent to $w$ in $G$ whose index in $\pi$ is the last among the edges incident to $w$ in $S_{G}$. If $i_{\pi}(e^{w}_{\operatorname{last}})>i_{\pi}(\{v,w\})$ (where $i_{\pi}(\{v,w\})$ is the index of edge $\{v,w\}$ in $\pi$), then $\{v,w\}$ replaces edge $e^{w}_{\operatorname{last}}$. Since both $G$ and $G^{\prime}$ are simple graphs, at most one edge incident to $w$ gets replaced by an edge in $E_{\operatorname{extra}}$ in $S_{G^{\prime}}$. This set of edge replacements leads to an edge edit distance of $2\Lambda$. ∎
The original sparsification algorithm in [solomon2021local] marks an arbitrary set of $\min(\deg(v),\Lambda)$ edges incident to each vertex $v$ and takes the subgraph consisting of all edges marked by both endpoints. In our setting, $\pi$ determines the arbitrary marking in our graphs. Hence, our sparsification procedure satisfies the below guarantee.
Theorem 6.2 (Theorem 3.3 in [solomon2021local]).
Let $G$ be a graph of arboricity at most $\alpha$ and $\Lambda\coloneqq 5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha$ for some $\eta\in(0,1]$. Suppose $H$ is obtained by marking an arbitrary set of $\min(\deg(v),\Lambda)$ edges incident to each vertex $v$ and taking the subgraph consisting of all edges marked by both endpoints. Then if $\mu(\cdot)$ denotes the size of a maximum matching of the input graph,
$\mu(H)\leq\mu(G)\leq(1+\eta)\mu(H).$
In particular, any $(\beta,\zeta)$-approximate maximum matching for $H$ is an $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$. Moreover, this holds for $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ where $\pi\in P_{\binom{n}{2}}$ is a total order over unordered vertex pairs.
6.2 Proof of Theorem 1.8
6.2.1 Node-DP Maximum Matching with Public Arboricity Bound
In this section, we design a node-DP algorithm to output an implicit matching in the central model using the sparsification techniques derived in Section 6.1 and any edge DP algorithm (e.g. Theorem 1.3) that outputs implicit solutions in the $\varepsilon$-DP setting.
Theorem 6.3.
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, there is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log(n)}{\eta\varepsilon}\right)$and (ii) if $\tilde{\alpha}\geq\alpha$, the implicit solution is at least $\frac{1}{2+\eta}$ the size of a maximum matching.
We emphasize that our privacy guarantees always hold but the utility guarantee is dependent on the public bound $\tilde{\alpha}$.
Proof Sketch.
The algorithm first discards edges according to Theorem 6.2 until there are at most $\Lambda=O(\tilde{\alpha}/\eta)$ edges incident to each vertex. This ensure that the edge-edit distance of node-neighboring graphs is at most $2\Lambda$ (Proposition 6.1). We can thus run any $(\frac{\varepsilon}{2\Lambda})$-edge DP algorithm to ensure $\varepsilon$-node DP. ∎
6.2.2 Removing the Assumption of Public Bound
In practice, it is not always possible to obtain a good public bound on the arboricity. However, we can always compute an $\varepsilon$-node DP estimate of the arboricity up to $O(\log(n)/\varepsilon)$ additive error since the sensitivity of the arboricity is $1$ even for node-neighboring graphs. This incurs a slight blow-up in $b$ and accounting for this blowup leads to a proof of Theorem 1.8.
7 Improved Node-DP Implicit Bipartite Matchings
In this section, we restate and prove Theorem 1.10. See 1.10
As in [hsu2014private], recall that we have a bipartite graph $G=(V_{L}\cup V_{R},E)$, where we think of the left nodes in $V_{L}$ as items and the right nodes in $V_{R}$ as bidders. Let’s say there are $n=|V_{R}|$ bidders and $k=|V_{L}|$ items. We define two of these bipartite graphs to be neighbors if they differ by a single bidder and all edges incident to the bidder. Our goal is to implicitly output an allocation of items to bidders such that each item is allocated to at most $s$ bidders and each bidder is allocated at most $1$ item, while guaranteeing differential privacy for the output.
Our algorithm for node-private bipartite matching follows the same template as that of [hsu2014private], based on a deferred acceptance type algorithm from [kelso1982job]. For each item, [kelso1982job] runs an ascending price auction where at every iteration, each item $u$ has a price $p_{u}$. Then, over a sequence of $T$ rounds, the algorithm processes the bidders in some publicly known order and each one bids on their cheapest neighboring item which has price less than $1$ (each edge indicates a potential maximum utility value of $1$ for the item). At any moment, the $s$ most recent bidders for an item are tentatively matched to that item, and all earlier bidders for it become unmatched.
In the private implementation of this algorithm in [hsu2014private], they keep private continual counters for the number of bidders matched with each item which they continually output. They additionally output the sequence of prices for each item. Using this information, each bidder can reconstruct the cheapest neighboring item at the given prices. They then send the bit 1 to the appropriate counter, and store the reading of the counter when they made the bid. When the counter indicates that $s$ bids have occurred since their initial bid, the bidder knows that they have become unmatched. The final matching which is implicitly output is simply the set of edges which have not been unmatched.
The main difference in our algorithm is in implementing the counters. In the private algorithm of [hsu2014private], each of these counters need to be accurate at each of the $nT$ iterations. This causes significant privacy loss. Our algorithm has a different implementation (of the same basic algorithm), where we use the fact that only the counts at the end of each of the $T$ iterations are important. Moreover, our implementation does not need the counts themselves, but just needs to check whether they are above the supply $s$. This allows us to use the Multidimensional AboveThreshold Mechanism (Algorithm 1) to get further improvements.
1 Input: Graph $G=(V,E)$, approximation factor $\eta\in(0,1)$, privacy parameter $\varepsilon>0$, matching parameter $b=\Omega(\log(n)/(\eta^{4}\varepsilon))$.
2 Output: An $\varepsilon$-node differentially private implicit $(1+\eta)$-approximate $b$-matching.
Initialization : $T\leftarrow 73/\eta^{2}$, $\varepsilon^{\prime}\leftarrow\varepsilon\eta/3T$, $\text{start}_{u}\leftarrow 0$, $\text{end}_{u}\leftarrow nT$, $p_{u}\leftarrow\eta$, and $c_{u}\leftarrow 0$ for each $u\in V_{L}$.
3 for $t\leftarrow 1$ to $T$ do
4       foreach item $u\in V_{L}$ do
5             Initialize noisy threshold $\tilde{t}_{u}\leftarrow(p_{u}/\eta)\cdot s+\mathrm{Lap}(2/\varepsilon^{\prime})$
6
      // Each bidder $v$ proposes to the neighbor $u$ with lowest price
7       foreach bidder $v\in V_{R}$ do
8             if $v$ is not matched then
9                   Choose a neighbor $u$ of $v$ with smallest $p_{u}$, if any exist, breaking ties arbitrarily
10                   if $p_{u}\leq 1$ then
11                         Denote bidder $v$ as (temporarily) matched with item $u$ for timestep $t$
12                         $c_{u}\leftarrow c_{u}+1$
13
14
15
16      foreach item $u\in V_{L}$ do
17             Release $p_{u}$
18             if $p_{u}\leq 1$ and $c_{u}+\mathrm{Lap}(4/\varepsilon^{\prime})\geq\tilde{t}_{u}$ then
19                   $p_{u}\leftarrow p_{u}+\eta$
20                   Re-initialize the noisy threshold $\tilde{t}_{u}\leftarrow(p_{u}/\eta)\cdot s+\mathrm{Lap}(2/\varepsilon^{\prime})$
21
22
      // Only the most recent (approximately) $s$ bidders for each item $u$ remain matched with $u$
23       foreach item $u\in V_{L}$ do
24             Release startu
25             $\tilde{s}_{u}\leftarrow s+\mathrm{Lap}(2/\varepsilon^{\prime})-18c\log(n)/\varepsilon^{\prime}$
26             Let $M_{u}$ denote the number of bidders matched with $u$, between timesteps $\text{start}_{u}$ and $\text{end}_{u}$
27             while $M_{u}+\mathrm{Lap}(4/\varepsilon^{\prime})\geq\tilde{s}_{u}$ do
28                   if a bidder $v$ matched with item $u$ at timestep $\text{start}_{u}$ then
29                         Mark bidder $v$ as unmatched
30
31                  $\text{start}_{u}\leftarrow\text{start}_{u}+1$ and update $M_{u}$ accordingly
32
33
      // Terminate early if the algorithm is close to convergence
34       Let $c_{0}$ denote the number of bids made in this iteration
35       if $c_{0}+\mathrm{Lap}(1/\varepsilon^{\prime})\leq\eta\cdot\mathrm{OPT}/3-3c\log(n)/\varepsilon^{\prime}$ then
36             Terminate the algorithm
37
38
Algorithm 5 Node-Differentially Private $(1+\eta)$-Approximate Maximum $b$-Matching
7.1 Proof of Theorem 1.10
The pseudocode for Theorem 1.10 is presented in Algorithm 5. Following the structure of previous sections, we partition the proof of Theorem 1.10 into its privacy and utility proofs. We first present the privacy proof in Section 7.1.1 and then the approximation proof in Section 7.1.2. Finally, we bring the arguments together in Section 7.1.3 to complete the proof of Theorem 1.10.
7.1.1 Privacy Proof
First, we describe the implicit output and how each bidder reconstructs who they are matched with. Like in the previous algorithm given in [hsu2014private], for each of the $nT$ iterations, the price of each item is outputted, creating a public sequence of prices for each item over the $nT$ iterations. At each iteration, the outputted prices can be used by the bidder to determine their cheapest neighbor. The algorithm also outputs $\text{start}_{u}$ and $\text{end}_{u}$ for each item $u$ at the end of each of the $T$ rounds. This indicates that only bidders which were matched with item $u$ between iteration $\text{start}_{u}$ and $\text{end}_{u}$ are matched with $u$.
Next, we show that the output of the algorithm is $\varepsilon$-node differentially private. On a high level, the proof follows by showing that the algorithm consists of (concurrent) compositions of instances of MAT and the Laplace mechanism. The pseudocode for our algorithm is given in Algorithm 5 which uses a constant $c\geq 1$ which is the constant used in the high probability $1-\frac{1}{n^{c}}$ bound.
Theorem 7.1.
Algorithm 5 is $\varepsilon$-node differentially private.
Proof.
Our algorithm releases the $p_{u}$ and startu of every item for each of the $T$ iterations. Hence, we show that these releases are $\varepsilon$-node differentially private. We first show that Algorithm 5 to Algorithm 5 can be seen as $73/\eta^{2}$ instances of the Multidimensional AboveThreshold mechanism (Algorithm 1) with privacy parameter $\varepsilon^{\prime}$ and queries $c_{u}$ for $u\in V_{L}$. Observe that each bidder $v\in V_{R}$ bids at most $T$ times, once per each of the $T$ iterations. As a result, the $\ell_{1}$ sensitivities of the count queries $c_{u}$ is at most $T$ since one additional bidder will bid on any item at most $T$ times. Furthermore, the noisy threshold $\tilde{t}_{u}$ is used at most $1/\eta$ times since after the prices exceed $1$, they are updated but no longer used. For each update of the threshold, the outputs are $T\varepsilon^{\prime}$-differentially private, so the entirety of the outputs of Algorithm 5 to Algorithm 5 is $T\varepsilon^{\prime}/\eta$-differentially private.
Similarly, for each of the $T$ iterations, we will show that Algorithm 5 to Algorithm 5 can be seen as an instance of the Multidimensional AboveThreshold mechanism with privacy parameter $\varepsilon^{\prime}$ and queries $M_{u}$ for $u\in V_{L}$. Recall that $M_{u}$ is the number of bidders matched with $u$ between $\text{start}_{u}$ and $\text{end}_{u}$. Since each bidder can only match with $1$ item at a time, the sensitivity of the queries is $1$, so each iteration of Algorithm 5 to Algorithm 5 is $\varepsilon^{\prime}$-differentially private. Applying the concurrent composition theorem over $T$ iterations, we have that Algorithm 5 to Algorithm 5is $T\varepsilon^{\prime}$-differentially private.
Finally, for each of the $T$ rounds, we have that Algorithms 5, 5 and 5 can be analyzed as an instance of the Laplace mechanism with privacy parameter $\varepsilon^{\prime}$ and query $c_{0}$. In a given iteration, each bidder makes only a single bid so the sensitivity of the query is $1$, implying that this part is $\varepsilon^{\prime}$-differentially private. Applying the concurrent composition theorem, we have that Algorithms 5, 5 and 5 is $T\varepsilon^{\prime}$-differentially private. Finally, by concurrent composition, the entire algorithm is $T\varepsilon^{\prime}+T\varepsilon^{\prime}+T\varepsilon^{\prime}/\eta\leq\varepsilon$-differentially private. ∎
7.1.2 Utility Proof
On a high level, the utility proof shows that $T$ iterations of the deferred acceptance algorithm suffices to match each bidder with a suitable item. Slightly more formally, we say a bidder is satisfied if they are matched with an $\eta$-approximate favorite good (i.e., $w_{\mu(v),v}-p_{\mu(v)}\geq w_{u,v}-p_{u}-\eta$ for all items $u$ where $\mu(v)$ indicates the item the bidder is matched with). Such a notion has been referred to, in the matching literature, as “happy” bidders [ALT21, LKK23]. We show below that at least a $(1-\eta)$-fraction of the bidders are satisfied, which directly leads to our approximation bound.
Lemma 7.2.
Assume that each Laplace random variable satisfies $|\mathrm{Lap}(\beta)|\leq 3c\beta\log(n)$ and $s\geq 72c\log(n)/(\eta\varepsilon^{\prime})$ for constant $c\geq 1$. Then, we have that at most $\eta\cdot\text{OPT}$ bidders are unsatisfied at termination.
Proof.
Observe that the number of unsatisfied bidders is exactly the number of bidders who were unmatched (Algorithm 5) by their item in the final round. We will first prove the claim when the algorithm terminates early in Algorithm 5. We will then show that the algorithm always terminates early under our assumptions on $s$ and our choice of $T=73/\eta^{2}$. Combining the two gives the desired claim.
If the algorithm terminates early, then we have $c_{0}+\mathrm{Lap}(1/\varepsilon^{\prime})\leq\eta\cdot\text{OPT}/3-3c\log(n)/\varepsilon^{\prime}$ in Algorithm 5, implying that the number of bids at the final round is at most $Q\coloneqq\eta\cdot\text{OPT}/3$ by our assumption that $\mathrm{Lap}(1/\varepsilon^{\prime})\leq 3c\log(n)/\varepsilon^{\prime}$. Let $N$ be the number of items which unmatched (Algorithm 5) with some bidder in this iteration. Since each such item is matched with at least $s-36c\log(n)/\varepsilon^{\prime}$ bidders at the end of the round (due to Algorithm 5), the total number of bidders who matched with these goods at the beginning of the round must be at least
$(s-36c\log(n)/\varepsilon^{\prime})N-Q.$
Next, observe that at most OPT bidders can be matched at the same time, by definition of OPT. Combining with the above inequality, we have that
$N\leq(\text{OPT}+Q)/(s-36c\log(n)/\varepsilon^{\prime}).$
Thus, the total number of bidders who were unmatched is at most
$sN-[(s-36c\log(n)/\varepsilon^{\prime})N-Q]$ $=36cN\log(n)/\varepsilon^{\prime}+Q$
$\leq\frac{36c\log(n)/\varepsilon^{\prime}}{s-36c\log(n)/\varepsilon^{\prime}}\cdot\left(\text{OPT}+Q\right)+Q$
$\leq\frac{36c\log(n)/\varepsilon^{\prime}}{s-36c\log(n)/\varepsilon^{\prime}}\cdot\left(\text{OPT}+\frac{\eta\cdot\text{OPT}}{3}\right)+Q.$
For $s\geq(1+2/\eta)72c\log(n)/\varepsilon^{\prime}=\Theta(\log(n)/\varepsilon\eta^{4})$, the above expression is upper bounded by $\eta\cdot\text{OPT}$, as desired.
If the algorithm doesn’t terminate early, then we have $c_{0}+\mathrm{Lap}(1/\varepsilon^{\prime})>\eta\cdot\text{OPT}/3-3c\log(n)/\varepsilon^{\prime}$ in Algorithm 5 for each iteration of the algorithm. This implies that the number of bids at each of the $T$ iterations of the algorithm is at least $\eta\cdot\text{OPT}/3-36c\log(n)/\varepsilon^{\prime}$. This implies that the total number of bids over all the iterations is lower bounded by
$B\geq T\cdot(Q-36c\log(n)/\varepsilon^{\prime}).$ (1)
As before, we have that at most OPT bidders can be matched at the same time. There are at most OPT bids on the under-demanded goods, since bidders are never unmatched with these goods. Furthermore, each of the over-demanded goods are matched with at least $s-36c\log(n)/\varepsilon^{\prime}$ bidders, so there are at most $\text{OPT}/(s-36c\log(n)/\varepsilon^{\prime})$ bidders. Since each such good takes at most $s+36c\log(n)/\varepsilon^{\prime}$ bids at each of the $1/\eta$ price levels, the total number of bids is thus upper bounded by
$B\leq\text{OPT}+\frac{\text{OPT}}{\eta}\left(\frac{s+36c\log(n)/\varepsilon^{\prime}}{s-36c\log(n)/\varepsilon^{\prime}}\right)\leq\frac{6\cdot\text{OPT}}{\eta},$ (2)
where the second inequality holds since $s\geq 72c\log(n)/\varepsilon^{\prime}$.
Combining our two estimates in Eqs. 1 and 2, we have
$T\cdot(Q-36c\log(n)/\varepsilon^{\prime})\leq B\leq\frac{6\cdot\text{OPT}}{\eta},$
implying that
$T\leq\frac{6\cdot\text{OPT}}{\eta}\cdot\left(\frac{1}{\eta\cdot\text{OPT}/3-36c\log(n)/\varepsilon^{\prime}}\right)\leq\frac{72}{\eta^{2}},$
where we have used that $\eta\cdot\text{OPT}\geq\eta\cdot s\geq 144c\log(n)/\varepsilon^{\prime}$ if there are at least $s$ edges; otherwise, if there are less than $s$ edges, the returned matching will equal the number of edges on the first iteration. Thus, this is a contradiction since $T=\frac{73}{\eta^{2}}$, so we can conclude that the algorithm must terminate early. ∎
7.1.3 Putting it Together
Proof of Theorem 1.10.
First, observe that for any Laplace random variable $\mathrm{Lap}(\beta)$, we have that $\mathrm{Lap}(\beta)\leq 3c\beta\log(n)$ with probability at least $1-1/n^{3c}$. There are $O(n^{2})$ total Laplace random variables in the algorithm, so a union bound implies each of them satisfies the concentration with probability at least $1-1/n^{c}$. We condition on this event for the remainder of the analysis. In particular, we have that at most $s$ bidders are matched with each item $u$, since at most $s$ bidders between $\text{start}_{u}$ and $\text{end}_{u}$ are matched with $u$ in the while loop starting in Algorithm 5. Now, we can start the analysis.
For each edge $(u,v)\in E$, set $w_{u,v}=1$; set $w_{u,v}=0$ otherwise. Let $\mu$ denote the matching (implicitly) output by the algorithm, and consider the optimal matching $\mu^{*}$. For each matched bidder $v$, we have
$w_{\mu(v),v}-p_{\mu(v)}\geq w_{\mu^{*}(v),v}-p_{\mu^{*}(v)}-\eta.$
Since there are at most OPT such bidders (call them $S$), summing the above over all $S$ gives
$\sum_{v\in S}[w_{\mu(v),v}-p_{\mu(v)}]\geq\sum_{v\in S}[w_{\mu^{*}(v),v}-p_{\mu^{*}(v)}]-\eta\cdot\text{OPT}.$
Let $N_{u},N_{u}^{*}$ be the number of times $u$ is respectively matched in $\mu,\mu^{*}$. Then rearranging gives
$\sum_{v\in S}[w_{\mu^{*}(v),v}-w_{\mu(v),v}]\leq\sum_{u\in V_{L}}[p^{*}_{u}\cdot N_{u}^{*}-p_{u}\cdot N_{u}]-\eta\cdot\text{OPT}.$
Next, observe that if a good $u$ has $p_{u}>0$, this means that at least $s-18c\log(n)/\varepsilon^{\prime}$ bidders were (temporarily) matched with it, due to Algorithm 5. This directly implies that the number of bidders matched with $u$ at the termination is at least $s-36c\log(n)/\varepsilon^{\prime}$, because goods only unmatch with bidders until they are matched with $s-36c\log(n)/\varepsilon^{\prime}$ bidders (Line 5). Thus, there can be at most $\text{OPT}/(s-36c\log(n)/\varepsilon^{\prime})$ goods with $p_{u}>0$. For each of these goods, we have $p^{*}_{u}\cdot N_{u}^{*}-p_{u}\cdot N_{u}\leq s$, so we have
$\sum_{v\in S}[w_{\mu^{*}(v),v}-w_{\mu(v),v}]\leq\frac{\text{OPT}\cdot s}{s-36c\log(n)/\varepsilon^{\prime}}-\eta\cdot\text{OPT}.$
Finally, Lemma 7.2 implies that at most $\eta\cdot\text{OPT}$ bidders are not in $S$. Summing over all bidders, we have
$\sum_{v\in V_{R}}[w_{\mu^{*}(v),v}-w_{\mu(v),v}]\leq\frac{\text{OPT}\cdot s}{s-36c\log(n)/\varepsilon^{\prime}}+\eta\cdot\text{OPT}-\eta\cdot\text{OPT}.$
Since we have $s\geq(1+2/\eta)\cdot 72c\log(n)/\varepsilon^{\prime}$, the first term on the right hand side is at most $2\eta\cdot\text{OPT}$. Scaling down $\eta$ by a constant factor $1/2$ gives the desired result with probability at least $1-\frac{1}{n^{c}}$ for constant $c\geq 1$. ∎
8 Arbitrary Edge-Order Continual Release Implicit Matchings
We now give algorithms for matchings in the continual release model. We give algorithms that satisfy edge-DP and node-DP in two different types of input streams. Specifically, we consider two types of insertion-only streams where nodes and edges can be inserted but not deleted: 1) arbitrary-order edge arrival streams and 2) adjacency-list order streams.
In this section, we focus on arbitrary-order edge arrival streams, where edge insertions arrive in an arbitrary order, and show that our edge and node-DP implicit matching algorithms can be implemented in this model. Our results for the second stream type (adjacency-list order streams) is presented in Section 9. To avoid redundancy, we present the edge-DP algorithm (Theorem 1.11), which is essentially a lazy-update version of our static algorithm (Algorithm 2), and then briefly discuss the minor changes needed to obtain a node-DP algorithm (Theorem 1.12). Our results uses a variant of the standard SVT tool that allows us to answer “above” $c$ times, which we recall in Section A.3 for completeness.
See 1.11
Implementing the node-DP arboricity sparsifier of Section 6.2.1 in the continual release model leads to Theorem 1.12, our main continual release node-DP result restated below.
See 1.12
8.1 Proof of Theorem 1.11
We begin with the pseudocode of the $\varepsilon$-edge DP algorithm in Algorithm 6. Then, we prove privacy and utility separately in Sections 8.1.1 and 8.1.2, respectively. The proof of privacy is non-trivial since we cannot simply apply composition. Instead, we directly argue by the definition of privacy.
1
Input: Arbitrary edge-order stream $S$, privacy parameter $\varepsilon>0$, approximation parameter $\eta\in(0,1]$,
Output: An $\varepsilon$-node differentially private implicit $b$-matching after each time stamp.
2 $j_{1}\leftarrow 0$
3 $c\leftarrow a\eta^{-1}\log(n)$
4 Initialize class $\textsc{EstimateSVT}\leftarrow\operatorname{SVT}(\nicefrac{{\varepsilon}}{{3}},1,c)$   (Algorithm 10)
5 for edge $e_{t}\in S$ do
       // $\nu(G_{t})$ denotes the size of the maximum matching of the dynamic graph $G_{t}$ at time $t$.
6       while EstimateSVT.ProcessQuery$(\nu(G_{t}),(1+\eta)^{j_{t}})$ is “above” do
7             $j_{t}\leftarrow j_{t}+1$
8
9      if $t=1$ or $j_{t}>j_{t-1}$ then
10             $\operatorname{solution}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP implicit $b$-matching using Algorithm 2 on input $G_{t}$ (Theorem 1.3)
11             $\operatorname{estimate}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP estimate of the size of the current maximum matching
12      $j_{t+1}\leftarrow j_{t}$
13       Output solution, estimate
Algorithm 6 Arbitrary Edge-Order Continual Release Edge-DP Matching
8.1.1 Privacy Proof
Lemma 8.1.
Algorithm 6 is $\varepsilon$-edge DP.
Proof of Lemma 8.1.
Let $(j_{t},\operatorname{solution}_{t},\operatorname{estimate}_{t})_{t=1}^{T}$ denote the random sequence of outputs from Algorithm 6 corresponding to the SVT estimate of the maximum matching in the current graph $G_{t}$, the implicit $b$-matching solution for timestamp $t$, and the estimate of the largest matching of $G_{t}$ contained within $\operatorname{solution}_{t}$. Similarly, let $(\widetilde{j}_{t},\widetilde{\operatorname{solution}}_{t},\widetilde{\operatorname{estimate}}_{t})_{t=1}^{T}$ be the random sequence of outputs on an edge-neighboring stream. We may assume that $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$.
Fix any deterministic sequence $(J_{t},B_{t},M_{t})_{t=1}^{T}$ of possible outputs. We have
$\Pr\left[(j_{t},\operatorname{solution}_{t},\operatorname{estimate}_{t})_{t=1}^{T}=(J_{t},B_{t},M_{t})_{t=1}^{T}\right]$
$=\prod_{t=1}^{T}\Pr[j_{t}=J_{t}\mid<t]\cdot\Pr[\operatorname{solution}_{t}=B_{t}\mid j_{t}=J_{t},<t]$
$\qquad\cdot\Pr[\operatorname{estimate}_{t}=M_{t}\mid\operatorname{solution}_{t}=B_{t},j_{t}=J_{t},<t]\,.$
Here the notation $<t$ is a shorthand that denotes the event that $(j_{\tau},\operatorname{solution}_{\tau},\operatorname{estimate}_{\tau})_{\tau=1}^{t-1}=(J_{\tau},B_{\tau},M_{\tau})_{\tau=1}^{t-1}$ for $t\geq 2$ and the trivial event of probability $1$ if $t=1$. Thus our goal is to bound the product of ratios
$\left(\prod_{t=1}^{T}\frac{\Pr[j_{t}=J_{t}\mid<t]}{\Pr[\tilde{j}_{t}=J_{t}\mid<t]}\right)\cdot\left(\prod_{t=1}^{T}\frac{\Pr[\operatorname{solution}_{t}=B_{t}\mid j_{t}=J_{t},<t]}{\Pr[\tilde{\operatorname{solution}}_{t}=B_{t}\mid\tilde{j}_{t}=J_{t},\tilde{<}t]}\right)$
$\qquad\cdot\left(\prod_{t=1}^{T}\frac{\Pr[\operatorname{estimate}_{t}=M_{t}\mid\operatorname{solution}_{t}=B_{t},j_{t}=J_{t},<t]}{\Pr[\tilde{\operatorname{estimate}}_{t}=M_{t}\mid\tilde{\operatorname{solution}}_{t}=B_{t},\tilde{j}_{t}=J_{t},\tilde{<}t]}\right)\,.$
Now, conditioned on the event that $j_{t-1}=J_{t-1}$, it holds that $j_{t}$ is independent of $\operatorname{solution}_{t-1}$ and $\operatorname{estimate}_{t-1}$. Hence the first product is at most $e^{\varepsilon/3}$ by the privacy guarantees of the SVT (Theorem A.3). For the second and third products, remark that given $j_{t}=J_{t-1}=\tilde{j}_{t}$, then $\operatorname{solution}_{t}=B_{t-1}=\widetilde{\operatorname{solution}}_{t}$ with probability $1$ and similarly for $\operatorname{estimate}_{t}$. Since $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$, at most $c=a\log(n)/\eta$ of the ratios from the second and third products are not 1. We can bound each of these ratios by $e^{\varepsilon/3c}$ using the individual privacy guarantees of Algorithm 2 (Theorem 1.3) and the Laplace mechanism. This concludes the proof. ∎
8.1.2 Utility Proof
Lemma 8.2.
Algorithm 6 outputs a sequence of implicit $b$-matchings with the following guarantee with probability at least $1-1/\operatorname{poly}(n)$: (i) $b=O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
Proof.
(i) follow directly from Theorem 1.3 and the fact that we require each call to Algorithm 2 to satisfy ($\nicefrac{{\varepsilon}}{{3c}}$)-privacy for $c=a\log(n)/\eta$.
To see (ii), we first observe that this certainly holds at timestamps when $\operatorname{solution}$ is updated. By the guarantees of SVT, $(1+\eta)^{j_{t}}$ is a $(1+\eta,O(c\log(n)/\varepsilon))$-approximate estimate of size of the current maximum matching. Hence we incur at most this much error in between updates. ∎
Combining Lemma 8.1 and Lemma 8.2 yields the proof of Theorem 1.11. The pseudocode for Theorem 1.11 is given in Algorithm 6.
8.2 Proof of Theorem 1.12
8.2.1 Extension to Node-Differential Privacy with Public Arboricity Bound
We now describe the necessary modifications to adapt our edge DP continual release algorithm to the challenging node DP setting. Similary to Section 6.2.1, we first present an algorithm assuming a public bound $\tilde{\alpha}$ on the arboricity.
Theorem 8.3.
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, Algorithm 7 is an $\varepsilon$-node DP algorithm in the arbitrary order continual release model that outputs implicit $b$-matchings. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\eta^{2}\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
The main task is to implement the arboricity matching sparsifier (Theorem 6.2) in an arbitrary edge-order stream. Note that there is a natural total order of vertex pairs given by the arrival order of edges. Thus we can simply discard edges where one of its endpoints has already seen more than some threshold $\Lambda$ incident edges (Algorithm 7 of Algorithm 7).
Let the sparsified graph $H$ be obtained from the input graph $G$ as described above. Write $c=O(\log(n)/\eta)$ be the maximum SVT budget for “above” queries (Algorithm 10). Our algorithm proceeds as following for each edge: Similar to Algorithm 6 of Algorithm 6, we use an SVT instance to check when to update our solution. Since the maximum matching size sensitivity is 1 for node-neighboring graphs, this step requires no change from the edge-DP algorithm. At timestamps where we must update the solution, we run a static $O(\nicefrac{{\varepsilon}}{{c\Lambda}})$-edge DP implicit matching algorithm (Algorithm 2) on the sparsified graph $H$. This is similar to Algorithm 6 of Algorithm 7 except we need to execute the underlying algorithm with a smaller privacy parameter. Finally, we also estimate the size of the current maximum matching. Once again, this step is unchanged from Algorithm 6 of Algorithm 6. In particular, we note that we only use the sparsified graph $H$ for updating the solution and not for estimating the current maximum matching size.
Input: Arbitrary edge-order stream $S$, approximation parameter $\eta\in(0,1]$, public bound $\tilde{\alpha}>0$.
1 $\Lambda\leftarrow 5(1+\nicefrac{{5}}{{\eta}})2\tilde{\alpha}$
2 $H\leftarrow(V,\varnothing)$
3 $d_{v}\leftarrow 0$ for each vertex $v\in V$
4 $c\leftarrow a\log(n)/\eta$
5 for edge $e_{t}\in S$ do
6       if $\perp\neq e_{t}=\{u,v\}$ and $\max(d_{u},d_{v})<\Lambda$ then
7             Increment $d_{u},d_{v}$
8             $E[H]\leftarrow E[H]\cup\{e_{t}\}$
9
      // Check if $\nu(G_{t})$ has significantly increased using SVT with total privacy budget $\nicefrac{{\varepsilon}}{{3}}$ (Algorithm 6 of Algorithm 6)
       // If SVT is ‘‘above’’, compute $\operatorname{solution}$ with respect to $H_{t}$ with privacy budget $\nicefrac{{\varepsilon}}{{6c\Lambda}}$ (Algorithm 6 of Algorithm 6)
       // If SVT is ‘‘above’’, compute $\operatorname{estimate}$ with respect to $G_{t}$ with privacy budget $\nicefrac{{\varepsilon}}{{3c}}$ (Algorithm 6 of Algorithm 6)
10       Output $\operatorname{solution},\operatorname{estimate}$
Algorithm 7 Arbitrary Edge-Order Continual Release Node-DP Matching Algorithm
We sketch the proof of guarantees for Algorithm 7.
Sketch Proof of Theorem 8.3.
The privacy proof is identical that in the proof of Theorem 1.11. The utility guarantees follow similarly, with the exception that we use the approximation guarantees of Theorem 6.3 rather than Theorem 1.3. ∎
8.2.2 Removing the Assumption of Public Bound
Finally, we describe the modifications from Algorithm 7 to remove the assumption on a public bound $\tilde{\alpha}$. First, we compute $O(\log(n))$ sparsified graphs $H_{k}$ corresponding to setting the public parameter $\tilde{\alpha}=2^{k},k=1,2,\dots,\lceil\log_{2}(n)\rceil$. The SVT to estimate the current maximum matching size remains the same (we do not compute an estimate for each $k$). Next, we compute $O(\log(n))$ implicit $b$-matchings, one for each sparsified graph $H_{k}$. We also privately compute an upper bound $\tilde{\alpha}_{t}$ of the current arboricity of $G_{t}$ and choose the implicit solution with index $k^{\star}$ such that $2^{k^{\star}-1}<\tilde{\alpha}_{t}\leq 2^{k^{\star}}$. Accounting for the error in estimating the arboricity yields a proof of Theorem 1.12.
9 Adjacency-List Order Continual Release Implicit Matchings
In the adjacency-list order insertion streams, nodes arrive in an arbitrary order, and once a node arrives, all edges adjacent to the node arrive in an arbitrary order. Remark that in the adjacency-list order stream, all edges arrive twice, once per endpoint in the stream.
In this section, we give $\varepsilon$-DP algorithms for implicit matching in the continual release model with adjacency-list order streams. Our results for another stream type (arbitrary edge-order streams) are presented in Section 8. The adjacency-list order stream ensures each node that arrives in this model will be followed by its edges where the edges arrive in an arbitrary order. Our algorithm is a straightforward implementation of Algorithm 2 in the continual release model.
In particular, the nodes arrive in an arbitrary order and when a node arrives, it waits until all of its edges arrive and then performs the same proposal and response procedure as given in Algorithm 2. Since each node can contribute at most $1$ to the matching size, we have an additional additive error of $1$ in the maximal matching size in the continual release setting.
Below, we restate our final result (Theorem 1.13) about the privacy and utility guarantees of our continual release algorithm. See 1.13
Detailed Algorithm
We give our modified adjacency-list continual release algorithm in Algorithm 8. This algorithm takes a stream $S$ of updates consisting of node insertions and edge insertions. The $i$-th update in $S$ is denoted $u_{i}$ and it can either be a node update $v$ or an edge update $e_{i}$. In adjacency-list order streams, each node update is guaranteed to be followed by all edges adjacent to it; these edges arrive in an arbitrary order. The high level idea of our algorithm is for each node $v$ to implicitly announce the nodes it is matched to after we have seen all of the edges adjacent to it. Since each node can add at most $1$ additional edge to any matching, waiting for all edges to arrive for each node update will incur an additive error of at most $1$. After we have seen all edges adjacent to $v$ (more precisely, when we see the next node update), we run the exact same proposal procedure as given in Algorithm 2.
We first set the parameters used in our algorithm the same way that the parameters were set in Algorithm 2 in Algorithms 8 and 8. Then, we iterate through all nodes (Algorithm 8) to determine the noisy cutoff $\tilde{b}(u)$ for each node $u$ (Algorithm 8). Then, we initiate the set $K$ (initially empty) to be the set of adjacent edges for the most recent node update (Algorithm 8). The most recent node update is stored in $w$ (Algorithm 8). Then, $Q$ (initially an empty set) stores all nodes that we have seen so far; that is, $Q$ is used to determine whether an edge adjacent to the most recent node update $w$ is adjacent to a node that appears earlier in the stream (Algorithm 8).
For each update as it appears in the order of the stream (Algorithm 8), we first iterate through each subgraph index (Algorithm 8) to flip a coin with appropriate probability to determine whether the edge is included in the subgraph with index $r$. This procedure is equivalent to Algorithms 2 and 2 in Algorithm 2. Then, we check whether the update is a node update (Algorithm 8). If it is a node update, then we add $v$ to $Q$ (Algorithm 8), and if $w\neq\bot$, there was a previous node update stored in $w$ (Algorithm 8) and we process this node. We first check all nodes to see whether they have reached their matching capacity (Algorithms 8, 8, 8 and 8) using an identical procedure to Algorithms 2, 2, 2 and 2 in Algorithm 2.
Then, if $w$ has not satisfied the matching condition, then we iterate through all subgraph indices to find the smallest index that does not cause the matching for $w$ to exceed $b$. This procedure (Algorithms 8, 8, 8, 8, 8, 8 and 8) is identical to Algorithms 2, 2, 2, 2, 2, 2 and 2 in Algorithm 2 except for Algorithm 8. The only difference between Algorithm 8 in Algorithm 8 and Algorithm 2 in Algorithm 2 is that we check whether $v$ appears earlier in the ordering by checking $v\not\in Q$ and we check the coin flip for edge $\{w,v\}$ by checking $c(i,r)=\textsc{Heads}$. Thus these two lines are functionally identical. After we have updated the matching with node $w$’s match, we set $K$ to empty (Algorithm 8) and update $w$ to the new node $v$ (Algorithm 8).
Finally, if $u_{i}$ is instead an edge update $e_{i}$ (Algorithm 8), we add $e_{i}$ to $K$ to maintain the adjacency list of the most recent node update (Algorithm 8).
1
Input: Arbitrary order adjacency-list stream $S$, privacy parameter $\varepsilon>0$, matching parameter $b=\Omega(\log{n}/\varepsilon)$.
Output: An $\varepsilon$-locally edge differentially private implicit $b$-matching.
2
[…]
4 $\varepsilon^{\prime}\leftarrow\varepsilon/(2+1/\eta^{\prime})$
5
6for each node $u\in[n]$ in order do
7       $\tilde{b}(u)\leftarrow b-20\log(n)/\varepsilon^{\prime}+\text{Lap}(4/\varepsilon^{\prime})$
8
[…]
10 $w\leftarrow\bot$
11 $Q\leftarrow\varnothing$
12 for every update $u_{i}\in S$ do
13       for each subgraph index $r=0,\ldots,\lceil\log_{1+\eta^{\prime}}(n)\rceil$ do
14             Flip and release coin $c(i,r)$ which lands heads with probability $p_{r}=(1+\eta^{\prime})^{-r}$
15
16      if $u_{i}$ is a node update of node $v$ then
17             $Q\leftarrow Q\cup\{v\}$
18             if $w\neq\bot$ then
19                   for each node $u\in V$ which has not satisfied the matching condition do
20                         $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime})$
21                         if $M_{i}(u)+\nu_{i}(u)\geq\tilde{b}(u)$ then
22                               Output to transcript: node $u$ has reached their matching capacity
23
24
25                  if $w$ has not satisfied the matching condition then
26                         for subgraph index $r=0,\ldots,\lceil\log_{1+\eta^{\prime}}(n)\rceil$ do
27                               $W_{r}(w)=\{v:v\text{ active}\wedge v\in e_{i}\text{ where }e_{i}\in K\wedge v\not\in Q\wedge c(i,r)=\textsc{Heads}\}$
28                               $\widetilde{W}_{r}(v)=|W_{r}(v)|+\text{Lap}(2/\varepsilon^{\prime})$
29                        $\widetilde{M}_{i}(v)=M_{i}(v)+\text{Lap}(2/\varepsilon^{\prime})$
30                         $v$ computes smallest $r$ so that $\widetilde{M}_{i}(v)+\widetilde{W}_{r}(v)+c\log(n)/\varepsilon^{\prime}\leq b$, and matches with neighbors in $W_{r}$
31                         $v$ releases $r$
32
33            $K\leftarrow\varnothing$
34             $w\leftarrow v$
35
36      if $u_{i}$ is an edge update $e_{i}$ then
37             $K\leftarrow K\cup\{e_{i}\}$
38
Algorithm 8 Adjacency-List Order Continual Release Matching
9.1 Proof of Theorem 1.13
Proof.
We first prove that our continual release algorithm is $\varepsilon$-DP on the vector of outputs. First, our algorithm only outputs a new output each time a node update arrives. For every edge update, the algorithm outputs the same outputs as the last time the algorithm outputted a new output for a node update. Algorithm 8 implements all of the local randomizers used in Algorithm 2 in the following way. The order of the nodes that propose is given by the order of the node updates. Coins for edges are flipped in the same way as in Algorithm 2, the threshold for the matching condition is determined via an identical procedure, and finally, the proposal process is identical to Algorithm 2. Hence, the same set of local randomizers can be implemented as in Algorithm 2 and the continual release algorithm is $\varepsilon$-DP via concurrent composition.
The approximation proof also follows from the approximation guarantee for Algorithm 2 since the proposal procedure is identical except for the additive error. In the continual release model, there is an additive error of at most $1$ since for every edge update after a new node update, the node could be matched to an initial new edge update but is not matched until the final edge update is shown for that node. Since each node contributes at most $1$ to a matching, the additive error is at most $1$ for each update. ∎
10 Private Implicit $b$-Matchings
In this section, we extend our all of our results to the $(2+\eta)$-approximate $b$-matchings, including our node-DP results, which use a novel $b$-matching arboricity sparsifier. The key difference between the results in this section and our previous result is that our approximation bounds are given in terms of the maximum $b$-matching given large enough $b$. In particular, we satisfy the $b$-matching constraints (that our matching has degree less than $b$ for all vertices), with high probability, and output a $(2+O(\eta))$-maximum $b$-matching). Thus, our approximation is no longer a bicriteria approximation in this setting.
To obtain our $b$-matching results, we prove a stronger version of our matching result. Namely, we show that we can guarantee that our implicit output is at least the size of a $(2+\eta)$-approximate maximum $b^{\prime}$-matching for some $b^{\prime}<b$. In particular, $b$ has to be an additive factor of $\Omega\left(\frac{\ln(n)}{{\eta^{2}}\varepsilon}\right)$ greater than $b^{\prime}$ to ensure the returned solution is comparable to a $2$-approximate $b^{\prime}$-matching. Furthermore, if $b=\Omega\left(\frac{\ln(n)}{\eta^{3}\varepsilon}\right)$, then we get a $(2+\eta)$-approximate maximum $b$-matching.
10.1 $\varepsilon$-LEDP $b$-Matchings
Our revised algorithm is the same algorithm as Algorithm 2 except with a different lower bound on $b$ depending on what $b^{\prime}$-maximal matching we want to guarantee in our returned solution. This is our only change to the algorithm. When our lower bound on $b$ contains a factor of $1/\eta^{2}$, then we obtain a $(2+\eta)$-approximate maximum $b$-matching. We restate our modified algorithm below with the modified constants highlighted in blue. The constants do not affect our privacy proof; hence, we only give an updated version of our utility analysis following our pseudocode in Algorithm 2. Since the proof and algorithm only require minor changes for the generalization, we replicate our previous algorithm and proof and highlight our changes in blue.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, $b^{\prime}>0$ parameter for $b^{\prime}$-matching, matching parameter $b\geq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}}$, constant $\eta\in(0,1)$, and constant $c\geq 3$
Output: An $\varepsilon$-local edge differentially private implicit $b$-matching.
1 $\varepsilon^{\prime}\leftarrow\frac{\varepsilon}{2(1+2\eta)/\eta}$
2 for every pair of vertices $u\neq v\in V\times V$ and subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$  do
3       Flip and release coin $coin(u,v,r)$ which lands heads with probability $p_{r}=(1+\eta)^{-r}$
4
5for each node $u\in V$ do
6       $\tilde{b}(u)\leftarrow b-36c\ln(n)/\varepsilon^{\prime}+\text{Lap}(4/\varepsilon^{\prime})$
7       $M(u)\leftarrow\varnothing$
8       $A(u)\leftarrow\infty$
9
10for iteration $i=1$ to $n$, let $v=v_{i}$ do
       // Multidimensional-AboveThreshold for checking if each node has reached their matching threshold
11       for each node $u\in V$ such that $A(u)>i$ ($u$ has not satisfied its matching condition)  do
12             $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime})$
13             if $|M(u)|+\nu_{i}(u)\geq\tilde{b}(u)$  then
14                   $u$ releases $A(u)\leftarrow i$
15
16
      // If $v$ can still match with more edges, find an additional set to match $v$ with.
17       if $A(v)>i$ ($v$ has not satisfied its matching condition)  then
18             for subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$  do
19                   $W_{r}(v)=\{u:A(u)>i\wedge u\text{ later than $v$ in ordering }\land\{u,v\}\in E\wedge coin(u,v,r)=\textsc{Heads}\}$
20                   $|\widetilde{W}_{r}(v)|=|W_{r}(v)|+\text{Lap}(2/\varepsilon^{\prime})$
21
22            $|\widetilde{M}_{i}(v)|=|M(v)|+\text{Lap}(2/\varepsilon^{\prime})$
23             $v$ computes $r_{v}\leftarrow\min\{r:|\widetilde{M}_{i}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b\}$
24             $v$ releases $r_{v}$
25             for $u\in W_{r_{v}}(v)$ do
26                   $M(u)\leftarrow M(u)\cup\{v\}$
27                   $M(v)\leftarrow M(v)\cup\{u\}$
28
29
Algorithm 9 $\varepsilon$-LEDP Approximate Maximum $b$-Matching
We now restate and prove Theorem 1.4. See 1.4
The proof is implied by the following lemma which is a restatement with detailed constants.
Lemma 10.1.
For $\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}$, Algorithm 9 is $\varepsilon$-LEDP and outputs an implicit $b$-matching $M$ in the billboard model such that $|M|\geq\frac{1}{2}\operatorname{OPT}(b^{\prime})$ with probability at least $1-\frac{1}{n^{c}}$. Here $\operatorname{OPT}(b^{\prime})$ is the size of a maximum $b^{\prime}$-matching.
Proof.
For each iteration $i$, we will show that the number of nodes that $v=v_{i}$ matches with after iteration $i$ is either at least $b^{\prime}$, or every one of its unmatched neighbors is already matched to at least $b^{\prime}$ of their neighbors. Hence the final output satisfies the maximality condition of Lemma 4.2 and thus the approximation guarantees. We will show that this suffices to guarantee that our output is of comparable size to a maximal $b^{\prime}$-matching. Finally, we show that with high probability, we do not match with more than $b$ nodes.
First, all of the Laplace noise (Algorithms 9, 9, 9 and 9) are drawn from distributions with expectation 00. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
  • $b-36c\log(n)/\varepsilon^{\prime}\leq\tilde{b}(u)\leq b-24c\log(n)/\varepsilon^{\prime}$ for all $u\in V$,
  • $|\nu_{i}(u)|\leq{24c\ln(n)/\varepsilon^{\prime}}$ for all $i\in[n],u\in V$,
  • $|W_{r}(v)|-{6}c\log(n)/\varepsilon^{\prime}\leq|\widetilde{W}_{r}(v)|\leq|W_{r}(v)|+{6}c\log(n)/\varepsilon^{\prime}$ for all $v\in V,r\in[\lceil\log_{1+\eta}(n)\rceil]$, and
  • $|M(v)|-{6}c\log(n)/\varepsilon^{\prime}\leq|\widetilde{M}_{i}(v)|\leq|M(v)|+{6}c\log(n)/\varepsilon^{\prime}$ for all $v_{i}\in V$.
Let $\mathcal{E}$ denote the intersection of all the events above. By a union bound, $\mathcal{E}$ occurs with probability at least $1-\frac{1}{n^{2c}}$ for $c\geq 3$. We condition on $\mathcal{E}$ for the rest of the proof.
Case I: Suppose that Algorithm 9 does not execute, i.e. $v$ has already satisfied its matching condition as evaluated on Algorithms 9, 9 and 9. Then we know that $|M(v)|+\nu_{i}(v)\geq\tilde{b}(v)$ by definition. We then have that
$|M(v)|$ $\geq b-{60}c\ln(n)/\varepsilon^{\prime}$
$\geq{b}-\frac{72c\ln(n)}{\varepsilon^{\prime}}$
$={b}-\frac{8\cdot 72c\ln(n)}{\eta\varepsilon}$ (by our choice of $\varepsilon^{\prime}$ in Algorithm 9)
$\geq{b}-\frac{576c\ln(n)}{\eta\varepsilon}$
$>{b^{\prime}}\,.$ (by $b\geq b^{\prime}+\frac{576c\ln(n)}{\eta\varepsilon}$)
In other words, $v$ was already matched to $b^{\prime}$ neighbors from a previous iteration.
Case II.a: Now, assume that Algorithm 9 does execute. Let $W_{0}(v)$ be the set of unmatched neighbors of $v$ which appear later in the ordering. First, consider the case that $|W_{0}(v)|<\frac{12c\ln(n)}{\eta^{2}}$. We know that either $|M(v)|<b^{\prime}$ or $|M(v)|\geq b^{\prime}$. If $|M(v)|\geq b^{\prime}$, then $v$ is already matched to sufficiently many neighbors and there is nothing to prove. Otherwise, conditioned on $\mathcal{E}$,
$|W_{0}(v)|+3\cdot\frac{12c\ln(n)}{\varepsilon^{\prime}}+{b^{\prime}}$ $\leq\frac{12c\ln(n)}{\eta^{2}}+\frac{36c\ln(n)}{\varepsilon^{\prime}}+{b^{\prime}}$
$\leq\frac{12c\ln(n)}{\eta^{2}}+\frac{216c\ln(n)}{\eta\varepsilon}+{b^{\prime}}$
$\leq\frac{238c\ln(n)}{\eta^{2}\varepsilon}+{b^{\prime}}$
$\leq b$
and all of $W_{0}(v)$ is matched.
Case II.b.1: For the remainder of the proof, we assume $k\coloneqq|W_{0}(v)|\geq\frac{12c\ln(n)}{\eta^{2}}$. As a reminder, we are in the case that Algorithm 9 does execute. Let $r^{*}$ be the $r$ chosen in Algorithm 9. If the returned $r^{*}>\log_{1+\eta}(k\eta^{2}/(12c\ln{n}))$, we know that the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b$ did not hold for $r=\log_{1+\eta}\left(\frac{k\eta^{2}}{12c\ln{n}}\right)$. Note here we use $k\geq\frac{12c\ln(n)}{\eta^{2}}$ so that $r\geq 0$ is well-defined. Thus, we have that $p_{r}=\frac{12c\ln(n)}{k\eta^{2}}$ so the expected size of $W_{r}(v)$ is $\frac{12c\ln(n)}{\eta^{2}}$; that is, $\mathbb{E}[|W_{r}(v)|]=\frac{12c\ln(n)}{\eta^{2}}$. By a multiplicative Chernoff Bound (Theorem A.2), the true size of $W_{r}(v)$ is at most $18c\log(n)/\eta^{2}$ with probability at least $1-\frac{1}{n^{4c}}$. Since the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\log(n)/\varepsilon^{\prime}\leq b$ didn’t hold for such $r$, we have that
$|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+\frac{12c\ln(n)}{\varepsilon^{\prime}}$ $>b$
$|M(v)|+|W_{r}(v)|+\frac{6c\ln(n)}{\varepsilon^{\prime}}+\frac{6c\ln(n)}{\varepsilon^{\prime}}+\frac{12c\ln(n)}{\varepsilon^{\prime}}$ $>b$ (conditioning on $\mathcal{E}$)
$|M(v)|+|W_{r}(v)|+\frac{24c\ln(n)}{\varepsilon^{\prime}}$ $>b\,.$
In other words,
$|M(v)|$ $>b-|W_{r}(v)|-\frac{24c\ln(n)}{\varepsilon^{\prime}}$
$|M(v)|$ $>b-\frac{18c\ln(n)}{\eta^{2}}-\frac{24c\ln(n)}{\varepsilon^{\prime}}$ ( w.p. at least $1-\frac{1}{n^{4c}}$)
$|M(v)|$ $>b-\frac{18c\ln(n)}{\eta^{2}}-\frac{144c\ln(n)}{\eta\varepsilon}$ (by the choice of $\varepsilon^{\prime}$)
$|M(v)|$ $>b-\frac{162c\ln(n)}{\eta^{2}\varepsilon}$
$|M(v)|$ $>{b^{\prime}}\,.$ (since $b\geq{b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}}$)
In other words, $v$ was already matched from a previous iteration.
Case II.b.2: If $r^{*}\leq\log_{1+\eta}\left(\frac{k\eta^{2}}{12c\ln{n}}\right)$, there are two scenarios. Either $r^{*}=0$, then we match $v$ with all nodes which are still available so clearly the number of nodes $v$ is matched with is at least $b^{*}(v)$, or $r^{*}>1$. In the latter scenario, we first remark that $p_{r^{*}}\geq\frac{12c\ln(n)}{k\eta^{2}}$ so that
$\mathbb{E}[|W_{r^{*}-1}|]=(1+\eta)\mathbb{E}[|W_{r^{*}}|]\geq(1+\eta)\frac{12c\ln(n)}{\eta^{2}}\,.$
Thus we can apply a Chernoff bound on both $|W_{r^{*}-1}|,|W_{r^{*}}|$ and obtain high probability bounds. We also know that the threshold is exceeded for $r^{*}-1$, so we have $|\widetilde{M}(v)|+|\widetilde{W}_{r^{*}-1}(v)|+12c\ln(n)/\varepsilon^{\prime}\geq b$. Conditioning on $\mathcal{E}$, $|M(v)|+|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Thus by a Chernoff bound on $|W_{r^{*}-1}|$, we see that
$b-|M(v)|-\frac{36c\ln(n)}{\varepsilon^{\prime}}\leq|W_{r^{*}-1}|\leq(1+\eta)\mathbb{E}[|W_{r^{*}-1}|]=(1+\eta)^{2}\mathbb{E}[|W_{r^{*}}|]$
with probability $1-\frac{1}{n^{4c}}$. Now again by a Chernoff bound but now on $|W_{r^{*}}|$, we have
$|W_{r^{*}}|\geq\frac{(1-\eta)}{(1+\eta)^{2}}\left[b-|M(v)|-\frac{36c\ln(n)}{\varepsilon^{\prime}}\right]$
with probability $1-\frac{1}{n^{4c}}$.
Hence, if $|M(v)|<b^{\prime}$ (meaning it has not satisfied the matching criteria), it holds that
$|W_{r^{*}}(v)|$ $\geq\frac{(1-\eta)}{(1+\eta)^{2}}(b-|M(v)|-\frac{216c\ln(n)}{\eta\varepsilon})$
$\geq\frac{1-\eta}{(1+\eta)^{2}}\left(\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}-|M(v)|-\frac{216c\ln(n)}{\eta\varepsilon}\right)$
$\geq b^{\prime}-|M(v)|\,.$
It follows that $v$ is matched with at least $b^{\prime}$ neighbors at the end of this iteration.
Taking a union bound over all $r$ and $v\in V$, all Chernoff bounds hold regardless of the value of $r^{*}$ with probability at least $1-\frac{1}{n^{2c}}$ when $c\geq 3$. Taking into account the conditioning on $\mathcal{E}$, our algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for $c\geq 3$.
In summary, at iteration $i$, we ensure that 1) either $v=v_{i}$ was already matched to at least $b^{\prime}$ neighbors during prior iterations, 2) or it will be matched with a sufficient number of descendents to satisfy a total of at least $b^{\prime}$ matched neighbors after iteration $i$, 3) or it will be matched to all descendents that do not yet have $b^{\prime}$ matched neighbors. The only property to check is whether there are any ancestors $u$ of $v$ that are not matched to $v$ but have yet to achieve $b^{\prime}$ matched neighbors in the last case. But this cannot happen since $u$ would have matched to all descendents that did not yet have $b^{\prime}$ matched neighbors during its iteration and $v$ in particular did not have $b^{\prime}$ matched neighbors.
It remains only to check that we match each vertex to at most $b$ neighbors. To do so, it suffices to check that on the $i$-th iteration, each node which does not yet satisfy a matching condition can match with at least 1 more neighbor after Algorithm 9 and that $v=v_{i}$ is matched with at most $b$ neighbors at the end of the iteration (Algorithm 9). The first statement is guaranteed since any vertex $u$ not yet satisfying its matching condition satisfies
$|M(u)|$ $<\tilde{b}(u)-\nu_{i}(u)$
$|M(u)|$ $<b-24c\log(n)/\varepsilon^{\prime}+24c\log(n)/\varepsilon^{\prime}$ (conditioning on $\mathcal{E}$)
$|M(u)|$ $<b.$
If $v$ already satisfies its matching condition after Algorithm 9, the second statement is guaranteed to hold. Otherwise, $|M(v)|<b$ and we choose $r$ such that
$b$ $\geq|\tilde{M}_{i}(v)|+|\tilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}$
$\geq|M(v)|+|W_{r}(v)|-12c\ln(n)/\varepsilon^{\prime}+12c\ln(n)/\varepsilon^{\prime}$ (conditioning on $\mathcal{E}$)
$=|M(v)|+|W_{r}(v)|.$
Thus in this case, $v$ is also matched to at most $b$ neighbors, as desired. ∎
We now restate and prove Theorem 1.5. See 1.5 The proof follows from the following corollary of Lemma 10.1.
Corollary 10.2 ($(2+\eta)$-Approximate Maximum $b$-Matching).
When $b\geq\frac{576c\ln(n)}{\eta^{3}\varepsilon}$ for a constant $\eta\in(0,\nicefrac{{1}}{{2}})$, Algorithm 9 is $\varepsilon$-LEDP and outputs an implicit $(2+24\eta)$-maximum $b$-matching in the billboard model, with probability at least $1-\frac{1}{n^{c}}$.
Before proceeding with the proof, we recall that $b$-matchings have the following integral relaxation [schrijver2003combinatorial, Theorem 33.2], where $\partial(U)$ denotes the set of outgoing edges from $U\subseteq V$.
$\sum_{uv\in E}x(uv)$ $\leq b$ $\forall v\in V$
$\sum_{e\in E[U]}x(e)+\sum_{f\in F}x(f)$ $\leq\lfloor\frac{1}{2}(b|U|+|F|)\rfloor$ $\forall U\subseteq V,F\subseteq\partial(U)$
$x$ $\in[0,1]^{E}$
Note that it suffices to consider constraints of the second form only when $b|U|+|F|$ is odd, but we include the redundant constraints for the sake of simplicity.
Proof.
Choose $b^{\prime}$ such that
$b=(1+12\eta)b^{\prime}\geq(1+11\eta)b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}\,.$
Then we can apply Lemma 10.1 to deduce that Algorithm 9 outputs a $2$-approximate maximum $(\frac{b}{1+12\eta})$-matching.
Identify any $b$-matching $M$ as an indicator $\chi_{M}\in[0,1]^{E}$ of the $b$-matching polytope. Then $\frac{1}{1+12\eta}\chi_{M}$ is a feasible fractional solution of the $(\frac{b}{1+12\eta})$-matching polytope. Hence the size of a maximum $(\frac{b}{1+12\eta})$-matching is at least a $\frac{1}{1+12\eta}$ fraction of the size of a maximum $b$-matching. This concludes the proof. ∎
10.2 Node-Differentially Private $b$-Matchings
Similar to Section 6, we can implement our implicit $b$-matching algorithm in node-DP setting with the help of an arboricity sparsifier. However, this requires extending Solomon’s sparsifier [solomon2021local] to the setting of $b$-matchings. Doing so and repeating the steps in Section 6 yields a proof of Theorem 1.9. See 1.9
The rest of this section serves to prove the necessary sparsifier extension.
10.2.1 Bounded Arboricity Sparsifiers for $b$-Matchings
In order to extend the $b$-matching guarantees to the node-DP setting, we prove an extension of Solomon’s sparsifier (Theorem 6.2) to the case of $b$-matchings with an $O(b)$ additive increase in the degree threshold. This may be of independent interest beyond node differentially private matchings.
Theorem 10.3.
Let $G$ be a graph of arboricity at most $\alpha$ and $\Lambda_{b}\coloneqq 5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha+(b-1)$ for some $\eta\in(0,1]$. Suppose $H$ is obtained by marking an arbitrary set of $\min(\deg(v),\Lambda_{b})$ edges incident to each vertex $v$ and taking the subgraph consisting of all edges marked by both endpoints. Then if $\mu_{b}(\cdot)$ denotes the size of a maximum $b$-matching of the input graph,
$\mu_{b}(H)\leq\mu_{b}(G)\leq(1+\eta)\mu_{b}(H).$
In particular, any $(\beta,\zeta)$-approximate maximum $b$-matching for $H$ is an $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$. Moreover, this holds for $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda_{b})$ where $\pi\in P_{\binom{n}{2}}$ is a total order over unordered vertex pairs.
[…]
Proof.
We argue by induction on $b$. The base case of $b=1$ is guaranteed by Theorem 6.2. Suppose inductively the statement holds for $b-1$ for some $b\geq 2$ so that our goal is to extend it to the case $b$.
Let $M_{1}\cup\dots M_{b}$ be any $b$-matching of $G$ decomposed into $b$ disjoint matchings. Let $H$ denote a candidate sparsifier obtained by marking an arbitrary set $T(v)$ of at least $\min(\deg_{G}(v),\Lambda_{b})$ edges incident to each vertex $v$, and taking the subgraph $H=(V,E_{H})$ consisting of all edges marked by both endpoints, i.e. $E_{H}\coloneqq\{uv\in E(G):u\in T(v),v\in T(u)\}$.
Consider $G_{1}\coloneqq G-(M_{2}\cup\dots\cup M_{b})$ and note that $M_{1}$ is a matching in $G_{1}$. We claim that $E_{H}\cap E(G_{1})$ contains a matching of size at least $\frac{1}{1+\eta}|M_{1}|$. Indeed, let
$\forall v\in V,T_{1}(v)\coloneqq\{u\in T(v):uv\notin M_{2}\cup\dots\cup M_{b}\},\qquad E_{1}\coloneqq\{uv\in E:u\in T_{1}(v),v\in T_{1}(u)\}\,.$
Note that each $T_{1}(v)$ contains at least $\min(\deg_{G_{1}}(v),\Lambda_{1})$ neighbors of $v$ within $G_{1}$. Furthermore, we have $E_{1}\subseteq E_{H}$. Thus by the base case, $E_{1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum matching $M_{1}^{\prime}$ of $G_{1}$.
Remark that $M_{1}^{\prime}\cup M_{2}\cup\dots\cup M_{b}$ is a $b$-matching within $G$ by construction. Consider $G_{-1}\coloneqq G-M_{1}^{\prime}$, so that $M_{2}\cup\dots\cup M_{b}$ is a $(b-1)$-matching within $G_{-1}$. We claim that $E_{H}\cap E(G_{-1})$ contains a $(b-1)$-matching of size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. To prove this claim, we will rely on the induction hypothesis. Define
$\forall v\in V,T_{-1}(v)\coloneqq\{u\in T(v):uv\notin M_{1}^{\prime}\},\qquad E_{-1}\coloneqq\{uv\in E:u\in T_{-1}(v),v\in T_{-1}(u)\}\,.$
Then each $T_{-1}(v)$ contains at least $\min(d_{G_{-1}}(v),\Lambda_{b-1})$ neighbors of $v$ in $G_{-1}$. Furthermore, $E_{-1}\subseteq(E_{H}\setminus M_{1}^{\prime})$ by construction. Thus by the induction hypothesis, $E_{-1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum $(b-1)$-matching $N$ of $G_{-1}$.
We have identified disjoint edge subsets $M_{1}^{\prime},N\subseteq E_{H}$ such that $M_{1}^{\prime}$ is a matching in $G$ with size at least $\frac{1}{1+\eta}|M_{1}|$ and $N$ is a $(b-1)$-matching in $G-M_{1}^{\prime}$ with size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. Thus $M_{1}^{\prime}\cup N$ is a $b$-matching in $G$ with size at least
$|M_{1}^{\prime}\cup N|=|M_{1}^{\prime}|+|N|\geq\frac{1}{1+\eta}|M_{1}\cup\dots\cup M_{b}|\,.$
By the arbitrary choice of $M_{1}\cup\dots\cup M_{b}$, we conclude the proof. ∎
10.3 Arbitrary Edge-Order Continual Release $b$-Matchings
Similarly, we can extend our $b$-matching algorithms to the continual release setting. Here we state the guarantees for an $\varepsilon$-edge DP algorithm for arbitrary edge-order stream, but our algorithms also extend to the node-DP setting.
Theorem 10.4.
For $\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\Omega\left(\frac{\log^{2}(n)}{\eta^{3}\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm that outputs a sequence of implicit $b$-matchings in the billboard model such that each $b$-matching is a $(2+\eta)$-approximation with respect to the maximum $b^{\prime}$-matching size, with probability at least $1-\frac{1}{n^{c}}$.
11 Other Lower Bounds
11.1 Lower Bound for Implicit Matchings
We now give our lower bound for implicit solutions, which essentially matches our upper bound from Theorem 1.3. While it does not apply to all algorithms in the billboard model, it applies to the implicit solutions that our algorithms use (Definition 2.6).
Theorem 11.1.
Let $\gamma>0$ be arbitrary. Let $\mathcal{A}$ be an algorithm that satisfies $\varepsilon$-edge DP and outputs an implicit solution that has the size of a $\gamma$-approximate maximal matching with probability at least $1-\beta$, with $\beta\leq 1/(16e^{4\varepsilon})$. Then the degree of this implicit solution is at least $\Omega(\frac{1}{\varepsilon}\log(1/\beta))$ with probability at least $1/2$, regardless of the value of $\gamma$.
Note that this essentially matches Theorem 1.3 (up to constants) by setting $\beta=1/\mathrm{poly}(n)$. In the rest of this section we will prove Theorem 11.1. Recall that, given an input graph $H$, our implicit solutions essentially output a subset of vertices for each node $x$ whose intersection with $x$’s private adjacency list gives the nodes $x$ is matched to. We denote these public subsets of vertices by $\mathcal{S}=\{S_{x}\}_{x\in V}$ where $S_{x}$ is the public subset of vertices for node $x$. That is, $\mathcal{S}=\{S_{x}\}_{x\in V}$ is the implicit solution generated by $\mathcal{A}$. We say that $\mathcal{A}$ includes an edge $\{u,v\}$ if either $u\in S_{v}$ or $v\in S_{u}$. In other words, $\mathcal{A}$ includes an edge if that edge is an edge in $H(\mathcal{S})$.
Proof of Theorem 11.1.
Let $V=\{r,v_{1},v_{2},\dots,v_{n-1}\}$. Let $i\in\{2,3,\dots,n-1\}$, and let $G_{i}$ be a graph with just one edge $\{r,v_{i}\}$. Then $\mathcal{A}$ must include $\{r,v_{i}\}$ with probability at least $1-\beta$ in order to meet the utility guarantee of Theorem 11.1 (note that $\eta$ vanishes since in this case the maximal matching has size $1$). Now consider a graph $G$ with just one edge $\{r,v_{1}\}$. Since $G_{i}$ has distance $2$ from $G$, by the differential privacy guarantee we know that the probability that $\mathcal{A}$ includes $\{r,v_{i}\}$ when run on $G$ must be at least $1-e^{2\varepsilon}\beta$. Note that this is true for all $i\in\{2,3,\dots,n-1\}$.
To simplify notation, let $\mathcal{S}$ be the implicit solution output by $\mathcal{A}$. Now let $T$ be an arbitrary subset of $\{v_{2},v_{3},\dots,v_{n-1}\}$ of size $\frac{1}{2\varepsilon}\log(1/\beta)$. Then the above argument implies that the expected number of edges between $r$ and $T$ in $H(\mathcal{S})$ is at least $(1-e^{2\varepsilon}\beta)|T|$ when we run $\mathcal{A}$ on $G$, or equivalently the expected number of non-edges between $r$ and $T$ is at most $e^{2\varepsilon}\beta|T|$. So by Markov’s inequality,
Pr[number of non-edges from r to T in $\Pr\left[\text{number of non-edges from $r$ to $T$ in }H(\mathcal{S})\geq\frac{|T|}{2}\right]\leq\frac{2e^{2\varepsilon}\beta|T|}{|T|}=2e^{2\varepsilon}\beta.$
Let $Q$ denote the event that the number of non-edges from $r$ to $T$ in $H(\mathcal{S})$ is at least $\frac{|T|}{2}$. Let $G^{\prime}=(V,E^{\prime})$ be a different graph with the same vertex set but with $E^{\prime}=\{\{r,v_{i}\}:v_{i}\in T\}$. Since $G^{\prime}$ has distance $|T|$ from $G$, group privacy implies that the probability of $Q$ when we run $\mathcal{A}$ on $G^{\prime}$ is at most
$e^{|T|\varepsilon}2e^{2\varepsilon}\beta$ $=e^{\frac{1}{2\varepsilon}\log(\nicefrac{{1}}{{\beta}})\varepsilon}2e^{2\varepsilon}\beta=\beta^{-\frac{1}{2}}2e^{2\varepsilon}\beta=\beta^{\frac{1}{2}}2e^{2\varepsilon}\leq\frac{1}{2}.$
Thus with probability at most $\nicefrac{{1}}{{2}}$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at most $\nicefrac{{|T|}}{{2}}$ edges from $r$ to $T$. Thus with probability at least $\nicefrac{{1}}{{2}}$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at least $\nicefrac{{|T|}}{{2}}$ edges from $r$ to $T$. Since all of those edges are also edges of $G^{\prime}$, this means that the degree of the implicit solution is at least $\nicefrac{{|T|}}{{2}}=\Theta\left(\frac{1}{\varepsilon}\log(\nicefrac{{1}}{{\beta}})\right)$ with probability at least $\nicefrac{{1}}{{2}}$. ∎
11.2 Lower Bound for Node DP Matching Sparsifiers
In this section we show a strong lower bound against computing matching sparsifiers in the node DP setting. It is obviously impossible to release such a sparsifier explicitly under any reasonable privacy constraint, but we will show that node differential privacy rules out even releasing an implicit sparsifier. This is particularly interesting since despite this lower bound, our node DP algorithm makes crucial use of a matching sparsifier which it constructs. So our node DP algorithm, combined with this lower bound, shows that it is possible to create and use a matching sparsifier for node DP algorithms even though the sparsifier itself cannot be released without destroying privacy. This is the fundamental reason why we cannot design a local algorithm for matchings in the node DP setting; the obvious local version of our algorithm would by definition include the construction of the sparsifier in the transcript, and thus this lower bound shows that it cannot be both DP and have high utility.
Recall that a matching sparsifier of a graph $G=(V,E)$ is a subgraph $H$ of $G$ with the properties that a) the maximum degree of $H$ is small, and b) the maximum matching $H$ has size close to the maximum matching in $G$. Without privacy, Solomon [solomon2021local] showed that for any $\eta\in(0,1]$, it is possible to find an $H$ with maximum degree at most $O(\frac{1}{\eta}\alpha)$ (where $\alpha$ is the arboricity of $G$) and $\nu(G)\leq(1+\eta)\nu(H)$ (where recall that $\nu(\cdot)$ denotes the size of the maximum matching in the graph).
Since we obviously cannot release a sparsifier explicitly under any reasonable privacy constraint, we turn to implicit solutions. Formally, given a graph $G=(V,E)$, we want an $(\varepsilon,\delta)$-node DP algorithm which outputs an implicit matching sparsifier: a set of edges $E^{\prime}\subseteq\binom{V}{2}$. This set $E^{\prime}$ gives us the “true” matching sparsifier $E^{\prime}\cap E$. We claim that under node privacy, $E^{\prime}\cap E$ cannot have both small maximum degree and be a good approximation of the maximum matching.
Theorem 11.2.
Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-node DP algorithm which, given an input graph $G=(V,E)$, outputs an implicit matching sparsifier $E^{\prime}$. Then if $\nu(G)\leq(1+\eta)\cdot\text{E}[\nu(E^{\prime}\cap E)]$, the maximum degree of $E^{\prime}\cap E$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$ even if $G$ has arboricity $1$.
Proof.
Fix $V$ with $|V|=n$, and let $r\in V$. For any $S\subseteq V\setminus\{r\}$, let $G_{S}$ be the graph whose edge set consists of an edge from $r$ to all nodes in $S$ (i.e., a star from $r$ to $S$ and all other nodes having degree 00). Note that all such graphs have arboricity $1$. A similar symmetry argument to the other lower bounds implies that without loss of generality, we may assume that the probability that $\mathcal{A}$ includes some edge $\{r,v\}$ is the same for all $v\in S$ (we will denote this probability by $p_{S}$).
Fix some node $v\in S\setminus\{r\}$, and consider what happens when we run $\mathcal{A}$ on the graph $G_{\{v\}}$. Since $\text{E}[\nu(E^{\prime}\cap E)]\geq\frac{\nu(G_{\{v\}})}{1+\eta}$ by assumption, we know that $p_{\{v\}}=\Pr[\{r,v\}\in E^{\prime}]\geq\frac{1}{1+\eta}$.
Now consider the graph $G_{V\setminus\{r\}}$. Since $G_{\{v\}}$ and $G_{V\setminus\{r\}}$ are neighboring graphs (under node-differential privacy, since we simply changed the edges incident on $r$) and $\mathcal{A}$ is $(\varepsilon,\delta)$-node DP, we know that when we run $\mathcal{A}$ on $G_{V\setminus\{r\}}$ it must be the case that $\Pr[\{r,v\}\in E^{\prime}]\geq e^{-\varepsilon}p_{\{v\}}-\delta$. But now symmetry implies that this is true for all $v\in V\setminus\{r\}$, so we have that
$p_{V\setminus\{r\}}$ $\geq e^{-\varepsilon}p_{\{v\}}-\delta\geq\frac{1}{e^{\varepsilon}(1+\eta)}-\delta.$
Linearity of expectation then implies that the expected degree of $r$ in $E^{\prime}$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$. Since $E(G_{V\setminus\{r\}})$ consists of all edges from $r$ to all other nodes, this means that the expected maximum degree of $E^{\prime}\cap E$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$ as claimed. ∎
In particular, for the natural regime where $\varepsilon$ is a constant, $\delta\leq 1/n$, and the maximum matching approximation $(1+\eta)$ is a constant, Theorem 11.2 implies that the maximum degree must be $\Omega(n)$ rather than $O(1)$ for graphs of arboricity $1$.
[…]
Appendix A Deferred Preliminaries
A.1 Concentration Inequalities
Lemma A.1.
Given a random variable $X\sim Lap(b)$ drawn from a Laplace distribution with expectation 00, the probability $|X|>c\ln(n)$ is $n^{-\frac{c}{b}}$.
Theorem A.2 (Multiplicative Chernoff Bound).
Let $X=\sum_{i=1}^{n}X_{i}$ where each $X_{i}$ is a Bernoulli variable which takes value $1$ with probability $p_{i}$ and 00 with probability $1-p_{i}$. Let $\mu=\mathbb{E}[X]=\sum_{i=1}^{n}p_{i}$. Then, it holds:
  1. 1.
    Upper Tail: $\mathsf{Pr}(X\geq(1+\psi)\cdot\mu)\leq\exp\left(-\frac{\psi^{2}\mu}{2+\psi}\right)$ for all $\psi>0$;
  2. 2.
    Lower Tail: $\mathsf{Pr}(X\leq(1-\psi)\cdot\mu)\leq\exp\left(-\frac{\psi^{2}\mu}{3}\right)$ for all $10<\psi<$ 1.
A.2 Proof of Lemma 2.14
We now restate and prove Lemma 2.14. See 2.14
Proof.
Let $G_{1},G_{2}$ be neighboring graphs, and we will let $\Pr[\mathcal{M}(G_{1})=z]$, $\Pr[\mathcal{M}(G_{2})=z]$ denote the density functions of $\mathcal{M}(G_{1})$, $\mathcal{M}(G_{2})$ evaluated at $z\in\mathbb{R}^{k}$, by some abuse of notation. To prove $\epsilon$-differential privacy, we need to show that the ratio of $\Pr[\mathcal{M}(G_{i})=z]$ is upper bounded by $\exp(\epsilon)$, for any $z\in\mathbb{R}^{k}$.
First, we define some more notation. Let $\mathcal{M}_{i}(G_{1})$, $\mathcal{M}_{i}(G_{2})$ denote the output of the mechanism $\mathcal{M}$ on graphs $G_{1}$, $G_{2}$ when answering the $i^{th}$ (adaptive) query. Via some more abuse of notation, let $\Pr[\mathcal{M}_{i}(G_{1})=z_{i}|\mathcal{M}_{j}(G_{1})=z_{j}\text{ for }j\in[i-1]]$ and $\Pr[\mathcal{M}_{i}(G_{2})=z_{i}|\mathcal{M}_{j}(G_{2})=z_{j}\text{ for }j\in[i-1]]$ denote the conditional density functions of $\mathcal{M}_{i}(G_{1})$ and $\mathcal{M}_{i}(G_{2})$ evaluated at $z_{i}\in\mathbb{R}$, conditioned on the events $\mathcal{M}_{j}(G_{1})=z_{j}$ and $\mathcal{M}_{j}(G_{2})=z_{j}$ for $j=1,\ldots,i-1$.
Finally, fix $z\in\mathbb{R}^{k}$. We have the following:
$\frac{\Pr[\mathcal{M}(G_{1})=z]}{\Pr[\mathcal{M}(G_{2})=z]}$ $=\frac{\prod_{i=1}^{k}\Pr[\mathcal{M}_{i}(G_{1})=z_{i}|\mathcal{M}_{j}(G_{1})=z_{j}\text{ for }j\in[i-1]]}{\prod_{i=1}^{k}\Pr[\mathcal{M}_{i}(G_{1})=z_{i}|\mathcal{M}_{j}(G_{1})=z_{j}\text{ for }j\in[i-1]]}$ (3)
$=\frac{\prod_{i=1}^{k}\exp\left(-\frac{\epsilon|f_{i}(G_{1})-z_{i}|}{\Delta}\right)}{\prod_{i=1}^{k}\exp\left(-\frac{\epsilon|f_{i}(G_{2})-z_{i}|}{\Delta}\right)}$ (4)
$=\prod_{i=1}^{k}\exp\left(-\frac{\epsilon(|f_{i}(G_{1})-z_{i}|-|f_{i}(G_{2})-z_{i}|)}{\Delta}\right)$
$\leq\prod_{i=1}^{k}\exp\left(-\frac{\epsilon|f_{i}(G_{1})-f_{i}(G_{2})|}{\Delta}\right)$ (5)
$=\exp\left(-\frac{\epsilon\sum_{i=1}^{k}|f_{i}(G_{1})-f_{i}(G_{2})|}{\Delta}\right)$
$=\exp\left(-\frac{\epsilon\|f(G_{1})-f(G_{2})\|_{1}}{\Delta}\right)$
$\leq\exp(\epsilon).$ (6)
In the above, equality (3) follows by the chain rule for condition probabilities, equality (4) is just writing out the density function of the Laplace distribution since we are conditioning on the answers $\mathcal{M}_{j}(G_{1})$ and $\mathcal{M}_{j}(G_{2})$ for the previous queries, inequality (5) follows by the triangle inequality, and inequality (6) follows since $\Delta$ is the $\ell_{1}$-sensitivity of $f$. ∎
A.3 Multi-Response Sparse Vector Technique
[…]
We use the class $\textsc{SVT}(\varepsilon,\Delta,c)$ (Algorithm 10) where $\varepsilon$ is our privacy parameter, $\Delta$ is an upper bound on the maximum sensitivity of incoming queries, and $c$ is the maximum number of “above” queries we can make. The class provides a $\textsc{ProcessQuery}(query,threshold)$ function where $query$ is the query to SVT and $threshold$ is the threshold that we wish to check whether the query exceeds.
Theorem A.3 (Theorem 2 in [lyu2017understanding]).
Algorithm 10 is $\varepsilon$-DP.
We remark that the version of SVT we employ (Algorithm 10) does not require us to resample the noise for the thresholds (Algorithm 10) after each query but we do need to resample the noise (Algorithm 10) for the queries after each query.
1 Input: privacy budget $\varepsilon$, upper bound on query sensitivity $\Delta$, maximum allowed “above” answers $c$
2 Class SVT($\varepsilon,\Delta,c$)
3       $\varepsilon_{1},\varepsilon_{2}\leftarrow\nicefrac{{\varepsilon}}{{2}}$
4       $\rho\leftarrow\text{Lap}(\nicefrac{{\Delta}}{{\varepsilon_{1}}})$
5       $\operatorname{count}\leftarrow 0$
6
7       Function ProcessQuery($f_{t}(D),\tau_{t}$)
8             if $\operatorname{count}>c$ then
9                   return “abort”
10
11            if $f_{t}(D)+\text{Lap}({2c\Delta}/{\varepsilon_{2}})\geq\tau_{t}+\rho$ then
12                   return “above”
13                   $\operatorname{count}\leftarrow\operatorname{count}+1$
14
15            else
16                   return “below”
17
18
19
Algorithm 10 Sparse Vector Technique
Appendix B Lower Bound for Explicit Fractional Matchings
We now prove Theorem 1.2 which is restated below for convenience: in the basic 0/1 weight $(\varepsilon,\delta)$-edge DP model (Definition 2.7), the required error is essentially linear, even if we allow for the algorithm to output a fractional $b$-matching for very large $b$. This proof is essentially the same as the proof of Theorem 1.1, but we give it explicitly for completeness.
See 1.2
Given a graph $G=(V,E)$, let $\nu^{f}(G)$ denote the size of the largest fractional matching in $G$ Given a (randomized) algorithm $\mathcal{A}$ which outputs some fractional $b$-matching $X_{G}^{\mathcal{A}}$ when run on input graph $G=(V,E)$, i.e. $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$ is a random vector satisfying $\sum_{v\neq u}X_{G,uv}^{\mathcal{A}}\leq b$ for all $u\in V$ with probability 1, let $\nu_{\mathcal{A}}^{f}(G)=\nu^{f}((V,X_{G}^{\mathcal{A}}\big{|}_{E}))$ be the size of a maximum fractional matching in the graph restricted to edges that are in $G$.
We now begin our proof of Theorem 1.2. At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument we will instead use a symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that $\text{E}[X_{G,uv}^{\mathcal{A}}]$ for each non-edge $uv\notin E$ is identical. This makes it easy to argue that high-quality algorithms cannot satisfy differential privacy.
We begin with a description of our inputs and some more notation. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-differentially private algorithm which always returns a fractional $b$-matching $X_{G}^{\mathcal{A}}$ on $V$ as defined above. For $G\in\mathcal{M}$ and $uv\in\binom{V}{2}$, let $\text{E}[X_{G,uv}^{\mathcal{A}}]$ denote the expected fractional value that $\mathcal{A}(G)$ assigns to $uv$. Let $\nu_{\mathcal{A}}^{f}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}^{f}(G)]$ denote the worst-case expected utility of $\mathcal{A}$ over $\mathcal{M}$. Note that for $G\in\mathcal{M}$, every subset of $E(G)$ is a matching. By linearity of expectations, this is equivalent to saying that $\sum_{uv\in E(G)}{\text{E}[X_{G,uv}^{\mathcal{A}}]}\geq\nu_{\mathcal{A}}^{f}$ for all $G\in\mathcal{M}$.
With this notation in hand, we can now define and prove the main symmetry property.
Lemma B.1.
If $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-DP and has expected utility at least $\nu_{\mathcal{A}}^{f}$ over $\mathcal{M}$, then there is an algorithm $\mathcal{A}^{\prime}$ which also satisfies $(\varepsilon,\delta)$-DP and has $\nu_{\mathcal{A}^{\prime}}^{f}\geq\nu_{\mathcal{A}}^{f}$. Moreover $\mathcal{A}^{\prime}$ satisfies the following symmetry property: $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\text{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ for all $uv,ab\not\in E(G)$.
Proof.
Consider the following algorithm $\mathcal{A}^{\prime}$. We first choose a permutation $\pi$ of $V$ uniformly at random from the space $S_{V}$ of all permutations of $V$. Let $\pi(G)$ denote the graph isomorphic to $G$ obtained by applying this permutation, i.e., by creating an edge $\{\pi(u),\pi(v)\}$ for each edge $\{u,v\}\in E(G)$. We run $\mathcal{A}$ on $\pi(G)$, and let $\pi(H)$ denote the resulting fractional $b$-matching. We then “undo” $\pi$ to get the fractional $b$-matching $X_{G,uv}^{\mathcal{A}^{\prime}}=X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}$. We return $X_{G}^{\mathcal{A}^{\prime}}$.
We first claim that $\mathcal{A}^{\prime}$ is $(\varepsilon,\delta)$-DP. This is straightforward, since it simply runs $\mathcal{A}$ on a uniformly random permuted version of $G$, and then does some post-processing. Slightly more formally, for any neighboring graphs $G$ and $G^{\prime}$ and for any $\pi\in S_{V}$, the graphs $\pi(G)$ and $\pi(G^{\prime})$ are also neighboring graphs. Hence the differential privacy guarantee of $\mathcal{A}$ implies that running $\mathcal{A}$ on $\pi(G)$ and $\pi(G^{\prime})$ is $(\varepsilon,\delta)$-DP, and then “undoing” $\pi$ is simply post-processing. So $\mathcal{A}^{\prime}$ has the same privacy guarantees as $\mathcal{A}$.
Now we claim that $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$. Since $\pi$ was chosen uniformly at random, we know that $\pi(G)$ is distributed uniformly among $\mathcal{M}$. Hence for any $G\in\mathcal{M}$ we have that
$\text{E}[\nu_{\mathcal{A}^{\prime}}^{f}(G)]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\text{E}[\nu_{\mathcal{A}}^{f}(\pi(G))]\geq\frac{1}{n!}\sum_{\pi\in S_{V}}\nu_{\mathcal{A}}^{f}=\nu_{\mathcal{A}}^{f}\,.$
Since this is true for all $G\in\mathcal{M}$, we have that $\nu_{\mathcal{A}^{\prime}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}^{\prime}}(G)]\geq\nu_{\mathcal{A}}$, as claimed.
We now prove the symmetry property. Note that there are $n!$ permutations in $S_{V}$, but only $\frac{n!}{2^{n/2}(n/2)!}$ different perfect matchings (i.e., $|\mathcal{M}|=\frac{n!}{2^{n/2}(n/2)!}$). This is because $\pi(G)$ can equal $\pi^{\prime}(G)$ for many for many $\pi\neq\pi^{\prime}$. In particular, it is not hard to see that |$|\{\pi\in S_{V}:\pi(G)=G^{\prime}\}|=2^{n/2}(n/2)!$ for all $G,G^{\prime}\in\mathcal{M}$. Fix $G,G^{\prime}\in\mathcal{M}$, and let $\Pi(G,G^{\prime})=\{\pi\in S_{V}:\pi(G)=G^{\prime}\}$. It is not hard to see that if we draw a permutation $\pi$ uniformly at random from $\Pi(G,G^{\prime})$, then for any $\{u,v\}\not\in E(G)$, the pair $\{\pi(u),\pi(v)\}$ is uniformly distributed among the set $\{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})\}$. Thus we have that
$\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\text{E}[X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\pi\in\Pi(G,G^{\prime})}\text{E}[X_{G^{\prime},\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{ab\not\in E(G^{\prime})}\text{E}[X_{G^{\prime},ab}^{\mathcal{A}}].$
Hence $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ is actually independent of $u,v$, and thus for all $uv,ab\not\in E(G)$, we have that $\text{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\text{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ as claimed. ∎
Lemma B.1 implies that if we can prove an upper bound on the expected utility for DP algorithms that obey the symmetry property, then the same upper bound applies to all DP algorithms. We will now prove such a bound.
Lemma B.2.
Let $\mathcal{A}$ be an algorithm which is $(\varepsilon,\delta)$-DP, always outputs a fractional $b$-matching, and satisfies the symmetry property of Lemma B.1. Then
$\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n.$
Proof.
Fix some graph $G\in\mathcal{M}$, and let $uv,ab\in E(G)$. Let $G^{\prime}$ be the graph obtained by removing $uv$ and $ab$, and adding $ub$ and $av$. Note that $G$ and $G^{\prime}$ are distance $2$ away from each other.
For all $y\neq u,b$, note that $uy\notin E(G^{\prime})$. Thus it must be the case that $\text{E}[X_{G,uy}^{\mathcal{A}}]\leq b/(n-2)$, since otherwise the symmetry property of Lemma B.1 and linearity of expectations implies that the expected (fractional) degree of $u$ in the output of $\mathcal{A}(G)$ is at least $\sum_{y\neq u,b}\text{E}[X_{G,uy}^{\mathcal{A}}]>(n-2)\frac{b}{n-2}=b$, contradicting our assumption that $\mathcal{A}$ always outputs a fractional $b$-matching. Thus $\text{E}[X_{G^{\prime},uv}^{\mathcal{A}}]\leq b/(n-2)$. Since $G$ and $G^{\prime}$ are at distance $2$ from each other, this implies that
$\text{E}[X_{G,uv}^{\mathcal{A}}]$ $=\int_{0}^{1}x\cdot\Pr[X_{G,uv}^{\mathcal{A}}=x]\mathop{}\!\mathrm{d}x$
$\leq\int_{0}^{1}x\left(e^{2\varepsilon}\cdot\Pr[X_{G^{\prime},uv}^{\mathcal{A}}=x]+2\delta\right)\mathop{}\!\mathrm{d}x$
$=e^{2\varepsilon}\cdot\text{E}[X_{G^{\prime},uv}^{\mathcal{A}}]+\delta$
$\leq e^{2\varepsilon}\frac{b}{n-2}+\delta\,.$
Since this argument applies to all $uv\in E(G)$, we have that
$\text{E}[\nu_{\mathcal{A}}^{f}(G)]$ $=\sum_{uv\in E(G)}\text{E}[X_{G,uv}^{\mathcal{A}}]$
$\leq\sum_{uv\in E(G)}\left(e^{2\varepsilon}\frac{b}{n-2}+\delta\right)$
$\leq e^{2\varepsilon}\frac{bn}{2(n-2)}+\delta\frac{n}{2}.$
$\leq e^{2\varepsilon}b+\delta n.$
Since $\nu_{\mathcal{A}}^{f}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}^{f}(G)]$, this implies that $\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n$, as claimed. ∎
We can now use these two lemmas to prove Theorem 1.2.
Proof of Theorem 1.2.
The combination of Lemma B.1 and Lemma B.2 imply that any $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ which always outputs a fractional $b$-matching has $\nu_{\mathcal{A}}^{f}\leq e^{2\varepsilon}b+\delta n$. We also know that $\nu^{f}(G)=n/2$ for all $G\in\mathcal{M}$. Thus by the definition of an $(\alpha,\beta)$-approximation, it must be the case that $\alpha(e^{2\varepsilon}b+\delta n)+\beta\geq n/2$. ∎