### v3

# Differentially Private Matchings: Symmetry Lower Bounds, Arboricity Sparsifiers, and Public Vertex Subset Mechanism

[…]
$\AtEveryCite\localrefcontext[sorting=ynt$]
Differentially Private Matchings: Symmetry Lower Bounds, Arboricity Sparsifiers, and Public Vertex Subset Mechanism
Michael Dinitz Johns Hopkins University mdinitz@cs.jhu.edu George Z. Li Carnegie Mellon University gzli@andrew.cmu.edu Quanquan C. Liu Yale University quanquan.liu@yale.edu Felix Zhou Yale University felix.zhou@yale.edu
Computing matchings in graphs is a foundational algorithmic task. Despite extensive interest in differentially private (DP) graph analysis, work on privately computing matching solutions, rather than just their size, has been sparse. The sole prior work in the standard model of pure $\varepsilon$-differential privacy, by [hsu2014private, STOC’14 ], focused on allocations and was thus restricted to bipartite graphs. This paper presents a comprehensive study of differentially private algorithms for maximum matching and $b$-matching in general graphs, which also yields techniques that directly improve upon prior work in the bipartite setting. En route to solving these matching problems, we develop a set of novel techniques with broad applicability, including a new symmetry argument for DP lower bounds, the first private arboricity-based sparsifiers for node-DP, and the novel Public Vertex Subset Mechanism. We demonstrate the versatility of these tools by applying them to other DP problems, such as vertex cover [GuptaLMRT10, SODA’10], and beyond DP, such as low-sensitivity algorithms [varma2023average, SODA’21,SICOMP’23]. Our contributions include the following:
  • We present the first utility improvements for $\varepsilon$-differentially private bipartite matching since [hsu2014private].
  • We establish a strong lower bound on the error for computing explicit matching solutions, using a novel symmetry argument. This bound demonstrates a fundamental barrier, holding even when the algorithm is permitted to output non-edges. We further generalize this to a lower bound on graph sparsity for synthetic graph generation.
  • Motivated by our lower bound, we introduce a framework for implicit solutions and develop the Public Vertex Subset Mechanism. This new tool allows us to achieve tight (bicriteria) upper and lower bounds for the implicit matching problem.
  • We implement our primary algorithm for implicit matching in the local edge-DP (LEDP) model, achieving logarithmic round complexity.
  • To address the challenging node-DP setting, we introduce the first arboricity-based graph sparsifiers under differential privacy. We demonstrate the independent utility of these sparsifiers by applying them to the node-DP vertex cover and low-sensitivity matching problems.
  • We show that all of our algorithms can be implemented in the continual release model, providing strong guarantees under both edge- and node-privacy.
1 Introduction
Computing matchings is a foundational problem in graph algorithms with broad applications. While decades of research have yielded efficient non-private algorithms [edmonds1965maximum, HK71, MV80, feldman2024maximum, assadi2022semi, kapralov2013better, CHS09, Fischer20, neiman2015simple, BDL21, GP13, BFS12, FischerN18, balliu2021lower, konrad2024unconditional] across numerous computational models, many modern applications, from ad allocation, kidney exchange, to social recommendations, involve sensitive relational data [mehta2013online, BMW22, wang2018privacy, yi2016practical]. For such settings, differential privacy (DP) has emerged as the gold standard for simultaneously providing rigorous privacy guarantees and ensuring good quality solutions. Despite a surge in recent interest in DP graph algorithms and analytics [peng2025synthetic, AHS21, NRS07, GuptaLMRT10, BGM22, DMN23, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, RS16, raskhodnikova2016differentially, Upadhyay13], algorithms that privately compute the matching solution itself, rather than just its scalar value, remain underexplored. The foundational work of [hsu2014private, STOC’14 ] in the central $\varepsilon$-DP model is the sole prior work and was restricted to bipartite graphs for allocation problems. As in the non-private setting, general matching is a harder problem than bipartite matching and requires a new set of techniques.
This paper initiates a comprehensive study of differentially private algorithms for maximum matching and $b$-matching in general graphs; we also give the first utility improvements for the special case of $\varepsilon$-DP bipartite matching since [hsu2014private]. Along the way, we develop broadly useful techniques: a new symmetry argument yielding lower bounds for private explicit solutions, the first arboricity-based sparsifiers under node-DP (together with a novel impossibility proof for their public release), and the Public Vertex Subset Mechanism (PVSM) for privately selecting and announcing differentially private vertex subsets. We show how these tools lead to tight (bicriteria) upper and lower bounds for implicit matching, efficient implementations in the local edge DP (LEDP) model [qin17generating, IMC21locally, IMC21communication, DLRSSY22, ELRS22, MPSL25] with logarithmic communication rounds, and applications to other problems such as vertex cover. We also implement all of our matching algorithms in the continual release model under both edge- and node-privacy, achieving better guarantees compared to previous work.
A maximum matching solution is fundamentally a set of edges. However, releasing any set composed exclusively of edges from the input graph inherently violates differential privacy by directly revealing the existence of those sensitive edges. Therefore, any differentially private algorithm that outputs an explicit matching solution must be permitted to include “non-edges”; that is, given an input graph $G=(V,E)$, it must include edges not present in $E$ in order to ensure that released edges do not just contain those that actually exist. A common relaxation, therefore, is to permit the algorithm to output an explicit matching $H$ on the complete graph $K_{n}$, which may include “non-edges” (edges from $K_{n}\setminus E$). This model is best understood by viewing the input $G$ as a 0/1-weighted complete graph: edges in $E$ have weight 1, and non-edges have weight 0. Edge-DP in this context means neighboring graphs differ by a single edge’s weight (a 0/1 flip). The algorithm outputs an unweighted explicit matching $H$, and its utility is measured by summing the (private) weights of the edges it contains (i.e., its intersection with $E$, $|H\cap E|$). We compare this utility to the non-private maximum matching size, $\operatorname{OPT}$, in $G$. Formally, a $(\gamma,\beta)$-approximation outputs a matching $H$ whose total weight is at least $(\operatorname{OPT}/\gamma)-\beta$.
We prove a fundamental barrier to this approach using a novel symmetry argument. Our lower bound demonstrates two key implications: first, if the output $H$ is constrained to be a matching, it is impossible to achieve any reasonable utility. Second, even if we relax this to allow $H$ to be a $b$-matching (where the degree of every vertex in $H$ can be at most $b$), the degree $b$ must be large to guarantee non-trivial utility. Thus, our lower bound demonstrates a trade-off: to achieve high utility (i.e., include many 1-weight edges), the output $H$ must have a large degree $b$. If $b$ is small (e.g., $b=1$ for a true matching), the utility is necessarily low. The below theorem illustrates a simplified version of our lower bound:
Theorem 1.1 (Simplified Lower Bound for Explicit Solutions; See Theorem 2.1).
Let $\mathcal{A}$ satisfy $(\varepsilon,\delta)$-edge DP, where $\varepsilon=O(1)$ and $\delta=\operatorname{poly}(\nicefrac{{1}}{{n}})$, and always output a graph $H$ of maximum degree at most $b$. If $\mathcal{A}$ returns a $(\gamma,\beta)$-approximation (even in expectation) to the maximum matching on $0/1$-weighted inputs, then:
  • If $b=1$, then either $\gamma=\Omega(n)$ or $\beta=\Omega(n)$, or
  • If $b>1$, then $\beta=\Omega(n-\gamma\cdot b)$ or $\gamma=\Omega\left(\frac{n-\beta}{b}\right)$.
Theorem 1.1 shows that if we restrict $H$ to be a matching, then either the multiplicative or additive error must be $\Omega(n)$. It further shows that even if we allow $H$ to be a $b$-matching for constant $b$, the multiplicative or additive error must still be $\Omega(n)$. Thus, we cannot use the returned graph $H$ as an explicit solution and expect good utility.
Our lower bound thus motivates the study of implicit solutions for matching using the billboard model. The billboard model is a standard framework in differential privacy where a differentially private transcript (the “billboard”) is released publicly. While this public data alone is insufficient to generate a solution to the problem, each node (or “user”) can then privately decode its own high-utility solution (in the case of matching, its matched edges) by combining the public billboard with its own private adjacency list.
We study a natural form of implicit solutions (Definition 3.6) where, roughly speaking, each vertex can implicitly select a subset of its neighbors to be matched with. Even for this particular implicit solution, we show that any $\gamma$-approximate $\varepsilon$-edge DP maximum matching must contain at least one node matched with $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes. This lower bound is also proven using our symmetry argument.
Theorem 1.2 (Simplified Implicit Solution Lower Bound; See Theorem 5.3).
For any $\gamma>1$, a $\gamma$-approximate implicit solution in the billboard model to the maximum matching problem (Definition 3.6), produced by an $\varepsilon$-edge DP algorithm, contains at least one node matched to at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes, with probability at least $\nicefrac{{1}}{{2}}$.
We emphasize the critical difference between the solution types given in Theorem 1.1 and Theorem 1.2. Theorem 1.1 is a lower bound for explicit solutions in which the publicly released solution itself gives the matching. Theorem 1.2 instead gives a lower bound for implicit solutions that are locally decoded by each node using the public billboard in combination with its own private adjacency list. Hence, Theorem 1.2 is not a trivial consequence of Theorem 1.1, as the set of matched neighbors within an implicit solution is locally decoded by each vertex using the billboard and its adjacency list, rather than explicitly output by the algorithm.
Due to Theorem 1.2, any $\varepsilon$-DP implicit solution with good utility must be a $b$-matching with $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$. There are many settings in which we might want a private matching (or $b$-matching) in a general graph. For example, in online dating, an individual is often provided with a shortlist of top candidates (resulting in a $b$-matching). The matching should be kept private but individuals in every pair should know who they are matched with. We design an algorithm that produces implicit $(2+\eta)$-approximate $b$-matchings for any $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, which is tight against our lower bound for implicit solutions.
Theorem 1.3 (Simplified Implicit $b$-Matching; See Theorem 2.2).
Let $\varepsilon,\eta\in(0,1)$, and $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$. There is an $\varepsilon$-DP algorithm that, with high probability, outputs an implicit $(2+\eta)$-approximate $b$-matching in the billboard model.
This paper establishes a comprehensive set of upper and lower bounds for $\varepsilon$-DP maximum matching and $b$-matching; we present a complete set of novel results across both edge- and node-privacy in the central [DMNS06], local (LEDP) [IMC21locally, DLRSSY22], and continual release models [CSS11, DNPR10]. See Section 1.1 for a summary of our contributions. Our results are driven by three primary novel technical contributions that have broad-reaching applications beyond our paper to other graph algorithms as well.
Symmetry Argument for Graph DP Lower Bounds.
Our lower bound for explicit matchings is driven by a novel symmetry argument inspired by distributed graph algorithms. The core idea is to analyze a “symmetrized” algorithm $\mathcal{A}^{\prime}$ created by wrapping any $(\varepsilon,\delta)$-edge DP algorithm $\mathcal{A}$ as follows: on input $G$, $\mathcal{A}^{\prime}$ computes $\pi^{-1}(\mathcal{A}(\pi(G)))$ for a uniformly random vertex permutation $\pi$. This wrapper preserves $(\varepsilon,\delta)$-DP and worst-case expected utility. We then show that for certain graph problems, such as maximum matching, a large set of random vertex permutations must give equal utility guarantees by definition of DP, leading to high error.
We analyze this on the hard instance family $\mathcal{M}$ of all perfect matchings. When fractional matchings are allowed, symmetrization enforces that, for any $G\in\mathcal{M}$, the expected fractional weight assigned to every non-edge is identical. We then consider a pair of inputs $G_{1},G_{2}\in\mathcal{M}$ which differ by two edges (i.e., given edges $\{u,v\},\{a,b\}\in G_{1}$, create edges $\{u,b\}$ and $\{a,v\}$ in $G_{2}$). The DP guarantee bounds the expected weight on any true edge by $e^{4\varepsilon}$ times the average non-edge weight (plus a $\delta$ term). The degree bound on the returned matching then bounds this average non-edge weight by the algorithm’s degree budget $b$. Summing $\delta$ over the $n/2$ edges of the input matching yields our main lower bound: any $(\gamma,\beta)$-approximation outputting a graph of maximum degree $b$ must satisfy
[…]
For constant $\varepsilon$ and $\delta\leq 1/n$, this implies a stark trade-off: either $b=\Omega(n)$ or the multiplicative $\gamma$ or additive error $\beta$ must be large (potentially $\Omega(n)$ for constant $b$).
This symmetrization argument extends almost verbatim to the fractional “synthetic graph” relaxation (Definition 4.1), replacing the maximum degree $b$ with the average degree budget $\bar{b}=\frac{1}{n}\sum_{v}b_{v}$:
[…]
Thus, any DP algorithm aiming for low error must produce an output synthetic graph with linear average degree. This information-theoretic barrier demonstrates why dense outputs are unavoidable for explicit solutions, explaining the limitations of mechanisms like randomized response and formally motivating our focus on the billboard model with implicit solutions for our upper bounds. We believe our symmetry argument, used to establish multiple lower bounds in our paper, holds significant promise for proving lower bounds to a wider class of DP graph problems beyond matchings.
Public Vertex Subset Mechanism (PVSM).
At the core of our implicit algorithm in the billboard model is the Public Vertex Subset Mechanism (PVSM), a novel technique within graph DP of potential independent interest. Consider an input graph $G=(V,E)$ where each node wants to select a subset of its neighbors as part of the solution to a graph problem. The Public Vertex Subset Mechanism (PVSM) is our new graph DP tool that allows each node to implicitly select a subset of its neighbors. The mechanism’s public output, which is posted to the billboard, does not reveal the specific neighbors chosen. However, it is designed so that each neighbor can combine this public information with their private adjacency list to locally decode whether they were selected, thus ensuring the privacy of the individual edges is maintained.
The mechanism works as follows: first, for each pair of nodes $\{u,v\}$ and parameter $r$, a public coin is flipped with probability $p_{r}=(1+\eta)^{-r}$. These coin flips are all public, collectively defining a set of random subgraphs picked from the complete graph $K_{n}$. A “proposer” node $v$ privately selects an optimal parameter $r^{*}$ (using, e.g., the AboveThreshold mechanism) indicating a subset of neighbors to select. The selection of $r^{*}$ balances utility (capturing many neighbors) with a degree constraint on the output. Crucially, $v$ publicly releases only the index $r^{*}$ to the billboard.
Any “receiver” node $w$ can then locally decode its status using $r^{*}$ and the public coin flips. Node $w$ checks its private adjacency list to see if it is a neighbor of $v$. If it is, $w$ consults the public coin for the pair $\{v,w\}$ at level $r^{*}$. In the context of matchings, if the coin is heads, $w$ knows it has been selected by $v$, thus forming the implicit matched edge. This mechanism provides a general method for privately selecting and publicly encoding graph subsets, and we believe it can be readily adapted to other graph problems in the billboard model. This mechanism yields Theorem 1.3.
Arboricity-Based Node-DP Sparsifiers.
Beyond edge-DP, we also study node-DP, where neighboring input graphs differ by a single node and all its incident edges. We design the first arboricity-based sparsifiers tailored to matching. These sparsifiers (privately) reduce vertex degrees to $\widetilde{O}(\alpha)$ (where $\alpha$ is the arboricity), while maintaining a large enough ($b$-)matching in the sparsified graph. Our sparsifiers guarantee a property known as stability which ensures that the number of edges that differ from any (sparsified) neighboring graph is also bounded by $\widetilde{O}(\alpha)$. The sparsification reduces the edge sensitivity from a worst case $O(n)$ to $\widetilde{O}(\alpha)$. Then, we run our edge-DP algorithms on this sparsified graph to yield high-utility node-DP algorithms for maximum $b$-matching. This yields:
Theorem 1.4 (Simplified Node-DP $b^{\prime}$-Matching; See Theorem 2.6).
Let $\alpha$ denote the arboricity of the input graph $G=(V,E)$. We give a $\varepsilon$-node DP algorithm in the billboard model that, with high probability, outputs an implicit $b$-matching whose size is at least a $(2+\eta)$-approximate $b^{\prime}$-matching, for
$b=O\left(\frac{\alpha\log n}{\eta\varepsilon}+\frac{\log^{2}n}{\eta\varepsilon^{2}}+\frac{b^{\prime}\log n}{\varepsilon}\right).$
Prior work on node-DP sparsification has focused exclusively on reducing graphs using bounded maximum degree (and almost bounded maximum degree)  [BBS13, Day16, kasiviswanathan2013analyzing, wagaman2024time, CZ13]. Our work introduces the first node-DP sparsifier designed around a more general graph property: arboricity. This technical advancement is significant because the class of graphs with bounded arboricity (which includes, for example, all planar graphs and graphs from any minor-closed family) is substantially broader than the class of bounded-maximum-degree graphs.
We demonstrate the independent use of our novel arboricity-based sparsifier by applying it to develop a new node-DP algorithm for vertex cover as well as a new low-sensitivity matching algorithm [varma2023average, yoshida2021sensitivity, yoshida2026low]. We believe arboricity sparsification is of independent interest beyond DP matching.
Improvement on the Utility of [hsu2014private].
For bipartite graphs with supply $s=O\left(\frac{\log^{3.5}(n)}{\varepsilon}\right)$, [hsu2014private] design a node-DP $(1+\eta)$-approximate one-sided $s$-matching algorithm. As an extension of our techniques, we improve the bound to $s=O\left(\frac{\log n}{\varepsilon}\right)$ while maintaining the $(1+\eta)$-approximation guarantee (Theorem 2.8).
1.1 Summary of Results
We give the following set of results which we describe in more detail in Section 2. See Table 1 for a comprehensive summary of quantitative results.
  • Explicit solutions are provably hard. We prove a strong lower bound for computing explicit matching solutions under edge-DP via a new symmetry argument. The bound holds even if the algorithm is allowed to output non-edges and even when we only compare the returned $b$-matching to the size of a (non-private) maximum matching. We further generalize this to a lower bound on the sparsity of differentially private synthetic graph generation.
  • Implicit solutions via PVSM. Motivated by the lower bound, we work in the billboard model that allows for the release of implicit solutions. We introduce the Public Vertex Subset Mechanism (PVSM), a generic mechanism for differentially privately selecting public vertex subsets that are locally decodable by a node’s neighbors using only public coins and their own private adjacency lists. We give tight (bicriteria) upper and lower bounds for implicit maximum matching and $b$-matching in general graphs.
  • Local edge-DP (LEDP) in logarithmic rounds. We implement our implicit matching algorithm in the $\varepsilon$-LEDP model. Our sequential $\varepsilon$-LEDP billboard algorithm achieves optimal dependence on $b=O\left(\frac{\log n}{\varepsilon}\right)$ and a $(2+\eta)$-approximate maximum $b$-matching. We then give a distributed $\varepsilon$-LEDP algorithm that completes in $O(\log n)$ rounds using a proposer/receiver mechanism inspired by non-private distributed algorithms.
  • Node-DP via arboricity sparsification. To handle node-DP, we design the first arboricity-based sparsifiers under differential privacy, and use them to obtain implicit node-DP algorithms for matching in general graphs. We also show an impossibility for the public release of such sparsifiers. As another application of our sparsifier beyond matchings, we obtain a node-DP vertex cover algorithm.
  • Improving [hsu2014private]. We improve the bipartite node-DP guarantees of [hsu2014private] by decreasing $s$ (the one-sided degree parameter of the matching given in [hsu2014private]) from $s=O\left(\frac{\log^{3.5}(n)}{\varepsilon}\right)$ to $s=O\left(\frac{\log(n)}{\varepsilon}\right)$ while maintaining a $(1+\eta)$-approximate $s$-maximum matching.
  • Continual release. All of our results extend to the continual release model for both edge- and node-privacy, with accuracy guarantees that hold for every update.
1.2 Related Work
Private Allocations.
As discussed, the closest work to ours is the private allocation paper by [hsu2014private]. They give a node-differentially private algorithm that implicitly outputs an approximate maximum matching solution on bipartite graphs. The authors model any bipartite graph as a set of bidders (on one side) and goods (on the other). Then, they formulate a differentially private version of the ascending auction price algorithms of [kelso1982job] to compute Walrasian equilibrium prices of the goods (a set of prices ensuring optimal utility of buyers and no good is over-demanded). The algorithm releases differentially private counts of the number of bids each good has received as an implicit solution, using the private counting algorithms of [CSS11, DNPR10]. Assuming the supply $s$ satisfies $s=\Omega(\log^{3.5}(n)/\varepsilon)$, this yields a node-DP $(1+\eta)$-approximate $s$-matching.
Now, auction-based methods and the implicit solution concept of [hsu2014private] are inherently tied to bipartite graphs. While it is possible to reduce general matchings to a bipartite instance via random bipartitions [liu2024pointwise], this transformation incurs a factor of 2 in the multiplicative approximation factor and the obtained approximation ratio is only in expectation. This would imply a node-DP $(2+\eta)$-approximate $s$-matching in expectation for general graphs. Our novel solution concept allows us to handle general graphs, under edge-DP or node-DP assuming low arboricity, and obtain a $(2+\eta)$-approximate $s$-matching with high probability, while reducing the lower bound to $s=\Omega(\log(n)/\varepsilon)$. Moreover, specializing our techniques to the bipartite case (with arbitrary arboricity) recovers the node-DP $(1+\eta)$-approximation ratio of [hsu2014private], while reducing the bound on $s$ to $s=O(\log(n)/\varepsilon)$.
We remark that a followup work [hsu2016jointly] studies a class of private convex programming problems that capture fractional allocations with supply $s$ under approximate $(\varepsilon,\delta)$-differential privacy. They output an $s^{\prime}$-matching with $s^{\prime}>s$, but unlike our results, they compute a fractional allocation and guarantee that the sum of constraint violations ($\sum_{j}s_{j}^{\prime}-s_{j}$) is bounded, whereas our algorithm returns an integral solution while ensuring each constraint violation is bounded ($\max_{j}s_{j}^{\prime}-s_{j}$). Furthermore, our results are given for pure $\varepsilon$-DP (and $\varepsilon$-LEDP). Thus, neither the privacy nor the utility guarantees are comparable to ours.
Although [kannan2015pareto] also study private allocations, they do so in a relaxed notion of differential privacy called marginal differential privacy with the goal of computing approximate pareto optimal allocations. Thus both their privacy and utility guarantees differ from our setting.
Private Matching Size.
Differentially private estimations of the size of the maximum matching have been recently studied in various privacy models including the continual release model [dong2024joins, FHO21, wagaman2024time, raskhodnikova2024fully, epasto2024sublinear] and the sublinear model [BGM22]. In the standard central DP model, computing an approximate solution for the size of the maximum matching in the input graph with $O(\log(n)/\varepsilon)$ additive error is trivial since the sensitivity of the size of the maximum matching is $1$ and so we can use the Laplace mechanism to obtain such an estimate.
Implicit Solutions.
Implicit solutions in the graph setting were first introduced by [GuptaLMRT10] for various combinatorial optimization problems such as set cover and vertex cover. An implicit solution is one where the released information from a differentially private algorithm is $\varepsilon$-DP or $\varepsilon$-LEDP but the released solution does not immediately solve the problem. In particular, the released solution is an implicit solution from which each individual node can decode the explicit solution for itself. In the case of our matching algorithms, the explicit solution computed by every node is the set of edges in the matching that are adjacent to it.
Low-Sensitivity Matching.
Let $\mathcal{A}$ be a randomized graph algorithm that takes as input a graph $G$ and outputs a subset of vertex pairs $F_{G}\subseteq\binom{V}{2}$. Consider edge neighboring graphs $G\sim G^{\prime}$ that differ by a single edge and recall the 1-Wasserstein distance of the output distributions with respect to the symmetric difference metric is given by the following infimum over couplings $\Pi(F_{G},F_{G^{\prime}})$ of the output distributions:
$W_{1}(\mathcal{A}(G),\mathcal{A}(G^{\prime}))=W_{1}(F_{G},F_{G^{\prime}})=\inf_{\pi\in\Pi(F_{G},F_{G^{\prime}})}\left[|(F_{G}\setminus F_{G^{\prime}})\cup(F_{G^{\prime}}\setminus F_{G})|\right]\,.$
[varma2023average] defined the (worst-case) sensitivity of a matching algorithm $\mathcal{A}$ on an input graph $G$ to be
$\max_{e\in E(G)}W_{1}(\mathcal{A}(G),\mathcal{A}(G-e))\,.$
The matching algorithm with the best sensitivity bound for bounded-degree graphs is a $(1+\eta)$-approximation algorithm with sensitivity $\Delta^{O(1/\eta)}$ over graphs of maximum degree $\Delta$ due to [yoshida2026low].
For edge-neighboring graphs $G\sim G^{\prime}$ with arboricity at most $\alpha$, our stable matching sparsifier for bounded-arboricity graphs (Section 7.1.1) maps $G\mapsto H,G^{\prime}\mapsto H^{\prime}$ to edge-neighboring graphs $H\sim H^{\prime}$ with maximum degree $\Delta=O(\nicefrac{{\alpha}}{{\varepsilon}})$ while approximately preserving the maximum matching up a $(1+\eta)$-multiplicative factor. Running the [yoshida2026low] algorithm on the sparsified graph immediately yields a $(1+\eta)$-approximate matching algorithm with sensitivity $\alpha^{\tilde{O}(1/\eta)}$. For graph classes such as planar graphs with constant arboricity but maximum degree $n$, this improves the worst-case sensitivity from $\operatorname{poly}(n)$ to $O(1)$.
2 Technical Overview
We now give a technical overview for the main ideas behind our results and give the formal theorem statements for our upper and lower bounds.
2.1 Lower Bounds for Explicit Solutions (Section 4)
Lower Bounds via Symmetrization (Sections 4.1 and 4.2).
At a very high level, our lower bound for explicit matching solutions (Theorem 2.1) uses an argument similar to the “packing arguments” of [DMN23] for minimum cut. However, instead of using a packing/averaging argument, we use a symmetry argument reminiscent of symmetry arguments in distributed computing (e.g., [kuhn2016local]) and semidefinite programming (e.g., [dam2015semidefinite]). Namely, we argue that any DP algorithm for a collection of input graphs must be highly symmetric. Highly symmetric is defined in terms of the marginal probability of each non-existing edge; we show that the marginal probability of outputting any non-existent edge is identical.
Because it is impossible to differentially privately output any subgraph (including a matching) containing only true edges, we must allow our algorithm to output “edges” that are not true edges; in other words, our matching algorithm returns a matching where only some of the edges actually exist. This corresponds to changing the input graph $G=(V,E)$ to a weighted complete graph, where every $\{u,v\}\not\in E$ is assigned weight 0 in the complete graph and actual edges have weight $1$, and asking for a max-weight matching in this graph. We can even allow our algorithm to be bicriteria and return a $b$-matching [^1] Recall that in a $b$-matching every node has degree at most $b$, thus a traditional matching is a $1$-matching. rather than a true matching. In this case, we would want our returned $b$-matching to be of comparable size to a maximum matching, while minimizing the magnitude of $b$.
We construct a family $\mathcal{M}$ of graphs over $n$ vertices with the property that for every pair of vertices $u,v$, there are neighboring graphs $G\sim G^{\prime}\in\mathcal{M}$ where $uv$ is a matching edge in $G$ but a non-edge in $G^{\prime}$. Furthermore, we show that any algorithm with reasonable utility over $\mathcal{M}$ must have non-trivial probability of outputting $uv$ (matching edge) on input $G$. Under differential privacy constraints, the probability of outputting $uv$ (non-edge) on input $G^{\prime}$ must also be non-trivial. By the symmetry of non-edge probabilities, any DP algorithm that gives a solution with decent utility must essentially output $\Omega(n^{2})$ edges. In other words, any differentially private algorithm with reasonable approximation must output many non-edges in an explicit solution.
Using generalizations of the argument above, we can in fact prove a stronger lower bound on a relaxation of the explicit matching problem to synthetic graph generation. We then show that the specific lower bound on explicit matchings stated below follows as a corollary of this more general lower bound.
Theorem 2.1 (Lower Bound on Explicit Solutions).
Let $\mathcal{A}$ be an algorithm which satisfies $(\varepsilon,\delta)$-edge DP and always outputs a graph $H$ with maximum degree at most $b$. If $\mathcal{A}$ outputs a $(\gamma,\beta)$-approximation (even just in expectation) of the maximum matching of a $0/1$-weighted (Definition 3.7) input graph $G=(V,E)$, then
$e^{4\varepsilon}b+2\delta n\geq\frac{n}{2\gamma}-\beta.$
The proof of Theorem 2.1 is presented in Section 4. To interpret this theorem, think of $\delta\leq 1/n$ (so the $\delta$ term does not contribute anything meaningful) and small constant $\varepsilon$. Then, a few meaningful regimes of $b,\gamma,\beta$ include the following:
  • If $\gamma=1$ (so we are bounding additive loss), then $\beta\geq n/2-O(b)$. So if $b$ is at most $n/c$ for some reasonably large constant $c$, then the additive loss is still linear even though we are allowing our solution to have a quadratic number of total edges!
  • If $\beta=0$ (so we are bounding multiplicative loss), then we get that $\gamma\geq\Omega(n/b)$.
  • If $\gamma\leq\frac{\sqrt{n}}{2}$, then $\beta\geq\sqrt{n}-O(b)$. Thus, if $b\leq\sqrt{n}/c$ for some constant $c$, we still have polynomial additive loss even if we allow large polynomial multiplicative loss.
In addition to this lower bound, we also use our symmetry argument to prove the lower bounds for synthetic graph generation in Theorem 4.3 (which can be interpreted as a relaxation of explicit matchings), as well as the lower bound for implicit solutions given in Theorem 5.3.
2.2 Edge Privacy via Public Vertex Subset Mechanism (PVSM) (Sections 5 and 6)
Theorem 2.1 shows that we must find another way to return a subset of edges representing a matching. A natural approach, which we adopt through the rest of this paper, is to use “implicit” solutions in the sense of [GuptaLMRT10], which was later formalized by [hsu2014private] as the billboard model. In this model, there is a billboard of public information released by the coordinator or the nodes themselves. Using the billboard of public information and its own privately stored information, each node can privately determine which of its incident edges are in the solution (if any). So in this model, we output something (to the billboard) that is “locally decodable” at every node (where each node can also use its private information in the decoding). That is, the information released to the billboard is $\varepsilon$-differentially private and each node computes its own part of the solution by combining the public information on the billboard with its own private adjacency list. Henceforth, we refer to the information on the billboard as an “implicit solution”. We formally define this implicit solution representation in Definition 3.4; such a solution has been used previously in the context of vertex cover, set cover, and maximum matching in bipartite graphs [GuptaLMRT10, hsu2014private]. It is important to note that the billboard model is not a privacy model, but a solution release model.
Our results for edge privacy combine the billboard model of implicit solutions with a strong local version of privacy known as local edge differential privacy (LEDP) [IMC21locally, DLRSSY22]. In the LEDP model, nodes do not reveal their private information to anyone. Rather, nodes communicate with the curator over (possibly many) rounds, where in each round each node releases differentially private outputs which are accumulated in a global transcript. The transcript is publicly visible and must also be differentially private. It is very natural to combine the LEDP and billboard models by thinking of the billboard as this differentially private transcript, which is the approach we take.
Although both [hsu2014private] and our results are both in the billboard model for solution release, their algorithm hinges on the “binary mechanism” [CSS11], a mechanism for releasing a stream of $n$ private counting queries with only $O(\log n)$ error, rather than the $O(\operatorname{poly}(n))$ error attained through naive composition. This mechanism releases noisy partial sums such that each query can be answered by combining $O(\log n)$ partial sums. However, this requires a centralized curator who can compute the true partial sums before adding noise, which seems incompatible with the LEDP model.
LEDP algorithm (Section 5).
Our main algorithms (Theorem 2.2 and Theorem 2.3) take inspiration from the greedy algorithms and distributed algorithms for maximal matching, although our proofs also hold for $b$-matchings. In particular, the traditional greedy algorithm for maximal matching iterates through the nodes in the graph in an arbitrary order and matches the nodes greedily whenever there exists an unmatched node in the current iteration. However, this algorithm is clearly not private because each matched edge reveals the existence of a true edge in the graph. Thus, we must have some way of releasing public subsets of vertices that include both true edges and non-edges.
Instead, we introduce a novel DP proposal/receiver mechanism that allows nodes to propose a match set and receivers to accept the matches. We detail this proposal/receiver mechanism in Algorithm 2 and the guarantees are given in Theorem 2.2. Algorithm 2 proceeds in $n$ rounds using an arbitrary ordering of the nodes in the graph. We give as input a threshold $b$ which is a threshold for the number of edges in the matching each vertex is incident to. In the order provided by the ordering, each node $v$ privately checks using the Multidimensional AboveThreshold (MAT) technique introduced in [DLL23] whether it has reached its number of matched neighbors threshold. If it has not, then it asks all of its neighbors ordered after it whether they will match with it. This is the proposal phase of the algorithm. Each of these neighbors first determines using MAT whether they have exceeded their threshold for matching and then flips a coin (using progressively exponentially decreasing probability) to determine whether to be part of the set to match with $v$. Then, every vertex ordered after $v$ answers YES or NO depending on its coin flip. Intersecting the YES/NO responses with $v$’s private adjacency list results in $v$’s matches. This second portion of the algorithm is the receiver phase. All of these procedures can be done locally privately in the LEDP model.
Within this algorithm, we introduce the Public Vertex Subset Mechanism (PVSM) which may be of independent interest to other problems such as vertex cover, dominating set, or independent set. In this mechanism, each node releases a public subset of nodes obtained via coin flips. Specifically, for each pair of nodes, we flip a set of coins with appropriate probability parameterized by subgraph indices $r$. For index $r$, each coin is flipped with probability $(1+\eta)^{-r}$. Then, a node $v$ publicly releases a subgraph index $r$ to determine a public subset of vertices consisting of coin flips which landed heads for each pair $\{v,u\}$ containing $v$. The value of $r$ determines the size of this public subset of vertices (in expectation). Then, each node $w$ takes the public subsets released by all vertices and intersects them with their private adjacency lists to determine a private subset of vertices (consisting of neighbors of $w$) only $w$ knows about. Thus, the public subset can be written on the billboard and the private subset is used to determine each node’s implicit answer. We believe this mechanism will be helpful for other graph problems that release implicit solutions. We obtain our first upper bound result for matchings in Theorem 2.2.
Theorem 2.2 (LEDP $2$-Approximate Maximum $b^{\prime}$-Matching).
Let $\varepsilon\in(0,1)$, $\eta\in(0,1)$, and
$b\geq\frac{(1+\eta)^{2}}{1-\eta}\cdot b^{\prime}+\Omega\left(\frac{\log(n)}{\eta^{2}\varepsilon}\right)\,.$
Algorithm 2 is $\varepsilon$-LEDP and, with high probability, outputs an (implicit) $b$-matching in the billboard model that has the size of a $2$-approximate maximum $b^{\prime}$-matching.
There are two main regimes of interest, which are corollaries of Theorem 2.2:
  • If $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 is $\varepsilon$-LEDP and outputs a $b$-matching that has the size of a 2-approximate 1-matching (Corollary 5.1). Note this is a bicriteria approximation.
  • If $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$, Algorithm 2 is $\varepsilon$-LEDP and outputs a $(2+\eta)$-approximate $b$-matching (Theorem 5.2). This setting avoids the need for a bicriteria approximation.
$O(\log n)$-Round Algorithm (Section 6).
The above algorithm runs in $\Omega(n)$ rounds. To speed up our algorithm to run in polylogarithmic rounds (Theorem 2.3), we need to allow many vertices to simultaneously propose matches to their neighbors (rather than just one at a time). Intuitively, this can be done in ways similar to classical parallel and distributed algorithms for related problems like Boruvka’s parallel MST algorithm [Boruvka]. In these algorithms, nodes randomly choose whether to propose or to listen. Then, proposer nodes send proposals, receivers decide which of their proposals to accept, and receivers communicate the acceptances back to the proposers (see [DHIN19] for a recent example of this in distributed settings). However, the added LEDP constraint makes this process significantly more difficult. Note, for example, that even communicating “with neighbors” violates privacy, since the transcript would then reveal who the neighbors are! Hence, greater care must be taken regarding the messages each node transmits. Fortunately, by using techniques gleaned from our $n$-round algorithm, ideas from distributed algorithms, and improved analyses, our implementation goes through.
The crux of our $O(\log n)$ round distributed algorithm relies on our Public Vertex Subset Mechanism. We use this mechanism to release proposal sets on the part of the proposers and match sets on the part of the receivers. The proposers release public proposal sets that contain proposals to its neighbors to match. Using the Public Vertex Subset Mechanism, receivers can privately discern which nodes have proposed. Then, receivers also publish a public subset of nodes as their match sets. Using this mechanism, the proposers can then (privately) recognize which receivers accepted their proposals. Performing multiple rounds of this procedure leads to our desired Theorem 2.3. Our novel utility analyses depends on a novel noisy version of a charging argument, based on edge orientation, implicitly described in [khuller1994primal]. Our improved guarantees are given in Theorem 2.3.
Theorem 2.3 (Efficient 2-Approximate Maximum $b^{\prime}$-Matching).
Let $\varepsilon\in(0,1)$, $\eta\in(0,\tfrac{1}{2})$, and matching parameter $b^{\prime}\geq 1$ where $b$ satisfies
$b\;\geq\;\frac{(1+\eta)^{2}}{1-\eta}\,b^{\prime}\;+\;\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\,\varepsilon}\right).$
Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) $b$-matching in the billboard model whose size is at least a $2$-approximate maximum $b^{\prime}$-matching.
Similar to our sequential algorithm, we obtain the following results as special cases.
  • If $b=\Omega\left(\frac{\log^{2}(n)}{\varepsilon}\right)$, Algorithm 3 is $\varepsilon$-LEDP and outputs a $b$-matching that has the size of a 2-approximate 1-matching in $O(\log n)$ rounds (Corollary 6.1). This is a bicriteria approximation.
  • If $b=\Omega\left(\frac{\log^{2}(n)}{{\eta^{5}}\varepsilon}\right)$, Algorithm 3 is $\varepsilon$-LEDP and outputs a $(2+\eta)$-approximate $b$-matching in $O(\log n)$ rounds (Theorem 6.2). This avoids the need for a bicriteria approximation.
Implicit Solution Lower Bound (Section 5.4).
We complement our upper bounds with a lower bound which shows that even for the particular type of implicit solutions used in Theorem 2.2, our reliance on $b$ is necessary: if $b$ is sublogarithmic, then even an implicit solution cannot be a good approximation to the maximum matching in the graph. This proof also uses our symmetry argument. We start with a low-degree graph, and first use the DP guarantee over nearby graphs to argue that edges that do not exist must still be included with reasonably high probability (this is essentially a highly simplified version of the symmetry argument used in the previous lower bound). We then use group privacy to argue that there is a graph with logarithmic max degree (which is also logarithmically far from our starting graph) where many edges incident to the node of logarithmic degree must be included.
Theorem 2.4 (Informal Lower Bound on Implicit Solutions; See Theorem 5.3).
Let $\mathcal{A}$ be an algorithm which satisfies $\varepsilon$-edge DP and outputs an implicit solution of the type used in Theorem 2.2 which is the size of a $\gamma$-approximate maximal matching [^2] Given a graph $G=(V,E)$, a $b$-matching $F$ is said to be a $\gamma$-approximate maximal matching if there exists a maximal matching with size $M$ such that $F$ has size at least $M/\gamma$. If there is a maximal $b$-matching of size $M$ where $F$ has size at least $M/\gamma$, then we say that $F$ is a $\gamma$-approximate $b$-matching. with probability at least $1-\operatorname{poly}(\nicefrac{{1}}{{n}})$. Then, when decoded from the billboard, with probability at least $\nicefrac{{1}}{{2}}$, there is a node matched with at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes, regardless of the value of $\gamma>0$.
Interestingly, the value of $\eta$ does not actually affect this lower bound; a logarithmic loss in the degree is necessary for any multiplicative approximation.
2.3 Node Privacy via Arboricity Sparsification (Section 7)
In the node differential privacy setting, two graphs are considered neighbors if one can be obtained from the other by replacing all edges incident on any one node. This model is a significantly more challenging setting than edge privacy. We use a sparsifier based on the arboricity [^3] Recall that a graph has arboricity $\alpha$ if it can be partitioned into $\alpha$ edge-disjoint forests. of our input graph for our node-DP results.
Implicit Matchings (Sections 7.1 and 7.2).
To achieve node privacy in the central model, we combine our edge privacy results with an arboricity-based sparsifier inspired by the “matching sparsifier” of [solomon2021local]. Informally, a matching sparsifier of a graph $G$ is a subgraph $H$ such that 1) the maximum matching in $H$ is approximately as large as the maximum matching in $G$, and 2) the maximum degree of $H$ is at most $O(\alpha)$, where $\alpha$ is the arboricity of $G$. Note that the average degree in a graph with arboricity $\alpha$ is at most $O(\alpha)$, so such a sparsifier essentially turns the average degree into the maximum degree while approximately preserving the maximum matching.
We give an arboricity-based sparsifier that is stable, meaning that on node-neighboring input graphs, $G$ and $G^{\prime}$, the sparsified graphs, $S$ and $S^{\prime}$, will be $\widetilde{O}(\alpha)$ apart (in edge edit distance [^4] The edge edit distance between two graphs $S$ and $S^{\prime}$ is the number of edge additions and deletions to $S$ that will create $S^{\prime}$.). Our arboricity-based sparsifier ensures the maximum degree of any vertex is $\widetilde{O}(\alpha)$. Given an algorithm to compute such a stable sparsifier, we can intuitively combine this with edge privacy to get node privacy: since the maximum degree in the sparsifier is at most $\widetilde{O}(\alpha)$ and the edge edit distance to any neighboring graph is $\widetilde{O}(\alpha)$, basic group privacy implies that we can simply use our edge-private algorithm and incur an extra loss of at most an $\widetilde{O}(\alpha)$ factor. In other words, if the maximum degree is small and the edge edit distance to any neighboring graph is also small, then node privacy is approximately the same as edge privacy, and our arboricity-based sparsifier guarantees small maximum degree bounded by the $\alpha$ of the input graph (up to polylog factors).
If we could compute and publicly release a differentially private matching sparsifier, we would be done. Unfortunately, we show in Section 7.4 that this is impossible, by providing a strong lower bound against private matching sparsification (Theorem 2.5).
Theorem 2.5 (Sparsifier Public Release Lower Bound).
Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-node DP algorithm which, given an input graph $G=(V,E)$, outputs an implicit matching sparsifier $E^{\prime}$. Then, if $\operatorname{OPT}(G)\leq(1+\eta)\cdot\mathbb{E}[\operatorname{OPT}(E^{\prime}\cap E)]$, the maximum degree of $E^{\prime}\cap E$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$ even if $G$ has arboricity $1$.
Even though releasing sparsifiers does not satisfy DP, we show that we can use them in a node-DP algorithm. Intuitively, this is because given two node-neighboring graphs, computing a sparsifier of them can bring the distributions of the outputs “closer together”. In particular, although these sparsifiers cannot be released publicly, they decrease the “edge-distance” between node-neighboring graphs, where the edge-distance is the number of edges that differ between neighboring graphs. Our upper bound is given formally below.
Theorem 2.6 (Node-DP Approximate Maximum $b^{\prime}$-Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, $\alpha$ be the arboricity of the input graph, and
$b=\Omega\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta\varepsilon^{2}}+\frac{b^{\prime}\log(n)}{\varepsilon}\right)\,.$
There is an $\varepsilon$-node DP algorithm that, with high probability, outputs an (implicit) $b$-matching in the billboard model that has the size of a $\left(2+\eta\right)$-approximate maximum $b^{\prime}$-matching.
As a special case we have the following corollary:
  • If $b=\Omega\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta\varepsilon^{2}}\right)$, there is a $\varepsilon$-node DP algorithm that outputs a $b$-matching that has the size of a $(2+\eta)$-approximate 1-matching (Corollary 7.1).
Implicit Vertex Cover (Section 7.3).
As an independent application of our arboricity sparsifier, we design an algorithm for implicit node-DP vertex cover. Similar to matchings, it is impossible to explicitly release a small vertex subset that is guaranteed to cover all edges without blatantly revealing the (non-)existence of edges. Due to this, [GuptaLMRT10] designed an $\varepsilon$-edge DP implicit $O(\nicefrac{{1}}{{\varepsilon}})$-approximate vertex cover algorithm that releases an ordering of vertices where an edge is covered by its earlier endpoint in the ordering. Arboricity sparsification also (approximately) preserves vertex cover solutions. Thus, applying our arboricity sparsifier to reduce node-DP to edge-DP with a slight modification of the [GuptaLMRT10] algorithm yields the following result.
Theorem 2.7 (Node-DP Approximate Vertex Cover).
Let $\varepsilon\in(0,1)$ and $\alpha$ denote the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that outputs an implicit vertex cover (Definition 7.11) which is an $\widetilde{O}(\nicefrac{{\alpha}}{{\varepsilon^{2}}})$-approximation with probability $0.99$.
2.4 Utility Improvements in Bipartite Graphs over [hsu2014private] (Section 8)
Our techniques also lead to improved results for the setting of [hsu2014private]: bipartite graphs with node differential privacy in the billboard model. In this model, they view the nodes in the “left side” of the bipartition as goods and the “right side” nodes of the graph as bidders for these goods. Two such bipartite graphs are neighboring if one can be obtained from the other by adding/deleting any one bidder (and all its incident edges).
[hsu2014private] show guarantees of the following form. Suppose each item has supply $s$, meaning that each item can be matched with $s$ bidders (let us call these $s$-matchings); if $s=\Omega\left(\frac{\log^{3.5}n}{\varepsilon}\right)$, then they give an algorithm which (implicitly) outputs a $(1+\eta)$-maximum matching while guaranteeing differential privacy. [^5] They also consider the weighted version of the problem and further generalizations where each bidder has gross substitutes valuations. Our techniques also extend to these cases with improved results, but we limit our discussion here to matchings as that is the focus of this work. They also show that $s=\Omega(\nicefrac{{1}}{{\sqrt{\eta}}})$ is necessary to obtain any guarantees, leaving open the question of tightening these bounds. We substantially tighten the upper bound in Section 8, showing that $s=\Theta\left(\frac{\log{n}}{\varepsilon}\right)$ suffices:
Theorem 2.8.
If the supply is at least $s\geq\Omega\left(\frac{\log(n)}{\eta^{4}\varepsilon}\right)$, Algorithm 5 is $\varepsilon$-node DP and outputs a $(1+\eta)$-approximate maximum (one-sided) $s$-matching with high probability.
The starting point of our algorithm for node-private bipartite matchings is the same deferred acceptance type algorithm which [hsu2014private] base their private algorithm on. In their private implementation of the algorithm, they use continual counters [dwork2009complexity, CSS11] to keep track of the number of bidders each item is matched to. This is necessary for their implementation since their implicit solutions need to know the number of bidders an item is matched with at every iteration of the algorithm. We show that in a different private implementation of the algorithm, the counts of the number of bidders an item is matched to is only needed $O(1/\eta^{2})$ different times (where $\eta\in(0,1]$ is present in the $(1+\eta)$-approximation factor). Furthermore, a clever use of the Multidimensional AboveThreshold mechanism (Algorithm 1) to iteratively check if this count exceeds the threshold $s$ (where $s$ is the matching factor) suffices for the implementation, enabling us to substantially improve the error of the counts used.
2.5 Continual Release (Section 9)
We extend all of our results to the continual release model [DNPR10, CSS11] under two different types of streams: edge-order streams and adjacency-list streams, both commonly seen in non-private streaming literature (see e.g. [kallaugher2019complexity, mcgregor2016better, mcgregor2005finding]). In the graph continual release setting, edges are given as updates to the graph in a stream and the algorithm releases an output after each update in the stream. The continual release model requires the entire vector of outputs of the algorithm to be $\varepsilon$-differentially private. We consider edge-insertion streams where an update in the stream can either be $\bot$ (an empty update) or an insertion of an edge $\{u,v\}$. Edge-neighboring streams are two streams that differ in exactly one edge update. We also consider node-neighboring streams; node-neighboring streams are two streams that differ in all edges adjacent to any one node. We release a solution after every update over the course of $T=\operatorname{poly}(n)$ updates. [^6] Our algorithms extend to the case where $T=\omega(\operatorname{poly}(n))$ at the cost of factors of $\log(T)$. We focus on the case $T=\operatorname{poly}(n)$ as there are at most $O(n^{2})$ non-empty updates. The goal is to produce an accurate approximate solution for each update, with high probability.
The continual release setting is a more difficult setting than the static setting for several reasons. First, each piece of private data is used multiple times to produce multiple solutions, potentially leading to high error from composition. Second, depending on the algorithm, it may be possible to accumulate more errors as one releases more solutions (leading to compounding errors). Finally, for node-private algorithms, sparsification techniques need to be handled with more care since temporal edge updates can lead to sparsified solutions becoming unstable (causing the neighboring streams to become farther in edge-distance instead of closer). We solve all of these challenges to implement our matching algorithms in the continual release model.
We first adapt our LEDP and node-DP implicit $O(\log(n)/\varepsilon)$-matching algorithms to edge-order streams. These results are given in Corollary 9.1 and Theorem 2.10. The main idea behind our continual release algorithm is to use the sparse vector technique (SVT) to determine when to release a new solution at each timestep $t\in[T]$. Specifically, at timestep $t$, we check the current exact maximum matching size in the induced subgraph $G_{t}$ consisting of all updates up to and including $t$. If this matching size is greater than a $(1+\eta)$ factor (for some fixed $\eta\in(0,1]$) of our previously released solution, then we release a new solution. To release our new solution, we use our LEDP algorithm as a blackbox and pass $G_{t}$ into the algorithm. Then, we release the implicit solution that is the output of our LEDP algorithm. Since we can only increase our solution size $O(\log_{1+\eta}(n))$ times, we only accumulate an additional $O(\log(n)/\eta)$ factor in the error due to composition.
For our node-DP continual release algorithm for edge-order streams, we perform the same strategy as our edge-DP algorithm above except for one main change. We implement a stable version of our generalized matching sparsifier in Section 7 for edge-order streams in the continual release model. Then, for each update, we determine whether to keep it as part of our matching sparsifier. Using the sparsified set of edges, we run our SVT procedure to determine when to release a new matching and use our LEDP algorithm as a blackbox.
Finally, we adapt our LEDP matching algorithm to adjacency-list order streams in the continual release model. In adjacency-list order streams, updates consist of both vertices and edges where each edge shows up twice in the stream. Immediately following a vertex update, all edges adjacent to the vertex are given in an arbitrary order in the stream following the vertex update. Edge-neighboring adjacency-list order streams differ in exactly one edge update. We implement our LEDP algorithm in a straightforward manner in adjacency-list order streams. In particular, a node performs our proposal procedure once it sees all of its adjacent edges. Then, the node writes onto the blackboard the results of this proposal procedure. We maintain the same utility guarantees except for an additional additive error of $1$. Since a node must wait until it sees all of its adjacent edges before performing the proposal procedure, this error results from the most recent node update (where the node has yet to observe all of its adjacent edges).
Our results are given formally below. We give bicriteria approximate solutions where an $(\psi,\phi)$-approximation is one with a $\psi$-multiplicative error and $\phi$-additive error. Our approximations hold, with high probability, for every update.
Theorem 2.9.
For $\varrho,\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+O\left(\frac{\log^{2}(n)}{\varrho\eta^{2}\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching is a $\left(2+\varrho,\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximation with respect to the maximum $b^{\prime}$-matching size.
Similar to our static edge-DP algorithms, there are two regimes of particular interest.
  • If $b=\Omega\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)$, there is an $\varepsilon$-edge DP algorithm that outputs a sequence of implicit $b$-matchings with the size of a $\left(2+\varrho,\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximate 1-matching (Corollary 9.1).
  • If $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\varepsilon}\right)$, there is an $\varepsilon$-edge DP algorithm that outputs a sequence of implicit $b$-matchings that are $\left(2+\varrho,\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate $b$-matchings (Theorem 9.2). Note that we circumvent the need for a bicriteria approximation.
Our continual release algorithm can also be implemented under node-DP by implementing the appropriate arboricity sparsifier in the online setting.
Theorem 2.10.
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, $\alpha$ be the arboricity of the input graph, and
$b=\Omega\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{2}\varepsilon^{2}}+\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\,.$
There is an $\varepsilon$-node DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching has the size of a $\left(2+\eta,O\left(\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum $b^{\prime}$-matching.
As a special case, we have the following corollary:
  • If $b=\Omega\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{2}\varepsilon^{2}}\right)$, there is a $\varepsilon$-node DP algorithm that outputs a sequence of $b$-matchings that are the size of a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate 1-matching (Corollary 9.3).
The second type of stream we consider is the arbitrary adjacency-list order model [kallaugher2019complexity, goldreich2009proximity, czumaj2016relating, mcgregor2016better] in which nodes arrive in arbitrary order and once a node $v$ arrives as an update in the stream all edges adjacent to $v$ arrive in an arbitrary order as edge updates immediately after the arrival of $v$. In this model, we give the following results within Section 9.3.
Theorem 2.11.
For $\varepsilon\in(0,1)$ and $b=O(\log(n)/\varepsilon)$, Algorithm 8 is an $\varepsilon$-edge DP algorithm in the arbitrary adjacency-list continual release model that outputs a sequence of implicit $b$-matchings such that, with high probability, each $b$-matching has the same guarantee as Theorem 2.2 with the addition of an additive error of $1$ per update.
Our quantitative results are summarized in Table 1. For simplicity, we state all approximation guarantees and lower bounds with respect to the maximum matchings, whereas our algorithms can also produce approximations with respect to the optimal $b$-matching.
Model $b$ Approximation Bound Type Reference
$(\varepsilon,\delta)$-edge DP $O(1)$ $\Omega(n/b)$ Lower Bound Theorem 2.1
(explicit solutions)
$\varepsilon$-edge DP $\Omega(\log(n)/\varepsilon)$ $(2+\eta)$ Lower Bound Theorem 2.4
$\varepsilon$-LEDP ($n$ rounds) $O(\log(n)/\varepsilon)$ $2,(2+\eta)$ Upper Bound Corollary 5.1
Theorem 2.2
Theorem 5.2
$\varepsilon$-LEDP ($O(\log n)$ rounds) $O(\log^{2}(n)/\varepsilon)$ 22 Upper Bound Theorem 2.3
$\varepsilon$-node DP $O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+{\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}}\right)$ $\left(2+\eta\right)$ Upper Bound Corollary 7.1
$\varepsilon$-node DP (bipartite) $s=\Omega\left(\frac{\log(n)}{{\eta^{4}}\varepsilon}\right)$ $(1+\eta)$, $s$-matching Upper Bound Theorem 2.8
Edge-Order $\varepsilon$-Edge DP $O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Corollary 9.1
Continual Release
Edge-Order $\varepsilon$-Node DP $O\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+{\frac{\log^{3}(n)}{\eta^{3}\varepsilon^{2}}}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Theorem 2.10
Continual Release
Adj-List Edge $O(\log(n)/\varepsilon)$ $(2,1)$ Upper Bound Theorem 2.11
Continual Release
Table 1: Summary of results on differentially private matchings. All approximations are given in terms of the optimum maximum matching in the input graph. For our node-DP results, $\alpha$ denotes the arboricity of the graph. All results except the first row are with respect to implicit solutions while the first one is a lower bound for explicit solutions.
3 Preliminaries
3.1 Differential Privacy
We begin with the basic definitions of differential privacy for graphs. Two graphs $G$ and $G^{\prime}$ are said to be edge-neighboring if they differ in one edge, i.e., there are vertices $u,v\in V(G)=V(G^{\prime})$ such that $\{uv\}=(E(G)\setminus E(G^{\prime}))\cup(E(G^{\prime})\setminus E(G))$. They are said to be node-neighboring if they differ in all edges incident to a particular node. Differentially private algorithms on edge-neighboring and node-neighboring graphs satisfy edge-privacy [NRS07] and node-privacy [BBS13, CZ13, kasiviswanathan2013analyzing], respectively.
Definition 3.1 (Graph Differential Privacy [NRS07]).
Algorithm $\mathcal{A}(G)$ that takes as input a graph $G$ and outputs an object in $\text{Range}(\mathcal{A})$ is $(\varepsilon,\delta)$-edge (-node) differentially private ($(\varepsilon,\delta)$-edge (-node) DP) if for all $S\subseteq\text{Range}(\mathcal{A})$ and all edge- (node-)neighboring graphs $G$ and $G^{\prime}$,
$\Pr[\mathcal{A}(G)\in S]\leq e^{\varepsilon}\cdot\Pr[\mathcal{A}(G^{\prime})\in S]+\delta$
If $\delta=0$ in the above, then we drop it and simply refer to $\varepsilon$-edge (-node) differential privacy.
In this paper, we differentiate between the representation model which we will use to release solutions and the privacy model. The privacy model we use for our static algorithms is the local edge differential privacy (LEDP) model defined in [DLRSSY22]. We give the transcript-based definition, defined on $\varepsilon$-local randomizers, verbatim, below.
Definition 3.2 (Local Randomizer (LR) [DLRSSY22, kasiviswanathan2011can]).
An $\varepsilon$-local randomizer $R:\mathbf{a}\rightarrow\mathcal{Y}$ for node $v$ is an $\varepsilon$-edge DP algorithm that takes as input the set of its neighbors, $N(v)$, represented by an adjacency list $\mathbf{a}=(b_{1},\dots,b_{|N(v)|})$. In other words,
$\frac{1}{e^{\varepsilon}}\leq\frac{\mathsf{Pr}\left[R(\mathbf{a}^{\prime})\in Y\right]}{\mathsf{Pr}\left[R(\mathbf{a})\in Y\right]}\leq e^{\varepsilon}$
for all $\mathbf{a}$ and $\mathbf{a}^{\prime}$ where the symmetric difference is $1$ and all sets of outputs $Y\subseteq\mathcal{Y}$. The probability is taken over the random coins of $R$ (but not over the choice of the input).
Definition 3.3 (Local Edge Differential Privacy (LEDP) [DLRSSY22, JMN019]).
A transcript $\pi$ is a vector consisting of 5-tuples $(S^{t}_{U},S^{t}_{R},S^{t}_{\varepsilon},S^{t}_{\delta},S^{t}_{Y})$ – encoding the set of parties chosen, set of local randomizers assigned, set of randomizer privacy parameters, and set of randomized outputs produced – for each round $t$. Let $S_{\pi}$ be the collection of all transcripts and $S_{R}$ be the collection of all randomizers. Let $\mathtt{EOC}$ denote a special character indicating the end of computation. A protocol is an algorithm $\mathcal{A}:S_{\pi}\to(2^{[n]}\times 2^{S_{R}}\times 2^{\mathbb{R}^{\geq 0}}\times 2^{\mathbb{R}^{\geq 0}})\;\cup\{\mathtt{EOC}\}$ mapping transcripts to sets of parties, randomizers, and randomizer privacy parameters. The length of the transcript, as indexed by $t$, is its round complexity.
Given $\varepsilon\geq 0$, a randomized protocol $\mathcal{A}$ on (distributed) graph $G$ is $\varepsilon$-locally edge differentially private ($\varepsilon$-LEDP) if the algorithm that outputs the entire transcript generated by $\mathcal{A}$ is $\varepsilon$-edge differentially private on graph $G.$ If $t=1$, that is, if there is only one round, then $\mathcal{A}$ is called non-interactive. Otherwise, $\mathcal{A}$ is called interactive.
Now, we decouple the billboard model definition given in [hsu2014private] from the privacy model. Hence, the billboard model only acts as a solution release model. The billboard model takes as input a private graph and produces an implicit solution consisting of a public billboard. Then, each node can determine its own part of the solution using the public billboard and the private information about its adjacent neighbors. In this paper, our algorithms for producing the public billboard will be $\varepsilon$-LEDP.
Definition 3.4 (Billboard Model [hsu2014private]).
Given an input graph $G=(V,E)$, algorithms in the billboard model produces a public billboard $B$. Then, each node $v\in V$ in the graph deduces the portion of the explicit solution that $v$ participates in as a function of $B$ and its private neighborhood $N(v)$.
In particular, one can easily show that if every node processes the information contained in the public billboard using a deterministic algorithm, then explicitly revealing the solution contained at every node will be $\varepsilon$-edge DP (with respect to edge-neighboring graphs) except for the two nodes adjacent to the edge that differs between neighboring graphs. Morally speaking, such a set of explicit solutions does not leak any additional private information since the nodes that are endpoints to the edge that differs already know that this edge exists. Hence, we decouple the privacy definition from the solution release model and say an algorithm is $\varepsilon$-LEDP in the billboard model if the public billboard is produced via a $\varepsilon$-LEDP algorithm and the explicit solutions obtained by every node is via a deterministic algorithm at each node.
Lemma 3.5.
Given a public billboard produced from the billboard model (Definition 3.4), if each node produces their individual explicit solution using a (predetermined) deterministic algorithm then the produced explicit solutions are $\varepsilon$-edge DP except for the solutions produced by the endpoints of the edge that differs between edge-neighboring graphs.
Proof.
Given identical adjacency lists and an identical billboard, a deterministic algorithm will output identical solutions. Hence, by the definition of $\varepsilon$-edge DP, every node will produce identical explicit solutions except for the endpoints of the edge that differs between edge-neighboring graphs. ∎
We now define the implicit solutions that our algorithms will post to the billboard. Note that by the definition of $\varepsilon$-LEDP, the entire transcript is public and so, without loss of generality, is posted to the billboard. But the particular information that each node will use to produce its explicit solution is the following.
Definition 3.6 (Implicit Solution).
Given a graph $G=(V,E)$, an implicit solution is a collection $\mathcal{S}=\{S_{v}\}_{v\in V}$ where each $S_{v}\subseteq V$. An implicit solution $\mathcal{S}$ defines a graph $H(\mathcal{S})=(V,E^{\prime})$ where $E^{\prime}=\{\{u,v\}:u\in S_{v}\lor v\in S_{u}\}$. The degree of an implicit solution $\mathcal{S}$ is the maximum degree in the implicit graph $(V,E\cap E(H(\mathcal{S})))$, i.e., the maximum degree in the graph which is the intersection of $G$ and $H(\mathcal{S})$.
In other words, for each vertex $v$ we have a subset of nodes $S_{v}$. Think of $S_{v}$ as “potential matches” for $v$. Then, $H(\mathcal{S})$ is the graph obtained by adding an edge if either endpoint contains the other as a potential match. We obtain a third graph by intersecting the implicit graph with the true graph, and the maximum degree in this graph is the degree of the solution. From a billboard/LEDP perspective, this means that any node $v$ can locally decode an explicit solution from the implicit solution since it can perform the intersection of $H(\mathcal{S})$ with its own neighborhood.
Now, we define an additional data input model which we use in our proof of our lower bound.
Definition 3.7 ($0/1$-Weight Model [hsu2014private]).
Given a complete graph $G=(V,E)$, each edge in the graph is given a binary weight of 0 or $1$. Edge-neighbors are two graphs $G$ and $G^{\prime}$ where the weight of exactly one edge differs. Node-neighbors are two graphs where the weights of all edges adjacent to exactly one node differ.
Remark 3.8.
For a given implicit solution $\mathcal{S}$ (Definition 3.6), we remark that the graph $H(\mathcal{S})$ is a valid explicit solution in the 0/1-weight model (Definition 3.7). We define $\mathcal{S}$ as a collection of local neighborhoods to highlight the fact that our algorithms are local.
Suppose that $\mathcal{S}$ is the output of an edge-DP approximate maximum matching algorithm, then $H(\mathcal{S})$ would be a solution that attains the same approximation ratio in the 0/1-weight model. On one hand, our implicit matching algorithms (Theorem 2.2) ensure that the implicit graph $H(\mathcal{S})\cap G$ has $O(\log n)$ maximum degree. On the other hand, our explicit solution lower bound (Theorem 2.1) ensures that $H(\mathcal{S})$ must have a large $\Omega(n)$ maximum degree. This demonstrates that implicit solutions are necessary to obtain meaningful algorithmic solutions.
3.1.1 Continual Release
In this section, we define the continuous release model [DNPR10, CSS11]. We first define the concepts of edge-order and adjacency-list order streams and then edge-neighboring and node-neighboring streams.
Definition 3.9 (Edge-Order Graph Stream [wagaman2024time]).
In the edge-order continual release model, a graph stream $S\in\mathcal{S}^{T}$ of length $T$ is a $T$-element vector where the $t$-th element is an edge update $u_{t}=\{v,w,\mathtt{insert}\}$ (an edge insertion of edge $\{v,w\}$), or $\bot$ (an empty operation).
Definition 3.10 (Adjacency-List Order Graph Stream (adapted from [mcgregor2016better])).
In the adjacency-list order continual release model, a graph stream $S\in\mathcal{S}^{T}$ of length $T$ is a $T$-element vector where the $t$-th element is a node update $u_{t}=\{v\}$, an edge update $u_{t}=\{v,w,\mathtt{insert}\}$ (an edge insertion of edge $\{v,w\}$), or $\bot$ (an empty operation). Each node update is followed (in an arbitrary order) by all adjacent edges.
We use $G_{t}$ and $E_{t}$ to denote the graph induced by the set of updates in the stream $S$ up to and including update $t$. Now, we define neighboring streams as follows. Intuitively, two graph streams are edge neighbors if one can be obtained from the other by removing one edge update (replacing the edge update by an empty update in a single timestep); and they are node-neighbors if one can be obtained from the other via removing all edge updates incident to a particular vertex.
Definition 3.11 (Edge Neighboring Streams).
Two streams of updates, $S=[u_{1},\dots,u_{T}]$ and $S^{\prime}=[u^{\prime}_{1},\dots,u^{\prime}_{T}]$, are edge-neighboring if there exists exactly one timestamp $t^{*}\in[T]$ (containing an edge update in $S$ or $S^{\prime}$) where $u_{t^{*}}\neq u^{\prime}_{t^{*}}$ and for all $t\neq t^{*}\in[T]$, it holds that $u_{t}=u^{\prime}_{t}$. Streams may contain any number of empty updates, i.e. $u_{t}=\bot$. Without loss of generality, we assume for the updates $u_{t^{*}}$ and $u^{\prime}_{t^{*}}$, it holds that $u^{\prime}_{t^{*}}=\bot$ and $u_{t^{*}}=e_{t^{*}}$ is an edge insertion.
Definition 3.12 (Node Neighboring Streams).
Two streams of updates, $S=[u_{1},\dots,u_{T}]$ and $S^{\prime}=[u^{\prime}_{1},\dots,u^{\prime}_{T}]$, are node-neighboring if there exists exactly one vertex $v^{*}\in V$ where for all $t\in[T]$, $u_{t}\neq u^{\prime}_{t}$ only if $u_{t}$ or $u_{t}^{\prime}$ is an edge insertion of an edge adjacent to $v^{*}$. Streams may contain any number of empty updates, i.e. $u_{t}=\bot$. Without loss of generality, we assume for the updates $u_{t}\neq u^{\prime}_{t}$, it holds that $u^{\prime}_{t}=\bot$ and $u_{t}=e_{t}$ is an edge insertion of an edge adjacent to $v^{*}$.
We now define edge-privacy and node-privacy for edge-neighboring and node-neighboring streams, respectively.
Definition 3.13 (Edge Differential Privacy for Edge-Neighboring Streams).
Let $\varepsilon\in(0,1)$. An algorithm $\mathcal{A}(S):\mathcal{S}^{T}\rightarrow\mathcal{Y}^{T}$ that takes as input a graph stream $S\in\mathcal{S}^{T}$ is said to be $\varepsilon$-edge differentially private (DP) if for any pair of edge-neighboring graph streams $S,S^{\prime}$ (Definition 3.11) and for every $T$-sized vector of outcomes $Y\subseteq\text{Range}(\mathcal{A})$,
$\mathsf{Pr}\left[\mathcal{A}(S)\in Y\right]\leq e^{\varepsilon}\cdot\mathsf{Pr}\left[\mathcal{A}(S^{\prime})\in Y\right].$
Definition 3.14 (Node Differential Privacy for Node-Neighboring Streams).
Let $\varepsilon\in(0,1)$. An algorithm $\mathcal{A}(S):\mathcal{S}^{T}\rightarrow\mathcal{Y}^{T}$ that takes as input a graph stream $S\in\mathcal{S}^{T}$ is said to be $\varepsilon$-node differentially private (DP) if for any pair of node-neighboring graph streams $S,S^{\prime}$ (Definition 3.12) and for every $T$-sized vector of outcomes $Y\subseteq\text{Range}(\mathcal{A})$,
$\mathsf{Pr}\left[\mathcal{A}(S)\in Y\right]\leq e^{\varepsilon}\cdot\mathsf{Pr}\left[\mathcal{A}(S^{\prime})\in Y\right].$
3.2 Differential Privacy Tools
In this section, we state the privacy tools we use in our paper. The adaptive Laplace mechanism is a formalization of the Laplace mechanism for adaptive inputs that we employ in this work (and is used implicitly in previous works). For completeness, we include a proof of Lemma 3.15 in Section A.2.
Lemma 3.15 (Adaptive Laplace Mechanism; used implicitly in [wagaman2024time]).
Let $f_{1},\ldots,f_{k}$ with $f_{i}:\mathcal{G}\to\mathbb{R}$ be a sequence of adaptively chosen queries and let $f$ denote the vector $(f_{1},\ldots,f_{k})$. Suppose that the adaptive adversary gives the guarantee that the vector $f$ has $\ell_{1}$-sensitivity $\Delta$, regardless of the output of the mechanism. Then the Adaptive Laplace Mechanism $\mathcal{M}$ with vector-valued output $\tilde{f}(G)$ where $\tilde{f}_{i}(G)\coloneqq f_{i}(G)+\text{Lap}(\Delta/\varepsilon)$ for each query $f_{i}$ is $\varepsilon$-differentially private.
The Multidimensional AboveThreshold mechanism (Algorithm 1) is a generalization of the AboveThreshold mechanism [lyu2017understanding] which is traditionally used to privately answer sparse threshold queries.
Lemma 3.16 (Multidimensional AboveThreshold Mechanism [DLL23]).
Algorithm 1 is $\varepsilon$-LEDP.
1Input: Graph $G$, adaptive queries $\{\vec{f}_{1},\dots,\vec{f}_{n}\}$, threshold vector $\vec{T}$, privacy parameter $\varepsilon$, $\ell_{1}$-sensitivity $\Delta$.
2 Output: A sequence of responses $\{\vec{a}_{1},\dots,\vec{a}_{n}\}$ where $a_{i,j}$ indicates if $f_{i,j}(G)\geq\vec{T}_{j}$
1:for $j=1,\ldots,d$ do
2:  $\hat{T}_{j}\leftarrow\vec{T}_{j}+\text{Lap}(2\Delta/\varepsilon)$
3:end for
4:
5:for each query $\vec{f}_{i}\in\{\vec{f}_{1},\dots,\vec{f}_{n}\}$ do
6:  for $j=1,\ldots,d$ do
7:   Let $\nu_{i,j}\leftarrow\text{Lap}(4\Delta/\varepsilon)$
8:   if $f_{i,j}(G)+\nu_{i,j}\geq\hat{T}_{j}$ then
9:    Output $a_{i,j}=$ “above”
10:    Stop answering queries for coordinate $j$
11:   else
12:    Output $a_{i,j}=$ “below”
13:   end if
14:  end for
15:end for
Algorithm 1 Multidimensional AboveThreshold (MAT) [DLL23]
In the privacy analysis of our algorithms, we will often argue that each subroutine is DP and hence the whole algorithm is DP. However, since the access to private data is interactive, we will need some form of concurrent composition theorem, such as the one stated below.
Lemma 3.17 (Concurrent Composition Theorem [vadhan2021concurrent]).
If $k$ interactive mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are each $(\varepsilon,\delta)$-differentially private, then their concurrent composition is $\left(k\cdot\varepsilon,\frac{e^{k\varepsilon}-1}{e^{\varepsilon}-1}\cdot\delta\right)$-differentially private.
Finally, we use the following privacy amplification theorem for subsampling.
Lemma 3.18 (Privacy Amplification via Subsampling [li2012sampling]).
Let $\varepsilon\in(0,1]$. If elements from the private dataset are sampled with probability $p$ and we are given a $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ on the original dataset, then running $\mathcal{A}$ on the subsampled dataset gives a $(2p\varepsilon,p\cdot\delta)$-DP algorithm for $\varepsilon\in(0,1)$.
4 Lower Bounds for Explicit Solutions via Symmetrization
In this section, we introduce our novel symmetry argument technique for proving lower bounds for graph differential privacy. As previously mentioned, this is inspired by arguments from distributed computation. Connections between differential privacy and other areas have previously been studied [bun2023stability, dwork2015adaptive, beimel2022adaptive, feldman2020memorization]. We hope that our work can lead to more connections between DP and distributed algorithms, a direction so far underexplored.
Our goal is to prove Theorem 2.1 which shows, intuitively, that any explicit solution (when provided $0/1$ weights for the graph) for the maximum matching problem must have large error. More specifically, Theorem 2.1 as restated below states: in the basic 0/1 weight $(\varepsilon,\delta)$-edge DP model (Definition 3.7), the required error is essentially linear, even if we allow for the algorithm to output a $b$-matching for very large $b$.
See 2.1
We will prove a more general lower bound for synthetic graph generation (Definition 4.1), which can be interpreted as a relaxation of the explicit matchings. As we will see, this is a stronger result than Theorem 2.1 since the synthetic graph problem has less stringent constraints than matchings. In Section 4.1, we formally define this problem and prove its hardness. Then, we prove Theorem 2.1 as a straightforward corollary in Section 4.2.
4.1 Lower Bound on DP Synthetic Graph Generation
We now formalize our symmetry argument and use it to prove the following lower bound (Theorem 4.3), which is a relaxation of our explicit matching problem. We will use Theorem 4.3 to prove Theorem 2.1 in Section 4.2. Note that our symmetry argument is also used to prove the lower bound presented in Theorem 2.5.
Definition 4.1 (Synthetic Graph Generation).
Let $G=(V,E)$ be an input graph and $b\in[0,n]^{V}$ be vertex degree bounds. The corresponding synthetic graph problem is given by the following stochastic program. This program states: given (potentially randomized) variables $x_{uv}$, one for each pair of vertices $u\neq v\in V$, we want to maximize the expected sum of all variables corresponding to edges in $E$, subject to the constraints that the expected sum of all variables associated with any vertex $v$ is at most $b_{v}$.
$\max$ $\sum_{uv\in E(G)}\mathbb{E}[x_{uv}]$
$\sum_{u\neq v}\mathbb{E}[x_{uv}]\leq b_{v}$ $\forall v\in V$
$\mathbf{x}$ $\in[0,1]^{\binom{V}{2}}$
Synthetic graph generation (Definition 4.1) is interesting due to the fact that many LEDP mechanisms, such as randomized response, produce dense graph outputs. Such dense graph outputs are undesirable both due to the large error as well as slow processing times. Our lower bound below essentially states that such dense outputs are unavoidable if one wants to preserve differential privacy. To the best of our knowledge, while sparsity lower bounds for specific algorithm such as randomized response have been studied [epasto2024smooth] under pure $\varepsilon$-DP, Theorem 4.3 is the first algorithm-agnostic lower bound for synthetic graph generation under approximate $(\varepsilon,\delta)$-DP.
Remark 4.2.
In the case that $\mathbf{x}\in\{0,1\}^{\binom{V}{2}}$ is required to be integral, we can interpret Definition 4.1 as outputting a graph $H$ maximizing $\mathbb{E}[|E(H)\cap E(G)|]$ subject to degree constraints in expectation $\mathbb{E}[\deg_{H}(v)]\leq b_{v}$. This problem essentially asks for a synthetic graph with maximum overlap with a given input graph $G$ and has degree constraints on all vertices.
This can be viewed as a relaxation of the 0/1-weight matching problem in the case $E(G)$ is a matching (e.g., the maximum degree of $G$ is 1), in the sense that the output of any (randomized) 0/1-weight matching algorithm on such a graph $G$ is a feasible solution to the synthetic graph problem with input $G$.
Having defined the synthetic graph problem, we are ready to state the lower bound Theorem 4.3 below. In the below theorem, a solution $S\in[0,1]^{{V\choose 2}}$ is a $(\gamma,\beta)$-approximation of the solution to the Synthetic Graph Problem (Definition 4.1) if $\sum_{uv\in E}\mathbb{E}[S_{uv}]\geq\frac{|E|}{\gamma}-\beta$.
Theorem 4.3 (Lower Bound on Private Synthetic Graph Approximation).
Let $\mathcal{A}$ be an algorithm which satisfies $(\varepsilon,\delta)$-edge DP and always outputs a solution $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$ to the synthetic graph problem (Definition 4.1) parameterized by graph $G$ and degree bounds $b$ with average degree bound $\bar{b}\coloneqq\frac{1}{n}\sum_{v\in V}b_{v}$. If $\mathcal{A}$ outputs a $(\gamma,\beta)$-approximation (even in expectation), then $\gamma,\beta$ must satisfy
$e^{4\varepsilon}\bar{b}+2\delta n\geq\frac{n}{2\gamma}-\beta\,.$
4.1.1 Proof of Theorem 4.3
We now present our proof of Theorem 4.3. At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument, we will instead use a novel symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that $\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$ for each non-edge $uv\notin E$ is identical. This makes it easy to argue that high-utility algorithms cannot satisfy differential privacy.
We begin with a description of our inputs and some more notation. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-differentially private algorithm which always returns a fractional solution $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$ as defined above (Definition 4.1). For $G\in\mathcal{M}$ and $uv\in\binom{V}{2}$, let $\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$ denote the expected fractional value that $\mathcal{A}(G)$ assigns to $uv$.
With this notation in hand, we can now define and prove the main symmetry property.
Lemma 4.4.
If $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-DP and has expected utility at least $\nu_{\mathcal{A}}$ over $\mathcal{M}$, then there is an algorithm $\mathcal{A}^{\prime}$ which also satisfies $(\varepsilon,\delta)$-DP and has $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$. Moreover $\mathcal{A}^{\prime}$ satisfies the following symmetry property: $\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\mathbb{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ for all $uv,ab\not\in E(G)$ when executed on input $G\in\mathcal{M}$.
Proof.
Consider the following algorithm $\mathcal{A}^{\prime}$. We first choose a permutation $\pi$ of $V$ uniformly at random from the space $S_{V}$ of all permutations of $V$. Let $\pi(G)$ denote the graph isomorphic to $G$ obtained by applying this permutation, i.e., by creating an edge $\{\pi(u),\pi(v)\}$ for each edge $\{u,v\}\in E(G)$. We run $\mathcal{A}$ on $\pi(G)$, and let $X_{\pi(G)}^{\mathcal{A}}$ denote the resulting fractional solution for $\pi(G)$. We then “undo” $\pi$ to get the fractional solution $X_{G,uv}^{\mathcal{A}^{\prime}}=X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}$. We return $X_{G}^{\mathcal{A}^{\prime}}$.
We first claim that $\mathcal{A}^{\prime}$ is $(\varepsilon,\delta)$-DP. This can be seen since $\mathcal{A}^{\prime}$ simply runs $\mathcal{A}$ on a uniformly random permuted version of $G$, and then does some post-processing. More formally, for any neighboring graphs $G$ and $G^{\prime}$ and for any $\pi\in S_{V}$, the graphs $\pi(G)$ and $\pi(G^{\prime})$ are also neighboring graphs. Hence the differential privacy guarantee of $\mathcal{A}$ implies that running $\mathcal{A}$ on $\pi(G)$ and $\pi(G^{\prime})$ is $(\varepsilon,\delta)$-DP, and then “undoing” $\pi$ is simply post-processing. So $\mathcal{A}^{\prime}$ has the same privacy guarantees as $\mathcal{A}$.
Now we claim that $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$. Since $\pi$ was chosen uniformly at random, we know that $\pi(G)$ is distributed uniformly among $\mathcal{M}$. Hence for any $G\in\mathcal{M}$ we have that
$\mathbb{E}[\nu_{\mathcal{A}^{\prime}}(G)]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\mathbb{E}[\nu_{\mathcal{A}}(\pi(G))]\geq\frac{1}{n!}\sum_{\pi\in S_{V}}\nu_{\mathcal{A}}=\nu_{\mathcal{A}}\,.$
Since this is true for all $G\in\mathcal{M}$, we have that $\nu_{\mathcal{A}^{\prime}}\coloneqq\min_{G\in\mathcal{M}}\mathbb{E}[\nu_{\mathcal{A}^{\prime}}(G)]\geq\nu_{\mathcal{A}}$, as claimed.
We now prove the key symmetry property. Fix $G,G^{\prime}\in\mathcal{M}$, and let $\Pi(G,G^{\prime})=\{\pi\in S_{V}:\pi(G)=G^{\prime}\}$. Note that if we draw a permutation $\pi$ uniformly at random from $\Pi(G,G^{\prime})$, then for any $\{u,v\}\not\in E(G)$, the pair $\{\pi(u),\pi(v)\}$ is uniformly distributed among the set $\{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})\}$. Thus we have that
$\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\mathbb{E}[X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\pi\in\Pi(G,G^{\prime})}\mathbb{E}[X_{G^{\prime},\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{ab\not\in E(G^{\prime})}\mathbb{E}[X_{G^{\prime},ab}^{\mathcal{A}}].$
In particular, $\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ is actually independent of $u,v$, and thus for all $uv,ab\not\in E(G)$, we have that $\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\mathbb{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ as claimed. ∎
Lemma 4.4 implies that if we can prove an upper bound on the expected utility for DP algorithms that obey the symmetry property, then the same upper bound applies to all DP algorithms. We will now prove such a bound.
Lemma 4.5.
Let $\mathcal{A}$ be an algorithm which is $(\varepsilon,\delta)$-DP, always outputs a synthetic graph solution (Definition 4.1), and satisfies the symmetry property of Lemma 4.4. Then, for $\bar{b}\coloneqq\frac{1}{n}\sum_{v\in V}b_{v}$ and $n\geq 4$,
$\nu_{\mathcal{A}}\leq e^{4\varepsilon}\bar{b}+\delta n\,.$
Proof.
Fix some graph $G\in\mathcal{M}$, and let $uv,ab\in E(G)$. Let $G^{\prime}$ be the graph obtained by removing $uv$ and $ab$, and adding $ub$ and $av$. Note that $G$ and $G^{\prime}$ are distance $4$ away from each other.
For all $y\neq u,b$, note that $uy\notin E(G^{\prime})$. Thus it must be the case that $\mathbb{E}[X_{G^{\prime},uy}^{\mathcal{A}}]\leq b_{u}/(n-2)$, since otherwise the symmetry property of Lemma 4.4 and linearity of expectations implies that the expected (fractional) degree of $u$ in the output of $\mathcal{A}(G)$ is at least $\sum_{y\neq u,b}\mathbb{E}[X_{G^{\prime},uy}^{\mathcal{A}}]>(n-2)\frac{b_{u}}{n-2}=b_{u}$, contradicting our assumption that $\mathcal{A}$ always outputs a feasible solution. Since $G$ and $G^{\prime}$ are at distance $2$ from each other, the fact that $\mathbb{E}[X_{G^{\prime},uv}^{\mathcal{A}}]\leq b_{u}/(n-2)$ implies that
$\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$ $=\int_{0}^{1}\Pr[X_{G,uv}^{\mathcal{A}}\geq x]\mathop{}\!\mathrm{d}x$
$\leq\int_{0}^{1}\left(e^{2\varepsilon}\cdot\Pr[X_{G^{\prime},uv}^{\mathcal{A}}\geq x]+4\delta\right)\mathop{}\!\mathrm{d}x$
$=e^{4\varepsilon}\cdot\mathbb{E}[X_{G^{\prime},uv}^{\mathcal{A}}]+4\delta$
$\leq e^{4\varepsilon}\frac{b_{u}+b_{v}}{2(n-2)}+4\delta\,.$
Here the last inequality uses the fact that if $x\leq y,z$, then $x\leq\frac{1}{2}(y+z)$.
Since this argument applies to all $uv\in E(G)$, we have that
$\mathbb{E}[\nu_{\mathcal{A}}(G)]$ $=\sum_{uv\in E(G)}\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$
$\leq\sum_{uv\in E(G)}\left(e^{4\varepsilon}\frac{b_{u}+b_{v}}{2(n-2)}+4\delta\right)$
$=e^{4\varepsilon}\frac{\bar{b}n}{2(n-2)}+2\delta n.$
$\leq e^{4\varepsilon}\bar{b}+2\delta n.$ ($n\geq 4$)
Since $\nu_{\mathcal{A}}=\min_{G\in\mathcal{M}}\mathbb{E}[\nu_{\mathcal{A}}(G)]$, this implies that $\nu_{\mathcal{A}}\leq e^{4\varepsilon}\bar{b}+2\delta n$, as claimed. ∎
We can now use these two lemmas to prove Theorem 4.3.
Proof of Theorem 4.3.
The combination of Lemma 4.4 and Lemma 4.5 imply that any $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ which always outputs a fractional stochastic packing solution $X_{G}^{\mathcal{A}}$ satisfies
$\nu_{\mathcal{A}}\coloneqq\sum_{uv\in E(G)}\mathbb{E}[X_{G,uv}^{\mathcal{A}}]\leq e^{4\varepsilon}\bar{b}+2\delta n\,.$
We also know that $|E(G)|=n/2$ for all $G\in\mathcal{M}$. Thus by the definition of an $(\gamma,\beta)$-approximation, it must be the case that $e^{4\varepsilon}\bar{b}+2\delta n\geq\frac{n}{2\gamma}-\beta$. ∎
4.2 Lower Bound on Explicit Matching Solutions (Theorem 2.1)
We first explain how to prove Theorem 2.1 by applying Theorem 4.3. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Remark that Theorem 4.3 holds even if we restrict the input class to $\mathcal{M}$. As stated in Remark 4.2, over $\mathcal{M}$, any $(\gamma,\beta)$-approximation algorithm for explicit matchings (Definition 3.7) is a $(\gamma,\beta)$-approximation algorithm for the synthetic graph problem (Definition 4.1) as the objective functions coincide and any explicit matching solution is a valid synthetic graph approximation with uniform degree bounds $b_{v}=b$. Thus, Theorem 2.1 follows directly from Theorem 4.3.
5 LEDP Implicit $b$-Matching via Public Vertex Subset Mechanism
In this section, we state our main algorithm for maximal matching (Algorithm 2) and prove its guarantees in Theorem 2.2, which is restated below for readability. This algorithm uses our novel Public Vertex Subset Mechanism (PVSM) which may be applicable to other problems such as vertex cover, dominating set, or independent set. We demonstrate the versatility of PVSM in two different algorithms in this paper. We implement Algorithm 2 for various different models in the later sections. See 2.2
As an immediate corollary, Theorem 2.2 implies the following.
Corollary 5.1 (LEDP $2$-Approximate Maximum Matching).
For $\varepsilon\in(0,1)$, Algorithm 2 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$, in the billboard model, for $b=O(\log(n)/\varepsilon)$ that has the size of a $2$-approximate maximum matching.
As mentioned, when $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 outputs a $b$-matching that approximates 1-matchings, giving a bicriteria approximation. Moreover, we can avoid a bicriteria approximation when $b$ is slightly larger.
Theorem 5.2 (LEDP $(2+\eta)$-Approximate Maximum $b$-Matching).
For $\varepsilon\in(0,1)$ and $\eta\in(0,1/2)$, Algorithm 2 is an $\varepsilon$-LEDP algorithm that, with high probability, outputs an (implicit) $b$-matching, in the billboard model, for $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$ that has the size of a $(2+\eta)$-approximate maximum $b$-matching.
Our LEDP algorithm draws on a simple procedure for maximal non-private $b$-matching, described as follows. We take an arbitrary ordering on the vertices, and process them one by one. When considering the $i^{th}$ vertex $v_{i}$, let $b^{\prime}$ be the number of additional vertices $v_{i}$ can match with. We wish to find some subset of vertices of size $b^{\prime}$ from later in the ordering with which to match $v_{i}$. Some of these vertices may have already matched with $b$ vertices from previous iterations. We choose an arbitrary subset of $b^{\prime}$ vertices which have not already reached their limit from later in the ordering. We say that vertices which have reached their limit satisfy the matching condition. If there are fewer than $b^{\prime}$ vertices satisfying the matching condition, we match all of them with $v_{i}$. Then we move on to the next vertex in the order. It is clear that this procedure yields a maximal $b$-matching. In the following algorithms, we analyze a suitable privatization of this algorithm and show that given sufficiently large $b$, we are guaranteed to produce an implicit maximal ($1$-)matching. [^7] We remark that all of our algorithms can be straightforwardly adjusted to guarantee maximal $b$-matchings.
To complement our algorithm, we show in Section 5.4 that the bound on $b$ is essentially tight for our specific notion of implicit solution.
Theorem 5.3.
Let $\gamma>0$ be arbitrary. Let $\mathcal{A}$ be an algorithm that satisfies $\varepsilon$-edge DP and outputs an implicit solution that has the size of a $\gamma$-approximate maximal matching with probability at least $1-\beta$, with $\beta\leq 1/(16e^{4\varepsilon})$. Then the degree of this implicit solution is at least $\Omega(\frac{1}{\varepsilon}\log(1/\beta))$ with probability at least $1/2$, regardless of the value of $\gamma$.
Note that this essentially matches Corollary 5.1 (up to constants) by setting $\beta=1/\mathrm{poly}(n)$.
High-Level Overview.
Our approximation guarantees hold with probability $1-\frac{1}{n^{c}}$ for any constant $c\geq 3$. The constants in Algorithm 2 are given in terms of $c$ to achieve our high probability guarantee. At a high level, our $\varepsilon$-LEDP algorithm modifies our non-private procedure described above in the following way. The key part of the procedure that uses private information is the selection of the $b^{\prime}$ neighbors for each vertex $v$ to satisfy the matching condition. We cannot select an arbitrary set of these neighbors directly since this arbitrary selection would reveal the existence of an edge between $v$ and each selected neighbor. Thus, we must select an appropriate number of neighbors randomly and output a public vertex subset that also includes non-neighbors.
To solve the above challenge, we introduce a novel Public Vertex Subset Mechanism. We call a vertex that is proposing a set of vertices the proposer. The mechanism works by having a proposer who proposes a public set of vertices to match to. The public set of vertices contains vertices which are neighbors of the proposer and also non-neighbors of the proposer. To select a public subset of vertices, we flip a set of coins with appropriate probability for each pair of vertices in the graph. A total of $O(\log(n))$ coins are flipped for each pair; a coin is flipped with probability $1/(1+\eta)^{r}$ for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. Thus, these coins determine progressively smaller subsets of vertices. An edge is in the public set indexed by $r$ if the coin lands heads. It is necessary to have vertex subsets with different sizes since proposers need to choose an appropriately sized subset of vertices that simultaneously ensures it satisfies the matching condition and not does exceed the $b$ constraint, with high probability. The coin flips are public because they do not reveal the existence of any edge. Moreover, the intersection between the public subset of vertices with each node’s private knowledge of its adjacency list allows each node to know which of its neighbors are matched to it. This Public Vertex Subset Mechanism may be useful for other problems.
A proposer $v$ releases the proposal subset by releasing the index $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ that corresponds with the coin flips determining the set that $v$ wants to match to. We use the sparse vector technique on the size of the subset to determine which $r$ to release. In fact, we use a multidimensional version of the sparse vector technique, called the Multidimensional AboveThreshold (MAT) technique (Algorithm 1, [DLL23]) which is designed for SVT queries performed by all nodes of a graph. Once a proposer releases a public proposal subset, each of $v$’s neighbors, that have not met their matching conditions, can determine whether they are matched to $v$ using the public coin flips. We call $r$ the subgraph index that is released by each proposer. Thus, each node knows the set of vertices they are matched to using the transcript consisting of publicly released subgraph indices and the public releases of when each node satisfies their matching condition. Each node determines whether it has satisfied its matching condition using MAT.
Organization.
We first give a detailed description of our algorithm, including its pseudocode, in Section 5.1. The accompanying privacy and utility analysis follows in Section 5.2. Section 5.3 proves the special cases of our general result (Theorem 2.2). Finally, we show the complementary lower bound on $b$ in Section 5.4.
5.1 Detailed Algorithm Description
We now describe our algorithm, Algorithm 2, in detail. The algorithm is provided with a private graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, a matching parameter $b\geq\frac{576c\ln(n)}{\eta\cdot\varepsilon}+1$ and a constant $c\geq 1$ which is used in the high probability guarantee a privacy parameter $\varepsilon>0$, matching benchmark parameter $b^{\prime}\geq 1$, constant $\eta\in(0,1)$, constant $c\geq 3$ which is used in the high probability guarantee, and matching degree parameter $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}$ . We first set $\varepsilon^{\prime}$ which we use in our algorithm (Algorithm 2). Then, we flip our public set of coins. We flip $O(\log(n))$ coins for each pair of vertices $u\neq v\in V\times V$ (Algorithm 2). Each of the coins flips for pair $\{u,v\}$ is flipped with probability $1/(1+\eta)^{r}$ for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. The result of the coin flip is released and stored in $coin(u,v,r)$ (Algorithm 2).
We now determine the matching condition for each node $u\in V$ (Algorithm 2). To do so, we draw a noise variable from $\text{Lap}(4/\varepsilon^{\prime})$ and add it to $b$. This noise is used as part of MAT (Algorithm 1, Multidimensional AboveThreshold) for a noisy threshold. Next, we subtract $36c\ln(n)/\varepsilon^{\prime}$ to ensure that the matching condition does not exceed $b$ even if a large positive noise was drawn from $\text{Lap}(4/\varepsilon^{\prime})$. In other words, Algorithm 2 ensures that the matching condition is satisfied when $v$ is actually matched to a sufficiently large number of vertices, with high probability. We also initialize the private data structure $M(u)$ stored by $u$ that contains the set of $u$’s neighbors that $u$ is matched to. Note that in practice, $M(u)$ is not centrally stored but can be decoded by $u$ using information posted to the public billboard.
We then initialize a set $A(u)\leftarrow\infty$ for every $u\in V$ (Algorithm 2) indicating that no vertices have yet to satisfy their matching condition. We now iterate through all of the vertices one by one in an arbitrary order (Algorithm 2). During the $i$-th iteration, if there is any node $u\in V$ which has not satisfied its matching condition (Algorithm 2), we use MAT (Algorithm 1) to check whether it now satisfies the matching condition (Algorithms 2 and 2). If node $u$ now satisfies the matching condition, then node $u$ releases $A(u)\leftarrow i$ to indicate that $u$ satisfied its matching condition at iteration $i$ (Algorithm 2).
If the current node in the iteration, $v$, has not satisfied its matching condition (Algorithm 2), then we determine its proposal set using the Public Vertex Subset Mechanism (Algorithms 2 to 2). For each subgraph index $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ (Algorithm 2), we determine the private subset of vertices $W_{r}(v)$ using the public coin flips. Specifically, $W_{r}(v)$ contains the set of neighbors $u$ of $v$ that are still active, come after $v$ in the ordering, and where $coin(u,v,r)=\textsc{Heads}$ (Algorithm 2). The public set of nodes determined by index $r$ are all nodes $x$ where $coin(v,x,r)=\textsc{Heads}$. We then determine the noisy size of $|W_{r}(v)|$ by drawing a noise variable from $\text{Lap}(2/\varepsilon^{\prime})$ to be used in the adaptive Laplace mechanism (Algorithm 2). Finally, we also determine the noisy number of nodes currently matched to $v$ (Algorithm 2). Then, $v$ determines the smallest $r$ that satisfies the SVT check in Algorithm 2. Intuitively, this means we are finding the largest subset of neighbors of $v$ that does not exceed the $b$ bound, with high probability. Then, $v$ releases $r_{v}$ (Algorithm 2) and for each neighbor $u$ of $v$ in $W_{r_{v}}(v)$ (Algorithm 2), $u$ (privately) adds $v$ to $M(u)$ (Algorithm 2) and $v$ (privately) adds $u$ to $M(v)$ (Algorithm 2).
If Algorithm 2 processed the vertices in order $v_{1},\dots,v_{n}$, then $v_{i}$ can decode its matched neighbors using the following formula
$M(v_{i})=\{v_{j}:j<i\land A(v_{j})>j\land v_{i}\in W_{r_{v_{j}}}(v_{j})\}\cup W_{r_{v_{i}}}(v_{i})\,.$
Here $W_{r_{v_{i}}}(v_{i})=\varnothing$ by convention if $A(v_{i})\leq i$ (i.e. if $v_{i}$ reaches its matching condition before the $i$-th iteration).
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching benchmark parameter $b^{\prime}\geq 1$, constant $\eta\in(0,1)$, constant $c\geq 3$, matching degree parameter $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}$
Output: An $\varepsilon$-local edge differentially private implicit $b$-matching.
1 $\varepsilon^{\prime}\leftarrow\frac{\varepsilon}{2(1+2\eta)/\eta}$
2 for every pair of vertices $u\neq v\in V\times V$ and subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$ do
3   Flip and release coin $coin(u,v,r)$ which lands heads with probability $p_{r}=(1+\eta)^{-r}$
4
5for each node $u\in V$ do
6 $\tilde{b}(u)\leftarrow b-36c\ln(n)/\varepsilon^{\prime}+\text{Lap}(4/\varepsilon^{\prime})$
7 $M(u)\leftarrow\varnothing$
8 $A(u)\leftarrow\infty$
9
10for iteration $i=1$ to $n$, let $v=v_{i}$ do
 // Multidimensional-AboveThreshold for checking if each node has reached their matching threshold
11 for each node $u\in V$ such that $A(u)>i$ ($u$ has not satisfied its matching condition) do
12    $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime})$
13    if $|M(u)|+\nu_{i}(u)\geq\tilde{b}(u)$ then
14       $u$ releases $A(u)\leftarrow i$
15
16
 // If $v$ can still match with more edges, find an additional set to match $v$ with.
17 if $A(v)>i$ ($v$ has not satisfied its matching condition) then
18    for subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$ do
19       $W_{r}(v)=\{u:A(u)>i\wedge u\text{ later than $v$ in ordering }\land\{u,v\}\in E\wedge coin(u,v,r)=\textsc{Heads}\}$
20       $|\widetilde{W}_{r}(v)|=|W_{r}(v)|+\text{Lap}(2/\varepsilon^{\prime})$
21
22    $|\widetilde{M}_{i}(v)|=|M(v)|+\text{Lap}(2/\varepsilon^{\prime})$
23    $v$ computes $r_{v}\leftarrow\min\{r:|\widetilde{M}_{i}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b\}$
24    $v$ releases $r_{v}$
25    for $u\in W_{r_{v}}(v)$ do
26       $M(u)\leftarrow M(u)\cup\{v\}$
27       $M(v)\leftarrow M(v)\cup\{u\}$
28
29
Algorithm 2 $\varepsilon$-LEDP Approximate Maximum $b$-Matching
5.2 Analysis (Theorem 2.2)
The pseudocode for Theorem 2.2 can be found in Algorithm 2. We prove its privacy and approximation guarantees separately. In Section 5.2.1, we first prove that Algorithm 2 is $\varepsilon$-DP. The utility proofs can then be found in Section 5.2.2.
5.2.1 Privacy Guarantees
We now prove that our algorithm is $\varepsilon$-LEDP using our privacy mechanisms given in Section 3.2 and show that Algorithm 2 can be implemented using local randomizers. In our proof, we implement three different types of local randomizers that use private data and perform various instructions of our algorithm. Note that not all of the local randomizers in our algorithm releases the computed information publicly but the computed information from these local randomizer algorithms satisfy edge-privacy.
Lemma 5.4.
Algorithm 2 is $\varepsilon$-locally edge differentially private.
Proof.
The algorithm consists of three different types of local randomizers which use the private information. The first type of local randomizer checks at each iteration whether or not each node has already reached its matching capacity. The second local randomizer computes the estimated number of neighbors, $|\widetilde{W}_{r}(v)|$, which are still active and are ordered after the current node for each subgraph index $r$. The third computes a noisy estimate $|\widetilde{M}(v)|$ for $v=v_{i}$ at iteration $i$. We will argue that each of these local randomizers are individually differentially private. Then, by the concurrent composition theorem (Lemma 3.17), the entire algorithm is differentially private as no other parts of the algorithm use the private information.
First, we prove that the local randomizers checking whether or not each node has reached their matching capacity (Algorithms 2, 2, 2, 2, 2 and 2) is an instance of the Multidimensional AboveThreshold (MAT) mechanism with sensitivity $\Delta=2$. The vector of queries at each iteration $i$ is the number of other nodes $|M(u)|$ each node $u$ has already been matched to before iteration $i$. Fixing the outputs of the previous iterations, the addition or removal of a single edge can only affect $|M(u)|$ for two nodes, each by at most $1$. Thus, Algorithms 2, 2, 2, 2, 2 and 2 indeed implement an instance of the Multidimensional AboveThreshold mechanism. But then these operations are $\varepsilon^{\prime}$-differentially private by Lemma 3.16.
Next, we prove that for each subgraph index $r$, the procedures that produce noisy estimates of the size of $W_{r}(v)$ in Algorithms 2, 2 and 2 is implemented via local randomizers using the Adaptive Laplace Mechanism with sensitivity $\Delta=2$. Fixing the outputs of the other parts of the algorithm (which concurrent composition allows us to do), the addition or removal of a single edge $\{u,w\}$ only affects $|W_{r}(v_{i})|$ if $u=v_{i}$ or $w=v_{i}$, and it affects each $|W_{r}(v_{i})|$ by 1. Thus, we have shown above that for a fixed subgraph index $r$, all estimates of $|W_{r}(v_{i})|$ over the iterations $i$ are $\varepsilon^{\prime}$-differentially private. Since we sample each edge with probability $p_{r}$, privacy amplification (Lemma 3.18) implies that this step is actually $2p_{r}\varepsilon^{\prime}$-differentially private since $\varepsilon^{\prime}\leq 1$.
Third, we prove that producing noisy estimates of $|M(v_{i})|$ in Algorithm 2 is an instance of the Adaptive Laplace Mechanism with sensitivity $\Delta=2$. Fixing the outputs of the other parts of the algorithm (again, which concurrent composition allows us to do), the additional and removal of an edge $\{u,w\}$ can again only affect $M(v_{i})$ for the iterations $i$ where $v_{i}=u$ and $v_{i}=w$, each by at most 1. Thus, the sensitivity is indeed $\Delta=2$ even with the adaptive choice of queries, and the noisy estimates of $|M(v_{i})|$ are $\varepsilon^{\prime}$-differentially private by Lemma 3.15.
Finally, applying the concurrent composition theorem (Lemma 3.17) proves that the entire algorithm is $2\varepsilon^{\prime}\cdot\left(1+\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}p_{r}\right)$-differentially private. We can upper bound this by an infinite geometric series:
$2\varepsilon^{\prime}\cdot\left(1+\textstyle\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}p_{r}\right)$ $\leq 2\varepsilon^{\prime}\cdot\left(1+\textstyle\sum_{r=0}^{\infty}(1+\eta)^{-r}\right)=2\varepsilon^{\prime}\cdot\left(1+\frac{1+\eta}{\eta}\right).$
By the choice of $\varepsilon^{\prime}=\frac{\varepsilon}{2(1+2\eta)/\eta}$ (Algorithm 2), we can conclude that the algorithm is $\varepsilon$-LEDP. ∎
5.2.2 Utility
We first prove that the implicitly output $b$-matching is at least half the size of a maximum $b^{\prime}$-matching, with high probability.
Lemma 5.5.
Let $M\subseteq E$ be an edge subset satisfying the following maximality condition: every vertex is either matched to at least $b^{\prime}$ edges of $M$, or all its unmatched neighbors are matched to at least $b^{\prime}$ edges of $M$. Then if $\operatorname{OPT}$ denotes the size of a maximum $b^{\prime}$-matching, we have
$|M|\geq\frac{1}{2}\operatorname{OPT}\,.$
Proof.
Let $M^{*}$ denote an arbitrary $b^{\prime}$-matching. We would like to show that $\frac{|M|}{|M^{*}|}\geq\frac{1}{2}$. Since $\frac{|M\setminus M^{*}|}{|M^{*}\setminus M|}\leq\frac{|M|}{|M^{*}|}$, we may assume without loss of generality that $M\cap M^{*}=\varnothing$ after adjusting the $b^{\prime}$ values per vertex.
Let $S\subseteq V$ denote the set of nodes $v$ with at least $b^{\prime}(v)$ matched edges in $M$. We claim that every edge of $M^{*}$ must have an endpoint in $S$. Otherwise, there is an edge $uv$ such that $u,v$ are matched to less than $b^{\prime}(u),b^{\prime}(v)$ neighbors, respectively, but is not in $M$, contradicting the maximality condition.
Now, each $v\in S$ is incident to at most $b^{\prime}(v)$ edges in $M^{*}$, while every edge of $M$ is incident to at most $2$ edges of $S$. Moreover, each node in $S$ is incident to at least $b^{\prime}(v)$ edges of $M$ by construction. Thus by a double counting argument, $|M^{*}|\leq\sum_{v\in S}b^{\prime}(v)\leq 2|M|$ as desired. ∎
The utility proof of Theorem 2.2 is implied by the following lemma, which is a restatement with detailed constants.
Lemma 5.6.
For $\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}$, Algorithm 2 is $\varepsilon$-LEDP and outputs an implicit $b$-matching $M$ in the billboard model such that $|M|\geq\frac{1}{2}\operatorname{OPT}(b^{\prime})$ with probability at least $1-\frac{1}{n^{c}}$. Here $\operatorname{OPT}(b^{\prime})$ is the size of a maximum $b^{\prime}$-matching.
Proof.
For each iteration $i$, we will show that the number of nodes that $v=v_{i}$ matches with after iteration $i$ is either at least $b^{\prime}$, or every one of its unmatched neighbors is already matched to at least $b^{\prime}$ of their neighbors. Hence the final output satisfies the maximality condition of Lemma 5.5 and thus the approximation guarantees. We will show that this suffices to guarantee that our output is of comparable size to a maximal $b^{\prime}$-matching. Finally, we show that with high probability, we do not match with more than $b$ nodes.
First, all of the Laplace noise (Algorithms 2, 2, 2 and 2) are drawn from distributions with expectation 0. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
  • $b-36c\log(n)/\varepsilon^{\prime}\leq\tilde{b}(u)\leq b-24c\log(n)/\varepsilon^{\prime}$ for all $u\in V$,
  • $|\nu_{i}(u)|\leq{24c\ln(n)/\varepsilon^{\prime}}$ for all $i\in[n],u\in V$,
  • $|W_{r}(v)|-{6}c\log(n)/\varepsilon^{\prime}\leq|\widetilde{W}_{r}(v)|\leq|W_{r}(v)|+{6}c\log(n)/\varepsilon^{\prime}$ for all $v\in V,r\in[\lceil\log_{1+\eta}(n)\rceil]$, and
  • $|M(v)|-{6}c\log(n)/\varepsilon^{\prime}\leq|\widetilde{M}_{i}(v)|\leq|M(v)|+{6}c\log(n)/\varepsilon^{\prime}$ for all $v_{i}\in V$.
Let $\mathcal{E}$ denote the intersection of all the events above. By a union bound, $\mathcal{E}$ occurs with probability at least $1-\frac{1}{n^{2c}}$ for $c\geq 3$. We condition on $\mathcal{E}$ for the rest of the proof.
Case I: Suppose that Algorithm 2 does not execute, i.e. $v$ has already satisfied its matching condition as evaluated on Algorithms 2, 2 and 2. Then we know that $|M(v)|+\nu_{i}(v)\geq\tilde{b}(v)$ by definition. We then have that
$|M(v)|$ $\geq b-{60}c\ln(n)/\varepsilon^{\prime}$
$\geq{b}-\frac{72c\ln(n)}{\varepsilon^{\prime}}$
$={b}-\frac{8\cdot 72c\ln(n)}{\eta\varepsilon}$ (by our choice of $\varepsilon^{\prime}$ in Algorithm 2)
$\geq{b}-\frac{576c\ln(n)}{\eta\varepsilon}$
$>{b^{\prime}}\,.$ (by $b\geq b^{\prime}+\frac{576c\ln(n)}{\eta\varepsilon}$)
In other words, $v$ was already matched to $b^{\prime}$ neighbors from a previous iteration.
Case II.a: Now, assume that Algorithm 2 does execute. Let $W_{0}(v)$ be the set of unmatched neighbors of $v$ which appear later in the ordering. First, consider the case that $|W_{0}(v)|<\frac{12c\ln(n)}{\eta^{2}}$. We know that either $|M(v)|<b^{\prime}$ or $|M(v)|\geq b^{\prime}$. If $|M(v)|\geq b^{\prime}$, then $v$ is already matched to sufficiently many neighbors and there is nothing to prove. Otherwise, conditioned on $\mathcal{E}$,
$|W_{0}(v)|+3\cdot\frac{12c\ln(n)}{\varepsilon^{\prime}}+{b^{\prime}}$ $\leq\frac{12c\ln(n)}{\eta^{2}}+\frac{36c\ln(n)}{\varepsilon^{\prime}}+{b^{\prime}}$
$\leq\frac{12c\ln(n)}{\eta^{2}}+\frac{216c\ln(n)}{\eta\varepsilon}+{b^{\prime}}$
$\leq\frac{238c\ln(n)}{\eta^{2}\varepsilon}+{b^{\prime}}$
$\leq b$
and all of $W_{0}(v)$ is matched.
Case II.b.1: For the remainder of the proof, we assume $k\coloneqq|W_{0}(v)|\geq\frac{12c\ln(n)}{\eta^{2}}$. As a reminder, we are in the case that Algorithm 2 does execute. Let $r^{*}$ be the $r$ chosen in Algorithm 2. If the returned $r^{*}>\log_{1+\eta}(k\eta^{2}/(12c\ln{n}))$, we know that the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b$ did not hold for $r=\log_{1+\eta}\left(\frac{k\eta^{2}}{12c\ln{n}}\right)$. Note here we use $k\geq\frac{12c\ln(n)}{\eta^{2}}$ so that $r\geq 0$ is well-defined. Thus, we have that $p_{r}=\frac{12c\ln(n)}{k\eta^{2}}$ so the expected size of $W_{r}(v)$ is $\frac{12c\ln(n)}{\eta^{2}}$; that is, $\mathbb{E}[|W_{r}(v)|]=\frac{12c\ln(n)}{\eta^{2}}$. By a multiplicative Chernoff Bound (Theorem A.2), the true size of $W_{r}(v)$ is at most $18c\log(n)/\eta^{2}$ with probability at least $1-\frac{1}{n^{4c}}$. Since the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\log(n)/\varepsilon^{\prime}\leq b$ didn’t hold for such $r$, we have that
$|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+\frac{12c\ln(n)}{\varepsilon^{\prime}}$ $>b$
$|M(v)|+|W_{r}(v)|+\frac{6c\ln(n)}{\varepsilon^{\prime}}+\frac{6c\ln(n)}{\varepsilon^{\prime}}+\frac{12c\ln(n)}{\varepsilon^{\prime}}$ $>b$ (conditioning on $\mathcal{E}$)
$|M(v)|+|W_{r}(v)|+\frac{24c\ln(n)}{\varepsilon^{\prime}}$ $>b\,.$
In other words,
$|M(v)|$ $>b-|W_{r}(v)|-\frac{24c\ln(n)}{\varepsilon^{\prime}}$
$|M(v)|$ $>b-\frac{18c\ln(n)}{\eta^{2}}-\frac{24c\ln(n)}{\varepsilon^{\prime}}$ ( w.p. at least $1-\frac{1}{n^{4c}}$)
$|M(v)|$ $>b-\frac{18c\ln(n)}{\eta^{2}}-\frac{144c\ln(n)}{\eta\varepsilon}$ (by the choice of $\varepsilon^{\prime}$)
$|M(v)|$ $>b-\frac{162c\ln(n)}{\eta^{2}\varepsilon}$
$|M(v)|$ $>{b^{\prime}}\,.$ (since $b\geq{b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}}$)
In other words, $v$ was already matched from a previous iteration.
Case II.b.2: If $r^{*}\leq\log_{1+\eta}\left(\frac{k\eta^{2}}{12c\ln{n}}\right)$, there are two scenarios. Either $r^{*}=0$, then we match $v$ with all nodes which are still available so clearly the number of nodes $v$ is matched with is at least $b^{*}(v)$, or $r^{*}>1$. In the latter scenario, we first remark that $p_{r^{*}}\geq\frac{12c\ln(n)}{k\eta^{2}}$ so that
$\mathbb{E}[|W_{r^{*}-1}|]=(1+\eta)\mathbb{E}[|W_{r^{*}}|]\geq(1+\eta)\frac{12c\ln(n)}{\eta^{2}}\,.$
Thus we can apply a Chernoff bound on both $|W_{r^{*}-1}|,|W_{r^{*}}|$ and obtain high probability bounds. We also know that the threshold is exceeded for $r^{*}-1$, so we have $|\widetilde{M}(v)|+|\widetilde{W}_{r^{*}-1}(v)|+12c\ln(n)/\varepsilon^{\prime}\geq b$. Conditioning on $\mathcal{E}$, $|M(v)|+|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Thus by a Chernoff bound on $|W_{r^{*}-1}|$, we see that
$b-|M(v)|-\frac{36c\ln(n)}{\varepsilon^{\prime}}\leq|W_{r^{*}-1}|\leq(1+\eta)\mathbb{E}[|W_{r^{*}-1}|]=(1+\eta)^{2}\mathbb{E}[|W_{r^{*}}|]$
with probability $1-\frac{1}{n^{4c}}$. Now again by a Chernoff bound but now on $|W_{r^{*}}|$, we have
$|W_{r^{*}}|\geq\frac{(1-\eta)}{(1+\eta)^{2}}\left[b-|M(v)|-\frac{36c\ln(n)}{\varepsilon^{\prime}}\right]$
with probability $1-\frac{1}{n^{4c}}$.
Hence, if $|M(v)|<b^{\prime}$ (meaning it has not satisfied the matching criteria), it holds that
$|W_{r^{*}}(v)|$ $\geq\frac{(1-\eta)}{(1+\eta)^{2}}(b-|M(v)|-\frac{216c\ln(n)}{\eta\varepsilon})$
$\geq\frac{1-\eta}{(1+\eta)^{2}}\left(\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}-|M(v)|-\frac{216c\ln(n)}{\eta\varepsilon}\right)$
$\geq b^{\prime}-|M(v)|\,.$
It follows that $v$ is matched with at least $b^{\prime}$ neighbors at the end of this iteration.
Taking a union bound over all $r$ and $v\in V$, all Chernoff bounds hold regardless of the value of $r^{*}$ with probability at least $1-\frac{1}{n^{2c}}$ when $c\geq 3$. Taking into account the conditioning on $\mathcal{E}$, our algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for $c\geq 3$.
In summary, at iteration $i$, we ensure that 1) either $v=v_{i}$ was already matched to at least $b^{\prime}$ neighbors during prior iterations, 2) or it will be matched with a sufficient number of descendents to satisfy a total of at least $b^{\prime}$ matched neighbors after iteration $i$, 3) or it will be matched to all descendents that do not yet have $b^{\prime}$ matched neighbors. The only property to check is whether there are any ancestors $u$ of $v$ that are not matched to $v$ but have yet to achieve $b^{\prime}$ matched neighbors in the last case. But this cannot happen since $u$ would have matched to all descendents that did not yet have $b^{\prime}$ matched neighbors during its iteration and $v$ in particular did not have $b^{\prime}$ matched neighbors.
It remains only to check that we match each vertex to at most $b$ neighbors. To do so, it suffices to check that on the $i$-th iteration, each node which does not yet satisfy a matching condition can match with at least 1 more neighbor after Algorithm 2 and that $v=v_{i}$ is matched with at most $b$ neighbors at the end of the iteration (Algorithm 2). The first statement is guaranteed since any vertex $u$ not yet satisfying its matching condition satisfies
$|M(u)|$ $<\tilde{b}(u)-\nu_{i}(u)$
$|M(u)|$ $<b-24c\log(n)/\varepsilon^{\prime}+24c\log(n)/\varepsilon^{\prime}$ (conditioning on $\mathcal{E}$)
$|M(u)|$ $<b.$
If $v$ already satisfies its matching condition after Algorithm 2, the second statement is guaranteed to hold. Otherwise, $|M(v)|<b$ and we choose $r$ such that
$b$ $\geq|\tilde{M}_{i}(v)|+|\tilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}$
$\geq|M(v)|+|W_{r}(v)|-12c\ln(n)/\varepsilon^{\prime}+12c\ln(n)/\varepsilon^{\prime}$ (conditioning on $\mathcal{E}$)
$=|M(v)|+|W_{r}(v)|.$
Thus in this case, $v$ is also matched to at most $b$ neighbors, as desired. ∎
5.3 Corollaries
By choosing a fixed constant $\eta=\nicefrac{{1}}{{2}}$, matching benchmark parameter $b^{\prime}=1$, and matching degree parameter $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, Theorem 2.2 immediately yields a $b$-matching that is the size of a $2$-approximate 11-matching. As stated in Theorem 5.2, when $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$ is slightly larger, we avoid the need for bicriteria approximation and our algorithm yields a $(2+\eta)$-approximate $b$-matching. The proof follows from the following corollary of Lemma 5.6.
Corollary 5.7 ($(2+\eta)$-Approximate Maximum $b$-Matching).
When $b\geq\frac{576c\ln(n)}{\eta^{3}\varepsilon}$ for a constant $\eta\in(0,\nicefrac{{1}}{{2}})$, Algorithm 2 is $\varepsilon$-LEDP and outputs an implicit $(2+24\eta)$-maximum $b$-matching in the billboard model, with probability at least $1-\frac{1}{n^{c}}$.
Before proceeding with the proof, we recall that $b$-matchings have the following integral linear programming relaxation [schrijver2003combinatorial, Theorem 33.2], where $\partial(U)$ denotes the set of outgoing edges from $U\subseteq V$.
$\sum_{uv\in E}x(uv)$ $\leq b$ $\forall v\in V$
$\sum_{e\in E[U]}x(e)+\sum_{f\in F}x(f)$ $\leq\lfloor\frac{1}{2}(b|U|+|F|)\rfloor$ $\forall U\subseteq V,F\subseteq\partial(U)$
$x$ $\in[0,1]^{E}$
Note that it suffices to consider constraints of the second form only when $b|U|+|F|$ is odd, but we include the redundant constraints for the sake of simplicity.
Proof.
Choose $b^{\prime}$ such that
$b=(1+12\eta)b^{\prime}\geq(1+11\eta)b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}\,.$
Then we can apply Lemma 5.6 to deduce that Algorithm 2 outputs a $2$-approximate maximum $(\frac{b}{1+12\eta})$-matching.
Identify any $b$-matching $M$ as an indicator $\chi_{M}\in[0,1]^{E}$ of the $b$-matching polytope. Then $\frac{1}{1+12\eta}\chi_{M}$ is a feasible fractional solution of the $(\frac{b}{1+12\eta})$-matching polytope. Hence the size of a maximum $(\frac{b}{1+12\eta})$-matching is at least a $\frac{1}{1+12\eta}$ fraction of the size of a maximum $b$-matching. This concludes the proof. ∎
5.4 Lower Bound for Implicit Matchings
We now prove Theorem 5.3, our lower bound for implicit solutions, which essentially matches our upper bound from Corollary 5.1. While it does not apply to all algorithms in the billboard model, it applies to the implicit solutions that our algorithms use (Definition 3.6).
Recall that, given an input graph $H$, our implicit solutions essentially output a subset of vertices for each node $x$ whose intersection with $x$’s private adjacency list gives the nodes $x$ is matched to. We denote these public subsets of vertices by $\mathcal{S}=\{S_{x}\}_{x\in V}$ where $S_{x}$ is the public subset of vertices for node $x$. That is, $\mathcal{S}=\{S_{x}\}_{x\in V}$ is the implicit solution generated by $\mathcal{A}$. We say that $\mathcal{A}$ includes an edge $\{u,v\}$ if either $u\in S_{v}$ or $v\in S_{u}$. In other words, $\mathcal{A}$ includes an edge if that edge is an edge in $H(\mathcal{S})$.
Proof of Theorem 5.3.
Let $V=\{r,v_{1},v_{2},\dots,v_{n-1}\}$. Let $i\in\{2,3,\dots,n-1\}$, and let $G_{i}$ be a graph with just one edge $\{r,v_{i}\}$. Then $\mathcal{A}$ must include $\{r,v_{i}\}$ with probability at least $1-\beta$ in order to meet the utility guarantee of Theorem 5.3 (note that $\eta$ vanishes since in this case the maximal matching has size $1$). Now consider a graph $G$ with just one edge $\{r,v_{1}\}$. Since $G_{i}$ has distance $2$ from $G$, by the differential privacy guarantee we know that the probability that $\mathcal{A}$ includes $\{r,v_{i}\}$ when run on $G$ must be at least $1-e^{2\varepsilon}\beta$. Note that this is true for all $i\in\{2,3,\dots,n-1\}$.
To simplify notation, let $\mathcal{S}$ be the implicit solution output by $\mathcal{A}$. Now let $T$ be an arbitrary subset of $\{v_{2},v_{3},\dots,v_{n-1}\}$ of size $\frac{1}{2\varepsilon}\log(1/\beta)$. Then the above argument implies that the expected number of edges between $r$ and $T$ in $H(\mathcal{S})$ is at least $(1-e^{2\varepsilon}\beta)|T|$ when we run $\mathcal{A}$ on $G$, or equivalently the expected number of non-edges between $r$ and $T$ is at most $e^{2\varepsilon}\beta|T|$. So by Markov’s inequality,
[…]
Let $Q$ denote the event that the number of non-edges from $r$ to $T$ in $H(\mathcal{S})$ is at least $\frac{|T|}{2}$. Let $G^{\prime}=(V,E^{\prime})$ be a different graph with the same vertex set but with $E^{\prime}=\{\{r,v_{i}\}:v_{i}\in T\}$. Since $G^{\prime}$ has distance $|T|$ from $G$, group privacy implies that the probability of $Q$ when we run $\mathcal{A}$ on $G^{\prime}$ is at most
$e^{|T|\varepsilon}2e^{2\varepsilon}\beta$ $=e^{\frac{1}{2\varepsilon}\log(\nicefrac{{1}}{{\beta}})\varepsilon}2e^{2\varepsilon}\beta=\beta^{-\frac{1}{2}}2e^{2\varepsilon}\beta=\beta^{\frac{1}{2}}2e^{2\varepsilon}\leq\frac{1}{2}.$
Thus with probability at most $\nicefrac{{1}}{{2}}$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at most $\nicefrac{{|T|}}{{2}}$ edges from $r$ to $T$. Thus with probability at least $\nicefrac{{1}}{{2}}$, when we run $\mathcal{A}$ on $G^{\prime}$ the implicit solution we get includes at least $\nicefrac{{|T|}}{{2}}$ edges from $r$ to $T$. Since all of those edges are also edges of $G^{\prime}$, this means that the degree of the implicit solution is at least $\nicefrac{{|T|}}{{2}}=\Theta\left(\frac{1}{\varepsilon}\log(\nicefrac{{1}}{{\beta}})\right)$ with probability at least $\nicefrac{{1}}{{2}}$. ∎
6 $O(\log n)$-Round LEDP Implicit $b$-Matching via PVSM
In this section, we present a distributed implementation of our matching algorithm that uses $O(\log n)$ rounds in the LEDP model. Specifically, we prove Theorem 2.3, restated below.
See 2.3
As an immediately corollary for the case $b^{\prime}=1$, we obtain the following.
Corollary 6.1.
Let $\varepsilon\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{\varepsilon}\right)$. Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) $b$-matching in the billboard model whose size is at least a $2$-approximate maximum $1$-matching.
Similar to the sequential algorithm, for $b$ sufficiently large, we can ensure an approximation guarantee with respect to the maximum $b$-matching.
Theorem 6.2.
Let $\varepsilon\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{{\eta^{5}}\varepsilon}\right)$. Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) $b$-matching in the billboard model whose size is at least a $(2+\eta)$-approximate maximum $b$-matching.
High-Level Overview.
Our algorithm performs multiple rounds of matching where in each round some nodes are proposers and others are receivers. A node participates in a round if it has not satisfied its matching condition; a node that participates in the current round is an active node. The proposers are chosen randomly and propose to a set of nodes to match. We call the set of nodes each node proposes to as the proposal set. Receivers are active nodes that are not chosen as proposers and receive the proposers’ proposals. Then, each receiver chooses a subset of proposals to accept. The algorithm continues until all nodes satisfy the matching condition.
The algorithm we use to prove Theorem 2.3 modifies Algorithm 2. Our algorithm (the pseudocode is given in Algorithm 3) modifies Algorithm 2 in multiple ways and we briefly comment on the three main changes here. First, we perform several rounds of matching with multiple proposers in each round. Each round uses a fresh set of coin flips. A node may propose in more than one round because many proposers may propose to the same set of receivers; then, many of the proposers will remain unmatched and will need to propose again in a future round. Second, each node that has not satisfied its matching condition determines whether it is a proposer with $1/2$ probability. Third, when all proposers have released their proposal sets, the receivers release their match sets which chooses among the proposers who proposed to them. Then, only the pairs that exist in both the proposal and match sets will be matched.
We show that $O(\log(n))$ rounds are sufficient, with high probability, to obtain a matching that satisfies Theorem 2.3. Notably, we use our Public Vertex Subset Mechanism that we developed in Section 5 as the main routine for both our proposal and match sets.
We give some intuition about why we need to make these changes for the distributed version of our algorithm. We need a fresh set of coin flips for each round because a proposer may participate in multiple rounds. An intuitive reason for why this is the case is due to the fact that some (unlucky) proposers may not be matched to most of the receivers in their proposal sets, in the current round. Hence, a proposer would need to choose another proposal set in the next round. In order to ensure that this new proposal set does not depend on the matched vertices of the previous round, we must flip a new set of coins. Second, because proposers release their sets simultaneously, we cannot have a node simultaneously propose and receive; if a node simultaneously proposes and receives, we do not have a way to ensure that the matching thresholds are not exceeded. Thus, we have a two-round process where in the first synchronous round, proposers first propose and then in the next round, receivers decide the matches.
6.1 Detailed Algorithm Description
We now describe our pseudocode for Algorithm 3 in detail and then prove its properties. We use a number of constants that guarantees our $\Theta(\log n)$ round algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$. Our algorithm returns a $(2+\eta)$-approximate $b^{\prime}$-matching with high probability when the degree parameter $b$ satisfies the bound given in Theorem 2.3.
In our pseudocode given in Algorithm 3, we define $M(u)$ to be the set of nodes currently matched to $u\in V$. For each round $i$, we let $V_{i}$ be the set of active vertices which have not reached their matching $b^{\prime}$-condition in round $i$. Our algorithm uses a subroutine (Algorithm 4) which takes a vertex $v$, a set of vertices $S$ from which to select a proposal or matching set, and returns a subgraph index which determines a subset of vertices based on public i.i.d. coin flips. This algorithm implements a more specific version of our Public Vertex Subset Mechanism.
Algorithm 3 takes as input a graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, a matching benchmark parameter $b^{\prime}\geq 1$, and a degree parameter $b$. The algorithm returns an $\varepsilon$-LEDP implicit $b$-matching. First, we set some additional approximation and privacy parameters in Algorithm 3. We then iterate through every node in Algorithm 3 simultaneously to determine the noisy threshold of every node (Algorithm 3). This threshold is used to determine if a vertex satisfies its matching condition. In particular, if the estimated number of possible matches for the node is greater than the threshold, then with high probability, the node is matched to at least $b^{\prime}$ neighbors. The threshold is set in Algorithm 3 using Laplace noise. We add Laplace noise to the threshold as part of an instance of the Multidimensional AboveThreshold (MAT) technique (Lemma 3.16). This is in turn used to determine for every node whether the number of matches exceeds this threshold. We also initialize an empty set for each $u\in V$, denoted $M(u)$, that contains the set of nodes $v$ is matched to (Algorithm 3).
$V_{i}$ contains the set of remaining active nodes in round $i$. Initially, in the first round, all nodes are active (Algorithm 3). We proceed through $O(\log n)$ rounds of matching (Algorithm 3). For each node, we first check using the MAT (simultaneously in Algorithm 3) whether its matching condition has been met. To do so, we add Laplace noise from the appropriate distribution (Algorithm 3) to the size of $u$’s current matches (Algorithm 3). If this noisy size exceeds the noisy threshold, we output that node $u$ has satisfied its matching condition and we remove $u$ from $V_{i}$ (Algorithm 3). Next, we flip the coins for round $i$. These coin flips are used to determine the proposal and match sets. Recall that we produce coin flips for each pair of nodes in the graph. Then, the coin flips are used to determine an implicit set of edges used to match nodes. Although the coin flips are public, only the endpoints of each existing edge knows whether that edge is added to a proposal or match set. All coin flips are done simultaneously and are performed by the curator (Algorithm 3). We flip a coin for each unique pair of vertices and for each $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$. The probability that the coin lands Heads is determined by $r$. Specifically, the coin for the $(i,j,r)$ tuple, denoted $coin(i,j,r)$, is Heads with probability $(1+\eta)^{-r}$ (Algorithm 3). This ensures that the $(r+1)$-th set is, in expectation, a factor of $\frac{1}{1+\eta}$ smaller than the $r$-th set.
We then flip another set of coins to determine which nodes are proposers (Algorithm 3) in round $i$. Each node is a proposer with $1/2$ probability; the result of the coin flip is stored in $a_{i}(v)$ for each $v\in V$. We only select proposers from the set of active vertices, $V_{i}$. For each selected proposer, we simultaneously call the procedure PrivateSubgraph with the inputs $G,\varepsilon^{\prime},\eta,b,$ $w,V_{i}\setminus P_{i}$ and coincoin (Algorithms 3 and 3). The pseudocode for PrivateSubgraph is given in Algorithm 4. The function takes as input a graph $G$, a privacy parameter $\varepsilon^{\prime}$, an approximation parameter $\eta$, the degree cap $b$, a vertex subset $S$, a vertex $w\in V$, and all public coin flips coincoin. The function then iterates through all possible subgraph indices $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ (Algorithm 4). For each index, we determine the set of nodes $u$ that satisfy the following conditions: $u\in S$, the edge $\{w,u\}$ exists, and the coin flip $coin(u,w,r)$ is Heads. This set of nodes is labeled $W_{r}(w)$ (Algorithm 4). We then add Laplace noise to the size of this set to obtain a noisy estimate for the size of $W_{r}(w)$ (Algorithm 4). We use the Adaptive Laplace Mechanism to determine the smallest $r$ (most number of nodes $w$ can propose to) that does not exceed $b$. To do this, we add Laplace noise to the size of the set of nodes matched to $w$ (Algorithm 4) and find the smallest $r$ such that the sum of the noisy proposal set size, $|\widetilde{W}_{r}(w)|$, and the noisy matched set size, $|\widetilde{M}(w)|$, plus $27c\log(n)/\varepsilon^{\prime}$ does not exceed $b$ (Algorithm 4). The term $27c\log(n)/\varepsilon^{\prime}$ is added to ensure that $|W_{r}(w)|+|M(w)|$ does not exceed our $b$ bound (by drawing negative noises), with high probability. The procedure releases the smallest subgraph index (Algorithm 4) satisfying Algorithm 4.
We save the released index from PrivateSubgraph in $r_{w}$. Then, $w$ releases $r_{w}$ (Algorithm 3). After the proposers release their subgraph indices, the receivers then determine their match sets. We iterate through all receivers simultaneously and for each receiver $w$, the receiver $w$ computes the set of proposers that proposed to it, denoted as $R$ (Algorithm 3). The receiver can privately compute this set since they know which edges are incident to it, which of their neighbors are proposers, the coin flips of each of the incident edges, and the released subgraph indices of their neighbors. Using the publicly released subgraph indices $r_{v}$ of each neighbor $v$, receiver $w$ can then check $coin(v,w,r_{v})$ to see if edge $\{v,w\}$ is included in proposer $v$’s proposal set. Using $R$, receiver $w$ then computes the match set by calling PrivateSubgraph (Algorithm 3). The receiver releases the subgraph index associated with the match set (Algorithm 3).
The final steps compute the new edges that are in the matching that each node stores privately. We iterate through all pairs of proposers and receivers simultaneously (Algorithm 3). For each pair, $v$ checks whether it is in $W_{r_{u}}(u)$ by checking if $coin(v,u,r_{u})=$ Heads and vice versa for $w$ (Algorithm 3). Then, if the pair is in both the proposal and match sets, $v,u$ matches with each other and $v$ adds $u$ to $M(v)$ and $u$ adds $v$ to $M(u)$ (Algorithm 3). The sets $M(u)$ and $M(v)$ are stored privately but they are computed using the public transcript; hence, the release transcript is the implicit solution that allows each node to know and privately store which nodes it is matched to.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching benchmark parameter $b^{\prime}\geq 1$, constant $c\geq 1$
Output: An $\varepsilon$-local edge differentially private ($\varepsilon$-LEDP) implicit $b$-matching
1
2Let $\eta\leftarrow 1/2,\varepsilon^{\prime}\leftarrow\varepsilon/(3072c\log_{16/15}(n)),\varepsilon^{\prime\prime}\leftarrow\varepsilon/3$
3 Set parameter $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{518\log_{1+\eta}(n)}{\eta^{4}\,\varepsilon^{\prime}}$
4 for each node $u\in V$ (simultaneously) do
5 $\tilde{b}(u)\leftarrow b-259c\log_{1+\eta}(n)/\varepsilon^{\prime}+\mathrm{Lap}(4/\varepsilon^{\prime\prime})$
6 $M(u)\leftarrow\varnothing$
7 end for
[…]
9 for round $i=1$ to $\lceil 512c\cdot\log_{16/15}(n)\rceil$ do
10 for each node $u\in V_{i}$ (simultaneously) do
11    $\nu_{i}(u)\leftarrow\mathrm{Lap}(8/\varepsilon^{\prime\prime})$
12    if $|M(u)|+\nu_{i}(u)\geq\widetilde{b}(u)$ then
13       Release: node $u$ has satisfied matching condition and remove $u$ from $V_{i}$
14
15      end if
16
17   end for
18
19 for each tuple $(i,j,r)$ where $i\in[n]$, $j\in\{i+1,\dots,n\}$ and $r\in\{0,\ldots,\lceil\log_{1+\eta}(n)\rceil\}$ do
20      Flip and release $coin_{m}(i,j,r)$ and $coin_{p}(i,j,r)$ which each lands Heads with probability $p_{r}=(1+\eta)^{-r}$
21
22   end for
23 $P_{i}\leftarrow\{v:a_{i}(v)=\textsc{Heads}\wedge v\in V_{i}\}$ where $a_{i}(v)$ is Heads with probability $p=1/2$
24
25 for $w\in P_{i}$ simultaneously do
[…]
27    Release $r_{w}$
28
29   end for
30
31 for $w\in V_{i}\setminus P_{i}$ simultaneously do
32    $R\leftarrow\{u\mid w\in W_{r_{u}}(u)\wedge u\in P_{i}\}$
[…]
34    Release $r_{w}$
35
36   end for
37 for $v\in P_{i}$ and $u\in V_{i}\setminus P_{i}$ (simultaneously) do
38    if $v\in W_{r_{u}}(u)$ and $u\in W_{r_{v}}(v)$ then
39       $v,u$ matches; add $u$ to $M(v)$ and $v$ to $M(u)$
40
41      end if
42
43   end for
44 $V_{i+1}\leftarrow V_{i}$
45 end for
Algorithm 3 $O(\log n)$-Round $\varepsilon$-LEDP Approximate Maximum $b^{\prime}$-Matching
Input: Graph $G=(V,E)$, privacy parameters $\varepsilon^{\prime}$, approximation parameter $\eta$, degree cap $b$, vertex set $S$, vertex $v\in V$, and public coin flips coincoin
Output: Release subgraph index $r$
1
2FunctionPrivateSubgraph($G,\varepsilon^{\prime},\eta,b,w,S,coin$) begin
3 for subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$ do
4    $W_{r}(w)\leftarrow\{u\mid u\in S\wedge\{w,u\}\in E\wedge coin(u,w,r)=\textsc{Heads}\}$
5    $|\widetilde{W}_{r}(w)|\leftarrow|W_{r}(w)|+\text{Lap}(4/\varepsilon^{\prime})$
6   end for
7 $|\widetilde{M}_{i}(w)|\leftarrow|M(w)|+\text{Lap}(2/\varepsilon^{\prime})$
8 $r\leftarrow\min\left(\{r:|\widetilde{M}_{i}(w)|+|\widetilde{W}_{r}(w)|+27c\log(n)/\varepsilon^{\prime}\leq b\}\right)$
9   Return $r$
10 end
11
Algorithm 4 Private Subgraph Release
6.2 Analysis (Theorem 2.3)
The pseudocode for Theorem 2.3 is presented in Algorithm 3. We prove its privacy and utility guarantees separately. We first prove that it satisfies $\varepsilon$-DP in Section 6.2.1 and then prove its approximation guarantees in Section 6.2.2.
6.2.1 Privacy Guarantees
Our privacy proof follows a similar flavor to the $b$-matching privacy proof given in the previous section except for one main difference. While the adaptive Laplace mechanism was called twice per node in its use in the previous algorithm, this is not the case in our distributed algorithm. Suppose we are given neighbor graphs $G$ and $G^{\prime}$ with edge $\{u,v\}$ that differs between the two. Nodes $u$ and/or $v$ can propose a set many times during the course of the algorithm since their proposed sets are not guaranteed to be matched. If they are particularly unlucky, they could propose a set during every round of the algorithm. This means that by composition, we have privacy loss proportional to the number of rounds. We formally prove the privacy of our algorithm below.
Lemma 6.3.
Algorithm 3 is $\varepsilon$-LEDP.
Proof.
To prove that Algorithm 3 is $\varepsilon$-LEDP, we show that our algorithm can be implemented using local randomizers. Each node $v$ in the algorithm only releases a privatized set of information in the following Algorithms 3, 3 and 3: when $v$ has satisfied its $b^{\prime}$-matching condition (Algorithm 3), when $v$ releases a proposal subgraph index (Algorithm 3), and when $v$ releases a match set index (Algorithm 3).
Each node implements three different local randomizers for releasing each of the aforementioned three types of information. Our local randomizers are created for each type of private information used to determine the released output. Namely, we use a local randomizer to determine when a node has satisfied the $b^{\prime}$-matching condition (Algorithm 3), a local randomizer for computing the noisy proposal or match sets in Algorithm 4 of Algorithm 4, and a local randomizer for determining the noisy set of matched edges in Algorithm 4 of Algorithm 4. The releases in Algorithms 3 and 3 solely depend on the noisy proposal/match sets and the noisy set of matched edges. Then, by the concurrent composition theorem, the entire algorithm is differentially private since no other parts of the algorithm use the private graph information.
First, before we dive into the privacy components, we note that the curator flips coins and releases the result of the coins (Algorithms 3 and 3); these coin flips do not lose any privacy since the coins are not tied to private information. The coin flips are performed for each public pair of distinct nodes. Our proof follows the privacy proof of Lemma 5.4 except we need to account for multiple uses of the adaptive Laplace mechanism for each node.
The first set of local randomizers for all nodes $u\in V$ that releases whether $u$ has satisfied its matching condition can be implemented using an instance of the Multidimensional AboveThreshold (MAT) mechanism with sensitivity $2$. We show how Algorithms 3, 3, 3, 3, 3 and 3 can be implemented using MAT. First, our vectors of queries at each round $i$ is an $n$-length vector which contains the number of nodes each node $u\in V$ has already matched to. Specifically, the vector of queries at each round $i$ is the number of other nodes, $|M(u)|$, each node $u$ has already been matched to before round $i$. Conditioning on the outputs of the previous iterations, the addition or removal of a single edge can only affect $M(u)$ for two nodes, each by at most $1$. Hence, the sensitivity of the vector of queries is $2$; furthermore, the sensitivity of each vector of queries for each of the rounds is $2$. This means that MAT can be implemented with $\Delta_{M}=2$. Thus, the first set of local randomizers is $\varepsilon^{\prime\prime}$-differentially private by Lemma 3.16.
The second type of local randomizer computes the $|\widetilde{W}_{r}(w)|$ values for each $w\in V$ that is a proposer or receiver. As before, each call of Algorithm 4 can be implemented as a local randomizer using the Adaptive Laplace Mechanism with sensitivity $\Delta=1$; the sensitivity is $1$ because on neighboring adjacency lists, $|W_{r}(w)|$ differs by at most $1$. In fact, given that each edge is selected with probability $p_{r}=(1+\eta)^{-r}$, we can amplify the privacy guarantee for this local randomizer in terms of $p$ using Lemma 3.18. For a given randomizer and parameter $r$, by Lemma 3.18, the randomizer is $\frac{2}{(1+\eta)^{r}}\cdot\frac{\varepsilon^{\prime}}{4}=\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}$-DP. Then, the sum of the privacy parameters of all calls to this second type of local randomizer on node $w$ is
$\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}=\frac{\varepsilon^{\prime}(\eta n+n-1)}{2\eta\cdot n}\leq\frac{\varepsilon^{\prime}\cdot(\eta+1)}{2\eta}=\frac{3\varepsilon^{\prime}}{2}.$
The last local randomizer implementation is for estimating $|\widetilde{M}(w)|$ in Algorithm 4; as before it is an instance of the Adaptive Laplace Mechanism with sensitivity $1$. Given two neighboring adjacency lists and conditioning on all previous local randomizer outputs, the number of nodes matched to $w$ differs by $1$. Hence, each call of Algorithm 4 can be implemented using a $\frac{\varepsilon^{\prime}}{2}$-local randomizer by Lemma 3.15.
Finally, applying the concurrent composition theorem (Lemma 3.17) proves the privacy guarantee of the entire algorithm. All calls to our MAT local randomizers is $\varepsilon^{\prime\prime}$-DP which simplifies to $\varepsilon/3$-DP by our setting of $\varepsilon^{\prime\prime}$. Then, all calls to our second type of local randomizers incur a privacy loss of at most
$2\cdot\frac{3\varepsilon^{\prime}}{2}\cdot 512c\log_{16/15}(n)=3\cdot\frac{\varepsilon}{3072\log_{16/15}(n)}\cdot 512c\log_{16/15}(n)=\frac{\varepsilon}{2}\,.$
All calls to our third type of local randomizer incur a privacy loss of at most
$2\cdot\frac{\varepsilon^{\prime}}{2}\cdot 512c\log_{1+\eta}(n)=2\cdot\frac{\varepsilon}{2\cdot 3072c\log_{16/15}(n)}\cdot 512c\log_{1+\eta}(n)=\frac{\varepsilon}{6}\,.$
Thus, all calls to all local randomizers give $\frac{\varepsilon}{3}+\frac{\varepsilon}{6}+\frac{\varepsilon}{2}=\varepsilon$-DP. Hence, our algorithm is $\varepsilon$-LEDP since we can implement our algorithm using local randomizers and our produced transcript preserves $\varepsilon$-differential privacy. ∎
6.2.2 Utility and Number of Rounds
The goal of this section is to prove the following lemma, which gives the formal approximation guarantees and round complexity of Algorithm 3.
Lemma 6.4.
For any $\varepsilon\in(0,1)$, Algorithm 3 returns a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum $b^{\prime}$-matching, with high probability, when $b$ satisfies the lower bound in Theorem 2.3, in $O(\log(n))$ rounds.
Towards proving Lemma 6.4, we first prove the following lemma. Recall that our algorithm randomly divides the population into (roughly) half proposers and half receivers. Then, the proposers first propose their proposal sets, and then the receivers respond with their match sets for each of $32c\cdot\lceil\log_{16/15}(n)\rceil$ rounds. The following crucial lemma gives the “progress” of the algorithm after each round. Namely, to ensure we produce a maximal $b^{\prime}$-matching in $O(\log n)$ rounds, with high probability, roughly a constant fraction of the nodes would need to exceed their $b^{\prime}$-matching condition every round; hence, these nodes would not participate in future matching rounds. For convenience in the below proofs, when we write $\log(n)$, we mean $\log_{1+\eta}(n)$.
In the analysis below, we call the set of nodes that have not exceeded their $b^{\prime}$-matching condition and have a non-zero number of neighbors who have not reached their $b^{\prime}$-matching condition, the hopeful nodes. A node that is not hopeful is called unhopeful.
First, note that every active node in any round $i$ is either a proposer or receiver by definition. We first prove the following lemma, which states that a vertex $w$ will choose $r_{w}=0$ with high probability, as a proposer or receiver if its degree is less than $27c\log(n)/\varepsilon^{\prime}$.
Lemma 6.5.
If a node $w\in V_{i}$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$ among the hopeful nodes in round $i$, then it will choose $r_{w}=0$ with probability at least $1-\frac{1}{n^{5c}}$.
Proof.
Suppose $X\sim\text{Lap}(4/\varepsilon^{\prime})$ and $Y\sim\text{Lap}(2/\varepsilon^{\prime})$ are the noise random variables chosen in Algorithms 4 and 4, respectively. Since $w$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$, it will pick $r_{w}=0$ if $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ by our setting of $b$ in Algorithm 3 of Algorithm 3 and Algorithm 4 of Algorithm 4. By Lemma A.1, $X\leq 24c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$ and $Y\leq 12c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$. Then, $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{5c}}$ when $n\geq 2$. ∎
We now prove an additional lemma about the size of any proposal or match set an active vertex in round $i$ will choose. In particular, if a vertex is active in round $i$, then it will propose a proposal set or match set of size at least $32c\log(n)/\varepsilon^{\prime}$, with high probability.
Lemma 6.6.
If vertex $w\in V_{i}$ is active in round $i$, then $w$ will pick $r_{w}=0$ or an $r_{w}$ where $|W_{r_{w}}(w)|\geq 20c\log(n)/\varepsilon^{\prime}$ with probability at least $1-O\left(\frac{1}{n^{2c}}\right)$.
Proof.
If $w$ is still active in round $i$, this means that $w$ did not meet its $b^{\prime}$-matching condition in round i1i-1. Hence, by Algorithm 3 of Algorithm 3 and Lemma A.1, it holds that $|M(w)|<b-259c\log(n)/\varepsilon^{\prime}+\text{Lap}(4/\varepsilon^{\prime\prime})-\text{Lap}(8/\varepsilon^{\prime\prime})$. By Lemma A.1 and the setting of $\varepsilon^{\prime\prime}$, we can bound $b-439c\log(n)/\varepsilon^{\prime}\leq|M(w)|\leq b-79c\log(n)/\varepsilon^{\prime}$ with probability at least $1-O\left(\frac{1}{n^{2c}}\right)$. Since our proposal/match selection guarantees $|\widetilde{M}(w)|+|\widetilde{W}_{r}(w)|+27c\log(n)/\varepsilon^{\prime}\leq b$, there is sufficient slack to choose an index $r$ with $\mathbb{E}[|W_{r}(w)|]=40c\log(n)/\varepsilon^{\prime}$. By a Chernoff bound and independence of the public coins, with probability at least $1-\frac{1}{n^{5c}}$ we have $|W_{r_{w}}(w)|\geq 20c\log(n)/\varepsilon^{\prime}$, or else $r_{w}=0$ if the expectation is smaller (in which case all available edges are taken). ∎
Using our above lemmas, we will now prove our main lemma that enough progress is made in each round of our algorithm such that we find a maximal matching at the end of all of our rounds. Specifically, this lemma shows that with a large enough constant probability, a constant fraction of the remaining edges in the graph become unhopeful. An edge is unhopeful if at least one of its endpoints is unhopeful.
Lemma 6.7.
In each round $i$ of Algorithm 3 with $H$ hopeful edges, at least $H/16$ edges become unhopeful with probability at least $1/16$.
Proof.
Let $H$ be the set of hopeful edges in round $i$ where a hopeful edge is one where both endpoints of the edge are hopeful. We show that in each of $32c\log_{16/15}(n)$ rounds, a constant fraction of hopeful edges become unhopeful with high constant probability.
By definition of a hopeful edge, $e=\{u,v\}$, if either $u$ or $v$ becomes a proposer then they will propose a non-empty set (if they have at least one receiver neighbor). The marginal probability that any hopeful edge has one endpoint become the proposer and the other remains a receiver is $1/4$ by Algorithm 3. We use an orientation charging scheme as follows: in the induced graph on the endpoints of $H$, orient all hopeful edges from lower to higher degree. A vertex is good if at least $1/3$ of its incident hopeful edges are oriented into it; an oriented edge into a good vertex is good. A standard charging shows at least $H/2$ edges are good.
Fix a good, hopeful edge $e=(u,v)$ oriented from $u$ to $v$. The marginal probability that $u$ is a proposer and $v$ is a receiver is $1/4$. By Lemma 6.6, the marginal probability that $u$ proposes $v$ is at least $\min\!\left(1,\frac{20c\log(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$. The expected number of proposals received by $v$ is therefore at least $\min\!\left(\deg(v),\frac{20c\log(n)}{4\varepsilon^{\prime}}\right)$, so by a Chernoff bound, with probability at least $1-\frac{1}{n^{2c}}$ the receiver $v$ either accepts one of these proposals or reaches its $b^{\prime}$-matching condition in this round. Hence, with probability at least $1/4\cdot(1-1/n^{2c})$, a good, hopeful edge becomes unhopeful in this round. Taking expectations and applying Markov’s inequality as in the standard analysis yields the claim that at least $H/16$ edges become unhopeful with probability at least $1/16$ (for $n$ large enough). ∎
Using all of the above lemmas, we prove Lemma 6.4, which states the final utility guarantee of our algorithm: It is guaranteed to return a maximal matching in $O(\log n)$ rounds, with high probability.
Proof of Lemma 6.4.
By Lemma 6.7, in each round $i$, at least $1/16$ of the hopeful edges become unhopeful with probability at least $1/16$. In $32\cdot 16c\log(n)$ rounds, the expected number of rounds for which at least $1/16$ of the hopeful edges become unhopeful is $32c\log(n)$ rounds. By a Chernoff bound, the probability that fewer than $16c\log(n)$ rounds are successful is at most $n^{-4c}$. Hence, in $O(\log n)$ rounds, all edges become unhopeful.
We now show that if all edges become unhopeful, then all nodes are either matched to at least $b^{\prime}$ neighbors or all of their neighbors are matched to at least $b^{\prime}$ neighbors. An edge becomes unhopeful if at least one of its endpoints satisfies the $b^{\prime}$-matching condition. By Lemma A.1, a node satisfies the matching condition if and only if $|M(u)|\geq b^{\prime}$ with probability at least $1-\frac{1}{n^{2c}}$. Hence, if a node is adjacent to all unhopeful edges, then it either satisfies the matching condition and is matched to at least $b^{\prime}$ neighbors with high probability, or it is adjacent to endpoints which are matched with at least $b^{\prime}$ neighbors with high probability.
Finally, by Lemma A.1, $|M(u)|\geq b$ with probability at most $1/n^{2c}$, and this is ruled out by the choice of $r$ in Algorithm 4. By Lemma 6.6 and a union bound over all failure events, with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$, we output a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum $b^{\prime}$-matching in $O(\log n)$ rounds provided that $b$ satisfies the lower bound in Theorem 2.3.
The number of rounds of our algorithm is determined by Algorithm 3 (which is $O(\log n)$) since each round contains a constant number of synchronization points and all nodes (proposers and receivers) perform their instructions simultaneously. ∎
Combining Lemma 6.3 and Lemma 6.4 yields the proof of Theorem 2.3.
6.3 Corollaries
By choosing a fixed constant $\eta=\nicefrac{{1}}{{2}}$, benchmark parameter $b^{\prime}=1$, and matching degree $b=\Omega\left(\frac{\log^{2}(n)}{\varepsilon}\right)$, Theorem 2.3 immediately implies Corollary 6.1
In order to obtain a non-bicriteria approximation when $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{5}\varepsilon}\right)$, we use the same technique as the sequential edge-DP algorithm, but this time we execute the $O(\log n)$-round algorithm (Theorem 2.3) with $b^{\prime}=\frac{b}{1+O(\eta)}$ instead of the sequential algorithm. The proof is identical to that of Corollary 5.7. This yields a proof of Theorem 6.2.
7 Node Differential Privacy via Arboricity Sparsification
Node-DP Matchings.
Using sparsification techniques, we demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. In particular, we prove Theorem 2.6, which we restate below for convenience. See 2.6
As a special case, we have the following corollary
Corollary 7.1 (Node-DP Approximate Maximum Matching).
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm that, with high probability, outputs an (implicit) matching with degree at most $b$ in the billboard model for $b=O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+\frac{\log^{2}(n)}{\eta\varepsilon^{2}}\right)$ that has the size of a $(2+\eta)$-approximate maximum matching.
Node-DP Vertex Cover.
To demonstrate the applicability of our techniques, we also design an implicit node-DP vertex cover algorithm using arboricity sparsification. Its guarantees are summarized in Theorem 2.7, which we repeat below. See 2.7
The class of bounded arboricity [^8] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree n1n-1 and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. For simplicity of presentation, we first design an algorithm assuming a public bound $\tilde{\alpha}$. Then our algorithm is always private but the approximation guarantees hold when $\tilde{\alpha}$ upper bounds the arboricity of the input graph. When such a public bound is unavailable, we show that privately estimating $\alpha$ yields the same guarantees up to logarithmic factors.
The key idea is that a judicious choice of a sparsification algorithm reduces the edge edit distance between node-neighboring graphs to some factor $\Lambda=O(\alpha)$. Such sparsification is useful since, in the worst case, the edge edit distance pre-sparsification can be $\Omega(n)$. By group privacy, it then suffices to run any edge-DP algorithm with privacy parameter $\nicefrac{{\varepsilon}}{{\Lambda}}$ after sparsification to achieve node-privacy.
Interestingly, in Section 7.4, we show that actually releasing such a sparsifier is impossible under privacy constraints (Theorem 2.5). However, we can still use them as subroutines in our algorithms!
Organization.
The rest of the section is organized as follows. In Section 7.1, we tackle node-DP matchings as a warm-up before addressing general $b$-matching. This entails adapting a matching sparsifier to the DP setting in Section 7.1.1, allowing us to reduce node-DP matching to edge-DP matching for bounded arboricity graphs. We treat the more general case of $b$-matchings in Section 7.2. This requires extending our matching sparsifier to a $b$-matching sparsifier in Section 7.2.1, which may be of independent interest beyond our work. To further illustrate the power of arboricity sparsification, we further design a node-DP vertex cover algorithm in Section 7.3. This includes a vertex cover sparsifier in Section 7.3.1, which allows us to reduce node-DP vertex cover to edge-DP vertex cover, similar to matchings. Finally, we present lower bounds for releasing our sparsifiers in Section 7.4.
7.1 Node-DP Implicit Matching
In this section, we design a node-DP matching algorithm via arboricity sparsification.
7.1.1 Bounded Arboricity Matching Sparsifier
For our bounded arboricity graphs, we take inspiration from the bounded arboricity sparsifier of [solomon2021local]. A closely related line of work is that of edge degree constrained subgraphs (EDCS) [assadi2019coresets, bernstein2015fully, bernstein2016faster]. We modify the sparsifier from [solomon2021local] to show Proposition 7.2, which states that node-neighboring graphs have small edge edit distance post-sparsification.
Matching Sparsifier.
Our sparsification algorithm $\textsc{ContractionSparsify}_{\pi}$ proceeds as follows. Given an ordering $\pi\in P_{\binom{n}{2}}$ over unordered vertex pairs, a graph $G$, and a degree threshold $\Lambda$, each vertex $v$ marks the first $\min(\deg_{G}(v),\Lambda)$ incident edges with respect to $\pi$. Then, $H$ is obtained from $G$ by taking all vertices of $G$ as well as edges that were marked by both endpoints. In the central model, we can take $\pi$ to be the lexicographic ordering of edges $\{u,v\}$. Thus, we omit the subscript $\pi$ in the below analyses with the understanding that there is a fixed underlying ordering.
Proposition 7.2.
Let $\pi\in P_{\binom{n}{2}}$ be a total ordering over unordered vertex pairs and $G\sim G^{\prime}$ be node neighboring graphs. Then the edge edit distance between $H\coloneqq\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ and $H^{\prime}\coloneqq\textsc{ContractionSparsify}_{\pi}(G^{\prime},\Lambda)$ is at most $2\Lambda$.
Proof.
Let $S_{G}$ and $S_{G^{\prime}}$ be the sparsified graphs of $G$ and $G^{\prime}$, respectively. Suppose without loss of generality that $G^{\prime}$ contains $E_{\operatorname{extra}}$ additional edges incident to vertex $v$ and $v$ has degree 0 in graph $G$. Then, for each edge $\{v,w\}\in E_{\operatorname{extra}}$, let $e^{w}_{\operatorname{last}}$ be the edge adjacent to $w$ in $G$ whose index in $\pi$ is the last among the edges incident to $w$ in $S_{G}$. If $i_{\pi}(e^{w}_{\operatorname{last}})>i_{\pi}(\{v,w\})$ (where $i_{\pi}(\{v,w\})$ is the index of edge $\{v,w\}$ in $\pi$), then $\{v,w\}$ replaces edge $e^{w}_{\operatorname{last}}$. Since both $G$ and $G^{\prime}$ are simple graphs, at most one edge incident to $w$ gets replaced by an edge in $E_{\operatorname{extra}}$ in $S_{G^{\prime}}$. This set of edge replacements leads to an edge edit distance of $2\Lambda$. ∎
The original sparsification algorithm in [solomon2021local] marks an arbitrary set of $\min(\deg(v),\Lambda)$ edges incident to each vertex $v$ and takes the subgraph consisting of all edges marked by both endpoints. In our setting, $\pi$ determines the arbitrary marking in our graphs. Hence, our sparsification procedure satisfies the below guarantee.
Theorem 7.3 (Theorem 3.3 in [solomon2021local]).
Let $G$ be a graph of arboricity at most $\alpha$ and $\Lambda\coloneqq 5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha$ for some $\eta\in(0,1]$. Suppose $H$ is obtained by marking an arbitrary set of $\min(\deg(v),\Lambda)$ edges incident to each vertex $v$ and taking the subgraph consisting of all edges marked by both endpoints. Then if $\mu(\cdot)$ denotes the size of a maximum matching of the input graph,
$\mu(H)\leq\mu(G)\leq(1+\eta)\mu(H).$
In particular, any $(\beta,\zeta)$-approximate maximum matching for $H$ is an $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$. Moreover, this holds for $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ where $\pi\in P_{\binom{n}{2}}$ is a total order over unordered vertex pairs.
7.1.2 Algorithm
In this section, we design a node-DP algorithm to output an implicit matching in the central model assuming a given public bound $\tilde{\alpha}$ on the arboricity $\alpha$ using the sparsification techniques derived in Section 7.1.1 and any edge DP algorithm (e.g. Theorem 2.2) that outputs implicit solutions in the $\varepsilon$-DP setting.
Theorem 7.4.
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, there is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log(n)}{\eta\varepsilon}\right)$and (ii) if $\tilde{\alpha}\geq\alpha$, the implicit solution is at least $\frac{1}{2+\eta}$ the size of a maximum matching.
We emphasize that our privacy guarantees always hold but the utility guarantee is dependent on the public bound $\tilde{\alpha}$.
Proof Sketch.
The algorithm first discards edges according to Theorem 7.3 until there are at most $\Lambda=O(\tilde{\alpha}/\eta)$ edges incident to each vertex. This ensure that the edge-edit distance of node-neighboring graphs is at most $2\Lambda$ (Proposition 7.2). We can thus run our $(\frac{\varepsilon}{2\Lambda})$-edge DP Algorithm 2 (Theorem 2.2) to ensure $\varepsilon$-node DP. ∎
Public Arboricity Bound.
In practice, it is not always possible to obtain a good public bound on the arboricity. However, we can always compute an $\varepsilon$-node DP estimate of the arboricity up to $O(\log(n)/\varepsilon)$ additive error since the sensitivity of the arboricity is $1$ even for node-neighboring graphs. This incurs a slight blow-up in $b$ and accounting for this blowup leads to a proof of Corollary 7.1.
7.2 Node-DP Implicit $b$-Matching (Theorem 2.6)
In this section, we design a node-DP $b$-matching algorithm via arboricity sparsification.
7.2.1 Bounded Arboricity $b$-Matching Sparsifier
In order to extend the $b$-matching guarantees to the node-DP setting, we prove an extension of Solomon’s sparsifier (Theorem 7.3) to the case of $b$-matchings with an $O(b)$ additive increase in the degree threshold. This may be of independent interest beyond node differentially private matchings.
$b$-Matching Sparsifier.
Our $b$-Matching sparsifier is identical to our matching sparsifier given in Section 7.1.1, with the exception that the degree threshold $\Lambda$ is adjusted as below.
Theorem 7.5.
Let $G$ be a graph of arboricity at most $\alpha$ and $\Lambda_{b}\coloneqq 5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha+(b-1)$ for some $\eta\in(0,1]$. Suppose $H$ is obtained by marking an arbitrary set of $\min(\deg(v),\Lambda_{b})$ edges incident to each vertex $v$ and taking the subgraph consisting of all edges marked by both endpoints. Then if $\mu_{b}(\cdot)$ denotes the size of a maximum $b$-matching of the input graph,
$\mu_{b}(H)\leq\mu_{b}(G)\leq(1+\eta)\mu_{b}(H).$
In particular, any $(\beta,\zeta)$-approximate maximum $b$-matching for $H$ is an $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$. Moreover, this holds for $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda_{b})$ where $\pi\in P_{\binom{n}{2}}$ is a total order over unordered vertex pairs.
Let $\Lambda_{b}$ be as in Theorem 7.5, $T(v)\subseteq V$ denote some subset of neighbors of $v$ with size at least $\min(\deg_{G}(v),\Lambda_{b})$ for each $v\in V$, and $E_{H}\coloneqq\{uv\in E(G):u\in T(v),v\in T(u)\}$. Then [solomon2021local] showed that $E_{H}$ contains a large $b$-matching for the case of $b=1$. We extend the result to arbitrary $1\leq b\leq n$.
Proof.
We argue by induction on $b$. The base case of $b=1$ is guaranteed by Theorem 7.3. Suppose inductively the statement holds for b1b-1 for some $b\geq 2$ so that our goal is to extend it to the case $b$.
Let $M_{1}\cup\dots M_{b}$ be any $b$-matching of $G$ decomposed into $b$ disjoint matchings. Let $H$ denote a candidate sparsifier obtained by marking an arbitrary set $T(v)$ of at least $\min(\deg_{G}(v),\Lambda_{b})$ edges incident to each vertex $v$, and taking the subgraph $H=(V,E_{H})$ consisting of all edges marked by both endpoints, i.e. $E_{H}\coloneqq\{uv\in E(G):u\in T(v),v\in T(u)\}$.
Consider $G_{1}\coloneqq G-(M_{2}\cup\dots\cup M_{b})$ and note that $M_{1}$ is a matching in $G_{1}$. We claim that $E_{H}\cap E(G_{1})$ contains a matching of size at least $\frac{1}{1+\eta}|M_{1}|$. Indeed, let
$\forall v\in V,T_{1}(v)\coloneqq\{u\in T(v):uv\notin M_{2}\cup\dots\cup M_{b}\},\qquad E_{1}\coloneqq\{uv\in E:u\in T_{1}(v),v\in T_{1}(u)\}\,.$
Note that each $T_{1}(v)$ contains at least $\min(\deg_{G_{1}}(v),\Lambda_{1})$ neighbors of $v$ within $G_{1}$. Furthermore, we have $E_{1}\subseteq E_{H}$. Thus by the base case, $E_{1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum matching $M_{1}^{\prime}$ of $G_{1}$.
Remark that $M_{1}^{\prime}\cup M_{2}\cup\dots\cup M_{b}$ is a $b$-matching within $G$ by construction. Consider $G_{-1}\coloneqq G-M_{1}^{\prime}$, so that $M_{2}\cup\dots\cup M_{b}$ is a (b1)(b-1)-matching within $G_{-1}$. We claim that $E_{H}\cap E(G_{-1})$ contains a (b1)(b-1)-matching of size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. To prove this claim, we will rely on the induction hypothesis. Define
$\forall v\in V,T_{-1}(v)\coloneqq\{u\in T(v):uv\notin M_{1}^{\prime}\},\qquad E_{-1}\coloneqq\{uv\in E:u\in T_{-1}(v),v\in T_{-1}(u)\}\,.$
Then each $T_{-1}(v)$ contains at least $\min(d_{G_{-1}}(v),\Lambda_{b-1})$ neighbors of $v$ in $G_{-1}$. Furthermore, $E_{-1}\subseteq(E_{H}\setminus M_{1}^{\prime})$ by construction. Thus by the induction hypothesis, $E_{-1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum (b1)(b-1)-matching NN of $G_{-1}$.
We have identified disjoint edge subsets $M_{1}^{\prime},N\subseteq E_{H}$ such that $M_{1}^{\prime}$ is a matching in $G$ with size at least $\frac{1}{1+\eta}|M_{1}|$ and $N$ is a (b1)(b-1)-matching in $G-M_{1}^{\prime}$ with size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. Thus $M_{1}^{\prime}\cup N$ is a $b$-matching in $G$ with size at least
$|M_{1}^{\prime}\cup N|=|M_{1}^{\prime}|+|N|\geq\frac{1}{1+\eta}|M_{1}\cup\dots\cup M_{b}|\,.$
By the arbitrary choice of $M_{1}\cup\dots\cup M_{b}$, we conclude the proof. ∎
7.2.2 Algorithm
Similar to Section 7, we can implement our implicit $b$-matching algorithm in node-DP setting with the help of the $b$-matching arboricity sparsifier we designed in Section 7.2.1. Then, repeating the steps in Section 7.1 yields a proof of Theorem 2.6.
7.3 Node-DP Implicit Vertex Cover (Theorem 2.7)
In this section, we design a node-DP vertex cover algorithm via arboricity sparsification.
7.3.1 Bounded Arboricity Vertex Cover Sparsifier
Towards proving our node-DP vertex cover algorithm (Theorem 2.7), we describe the arboricity sparsifier for vertex cover.
Let $G_{\leq}^{\Lambda}$ denote the induced subgraph on all vertices with degree less than or equal to some bound $\Lambda$ (which we will set) and $V_{>}^{\Lambda}$ denotes the set of vertices with degree greater than $\Lambda$. [solomon2021local] showed that the union of any $\beta$-approximate vertex cover for $G_{\leq}^{\Lambda}$ and the set of vertices $V_{>}^{\Lambda}$ yields a $(\beta+\eta)$-approximate vertex cover for $G$ for any constant $\eta>0$. We extend the ideas of [solomon2021local] to more general sparsifiers for vertex cover. Our argument holds in greater generality at the cost of an additional additive error of $\operatorname{OPT}(G)$ in the approximation guarantees.
Vertex Cover Sparsifier.
Our vertex cover sparsifier is identical to our matching sparsifier given in Section 7.1.1. Specifically, each vertex keeps the first $\min(\deg_{G}(v),\Lambda)$ incident edges with respect to some fixed ordering $\pi$.
Theorem 7.6.
Let $\eta>0$, $G$ be a graph of arboricity at most $\alpha$, and set
$\Lambda:=(1+\nicefrac{{1}}{{\eta}})\cdot 2\alpha.$
Let $H\preceq G$ be any subgraph of $G$ such that $G_{\leq}^{\Lambda}\preceq H\preceq G$. Here $\preceq$ denotes the subgraph relation. If $C$ is a $(\beta,\zeta)$-approximate vertex cover of $H$, then $C\cup V_{>}^{\Lambda}$ is a cover of $G$ with cardinality at most
$(1+\beta+\eta)\operatorname{OPT}(G)+\zeta.$
Lemma 7.7 (Lemma 3.8 in [solomon2021local]).
Let $\eta>0$ and $G$ be a graph of arboricity at most $\alpha$ and set $\Lambda:=(1+\nicefrac{{1}}{{\eta}})\cdot 2\alpha$. If $C$ is any vertex cover for $G$, and $U_{>}:=V_{>}\setminus C$, then $|U_{>}|\leq\eta|C|$.
Remark 7.8.
Lemma 3.8 in [solomon2021local] holds as long as any induced subgraph of $G$ has average degree at most $2\alpha$, which is certainly the case for graphs of arboricity at most $\alpha$. Moreover, [solomon2021local] stated their result for
$V_{\geq}=\{v\in V:\deg(v)\geq\Lambda\},\qquad U_{\geq}=V_{\geq}\setminus C.$
But since $U_{>}\subseteq U_{\geq}$, the result clearly still holds. In order to remain consistent with our matching sparsifier, we state their result with $V_{>}$ and $U_{>}$.
Proof of Theorem 7.6.
Any edge in $H$ is covered by $C$, which is a $(\beta,\zeta)$-approximate vertex cover of $H$, and any edge that is not in $H$ is covered by $V_{>}^{\Lambda}$. Thus $C\cup V_{>}^{\Lambda}$ is indeed a vertex cover. By Lemma 7.7, $|V_{>}^{\Lambda}\setminus C^{*}|\leq\eta|C^{*}|$ where $C^{*}$ is a minimum vertex cover of $G$. Then,
$|C\cup V_{>}^{\Lambda}|$ $\leq|C|+|V_{>}^{\Lambda}|$ (1)
$\leq\beta\operatorname{OPT}(H)+\zeta+|V_{>}^{\Lambda}\cap C^{*}|+|V_{>}^{\Lambda}\setminus C^{*}|$ (2)
$\leq\beta\operatorname{OPT}(G)+\zeta+\operatorname{OPT}(G)+\eta\operatorname{OPT}(G)$ (3)
$=(1+\beta+\eta)\operatorname{OPT}(G)+\zeta.$ (4)
Eq. 3 follows by Lemma 7.7 and since $|V_{>}^{\Lambda}\cap C^{*}|\leq\operatorname{OPT}(G)$ due to $C^{*}$ being an optimal cover for $G$. ∎
Let $H$ denote our vertex cover sparsifier. Remark that we are guaranteed to have $G_{\leq}^{\Lambda}\subseteq H$ since if an edge $\{u,w\}$ has $\deg(u)\leq\Lambda$ and $\deg(w)\leq\Lambda$, then the edge is guaranteed to be in the first $\Lambda$ edges of the ordering $\pi$ for both of its endpoints.
7.3.2 Algorithm
We now present our node-DP vertex cover algorithm that releases an implicit vertex cover (Definition 7.11) using the sparsification techniques derived in Section 7.3.1. As a subroutine, we use the static edge DP implicit vertex cover algorithm (Theorem 7.10) of [GuptaLMRT10]. Similar to our node-DP matching algorithm, we first describe an algorithm assuming a public bound $\tilde{\alpha}$ on the arboricity $\alpha$.
Theorem 7.9.
Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, there is an $\varepsilon$-node DP algorithm that outputs an implicit vertex cover (Definition 7.11). Moreover, if $\tilde{\alpha}\geq\alpha$, the implicit solution is a $O(\nicefrac{{\tilde{\alpha}}}{{\varepsilon}})$-approximation with probability $0.99$.
Before proving Theorem 7.9, we first define the precise implicit solution released by our algorithm.
Value vs Solution.
The node-DP algorithm to estimate the value of a maximum matching immediately yields a node-DP algorithm for estimating the size of a minimum vertex cover. In fact, it suffices to compute the size of a greedy maximal matching, which has a coupled global sensitivity of 1 with respect to node neighboring graphs. This is then a $2$-approximation for the size of a minimum vertex cover. However, it may be desirable to output actual solutions to the vertex cover problem.
[GuptaLMRT10] demonstrated how to output an implicit solution through an ordering of the vertices. Specifically, the theorem below returns an implicit solution where every edge is considered an entity and can determine, using the released implicit solution, which of its endpoints covers it.
Theorem 7.10 (Theorem 5.1, Theorem 5.2 in [GuptaLMRT10]).
Fix $\beta\in(0,1)$. There is a polynomial time $\varepsilon$-edge DP algorithm EdgePrivateVC that outputs an implicit vertex cover with expected cardinality at most
$\left(2+\frac{16}{\varepsilon}\right)\operatorname{OPT}.$
By Markov’s inequality, running EdgePrivateVC guarantees a $O(\nicefrac{{1}}{{\varepsilon}})$-approximate solution with probability $0.999$.
Implicit Vertex Cover.
The implicit vertex cover from [GuptaLMRT10] is an ordering of the vertices where each edge is covered by the earlier vertex in the ordering. Such an implicit ordering allows for each edge to determine which of its endpoints covers it given the public ordering.
In our setting, if we run EdgePrivateVC on a sparsified graph $H$, the ordering we output is with respect to $H$ and some edges that were deleted from $G$ to obtain $H$ may not be covered by the intended solution. Thus we also output the threshold we input into the sparsifier so that an edge is first covered by a high-degree endpoint if it has one, and otherwise by the earlier vertex in the ordering.
Definition 7.11 (Implicit Vertex Cover).
An implicit solution is a pair
$\left(\pi,\Lambda\right)\in S_{n}\times\mathbb{Z}_{\geq 0}$
where $\pi$ is an ordering of the vertices and $\Lambda$ is the threshold used to perform the sparsification. Every edge is first covered by an endpoint with degree strictly greater than $\Lambda$ if it exists, and otherwise by the endpoint which appears earlier in $\pi$.
Remark 7.12 (Ordering-Only Implicit Vertex Cover).
We assume each edge is an entity that knows its endpoints as well as their (private) degrees. Ideally, we would utilize the exact same solution concept as [GuptaLMRT10], i.e., output an ordering only, which does not require each edge to have access to the degrees of its endpoints for decoding. However, we prove that any $(\varepsilon,\delta)$-node DP algorithm which outputs a [GuptaLMRT10] implicit solution necessarily incurs large $\Omega(n)$-error (Section 7.5).
With the precise notion of an implicit vertex cover in hand, we are now ready to prove Theorem 7.9.
Proof Sketch.
[…]
Public Arboricity Bound.
We can always compute an $\varepsilon$-node DP estimate of the arboricity up to $O(\nicefrac{{1}}{{\varepsilon}})$ additive error with probability $0.999$ since the sensitivity of the arboricity is $1$ even for node-neighboring graphs. This incurs a constant factor in the approximation ratio and leads to a proof of Theorem 2.7.
7.4 Lower Bound for Releasing Sparsifiers
In this section we show a strong lower bound against computing matching sparsifiers in the node DP setting. It is obviously impossible to release such a sparsifier explicitly under any reasonable privacy constraint, but we will show that node differential privacy rules out even releasing an implicit sparsifier. This is particularly interesting since despite this lower bound, our node DP algorithm makes crucial use of a matching sparsifier which it constructs. So our node DP algorithm, combined with this lower bound, shows that it is possible to create and use a matching sparsifier for node DP algorithms even though the sparsifier itself cannot be released without destroying privacy. This is the fundamental reason why we cannot design a local algorithm for matchings in the node DP setting; the obvious local version of our algorithm would by definition include the construction of the sparsifier in the transcript, and thus this lower bound shows that it cannot be both DP and have high utility.
Recall that a matching sparsifier of a graph $G=(V,E)$ is a subgraph $H$ of $G$ with the properties that a) the maximum degree of $H$ is small, and b) the maximum matching $H$ has size close to the maximum matching in $G$. Without privacy, Solomon [solomon2021local] showed that for any $\eta\in(0,1]$, it is possible to find an $H$ with maximum degree at most $O(\frac{1}{\eta}\alpha)$ (where $\alpha$ is the arboricity of $G$) and $\nu(G)\leq(1+\eta)\nu(H)$ (where recall that $\nu(\cdot)$ denotes the size of the maximum matching in the graph).
Since we obviously cannot release a sparsifier explicitly under any reasonable privacy constraint, we turn to implicit solutions. Formally, given a graph $G=(V,E)$, we want an $(\varepsilon,\delta)$-node DP algorithm which outputs an implicit matching sparsifier: a set of edges $E^{\prime}\subseteq\binom{V}{2}$. This set $E^{\prime}$ gives us the “true” matching sparsifier $E^{\prime}\cap E$. We claim that under node privacy, $E^{\prime}\cap E$ cannot have both small maximum degree and be a good approximation of the maximum matching.
See 2.5
Proof.
Fix $V$ with $|V|=n$, and let $r\in V$. For any $S\subseteq V\setminus\{r\}$, let $G_{S}$ be the graph whose edge set consists of an edge from $r$ to all nodes in $S$ (i.e., a star from $r$ to $S$ and all other nodes having degree 0). Note that all such graphs have arboricity $1$. A similar symmetry argument to the other lower bounds implies that without loss of generality, we may assume that the probability that $\mathcal{A}$ includes some edge $\{r,v\}$ is the same for all $v\in S$ (we will denote this probability by $p_{S}$).
Fix some node $v\in S\setminus\{r\}$, and consider what happens when we run $\mathcal{A}$ on the graph $G_{\{v\}}$. Since $\mathbb{E}[\nu(E^{\prime}\cap E)]\geq\frac{\nu(G_{\{v\}})}{1+\eta}$ by assumption, we know that $p_{\{v\}}=\Pr[\{r,v\}\in E^{\prime}]\geq\frac{1}{1+\eta}$.
Now consider the graph $G_{V\setminus\{r\}}$. Since $G_{\{v\}}$ and $G_{V\setminus\{r\}}$ are neighboring graphs (under node-differential privacy, since we simply changed the edges incident on $r$) and $\mathcal{A}$ is $(\varepsilon,\delta)$-node DP, we know that when we run $\mathcal{A}$ on $G_{V\setminus\{r\}}$ it must be the case that $\Pr[\{r,v\}\in E^{\prime}]\geq e^{-\varepsilon}p_{\{v\}}-\delta$. But now symmetry implies that this is true for all $v\in V\setminus\{r\}$, so we have that
$p_{V\setminus\{r\}}$ $\geq e^{-\varepsilon}p_{\{v\}}-\delta\geq\frac{1}{e^{\varepsilon}(1+\eta)}-\delta.$
Linearity of expectation then implies that the expected degree of $r$ in $E^{\prime}$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$. Since $E(G_{V\setminus\{r\}})$ consists of all edges from $r$ to all other nodes, this means that the expected maximum degree of $E^{\prime}\cap E$ is at least $\left(\frac{1}{e^{\varepsilon}(1+\eta)}-\delta\right)(n-1)$ as claimed. ∎
In particular, for the natural regime where $\varepsilon$ is a constant, $\delta\leq 1/n$, and the maximum matching approximation $(1+\eta)$ is a constant, Theorem 2.5 implies that the maximum degree must be $\Omega(n)$ rather than $O(1)$ for graphs of arboricity $1$.
7.5 Lower Bound for Ordering-Only Node-DP Implicit Vertex Cover
Let $G=(V,E)$ be a graph. As mentioned in Section 7.3, the implicit vertex cover used by [GuptaLMRT10] is an ordering. More formally, given an ordering $\pi$ of $V$, we let $G_{\pi}$ be the directed graph where for any edge $\{u,v\}\in E$ we direct it $(v,u)$ if $\pi(u)<\pi(v)$ and otherwise we direct it $(u,v)$. In other words, we direct it into the endpoint that is earlier in the ordering. Given $\pi$, let $S(\pi,G)=\{v\in V:\exists(u,v)\in E(G_{\pi})\}$. Then clearly $S(\pi,G)$ is a vertex cover of $G$, which we think of as the vertex cover defined implicitly by $\pi$.
Theorem 7.13.
Let $\mathcal{A}$ be an algorithm which satisfies $(\epsilon,\delta)$-node DP and outputs an ordering $\pi$ of $V$. Then the expected approximation ratio of $\mathcal{A}$ is at least $\frac{n}{4}e^{-\epsilon}\left(1-2\delta\right)$.
Proof.
Let $V$ be a set of $n$ nodes. For each $v\in V$, let $G_{v}$ be the star graph with vertex set $V$ with center node $v$. Let $G_{\emptyset}$ denote the empty graph with vertex set $V$. Note that $G_{\emptyset}$ and $G_{v}$ are neighboring graphs in the node DP setting for all $v\in V$.
Consider the distribution over orderings generated by running $\mathcal{A}$ on $G_{\emptyset}$. For every $v\in V$, let $p_{v}=\Pr_{\pi\sim\mathcal{A}(G_{\emptyset})}[\pi(v)\geq n/2]$ denote the probability that $v$ is in the second half of the nodes in the ordering. Then clearly $\sum_{v\in V}p_{v}=n/2$, since the expected number of nodes in the second half is exactly $n/2$ and is also (by linearity of expectations) equal to $\sum_{v\in V}p_{v}$. Thus there exists some node $v\in V$ such that $p_{v}\geq 1/2$.
Now consider $G_{v}$. Let $p^{\prime}_{v}=\Pr_{\pi\sim\mathcal{A}(G_{v})}[\pi(v)\leq n/2]$ be the same probability but when $\mathcal{A}$ is run on $G_{v}$. Then by node differential privacy, we get that $p_{v}\leq e^{\epsilon}p^{\prime}_{v}+\delta$, and hence $p^{\prime}_{v}\geq e^{-\epsilon}\left(p_{v}-\delta\right)\geq e^{-\epsilon}\left(\frac{1}{2}-\delta\right)$. Note that if $v$ is in the second half of $\pi$, then $|S(\pi,G_{v})|\geq n/2$, since every node in the first half of $\pi$ would be in $S(\pi,G_{v})$. Thus we have that
$\mathbb{E}_{\pi\sim\mathcal{A}(G_{v})}[|S(\pi,G_{v})|]$ $\geq\frac{n}{2}p^{\prime}_{v}\geq\frac{n}{2}e^{-\epsilon}\left(\frac{1}{2}-\delta\right)=\frac{n}{4}e^{-\epsilon}\left(1-2\delta\right).$
On the other hand, it is obvious that the optimal vertex cover in $G_{v}$ has size $1$ (it is just $\{v\}$). This implies the theorem. ∎
8 Improved Node-DP Implicit Bipartite Matchings
In this section, we restate and prove Theorem 2.8. See 2.8
8.1 Detailed Algorithm Description
As in [hsu2014private], recall that we have a bipartite graph $G=(V_{L}\cup V_{R},E)$, where we think of the left nodes in $V_{L}$ as items and the right nodes in $V_{R}$ as bidders. Let’s say there are $n=|V_{R}|$ bidders and $k=|V_{L}|$ items. We define two of these bipartite graphs to be neighbors if they differ by a single bidder and all edges incident to the bidder. Our goal is to implicitly output an allocation of items to bidders such that each item is allocated to at most $s$ bidders and each bidder is allocated at most $1$ item, while guaranteeing differential privacy for the output.
Our algorithm for node-private bipartite matching follows the same template as that of [hsu2014private], based on a deferred acceptance type algorithm from [kelso1982job]. For each item, [kelso1982job] runs an ascending price auction where at every iteration, each item $u$ has a price $p_{u}$. Then, over a sequence of $T$ rounds, the algorithm processes the bidders in some publicly known order and each one bids on their cheapest neighboring item which has price less than $1$ (each edge indicates a potential maximum utility value of $1$ for the item). At any moment, the $s$ most recent bidders for an item are tentatively matched to that item, and all earlier bidders for it become unmatched.
In the private implementation of this algorithm in [hsu2014private], they keep private continual counters for the number of bidders matched with each item which they continually output. They additionally output the sequence of prices for each item. Using this information, each bidder can reconstruct the cheapest neighboring item at the given prices. They then send the bit 1 to the appropriate counter, and store the reading of the counter when they made the bid. When the counter indicates that $s$ bids have occurred since their initial bid, the bidder knows that they have become unmatched. The final matching which is implicitly output is simply the set of edges which have not been unmatched.
The main difference in our algorithm is in implementing the counters. In the private algorithm of [hsu2014private], each of these counters need to be accurate at each of the $nT$ iterations. This causes significant privacy loss. Our algorithm has a different implementation (of the same basic algorithm), where we use the fact that only the counts at the end of each of the $T$ iterations are important. Moreover, our implementation does not need the counts themselves, but just needs to check whether they are above the supply $s$. This allows us to use the Multidimensional AboveThreshold Mechanism (Algorithm 1) to get further improvements.
1Input: Graph $G=(V,E)$, approximation factor $\eta\in(0,1)$, privacy parameter $\varepsilon>0$, matching parameter $b=\Omega(\log(n)/(\eta^{4}\varepsilon))$.
2 Output: An $\varepsilon$-node differentially private implicit $(1+\eta)$-approximate $b$-matching.
Initialization : $T\leftarrow 73/\eta^{2}$, $\varepsilon^{\prime}\leftarrow\varepsilon\eta/3T$, $\text{start}_{u}\leftarrow 0$, $\text{end}_{u}\leftarrow nT$, $p_{u}\leftarrow\eta$, and $c_{u}\leftarrow 0$ for each $u\in V_{L}$.
3 for $t\leftarrow 1$ to $T$ do
4 foreach item $u\in V_{L}$ do
5      Initialize noisy threshold $\tilde{t}_{u}\leftarrow(p_{u}/\eta)\cdot s+\mathrm{Lap}(2/\varepsilon^{\prime})$
6
 // Each bidder $v$ proposes to the neighbor $u$ with lowest price
7 foreach bidder $v\in V_{R}$ do
8    if $v$ is not matched then
9         Choose a neighbor $u$ of $v$ with smallest $p_{u}$, if any exist, breaking ties arbitrarily
10       if $p_{u}\leq 1$ then
11            Denote bidder $v$ as (temporarily) matched with item $u$ for timestep $t$
12          $c_{u}\leftarrow c_{u}+1$
13
14
15
16 foreach item $u\in V_{L}$ do
17    Release $p_{u}$
18    if $p_{u}\leq 1$ and $c_{u}+\mathrm{Lap}(4/\varepsilon^{\prime})\geq\tilde{t}_{u}$ then
19       $p_{u}\leftarrow p_{u}+\eta$
20         Re-initialize the noisy threshold $\tilde{t}_{u}\leftarrow(p_{u}/\eta)\cdot s+\mathrm{Lap}(2/\varepsilon^{\prime})$
21
22
 // Only the most recent (approximately) $s$ bidders for each item $u$ remain matched with $u$
23 foreach item $u\in V_{L}$ do
24    Release startu
25    $\tilde{s}_{u}\leftarrow s+\mathrm{Lap}(2/\varepsilon^{\prime})-18c\log(n)/\varepsilon^{\prime}$
26      Let $M_{u}$ denote the number of bidders matched with $u$, between timesteps $\text{start}_{u}$ and $\text{end}_{u}$
27    while $M_{u}+\mathrm{Lap}(4/\varepsilon^{\prime})\geq\tilde{s}_{u}$ do
28       if a bidder $v$ matched with item $u$ at timestep $\text{start}_{u}$ then
29            Mark bidder $v$ as unmatched
30
31       $\text{start}_{u}\leftarrow\text{start}_{u}+1$ and update $M_{u}$ accordingly
32
33
 // Terminate early if the algorithm is close to convergence
34   Let $c_{0}$ denote the number of bids made in this iteration
35 if $c_{0}+\mathrm{Lap}(1/\varepsilon^{\prime})\leq\eta\cdot\mathrm{OPT}/3-3c\log(n)/\varepsilon^{\prime}$ then
36      Terminate the algorithm
37
38
Algorithm 5 Node-Differentially Private $(1+\eta)$-Approximate Maximum $b$-Matching
8.2 Analysis (Theorem 2.8)
The pseudocode for Theorem 2.8 is presented in Algorithm 5. Following the structure of previous sections, we partition the proof of Theorem 2.8 into its privacy and utility proofs. We first present the privacy proof in Section 8.2.1 and then the approximation proof in Section 8.2.2. Finally, we bring the arguments together in Section 8.2.3 to complete the proof of Theorem 2.8.
8.2.1 Privacy Proof
First, we describe the implicit output and how each bidder reconstructs who they are matched with. Like in the previous algorithm given in [hsu2014private], for each of the $nT$ iterations, the price of each item is outputted, creating a public sequence of prices for each item over the $nT$ iterations. At each iteration, the outputted prices can be used by the bidder to determine their cheapest neighbor. The algorithm also outputs $\text{start}_{u}$ and $\text{end}_{u}$ for each item $u$ at the end of each of the $T$ rounds. This indicates that only bidders which were matched with item $u$ between iteration $\text{start}_{u}$ and $\text{end}_{u}$ are matched with $u$.
Next, we show that the output of the algorithm is $\varepsilon$-node differentially private. On a high level, the proof follows by showing that the algorithm consists of (concurrent) compositions of instances of MAT and the Laplace mechanism. The pseudocode for our algorithm is given in Algorithm 5 which uses a constant $c\geq 1$ which is the constant used in the high probability $1-\frac{1}{n^{c}}$ bound.
Theorem 8.1.
Algorithm 5 is $\varepsilon$-node differentially private.
Proof.
Our algorithm releases the $p_{u}$ and startu of every item for each of the $T$ iterations. Hence, we show that these releases are $\varepsilon$-node differentially private. We first show that Algorithm 5 to Algorithm 5 can be seen as $73/\eta^{2}$ instances of the Multidimensional AboveThreshold mechanism (Algorithm 1) with privacy parameter $\varepsilon^{\prime}$ and queries $c_{u}$ for $u\in V_{L}$. Observe that each bidder $v\in V_{R}$ bids at most $T$ times, once per each of the $T$ iterations. As a result, the $\ell_{1}$ sensitivities of the count queries $c_{u}$ is at most $T$ since one additional bidder will bid on any item at most $T$ times. Furthermore, the noisy threshold $\tilde{t}_{u}$ is used at most $1/\eta$ times since after the prices exceed $1$, they are updated but no longer used. For each update of the threshold, the outputs are $T\varepsilon^{\prime}$-differentially private, so the entirety of the outputs of Algorithm 5 to Algorithm 5 is $T\varepsilon^{\prime}/\eta$-differentially private.
Similarly, for each of the $T$ iterations, we will show that Algorithm 5 to Algorithm 5 can be seen as an instance of the Multidimensional AboveThreshold mechanism with privacy parameter $\varepsilon^{\prime}$ and queries $M_{u}$ for $u\in V_{L}$. Recall that $M_{u}$ is the number of bidders matched with $u$ between $\text{start}_{u}$ and $\text{end}_{u}$. Since each bidder can only match with $1$ item at a time, the sensitivity of the queries is $1$, so each iteration of Algorithm 5 to Algorithm 5 is $\varepsilon^{\prime}$-differentially private. Applying the concurrent composition theorem over $T$ iterations, we have that Algorithm 5 to Algorithm 5is $T\varepsilon^{\prime}$-differentially private.
Finally, for each of the $T$ rounds, we have that Algorithms 5, 5 and 5 can be analyzed as an instance of the Laplace mechanism with privacy parameter $\varepsilon^{\prime}$ and query $c_{0}$. In a given iteration, each bidder makes only a single bid so the sensitivity of the query is $1$, implying that this part is $\varepsilon^{\prime}$-differentially private. Applying the concurrent composition theorem, we have that Algorithms 5, 5 and 5 is $T\varepsilon^{\prime}$-differentially private. Finally, by concurrent composition, the entire algorithm is $T\varepsilon^{\prime}+T\varepsilon^{\prime}+T\varepsilon^{\prime}/\eta\leq\varepsilon$-differentially private. ∎
8.2.2 Utility Proof
On a high level, the utility proof shows that $T$ iterations of the deferred acceptance algorithm suffices to match each bidder with a suitable item. Slightly more formally, we say a bidder is satisfied if they are matched with an $\eta$-approximate favorite good (i.e., $w_{\mu(v),v}-p_{\mu(v)}\geq w_{u,v}-p_{u}-\eta$ for all items $u$ where $\mu(v)$ indicates the item the bidder is matched with). Such a notion has been referred to, in the matching literature, as “happy” bidders [ALT21, LKK23]. We show below that at least a $(1-\eta)$-fraction of the bidders are satisfied, which directly leads to our approximation bound.
Lemma 8.2.
Assume that each Laplace random variable satisfies $|\mathrm{Lap}(\beta)|\leq 3c\beta\log(n)$ and $s\geq 72c\log(n)/(\eta\varepsilon^{\prime})$ for constant $c\geq 1$. Then, we have that at most $\eta\cdot\text{OPT}$ bidders are unsatisfied at termination.
Proof.
Observe that the number of unsatisfied bidders is exactly the number of bidders who were unmatched (Algorithm 5) by their item in the final round. We will first prove the claim when the algorithm terminates early in Algorithm 5. We will then show that the algorithm always terminates early under our assumptions on $s$ and our choice of $T=73/\eta^{2}$. Combining the two gives the desired claim.
If the algorithm terminates early, then we have $c_{0}+\mathrm{Lap}(1/\varepsilon^{\prime})\leq\eta\cdot\text{OPT}/3-3c\log(n)/\varepsilon^{\prime}$ in Algorithm 5, implying that the number of bids at the final round is at most $Q\coloneqq\eta\cdot\text{OPT}/3$ by our assumption that $\mathrm{Lap}(1/\varepsilon^{\prime})\leq 3c\log(n)/\varepsilon^{\prime}$. Let $N$ be the number of items which unmatched (Algorithm 5) with some bidder in this iteration. Since each such item is matched with at least $s-36c\log(n)/\varepsilon^{\prime}$ bidders at the end of the round (due to Algorithm 5), the total number of bidders who matched with these goods at the beginning of the round must be at least
$(s-36c\log(n)/\varepsilon^{\prime})N-Q.$
Next, observe that at most OPT bidders can be matched at the same time, by definition of OPT. Combining with the above inequality, we have that
$N\leq(\text{OPT}+Q)/(s-36c\log(n)/\varepsilon^{\prime}).$
Thus, the total number of bidders who were unmatched is at most
$sN-[(s-36c\log(n)/\varepsilon^{\prime})N-Q]$ $=36cN\log(n)/\varepsilon^{\prime}+Q$
$\leq\frac{36c\log(n)/\varepsilon^{\prime}}{s-36c\log(n)/\varepsilon^{\prime}}\cdot\left(\text{OPT}+Q\right)+Q$
$\leq\frac{36c\log(n)/\varepsilon^{\prime}}{s-36c\log(n)/\varepsilon^{\prime}}\cdot\left(\text{OPT}+\frac{\eta\cdot\text{OPT}}{3}\right)+Q.$
For $s\geq(1+2/\eta)72c\log(n)/\varepsilon^{\prime}=\Theta(\log(n)/\varepsilon\eta^{4})$, the above expression is upper bounded by $\eta\cdot\text{OPT}$, as desired.
If the algorithm doesn’t terminate early, then we have $c_{0}+\mathrm{Lap}(1/\varepsilon^{\prime})>\eta\cdot\text{OPT}/3-3c\log(n)/\varepsilon^{\prime}$ in Algorithm 5 for each iteration of the algorithm. This implies that the number of bids at each of the $T$ iterations of the algorithm is at least $\eta\cdot\text{OPT}/3-36c\log(n)/\varepsilon^{\prime}$. This implies that the total number of bids over all the iterations is lower bounded by
$B\geq T\cdot(Q-36c\log(n)/\varepsilon^{\prime}).$ (5)
As before, we have that at most OPT bidders can be matched at the same time. There are at most OPT bids on the under-demanded goods, since bidders are never unmatched with these goods. Furthermore, each of the over-demanded goods are matched with at least $s-36c\log(n)/\varepsilon^{\prime}$ bidders, so there are at most $\text{OPT}/(s-36c\log(n)/\varepsilon^{\prime})$ bidders. Since each such good takes at most $s+36c\log(n)/\varepsilon^{\prime}$ bids at each of the $1/\eta$ price levels, the total number of bids is thus upper bounded by
$B\leq\text{OPT}+\frac{\text{OPT}}{\eta}\left(\frac{s+36c\log(n)/\varepsilon^{\prime}}{s-36c\log(n)/\varepsilon^{\prime}}\right)\leq\frac{6\cdot\text{OPT}}{\eta},$ (6)
where the second inequality holds since $s\geq 72c\log(n)/\varepsilon^{\prime}$.
Combining our two estimates in Eqs. 5 and 6, we have
$T\cdot(Q-36c\log(n)/\varepsilon^{\prime})\leq B\leq\frac{6\cdot\text{OPT}}{\eta},$
implying that
$T\leq\frac{6\cdot\text{OPT}}{\eta}\cdot\left(\frac{1}{\eta\cdot\text{OPT}/3-36c\log(n)/\varepsilon^{\prime}}\right)\leq\frac{72}{\eta^{2}},$
where we have used that $\eta\cdot\text{OPT}\geq\eta\cdot s\geq 144c\log(n)/\varepsilon^{\prime}$ if there are at least $s$ edges; otherwise, if there are less than $s$ edges, the returned matching will equal the number of edges on the first iteration. Thus, this is a contradiction since $T=\frac{73}{\eta^{2}}$, so we can conclude that the algorithm must terminate early. ∎
8.2.3 Putting it Together
Proof of Theorem 2.8.
First, observe that for any Laplace random variable $\mathrm{Lap}(\beta)$, we have that $\mathrm{Lap}(\beta)\leq 3c\beta\log(n)$ with probability at least $1-1/n^{3c}$. There are $O(n^{2})$ total Laplace random variables in the algorithm, so a union bound implies each of them satisfies the concentration with probability at least $1-1/n^{c}$. We condition on this event for the remainder of the analysis. In particular, we have that at most $s$ bidders are matched with each item $u$, since at most $s$ bidders between $\text{start}_{u}$ and $\text{end}_{u}$ are matched with $u$ in the while loop starting in Algorithm 5. Now, we can start the analysis.
For each edge $(u,v)\in E$, set $w_{u,v}=1$; set $w_{u,v}=0$ otherwise. Let $\mu$ denote the matching (implicitly) output by the algorithm, and consider the optimal matching $\mu^{*}$. For each matched bidder $v$, we have
$w_{\mu(v),v}-p_{\mu(v)}\geq w_{\mu^{*}(v),v}-p_{\mu^{*}(v)}-\eta.$
Since there are at most OPT such bidders (call them $S$), summing the above over all $S$ gives
$\sum_{v\in S}[w_{\mu(v),v}-p_{\mu(v)}]\geq\sum_{v\in S}[w_{\mu^{*}(v),v}-p_{\mu^{*}(v)}]-\eta\cdot\text{OPT}.$
Let $N_{u},N_{u}^{*}$ be the number of times $u$ is respectively matched in $\mu,\mu^{*}$. Then rearranging gives
$\sum_{v\in S}[w_{\mu^{*}(v),v}-w_{\mu(v),v}]\leq\sum_{u\in V_{L}}[p^{*}_{u}\cdot N_{u}^{*}-p_{u}\cdot N_{u}]-\eta\cdot\text{OPT}.$
Next, observe that if a good $u$ has $p_{u}>0$, this means that at least $s-18c\log(n)/\varepsilon^{\prime}$ bidders were (temporarily) matched with it, due to Algorithm 5. This directly implies that the number of bidders matched with $u$ at the termination is at least $s-36c\log(n)/\varepsilon^{\prime}$, because goods only unmatch with bidders until they are matched with $s-36c\log(n)/\varepsilon^{\prime}$ bidders (Line 5). Thus, there can be at most $\text{OPT}/(s-36c\log(n)/\varepsilon^{\prime})$ goods with $p_{u}>0$. For each of these goods, we have $p^{*}_{u}\cdot N_{u}^{*}-p_{u}\cdot N_{u}\leq s$, so we have
$\sum_{v\in S}[w_{\mu^{*}(v),v}-w_{\mu(v),v}]\leq\frac{\text{OPT}\cdot s}{s-36c\log(n)/\varepsilon^{\prime}}-\eta\cdot\text{OPT}.$
Finally, Lemma 8.2 implies that at most $\eta\cdot\text{OPT}$ bidders are not in $S$. Summing over all bidders, we have
$\sum_{v\in V_{R}}[w_{\mu^{*}(v),v}-w_{\mu(v),v}]\leq\frac{\text{OPT}\cdot s}{s-36c\log(n)/\varepsilon^{\prime}}+\eta\cdot\text{OPT}-\eta\cdot\text{OPT}.$
Since we have $s\geq(1+2/\eta)\cdot 72c\log(n)/\varepsilon^{\prime}$, the first term on the right hand side is at most $2\eta\cdot\text{OPT}$. Scaling down $\eta$ by a constant factor $1/2$ gives the desired result with probability at least $1-\frac{1}{n^{c}}$ for constant $c\geq 1$. ∎
9 Continual Release of Implicit Matchings
We now give algorithms for matchings in the continual release model. We give algorithms that satisfy edge-DP and node-DP in two different types of input streams. Specifically, we consider two types of insertion-only streams where nodes and edges can be inserted but not deleted: 1) arbitrary-order edge arrival streams under both edge and node-DP and 2) adjacency-list order streams.
Edge-DP for Arbitrary Edge Order Streams.
We first focus on arbitrary-order edge arrival streams, where edge insertions arrive in an arbitrary order, and show that our edge and node-DP implicit matching algorithms can be implemented in this model. To avoid redundancy, we present the edge-DP algorithm (Theorem 2.9), which is essentially a lazy-update version of our static algorithm (Algorithm 2), and then briefly discuss the minor changes needed to obtain a node-DP algorithm (Theorem 2.10).
See 2.9
As a special case, we recover the following result for matchings.
Corollary 9.1.
For $\varrho\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with probability $1-\frac{1}{n^{c}}$, each $b$-matching is a $\left(2+\varrho,O\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximation with respect to the maximum 1-matching size.
Similar to Theorem 5.2, we can remove the need for a bicriteria approximation when the degree bound $b$ is slightly higher.
Theorem 9.2.
For $\eta\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching is a $\left(2+\eta,\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum $b$-matching.
Node-DP for Arbitrary Edge Order Streams.
Implementing the node-DP arboricity sparsifier of Section 7.1 in the continual release model leads to Theorem 2.10, our main continual release node-DP result restated below.
See 2.10 Similar to the edge-DP case, the case of $b^{\prime}=1$ leads to a special case with approximation guarantees with respect to the maximum 1-matching size.
Corollary 9.3.
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, $\alpha$ be the arboricity of the input graph, and $b=\Omega\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{2}\varepsilon^{2}}\right)$. There is an $\varepsilon$-node DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching has the size of a $\left(2+\eta,O\left(\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum $1$-matching.
Adjacency-List Order Streams.
In the adjacency-list order insertion streams, nodes arrive in an arbitrary order, and once a node arrives, all edges adjacent to the node arrive in an arbitrary order. Remark that in the adjacency-list order stream, all edges arrive twice, once per endpoint in the stream. See 2.11
Organization.
In Section 9.1, we present our $\varepsilon$-edge DP arbitrary edge-order continual release algorithm as well as the corollaries in Section 9.1.3. Then, we achieve stronger $\varepsilon$-node DP guarantees in Section 9.2. Finally, we present the details of our adjacency-list continual release algorithm in Section 9.3.
9.1 Edge-DP Arbitrary Edge-Order Continual Release Implicit $b$-Matchings
We begin with the pseudocode of the $\varepsilon$-edge DP algorithm in Algorithm 6. Then, we prove privacy and utility separately in Sections 9.1.1 and 9.1.2, respectively. The proof of privacy is non-trivial since we cannot simply apply composition. Instead, we directly argue by the definition of privacy.
Our results use a variant of the standard SVT tool that allows us to answer “above” $c$ times, which we recall in Section A.3 for completeness.
1
Input: Arbitrary edge-order stream $S$, privacy parameter $\varepsilon>0$, approximation parameter $\varrho\in(0,1]$,
Output: An $\varepsilon$-node differentially private implicit $b$-matching after each time stamp.
2 $j_{1}\leftarrow 0$
3 $c\leftarrow\log_{1+\varrho}(n)$
4 Initialize class $\textsc{EstimateSVT}\leftarrow\operatorname{SVT}(\nicefrac{{\varepsilon}}{{3}},1,c)$   (Algorithm 9)
5
61exfor edge $e_{t}\in S$ do
7
  1ex// $\nu(G_{t})$ denotes the size of the maximum matching of the dynamic graph $G_{t}$ at time $t$.
[…]
9    $j_{t}\leftarrow j_{t}+1$
10
11
12  1exif $t=1$ or $j_{t}>j_{t-1}$ then
13    $\operatorname{solution}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP implicit $b$-matching using Algorithm 2 on input $G_{t}$ (Theorem 2.2)
14    $\operatorname{estimate}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP estimate of the size of the current maximum $b^{\prime}$-matching
15
[…]
17   Output solution, estimate
Algorithm 6 Arbitrary Edge-Order Continual Release Edge-DP Implicit $b$-Matching
9.1.1 Privacy Proof
Lemma 9.4.
Algorithm 6 is $\varepsilon$-edge DP.
Proof of Lemma 9.4.
Let $(j_{t},\operatorname{solution}_{t},\operatorname{estimate}_{t})_{t=1}^{T}$ denote the random sequence of outputs from Algorithm 6 corresponding to the SVT estimate of the maximum matching in the current graph $G_{t}$, the implicit $b$-matching solution for timestamp $t$, and the estimate of the largest matching of $G_{t}$ contained within $\operatorname{solution}_{t}$. Similarly, let $(\widetilde{j}_{t},\widetilde{\operatorname{solution}}_{t},\widetilde{\operatorname{estimate}}_{t})_{t=1}^{T}$ be the random sequence of outputs on an edge-neighboring stream. We may assume that $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$.
Fix any deterministic sequence $(J_{t},B_{t},M_{t})_{t=1}^{T}$ of possible outputs. We have
$\Pr\left[(j_{t},\operatorname{solution}_{t},\operatorname{estimate}_{t})_{t=1}^{T}=(J_{t},B_{t},M_{t})_{t=1}^{T}\right]$
$=\prod_{t=1}^{T}\Pr[j_{t}=J_{t}\mid<t]\cdot\Pr[\operatorname{solution}_{t}=B_{t}\mid j_{t}=J_{t},<t]$
$\qquad\cdot\Pr[\operatorname{estimate}_{t}=M_{t}\mid\operatorname{solution}_{t}=B_{t},j_{t}=J_{t},<t]\,.$
Here the notation <t<t is a shorthand that denotes the event that $(j_{\tau},\operatorname{solution}_{\tau},\operatorname{estimate}_{\tau})_{\tau=1}^{t-1}=(J_{\tau},B_{\tau},M_{\tau})_{\tau=1}^{t-1}$ for $t\geq 2$ and the trivial event of probability $1$ if $t=1$. Thus our goal is to bound the product of ratios
$\left(\prod_{t=1}^{T}\frac{\Pr[j_{t}=J_{t}\mid<t]}{\Pr[\tilde{j}_{t}=J_{t}\mid<t]}\right)\cdot\left(\prod_{t=1}^{T}\frac{\Pr[\operatorname{solution}_{t}=B_{t}\mid j_{t}=J_{t},<t]}{\Pr[\tilde{\operatorname{solution}}_{t}=B_{t}\mid\tilde{j}_{t}=J_{t},\tilde{<}t]}\right)$
$\qquad\cdot\left(\prod_{t=1}^{T}\frac{\Pr[\operatorname{estimate}_{t}=M_{t}\mid\operatorname{solution}_{t}=B_{t},j_{t}=J_{t},<t]}{\Pr[\tilde{\operatorname{estimate}}_{t}=M_{t}\mid\tilde{\operatorname{solution}}_{t}=B_{t},\tilde{j}_{t}=J_{t},\tilde{<}t]}\right)\,.$
Now, conditioned on the event that $j_{t-1}=J_{t-1}$, it holds that $j_{t}$ is independent of $\operatorname{solution}_{t-1}$ and $\operatorname{estimate}_{t-1}$. Hence the first product is at most $e^{\varepsilon/3}$ by the privacy guarantees of the SVT (Theorem A.3). For the second and third products, remark that given $j_{t}=J_{t-1}=\tilde{j}_{t}$, then $\operatorname{solution}_{t}=B_{t-1}=\widetilde{\operatorname{solution}}_{t}$ with probability $1$ and similarly for $\operatorname{estimate}_{t}$. Since $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$, at most $c=a\log(n)/\eta$ of the ratios from the second and third products are not 1. We can bound each of these ratios by $e^{\varepsilon/3c}$ using the individual privacy guarantees of Algorithm 2 (Corollary 5.1) and the Laplace mechanism. This concludes the proof. ∎
9.1.2 Utility Proof
Lemma 9.5.
Let $\varepsilon\in(0,1)$, $\varrho,\eta\in(0,1)$, and $b\geq\frac{(1+\eta)^{2}}{1-\eta}\cdot b^{\prime}+O\left(\frac{\log^{2}(n)}{\varrho\eta^{2}\varepsilon}\right)$. Algorithm 6 outputs a sequence of implicit $b$-matchings with the following guarantee with probability at least $1-1/\operatorname{poly}(n)$: each implicit solution contains a $\left(2+\varrho,O\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximate maximum matching.
Proof.
The bound on $b$ follow directly from Theorem 2.2 and the fact that we require each call to Algorithm 2 to satisfy ($\nicefrac{{\varepsilon}}{{3c}}$)-privacy for $c=\log_{1+\varrho}(n)$.
To see the approximation guarantee, we first observe that this certainly holds at timestamps when $\operatorname{solution}$ is updated. By the guarantees of SVT, $(1+\varrho)^{j_{t}}$ is a $(1+\varrho,O(\log^{2}(n)/\varepsilon\varrho))$-approximate estimate of size of the current maximum matching. Hence we incur at most this much error in between updates. ∎
Combining Lemma 9.4 and Lemma 9.5 yields the proof of Corollary 9.1. The pseudocode for Corollary 9.1 is given in Algorithm 6.
9.1.3 Corollaries
As a special case of our general edge-DP algorithm (Theorem 2.9) when $\eta=\nicefrac{{1}}{{2}}$ and $b^{\prime}=1$, we recover the guarantees stated in Corollary 9.1.
In order to obtain a non-bicriteria approximation when $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\varepsilon}\right)$, we use the same technique as the static edge-DP algorithm of executing the general algorithm (Theorem 2.9) with $b^{\prime}=\frac{b}{1+O(\eta)}$. The proof is identical to that of Corollary 5.7. This yields a proof of Theorem 9.2.
9.2 Node-DP Arbitrary Edge-Order Continual Release Implicit $b$-Matchings
We now describe the necessary modifications to adapt our edge DP continual release algorithm to the challenging node DP setting. Similary to Section 7.1, we first present an algorithm assuming a public bound $\tilde{\alpha}$ on the arboricity.
Theorem 9.6.
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, Algorithm 7 is an $\varepsilon$-node DP algorithm in the arbitrary order continual release model that outputs implicit $b$-matchings. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
The main task is to implement the arboricity matching sparsifier (Theorem 7.3) in an arbitrary edge-order stream. Note that there is a natural total order of vertex pairs given by the arrival order of edges. Thus we can simply discard edges where one of its endpoints has already seen more than some threshold $\Lambda$ incident edges (Algorithm 7 of Algorithm 7).
Let the sparsified graph $H$ be obtained from the input graph $G$ as described above. Write $c=O(\log(n)/\eta)$ be the maximum SVT budget for “above” queries (Algorithm 9). Our algorithm proceeds as following for each edge: Similar to Algorithm 6 of Algorithm 6, we use an SVT instance to check when to update our solution. The only change is that $b^{\prime}$-matching sensitivity is $b^{\prime}$ for node-neighboring graphs. At timestamps where we must update the solution, we run a static $O(\nicefrac{{\varepsilon}}{{c\Lambda}})$-edge DP implicit matching algorithm (Algorithm 2) on the sparsified graph $H$. This is similar to Algorithm 6 of Algorithm 7 except we need to execute the underlying algorithm with a smaller privacy parameter. Finally, we also estimate the size of the current maximum matching. Once again, this step is unchanged from Algorithm 6 of Algorithm 6, with the slight exception that the sensitivity is now $b^{\prime}$. In particular, we note that we only use the sparsified graph $H$ for updating the solution and not for estimating the current maximum matching size.
Input: Arbitrary edge-order stream $S$, approximation parameter $\eta\in(0,1]$, public bound $\tilde{\alpha}>0$.
1 $\Lambda\leftarrow 5(1+\nicefrac{{5}}{{\eta}})2\tilde{\alpha}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}+(b^{\prime}-1)}$
2 $H\leftarrow(V,\varnothing)$
3 $d_{v}\leftarrow 0$ for each vertex $v\in V$
4 $c\leftarrow\log_{1+\eta}(n)$
5 for edge $e_{t}\in S$ do
6 if $\perp\neq e_{t}=\{u,v\}$ and $\max(d_{u},d_{v})<\Lambda$ then
7      Increment $d_{u},d_{v}$
8    $E[H]\leftarrow E[H]\cup\{e_{t}\}$
9
10
  1ex// Check if $\nu(G_{t})$ has significantly increased using SVT with sensitivity $b^{\prime}$ and total privacy budget $\nicefrac{{\varepsilon}}{{3}}$ (Algorithm 6 of Algorithm 6)
 // If SVT is ‘‘above’’, compute $\operatorname{solution}$ with respect to $H_{t}$ with privacy budget $\nicefrac{{\varepsilon}}{{6c\Lambda}}$ (Algorithm 6 of Algorithm 6)
 // If SVT is ‘‘above’’, compute $\operatorname{estimate}$ with respect to $G_{t}$ with privacy budget $\nicefrac{{\varepsilon}}{{3c}}$ (Algorithm 6 of Algorithm 6)
11   Output $\operatorname{solution},\operatorname{estimate}$
Algorithm 7 Arbitrary Edge-Order Continual Release Node-DP $b$-Matching Algorithm
We sketch the proof of guarantees for Algorithm 7.
Sketch Proof of Theorem 9.6.
The privacy proof is identical to that of Theorem 2.9. The utility guarantees follow similarly, with the exception that we use the approximation guarantees of Theorem 2.6 rather than Theorem 2.2. ∎
Public Arboricity Bound.
Finally, we describe the modifications from Algorithm 7 to remove the assumption on a public bound $\tilde{\alpha}$. First, we compute $O(\log(n))$ sparsified graphs $H_{k}$ corresponding to setting the public parameter $\tilde{\alpha}=2^{k},k=1,2,\dots,\lceil\log_{2}(n)\rceil$. The SVT to estimate the current maximum matching size remains the same (we do not compute an estimate for each $k$). Next, we compute $O(\log(n))$ implicit $b$-matchings, one for each sparsified graph $H_{k}$. We also privately compute an upper bound $\tilde{\alpha}_{t}$ of the current arboricity of $G_{t}$ and choose the implicit solution with index $k^{\star}$ such that $2^{k^{\star}-1}<\tilde{\alpha}_{t}\leq 2^{k^{\star}}$. Accounting for the error in estimating the arboricity yields a proof of Theorem 2.10.
9.3 Adjacency-List Order Continual Release Implicit Matchings
In this section, we give $\varepsilon$-DP algorithms for implicit matching in the continual release model with adjacency-list order streams. Our results for another stream type (arbitrary edge-order streams) are presented in Section 9.1. The adjacency-list order stream ensures each node that arrives in this model will be followed by its edges where the edges arrive in an arbitrary order. Our algorithm is a straightforward implementation of Algorithm 2 in the continual release model.
In particular, the nodes arrive in an arbitrary order and when a node arrives, it waits until all of its edges arrive and then performs the same proposal and response procedure as given in Algorithm 2. Since each node can contribute at most $1$ to the matching size, we have an additional additive error of $1$ in the maximal matching size in the continual release setting.
Detailed Algorithm.
We give our modified adjacency-list continual release algorithm in Algorithm 8. This algorithm takes a stream $S$ of updates consisting of node insertions and edge insertions. The $i$-th update in $S$ is denoted $u_{i}$ and it can either be a node update $v$ or an edge update $e_{i}$. In adjacency-list order streams, each node update is guaranteed to be followed by all edges adjacent to it; these edges arrive in an arbitrary order. The high level idea of our algorithm is for each node $v$ to implicitly announce the nodes it is matched to after we have seen all of the edges adjacent to it. Since each node can add at most $1$ additional edge to any matching, waiting for all edges to arrive for each node update will incur an additive error of at most $1$. After we have seen all edges adjacent to $v$ (more precisely, when we see the next node update), we run the exact same proposal procedure as given in Algorithm 2.
We first set the parameters used in our algorithm the same way that the parameters were set in Algorithm 2 in Algorithms 8 and 8. Then, we iterate through all nodes (Algorithm 8) to determine the noisy cutoff $\tilde{b}(u)$ for each node $u$ (Algorithm 8). Then, we initiate the set $K$ (initially empty) to be the set of adjacent edges for the most recent node update (Algorithm 8). The most recent node update is stored in $w$ (Algorithm 8). Then, $Q$ (initially an empty set) stores all nodes that we have seen so far; that is, $Q$ is used to determine whether an edge adjacent to the most recent node update $w$ is adjacent to a node that appears earlier in the stream (Algorithm 8).
For each update as it appears in the order of the stream (Algorithm 8), we first iterate through each subgraph index (Algorithm 8) to flip a coin with appropriate probability to determine whether the edge is included in the subgraph with index $r$. This procedure is equivalent to Algorithms 2 and 2 in Algorithm 2. Then, we check whether the update is a node update (Algorithm 8). If it is a node update, then we add $v$ to $Q$ (Algorithm 8), and if $w\neq\bot$, there was a previous node update stored in $w$ (Algorithm 8) and we process this node. We first check all nodes to see whether they have reached their matching capacity (Algorithms 8, 8, 8 and 8) using an identical procedure to Algorithms 2, 2, 2 and 2 in Algorithm 2.
Then, if $w$ has not satisfied the matching condition, then we iterate through all subgraph indices to find the smallest index that does not cause the matching for $w$ to exceed $b$. This procedure (Algorithms 8, 8, 8, 8, 8, 8 and 8) is identical to Algorithms 2, 2, 2, 2, 2, 2 and 2 in Algorithm 2 except for Algorithm 8. The only difference between Algorithm 8 in Algorithm 8 and Algorithm 2 in Algorithm 2 is that we check whether $v$ appears earlier in the ordering by checking $v\not\in Q$ and we check the coin flip for edge $\{w,v\}$ by checking $c(i,r)=\textsc{Heads}$. Thus these two lines are functionally identical. After we have updated the matching with node $w$’s match, we set $K$ to empty (Algorithm 8) and update $w$ to the new node $v$ (Algorithm 8).
Finally, if $u_{i}$ is instead an edge update $e_{i}$ (Algorithm 8), we add $e_{i}$ to $K$ to maintain the adjacency list of the most recent node update (Algorithm 8).
1
Input: Arbitrary order adjacency-list stream $S$, privacy parameter $\varepsilon>0$, matching parameter $b=\Omega(\log{n}/\varepsilon)$.
Output: An $\varepsilon$-locally edge differentially private implicit $b$-matching.
2
[…]
4 $\varepsilon^{\prime}\leftarrow\varepsilon/(2+1/\eta^{\prime})$
5
6for each node $u\in[n]$ in order do
7 $\tilde{b}(u)\leftarrow b-20\log(n)/\varepsilon^{\prime}+\text{Lap}(4/\varepsilon^{\prime})$
8
[…]
10 $w\leftarrow\bot$
11 $Q\leftarrow\varnothing$
12 for every update $u_{i}\in S$ do
13 for each subgraph index $r=0,\ldots,\lceil\log_{1+\eta^{\prime}}(n)\rceil$ do
14      Flip and release coin $c(i,r)$ which lands heads with probability $p_{r}=(1+\eta^{\prime})^{-r}$
15
16 if $u_{i}$ is a node update of node $v$ then
17    $Q\leftarrow Q\cup\{v\}$
18    if $w\neq\bot$ then
19       for each node $u\in V$ which has not satisfied the matching condition do
20          $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime})$
21          if $M_{i}(u)+\nu_{i}(u)\geq\tilde{b}(u)$ then
22               Output to transcript: node $u$ has reached their matching capacity
23
24
25       if $w$ has not satisfied the matching condition then
26          for subgraph index $r=0,\ldots,\lceil\log_{1+\eta^{\prime}}(n)\rceil$ do
27             $W_{r}(w)=\{v:v\text{ active}\wedge v\in e_{i}\text{ where }e_{i}\in K\wedge v\not\in Q\wedge c(i,r)=\textsc{Heads}\}$
28             $\widetilde{W}_{r}(v)=|W_{r}(v)|+\text{Lap}(2/\varepsilon^{\prime})$
29          $\widetilde{M}_{i}(v)=M_{i}(v)+\text{Lap}(2/\varepsilon^{\prime})$
30          $v$ computes smallest $r$ so that $\widetilde{M}_{i}(v)+\widetilde{W}_{r}(v)+c\log(n)/\varepsilon^{\prime}\leq b$, and matches with neighbors in $W_{r}$
31          $v$ releases $r$
32
33    $K\leftarrow\varnothing$
34    $w\leftarrow v$
35
36 if $u_{i}$ is an edge update $e_{i}$ then
37    $K\leftarrow K\cup\{e_{i}\}$
38
Algorithm 8 Adjacency-List Order Continual Release Matching
9.3.1 Proof of Theorem 2.11
Proof.
We first prove that our continual release algorithm is $\varepsilon$-DP on the vector of outputs. First, our algorithm only outputs a new output each time a node update arrives. For every edge update, the algorithm outputs the same outputs as the last time the algorithm outputted a new output for a node update. Algorithm 8 implements all of the local randomizers used in Algorithm 2 in the following way. The order of the nodes that propose is given by the order of the node updates. Coins for edges are flipped in the same way as in Algorithm 2, the threshold for the matching condition is determined via an identical procedure, and finally, the proposal process is identical to Algorithm 2. Hence, the same set of local randomizers can be implemented as in Algorithm 2 and the continual release algorithm is $\varepsilon$-DP via concurrent composition.
The approximation proof also follows from the approximation guarantee for Algorithm 2 since the proposal procedure is identical except for the additive error. In the continual release model, there is an additive error of at most $1$ since for every edge update after a new node update, the node could be matched to an initial new edge update but is not matched until the final edge update is shown for that node. Since each node contributes at most $1$ to a matching, the additive error is at most $1$ for each update. ∎
Acknowledgments
Felix Zhou acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). Quanquan C. Liu and Felix Zhou are supported by a Google Academic Research Award and NSF Grant #CCF-2453323.
Appendix A Deferred Preliminaries
A.1 Concentration Inequalities
Lemma A.1.
Given a random variable $X\sim Lap(b)$ drawn from a Laplace distribution with expectation 0, the probability $|X|>c\ln(n)$ is $n^{-\frac{c}{b}}$.
Theorem A.2 (Multiplicative Chernoff Bound).
Let $X=\sum_{i=1}^{n}X_{i}$ where each $X_{i}$ is a Bernoulli variable which takes value $1$ with probability $p_{i}$ and 0 with probability $1-p_{i}$. Let $\mu=\mathbb{E}[X]=\sum_{i=1}^{n}p_{i}$. Then, it holds:
  1. 1.
    Upper Tail: $\mathsf{Pr}(X\geq(1+\psi)\cdot\mu)\leq\exp\left(-\frac{\psi^{2}\mu}{2+\psi}\right)$ for all $\psi>0$;
  2. 2.
    Lower Tail: $\mathsf{Pr}(X\leq(1-\psi)\cdot\mu)\leq\exp\left(-\frac{\psi^{2}\mu}{3}\right)$ for all $0<\psi<1$.
A.2 Proof of Lemma 3.15
We now restate and prove Lemma 3.15. See 3.15
Proof.
Let $G_{1},G_{2}$ be neighboring graphs, and we will let $\Pr[\mathcal{M}(G_{1})=z]$, $\Pr[\mathcal{M}(G_{2})=z]$ denote the density functions of $\mathcal{M}(G_{1})$, $\mathcal{M}(G_{2})$ evaluated at $z\in\mathbb{R}^{k}$, by some abuse of notation. To prove $\epsilon$-differential privacy, we need to show that the ratio of $\Pr[\mathcal{M}(G_{i})=z]$ is upper bounded by $\exp(\epsilon)$, for any $z\in\mathbb{R}^{k}$.
First, we define some more notation. Let $\mathcal{M}_{i}(G_{1})$, $\mathcal{M}_{i}(G_{2})$ denote the output of the mechanism $\mathcal{M}$ on graphs $G_{1}$, $G_{2}$ when answering the $i^{th}$ (adaptive) query. Via some more abuse of notation, let $\Pr[\mathcal{M}_{i}(G_{1})=z_{i}|\mathcal{M}_{j}(G_{1})=z_{j}\text{ for }j\in[i-1]]$ and $\Pr[\mathcal{M}_{i}(G_{2})=z_{i}|\mathcal{M}_{j}(G_{2})=z_{j}\text{ for }j\in[i-1]]$ denote the conditional density functions of $\mathcal{M}_{i}(G_{1})$ and $\mathcal{M}_{i}(G_{2})$ evaluated at $z_{i}\in\mathbb{R}$, conditioned on the events $\mathcal{M}_{j}(G_{1})=z_{j}$ and $\mathcal{M}_{j}(G_{2})=z_{j}$ for $j=1,\ldots,i-1$.
Finally, fix $z\in\mathbb{R}^{k}$. We have the following:
$\frac{\Pr[\mathcal{M}(G_{1})=z]}{\Pr[\mathcal{M}(G_{2})=z]}$ $=\frac{\prod_{i=1}^{k}\Pr[\mathcal{M}_{i}(G_{1})=z_{i}|\mathcal{M}_{j}(G_{1})=z_{j}\text{ for }j\in[i-1]]}{\prod_{i=1}^{k}\Pr[\mathcal{M}_{i}(G_{1})=z_{i}|\mathcal{M}_{j}(G_{1})=z_{j}\text{ for }j\in[i-1]]}$ (7)
$=\frac{\prod_{i=1}^{k}\exp\left(-\frac{\epsilon|f_{i}(G_{1})-z_{i}|}{\Delta}\right)}{\prod_{i=1}^{k}\exp\left(-\frac{\epsilon|f_{i}(G_{2})-z_{i}|}{\Delta}\right)}$ (8)
$=\prod_{i=1}^{k}\exp\left(-\frac{\epsilon(|f_{i}(G_{1})-z_{i}|-|f_{i}(G_{2})-z_{i}|)}{\Delta}\right)$
$\leq\prod_{i=1}^{k}\exp\left(-\frac{\epsilon|f_{i}(G_{1})-f_{i}(G_{2})|}{\Delta}\right)$ (9)
$=\exp\left(-\frac{\epsilon\sum_{i=1}^{k}|f_{i}(G_{1})-f_{i}(G_{2})|}{\Delta}\right)$
$=\exp\left(-\frac{\epsilon\|f(G_{1})-f(G_{2})\|_{1}}{\Delta}\right)$
$\leq\exp(\epsilon).$ (10)
In the above, equality (7) follows by the chain rule for condition probabilities, equality (8) is just writing out the density function of the Laplace distribution since we are conditioning on the answers $\mathcal{M}_{j}(G_{1})$ and $\mathcal{M}_{j}(G_{2})$ for the previous queries, inequality (9) follows by the triangle inequality, and inequality (10) follows since $\Delta$ is the $\ell_{1}$-sensitivity of $f$. ∎
A.3 Multi-Response Sparse Vector Technique
We recall a variant of the sparse vector technique that enables multiple responses as introduced by [lyu2017understanding]. Let $D$ be an arbitrary (graph) dataset, $(f_{t},\tau_{t})$ a sequence of (possibly adaptive) query-threshold pairs, $\Delta$ an upper bound on the maximum sensitivity of all queries $f_{t}$, and an upper bound $c$ on the maximum number of queries to be answered “above”. Typically, the AboveThreshold algorithm stops running at the first instance of the input exceeding the threshold, but we use the variant where the input can exceed the threshold at most $c$ times where $c$ is a parameter passed into the function.
We use the class $\textsc{SVT}(\varepsilon,\Delta,c)$ (Algorithm 9) where $\varepsilon$ is our privacy parameter, $\Delta$ is an upper bound on the maximum sensitivity of incoming queries, and $c$ is the maximum number of “above” queries we can make. The class provides a $\textsc{ProcessQuery}(query,threshold)$ function where queryquery is the query to SVT and thresholdthreshold is the threshold that we wish to check whether the query exceeds.
Theorem A.3 (Theorem 2 in [lyu2017understanding]).
Algorithm 9 is $\varepsilon$-DP.
We remark that the version of SVT we employ (Algorithm 9) does not require us to resample the noise for the thresholds (Algorithm 9) after each query but we do need to resample the noise (Algorithm 9) for the queries after each query.
1Input: privacy budget $\varepsilon$, upper bound on query sensitivity $\Delta$, maximum allowed “above” answers $c$
2 Class SVT($\varepsilon,\Delta,c$)
3 $\varepsilon_{1},\varepsilon_{2}\leftarrow\nicefrac{{\varepsilon}}{{2}}$
4 $\rho\leftarrow\text{Lap}(\nicefrac{{\Delta}}{{\varepsilon_{1}}})$
5 $\operatorname{count}\leftarrow 0$
6
[…]
8    if $\operatorname{count}>c$ then
9       return “abort”
10
11    if $f_{t}(D)+\text{Lap}({2c\Delta}/{\varepsilon_{2}})\geq\tau_{t}+\rho$ then
12       return “above”
13       $\operatorname{count}\leftarrow\operatorname{count}+1$
14
15    else
16       return “below”
17
18
19
Algorithm 9 Sparse Vector Technique